### v3

# Task Aware Dreamer for Task Generalization in Reinforcement Learning

Abstract
A long-standing goal of reinforcement learning is to acquire agents that can learn on training tasks and generalize well on unseen tasks that may share a similar dynamic but with different reward functions. The ability to generalize across tasks is important as it determines an agent’s adaptability to real-world scenarios where reward mechanisms might vary. In this work, we first show that training a general world model can utilize similar structures in these tasks and help train more generalizable agents. Extending world models into the task generalization setting, we introduce a novel method named Task Aware Dreamer (TAD), which integrates reward-informed features to identify consistent latent characteristics across tasks. Within TAD, we compute the variational lower bound of sample data log-likelihood, which introduces a new term designed to differentiate tasks using their states, as the optimization objective of our reward-informed world models. To demonstrate the advantages of the reward-informed policy in TAD, we introduce a new metric called Task Distribution Relevance (TDR) which quantitatively measures the relevance of different tasks. For tasks exhibiting a high TDR, i.e., the tasks differ significantly, we illustrate that Markovian policies struggle to distinguish them, thus it is necessary to utilize reward-informed policies in TAD. Extensive experiments in both image-based and state-based tasks show that TAD can significantly improve the performance of handling different tasks simultaneously, especially for those with high TDR, and display a strong generalization ability to unseen tasks.
(a) Overview of this work
(b) Probabilistic graphical model designs
Figure 1: (a) An overview. Given a task distribution, we train the agent in training tasks and hope it to zero-shot generalize to testing tasks. For improving the generalization, we propose TAD, which utilizes $\Pi_{3}$ to encode all historical information for inferring the current task and novel reward-informed world models for capturing invariant latent features. (b) Probabilistic graphical model designs for the single-task setting (left) and the task-distribution setting (right). The latter inspires the design of reward-informed world models. Solid and dashed lines represent the generative process and the inference model, respectively.
1 Introduction
Deep Reinforcement Learning (DRL) has demonstrated significant advancements in diverse fields Mnih et al. (2016); Hafner et al. (2019a) , and a key factor in these accomplishments is an agent’s proficiency in assimilating lessons from dedicated training tasks. However, when the objective transitions from optimizing performance in specific tasks to achieving broad generalization across related tasks, new challenges emerge. One primary source of these challenges is the congruence in the underlying dynamics of different tasks, which exhibit distinct reward structures. To illustrate, two tasks might involve navigating a robot, but in one, the robot is rewarded for speed, while in the other, it is rewarded for caution. This tendency towards specialization harms the wider applicability of DRL in real-world scenarios. Developing generalizable agents that recognize and handle these subtle variations continues to be an area of keen interest and exploration in DRL.
For designing such generalizable agents, a promising pathway is to train a general world model Ha and Schmidhuber (2018); Hafner et al. (2019b) that helps the agent understand the world and make the decision. In this research, we first provide insight into how a unified dynamic model can help improve the sample efficiency in handling the distribution of tasks, especially by utilizing similar structures of trajectories sampled from different tasks (Theorem 1). Consequently, we propose a novel framework named Task Aware Dreamer (TAD), which extends existing world models into the task generalization setting. As current world models are primarily designed for the single-task setting (left of Fig. 1(b)), we consider the corresponding probabilistic graphical model of the task distribution setting (right of Fig. 1(b)) and extend single-task world model into novel reward-informed world models, which can capture learnable structures in the task distribution. To optimize these models, we compute the variational lower bound of the data log-likelihood while incorporating a novel term that categorizes tasks based on their states, serving as the primary training objective for the reward-informed world models.
Furthermore, we conduct detailed analyses of the components of TAD, which are simple but effective for handling task distribution. First, we point out that the reward-informed policy used in TAD is necessary for task generalization. While prior studies Ghosh et al. (2021) have highlighted that generalizing to unseen environments results in partial observability, rendering all Markovian policies suboptimal, they often overlook the relationship between the task distribution and the expressiveness of the chosen policy hypothesis set. To bridge this gap, we propose a novel metric of Task Distribution Relevance (TDR), which encapsulates the relevance of different tasks within the distribution through their optimal Q functions. Furthermore, we prove that both the commonly adopted hypothesis set $\Pi_{1}$ of Markovian policies and the set $\Pi_{2}$ of policies that encode historical states and actions, are sub-optimal under the task-distribution setting and struggle to generalize. This sub-optimality is proven to be related to TDR (Theorem 3), i.e., for task distributions with high TDR, the performance of these two policy hypotheses might degenerate significantly as they fail to identify the current task, a phenomenon we also demonstrate by experiments (Sec. 5.2). Consequently, the reward-informed policy hypothesis $\Pi_{3}$ utilized in TAD is necessary for handling task generalization. Besides the policy chosen, we also discuss the novel optimization term, i.e., categorizing tasks based on hidden state, in TAD and theoretically show that it can effectively reduce the gap between the policy return and the optimal return (Theorem 4).
Extensive experiments are performed on task distributions based on DeepMind control suite Tassa et al. (2018) and MuJoCo simulator Todorov et al. (2012), which are image-based and state-based respectively. Agents are trained on specific tasks and evaluated for their generalization capabilities on unseen evaluation tasks. Our findings corroborate our analyses, indicating that $\Pi_{1}$ and $\Pi_{2}$ falter in managing task distributions characterized by a high TDR. Contrastingly, TAD excels in discerning and simultaneously managing varied tasks, significantly surpassing benchmark performance. This underscores TAD’s impressive sample efficiency and adaptability in both image-based and state-based contexts. Additionally, our ablation studies spotlight TAD’s versatility, showcasing its prowess in dynamic generalization and handling cross-embodiment tasks. In summary, our contributions are:
  • We present a theoretical observation that a general world model can utilize similar structures of different tasks and improve the sample efficiency for training a generalizable agent for the task distribution.
  • Extending existing single-task world models into the task distribution setting, we then formulate a new algorithm, termed TAD, leveraging the variational lower bound for calculations and showcasing superior generalization capabilities in contrast to baselines.
  • To analyze TAD’s components, we analyze policy generalization across task distributions and introduce a new metric TDR, which quantifies the distribution relevance. Our findings indicate that Markovian policies tend to be sub-optimal, with performance gaps correlating with TDR. Further, we demonstrate that the new optimization term in TAD is effective for task generalization.
2 Related Work
2.1 Generalization in Reinforcement Learning
Contemporary RL methodologies consistently struggle when tasked with generalizing to new environments Song et al. (2019). Prior studies have crafted an array of training methodologies, such as loss regularization Cobbe et al. (2019); Wang et al. (2020), successor representation Touati and Ollivier (2021); Touati et al. (2022), network architecture design Lee et al. (2019); Raileanu and Fergus (2021), and data augmentation Raileanu et al. (2021); Hansen and Wang (2021). Despite this, there are few investigations into the mechanism behind policy generalization in unseen environments and its relationship with the distribution of those environments. Ghosh et al. took the initial step in demonstrating that generalizing to unseen environments introduces partial observability, thereby rendering deterministic Markovian policies sub-optimal. Also, some studies Lee et al. (2020); Ghosh et al. (2021) have experimentally indicated that stochastic or non-Markovian policies can yield superior generalization compared to deterministic Markovian policies. Nevertheless, expressive capabilities of differing hypothesis sets and their connection to the environment distribution remain understudied areas. This knowledge gap is significant for understanding current algorithms and the design of more robust algorithms with enhanced generalization capacities.
There are several topics closely related to generalization, like multi-task RL and meta RL. Multi-task RL Yang et al. (2020); Sodhani et al. (2021); Lee et al. (2022) primarily aims for the trained agent to excel across all training tasks. For boosting the generalization, Meta RL seeks to enable the trained agents to adapt to new tasks with as little data as possible, including gradient-based Finn et al. (2017) and context-based methods Duan et al. (2016); Rakelly et al. (2019). There are also some model-based methods Nagabandi et al. (2018); Clavera et al. (2018); Zhao et al. (2021) utilize learned models to boost the sample efficiency. Though some context-based methods like VariBAD Zintgraf et al. (2019) show zero-shot generalization ability, it is still significant to analyze the generalization in RL and design corresponding algorithms.
2.2 World Models
The concept of world models Ha and Schmidhuber (2018) aims to better learn environmental representations, which has potential advantages for generalization as it can capture invariant features across tasks. Classical methods utilize the Recurrent State Space Model (RSSM) Hafner et al. (2019b) for planning like PlaNet Hafner et al. (2019b) and policy learning like Dreamer Hafner et al. (2019a, 2020, 2023). Subsequent research explored reconstruction-free world models Deng et al. (2022), preserving representational capacity by avoiding reconstructing task-irrelevant elements. Also, techniques like temporal predictive coding Nguyen et al. (2021), cooperative reconstruction Fu et al. (2021), and Denoised MDP Wang et al. (2022) have been proposed for more effective task-relevant information encoding. Recent works Chen et al. (2021); Robine et al. (2023) integrate Transformer Vaswani et al. (2017) into world models for capturing long-term dependencies. Other studies propose using world models to extract environmental invariant features by learning from videos Seo et al. (2022) or exploration Sekar et al. (2020); Xu et al. (2022), and then fine-tuning to new tasks. However, most world models are designed for the single-task setting and struggle to manage multiple tasks without fine-tuning, limiting their effectiveness for zero-shot generalization to unseen tasks.
3 Preliminary
We consider the setting with a task distribution $\mathcal{T}$ of MDPs, where different tasks own the same dynamic and different rewards. Formally, each MDP $\mathcal{M}\sim\mathcal{T}$ is represented as $\mathcal{M}=(\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R}_{\mathcal{M}},\gamma)$. Here $\mathcal{S}$ and $\mathcal{A}$ denote the state and action spaces, respectively. For $\forall(s,a)\in\mathcal{S}\times\mathcal{A}$, $\mathcal{P}(\cdot|s,a)$ is a distribution over $\mathcal{S}$, representing the dynamic, $\mathcal{R}_{\mathcal{M}}(s,a)$ is the reward function of $\mathcal{M}$, and $\gamma\in(0,1)$ is a discount factor.
Although deterministic Markovian policy is well known to be optimal for handling the single task setting, some recent work Ghosh et al. (2021) has demonstrated that generalizing to unseen environments introduces partial observability, thus deterministic Markovian policy will be sub-optimal and we need to encode more historical information. Also, previous multi-task methods Rakelly et al. (2019); Zintgraf et al. (2019) will utilize all historical information to choose the action. Formally, given policy $\pi$, at each timestep $t$, the agent will use the whole history trajectory $(s_{0},a_{0},r_{0},s_{1},a_{1},...,s_{t})$ to sample action $a_{t}$, arrive at the next state $s_{t+1}\sim\mathcal{P}(\cdot|s_{t},a_{t})$, and get the current reward $r_{t}=\mathcal{R}_{\mathcal{M}}(s_{t},a_{t})$. The performance of policy $\pi$ in $\mathcal{M}$ is defined as the expected discounted return of the trajectory $\tau$: $J_{\mathcal{M}}(\pi)=\mathbb{E}_{\tau\sim\pi}\left[R(\tau)\triangleq\sum_{t=0}^{\infty}\gamma^{t}r_{t}\right]$, where $R(\tau)$ represents the discounted return of $\tau$. And we hope to find the optimal policy over $\mathcal{T}$, i.e., $\max_{\pi}\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}\left[J_{\mathcal{M}}(\pi)\right]$.
In practice, given the task distribution $\mathcal{T}$, we sample $M$ training tasks $\{\mathcal{M}_{m}\}_{m=1}^{M}$ for optimizing the agent, i.e., maximizing $\frac{1}{M}\sum_{m=1}^{M}J_{\mathcal{M}_{m}}(\pi)$. In the testing stage, we will sample $N$ unseen testing tasks $\{\mathcal{M}_{M+n}\}_{n=1}^{N}$ to evaluate its generalization ability, i.e., evaluating $\frac{1}{N}\sum_{n=1}^{N}J_{\mathcal{M}_{M+n}}(\pi)$.
4 Task Aware Dreamer
In this section, we propose a novel framework of Task Aware Dreamer (TAD) for handling task generalization, including the reward-informed world models and the optimization. Finally, we will provide theoretical analyses of designs in TAD.
4.1 Reward-Informed World Models
Our first observation is that model-based methods, in addition to proving to improve sample efficiency in the single-task setting Janner et al. (2019); Hafner et al. (2019a), are also effective in narrowing down the hypothesis space of the optimal Q function when considering task generalization:
Theorem 1 (Informally, a detailed version and proof are in Appendix B.1).
Given $M$ tasks $\{\mathcal{M}_{m}\}_{m=1}^{M}$ and corresponding dataset $D^{m}$ sampled from $\mathcal{M}_{m}$. We can consider three hypothesis classes of optimal Q function $\mathcal{H}_{1},\mathcal{H}_{2},\mathcal{H}_{3}$ as
  • $\mathcal{H}_{1}$ includes all optimal Q functions satisfying the Bellman optimality equation for all transitions in $\mathcal{D}_{m}$.
  • For any dynamics $p_{m}$ satisfying the dynamic transition in $\mathcal{D}_{m}$, $\mathcal{H}_{2}$ includes all optimal Q functions satisfying the Bellman optimality equation for all transitions consistent with $p_{m}$.
  • For any dynamics $p$ satisfying all dynamic transition in $\{\mathcal{D}_{m}\}_{m=1}^{M}$, $\mathcal{H}_{3}$ includes all optimal Q functions satisfying the Bellman optimality equation for all transitions consistent with $p$.
We have that $\mathcal{H}_{3}\subseteq\mathcal{H}_{2}\subseteq\mathcal{H}_{1}$.
Here $\mathcal{H}_{2}\subseteq\mathcal{H}_{1}$ utilizes the generalization ability of the trained world model, which is similar to the single-task setting. Moreover, $\mathcal{H}_{3}\subseteq\mathcal{H}_{2}$ is based on similar dynamic structures across different tasks, which benefits learning the general world model. In summary, this result inherits the previous result Young et al. (2022) where the reward function is unchanged and inspires us that training a general world model can narrow down the hypothesis of possible Q function as well as benefit training the agent.
As traditional world models Hafner et al. (2019b, a) mainly designed for the single-task setting, we first analyze the probabilistic graphical model of the task-distribution setting in Fig. 1(b) and further build reward-informed world models for handling the task distribution. As shown in Fig. 1(b), the reward does not only rely on the previous states and actions but also on the current task. Thus the joint distribution can be calculated as
$p(s_{1:T},o_{1:T},r_{1:T},a_{1:T-1},\mathcal{M})$ (1)
$=$ $p(\mathcal{M})\prod_{t=1}p(s_{t+1}|s_{t},a_{t})p(o_{t}|s_{t})p(r_{t}|s_{t},\mathcal{M}),$
and we choose the reward-informed inference model (analyses about the necessity of utilizing reward-informed policy in task generalization is in Sec. 4.3) to approximate state posteriors as
$q(s_{1:T}|o_{1:T},a_{1:T},r_{1:T})=\prod_{t=1}^{T}q(s_{t}|s_{t-1},a_{t-1},r_{t-1},o_{t}).$ (2)
Further, we can construct the variational lower bound of the log-likelihood on the data as
$\begin{split}&\ln p(o_{1:T},r_{1:T},\mathcal{M}|a_{1:T})\\ \geq&\sum_{t=1}^{T}\mathbb{E}_{q(s_{t}|o_{\leq t},a_{<t},r_{<t})}\left[\ln p(o_{t},r_{t}|\mathcal{M},s_{t})\right.\\ -&\left.\text{KL}\left(q(s_{t}|o_{\leq t},r_{<t},a_{<t})\|p(s_{t}|s_{t-1},a_{t-1})\right)\right]\\ +&\mathbb{E}_{q(s_{1:T}|o_{1:T},a_{1:T},r_{1:T})}\left[\ln p(\mathcal{M}|s_{1:T})\right].\end{split}$ (3)
It is a general form of the single task setting Hafner et al. (2019b) and the detailed derivation is in Appendix B.2. Here the first two terms in Eq.(3) are for reconstructing observations, predicting rewards, and inferring states, which are similar to the single-task setting. The last novel term is dedicated to predicting the current task from historical information, which is beneficial for improving the generalization as it encourages inferring the current task context, and we use one-hot task vectors to present $\mathcal{M}$ during training. Extending RSSM via Eq.(3), our reward-informed world models consist of:
Deterministic state model: $h_{t}=f(h_{t-1},s_{t-1},a_{t-1},r_{t-1}),$
Transition model: $p_{\theta}(s_{t}|h_{t}),$
Observation model: $p_{\theta}(o_{t}|h_{t},s_{t}),$
Reward model: $p_{\theta}(r_{t}|h_{t},s_{t}),$
Task model: $p_{\theta}(\mathcal{M}|h_{t},s_{t}).$
Here hidden state $h_{t}$ encodes historical states, actions, and rewards, by using gated recurrent unit (GRU) Chung et al. (2014) as $f(\cdot)=\text{GRU}(\cdot)$. Then the transition model, observation model, and reward model further predict state, observation, and reward respectively. Finally, the task model infers the current task to distinguish different tasks’ trajectories.
4.2 Optimization
Based on the above analyses, we now introduce the training of TAD in detail. Following previous work Hafner et al. (2019a), we adopt an alternating training approach between the reward-informed world models and the policy.
To balance different tasks, in the stage of collecting data, TAD utilizes different replay buffers $\{\mathcal{D}_{m}\}_{m=1}^{M}$ to store trajectories sampled from $\{\mathcal{M}_{m}\}_{m=1}^{M}$, respectively. Then in the training stage, TAD will sample data from each replay buffer and train the reward-informed world models via the objective, which follows Eq. (3), as
$\begin{split}&L_{\text{TAD}}=\sum_{i=1}^{M}\mathbb{E}_{q}\left[\sum_{t=1}^{T}\ln p_{\theta}(o_{t}^{i}|h_{t}^{i},s_{t}^{i})+\sum_{t=1}^{T}\ln p_{\theta}(r_{t}^{i}|h_{t}^{i},s_{t}^{i})\right.\\ &-\left.\sum_{t=1}^{T}D_{\text{KL}}(q(s_{t}^{i}|h_{t}^{i},o_{t}^{i})\|p_{\theta}(s_{t}^{i}|h_{t}^{i}))+\sum_{t=1}^{T}\ln p_{\theta}(\mathcal{M}_{i}|h_{t}^{i},s_{t}^{i})\right].\end{split}$ (4)
In Eq. (4), the first three items are similar to Dreamer for reconstructing observations, predicting rewards, and inferring states. The novel last term is calculated in Eq. (3) for encouraging the agent to encode task information into hidden states and infer the current task based on these hidden states, which benefits generalizing to unseen tasks (We demonstrate this in Sec. 4.3).
In terms of training the actor-critic, both of which are parameterized neural networks, we follow the actor-critic learning approach in Dreamer with a more powerful hypothesis set $\Pi_{3}$ (The necessity is shown in Sec. 4.3). When optimizing the policy, TAD samples a series of states from the replay buffer and starts from them to imagine trajectories via interacting with reward-informed world models, which is fixed at this time. Here reward-informed world models can capture invariant features, which are beneficial for the agent to train more effectively and gain better generalization. After obtaining imagined trajectories, the actor-critic can be optimized via maximizing the $\lambda$-return Schulman et al. (2015) and regressing the targets calculated by TD Sutton and Barto (2018), respectively.
Algorithm 1 Task Aware Dreamer (TAD)
0:  $M$ training tasks $\{\mathcal{M}_{m}\}_{m=1}^{M}$, $M$ replay buffers $\{\mathcal{D}_{m}\}_{m=1}^{M}$, $N$ testing tasks $\{\mathcal{M}_{M+n}\}_{n=1}^{N}$, initialize parameters of world models, the policy, and the critic.
1:  while not converge do
2:     for update $\text{update step}=1,2,...,U$ do
3:        Sample $o$-$a$-$r$ pairs $\{(o_{t}^{i},a_{t}^{i},r_{t}^{i})_{t=1}^{T}\}$ form each replay buffer $\mathcal{D}_{i},i=1,2,...,M$
4:        Calculate the deterministic state $h$ and further calculate model states $s$.
5:        Update the world models via optimizing Eq. (4).
6:        Collect imagined trajectories from each $s$ via the policy and the world models and use these imagined trajectories to update the policy and the critic.
7:     end for
8:     Collect trajectories from $\mathcal{M}_{m}(m=1,2,...,M)$ and store them into the replay buffer $\mathcal{D}_{m}$.
9:  end while
10:  Evaluate the agent in testing environments $\{\mathcal{M}_{M+n}\}$.
4.3 Theoretical Analyses
In this part, we will provide some analyses to show that components of TAD are simple but effective for task generalization.
Is it necessary to use policies that utilize all historical information for task generalization?
Below we will introduce $3$ types of widely used policy hypotheses and show that $\Pi_{3}$, used in TAD, is necessary for handling the task generalization.
  1. 1.
    Markovian policy set $\Pi_{1}$ Sutton and Barto (2018); Yarats et al. (2021), i.e., $\Pi_{1}=\{\pi|\pi:\mathcal{S}\rightarrow\Delta(\mathcal{A})\}$, here $\Delta(\mathcal{A})$ represents a distribution over $\mathcal{A}$, which is widely used and optimal for the single-task setting;
  2. 2.
    $\mathcal{S}$-$\mathcal{A}$ memorized policy set $\Pi_{2}$ Hafner et al. (2019a, 2020); Lee et al. (2020), i.e., $\Pi_{2}=\{\pi|\pi:\mathcal{H}\rightarrow\Delta(\mathcal{A})\}$, here $\mathcal{H}=\cup_{t=1}^{\infty}\mathcal{H}_{t},\mathcal{H}_{t}=(\mathcal{S}\times\mathcal{A})^{t-1}\times\mathcal{S}$;
  3. 3.
    $\mathcal{S}$-$\mathcal{A}$$-\mathcal{R}$ memorized policy set $\Pi_{3}$ Rakelly et al. (2019); Zintgraf et al. (2019), i.e., $\Pi_{3}=\{\pi|\pi:\mathcal{L}\rightarrow\Delta(\mathcal{A})\}$, here $\mathcal{L}=\cup_{t=1}^{\infty}\mathcal{L}_{t},\mathcal{L}_{t}=(\mathcal{S}\times\mathcal{A}\times\mathbb{R})^{t-1}\times\mathcal{S}$.
As illustrated in Fig. 1(a), naturally $\Pi_{1}\subseteq\Pi_{2}\subseteq\Pi_{3}$. Next, we will analyze their expressive ability under task distribution $\mathcal{T}$. Denote $J_{\mathcal{T}}^{*}\triangleq\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}\left[\max_{\pi}J_{\mathcal{M}}(\pi)\right]$ as the optimal return under $\mathcal{T}$, and $J_{\mathcal{T}}^{i}\triangleq\max_{\pi\in\Pi_{i}}\left[\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}J_{\mathcal{M}}(\pi)\right]$ as the optimal return for $\Pi_{i}(i=1,2,3)$ under $\mathcal{T}$. Our first main result shows that, although $\Pi_{1}\subseteq\Pi_{2}$, they own the same expressive ability, i.e., $J_{\mathcal{T}}^{1}=J_{\mathcal{T}}^{2}$, and are both sub-optimal, summarized as:
Theorem 2 (Sub-Optimality of $\Pi_{1},\Pi_{2}$. Proof in Appendix B.3).
We set ¯$\bar{\mathcal{M}}=(\mathcal{S},\mathcal{A},\mathcal{P},\bar{\mathcal{R}},\gamma)$, here ¯$\bar{\mathcal{R}}=\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}[\mathcal{R}_{\mathcal{M}}]$. For $\forall\pi\in\Pi_{2}$, we have $\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}[J_{\mathcal{M}}(\pi)]=J_{\bar{\mathcal{M}}}(\pi)$ and further $J_{\mathcal{T}}^{1}=J_{\mathcal{T}}^{2}\leq J_{\mathcal{T}}^{*}$.
Theorem 2 reveals that the cumulative returns of policies in $\Pi_{1}$ and $\Pi_{2}$ are the same as their returns in the “average” MDP $\bar{\mathcal{M}}$, where the reward function is the average of reward functions in different tasks. Also, as $\Pi_{1}$ and $\Pi_{2}$ only choose actions via current state or historical state-action pairs, they cannot distinguish different tasks and are both sub-optimal.
To quantitatively analyze the characteristic of $\mathcal{T}$ and the gap between $J_{\mathcal{T}}^{1},J_{\mathcal{T}}^{2}$ and $J_{\mathcal{T}}^{*}$, we propose a novel metric named Task Distribution Relevance (TDR) of the distribution $\mathcal{T}$ as
Definition 1 (Task Distribution Relevance).
For any task distribution $\mathcal{T}$ and state $s$, the Task Distribution Relevance of $\mathcal{T}$ and $s$ is defined as
[…]
Intuitively, TDR describes the relevance of $\mathcal{T}$ via optimal $Q$ functions, which determine the distribution of optimal actions in corresponding tasks. Based on TDR, we can bound the gap:
Theorem 3 (Proof in Appendix B.4).
Assume $\pi_{\mathcal{M}}^{*}=\mathop{\arg\max}_{\pi}J_{\mathcal{M}}(\pi)$, for $\forall\pi\in\Pi_{1}$, we have
[…]
Theorem 3 demonstrates that the gap between $J_{\mathcal{T}}^{1},J_{\mathcal{T}}^{2}$ and $J_{\mathcal{T}}^{*}$ is related to the TDR of $\mathcal{T}$. When considering $\mathcal{T}$ with high TDR, i.e., the optimal $Q$ values in different tasks differ greatly, the performance of $\Pi_{1}$ and $\Pi_{2}$ will be extremely poor since they cannot differentiate different tasks and their expressive abilities are significantly limited. This conclusion is further verified empirically in the experimental section. Moreover, we can show that $J_{\mathcal{T}}^{1},J_{\mathcal{T}}^{2}$ might be arbitrarily small when $J_{\mathcal{T}}^{3}$ might be arbitrarily close to $J_{\mathcal{T}}^{*}$ (More details are in Appendix B.5), which demonstrates that $\Pi_{3}$ owns stronger expressive ability. Consequently, it is necessary to utilize $\Pi_{3}$ to distinguish different tasks for enhancing expression ability and generalization over the task distribution.
Is optimizing $p_{\theta}(\mathcal{M}|h_{t},s_{t})$ helpful for task generalization?
Now we will show that, although $p_{\theta}(\mathcal{M}|h_{t},s_{t})$ is a simple term, optimizing it can be effective for reaching the generalizable agent in the task distribution.
As the input space of $\Pi_{3}$ is $\mathcal{L}$, i.e., all partial trajectories, our major result analyzes the relation between the policy $\pi\in\Pi_{3}$ and the task posterior $p(\mathcal{M}|l),l\in\mathcal{L}$. Notice that $p_{\theta}(\mathcal{M}|h_{t},s_{t})$ can be regard a deformation of $p(\mathcal{M}|l),l\in\mathcal{L}$ as $h_{t}$ encodes all historical information.
Theorem 4 (Informally, detailed analyses and proof are in Appendix B.6).
For any policy $\pi\in\Pi_{3}$, we have
$\begin{split}&J_{\mathcal{T}}^{*}-\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}\left[J_{\mathcal{M}}(\pi)\right]\\ =&\frac{1}{1-\gamma}\int_{\mathcal{L}}p(l)dl\left[\int p(\mathcal{M}|l)\max_{a}Q_{\mathcal{M}}^{*}(l,a)d\mathcal{M}\right.\\ &\qquad\qquad\qquad-\left.\max_{a}\int p(\mathcal{M}|l)Q_{\mathcal{M}}^{*}(l,a)d\mathcal{M}\right],\end{split}$ (7)
here $p(l)$ is a distribution of $\mathcal{L}$ related to $\mathcal{T},\pi$ and $p(\mathcal{M}|l)$ is the task posterior related to $\pi$.
Consequently, maximizing $p(\mathcal{M}|l)$, i.e., making the distribution of $p(\mathcal{M}|l)$ to be closer to some Dirac distribution, can significantly reduce the right part of Eq. (7), thus is effective for improving the generalization ability of $\pi$. More details and discussion are also provided in Appendix B.6.
Tasks Hypothesis Walker-stand&\&&walk Acrobot-swingup&\&&sparse Reacher-easy&\&&hard Walker-walk&\&&prostrate
CURL $\Pi_{1}$ 563.4 $\pm$ 16.0 11.0 $\pm$ 1.8 552.5 $\pm$ 88.9 441.0 $\pm$ 7.7
SAC+AE $\Pi_{1}$ 683.8 $\pm$ 164.3 11.7 $\pm$ 5.4 669.5 $\pm$ 99.4 445.8 $\pm$ 2.6
PlaNet $\Pi_{2}$ 990.6 $\pm$ 3.8 12.2 $\pm$ 9.8 341.0 $\pm$ 228.3 445.4 $\pm$ 5.1
Dreamer $\Pi_{2}$ 992.3 $\pm$ 1.7 159.9 $\pm$ 108.4 506.1 $\pm$ 382.9 449.7 $\pm$ 12.5
TAD $\Pi_{3}$ 993.9 $\pm$ 1.0 217.3 $\pm$ 107.1 823.3 $\pm$ 237.4 913.7 $\pm$ 11.5
[…]
Tasks Hypothesis Cheetah__\__speed Pendulum__\__angle Walker__\__speed
Train Test Train Test Train Test
CURL $\Pi_{1}$ 211.7 $\pm$ 13.7 57.4 $\pm$ 26.6 137.1 $\pm$ 8.6 57.4 $\pm$ 21.8 127.0 $\pm$ 33.7 77.5 $\pm$ 11.5
SAC+AE $\Pi_{1}$ 182.2 $\pm$ 7.6 115.2 $\pm$ 10.1 144.0 $\pm$ 4.5 68.6 $\pm$ 3.7 136.8 $\pm$ 34.4 27.5 $\pm$ 10.5
PlaNet $\Pi_{2}$ 176.6 $\pm$ 25.9 83.0 $\pm$ 52.2 99.8 $\pm$ 28.3 80.5 $\pm$ 24.2 173.9 $\pm$ 19.3 58.4 $\pm$ 23.7
Dreamer $\Pi_{2}$ 250.2 $\pm$ 9.6 3.0 $\pm$ 2.2 92.2 $\pm$ 19.9 79.6 $\pm$ 14.8 197.6 $\pm$ 24.6 10.0 $\pm$ 6.5
TAD $\Pi_{3}$ 951.9 $\pm$ 3.3 876.9 $\pm$ 51.1 344.8 $\pm$ 2.8 132.6 $\pm$ 14.2 241.2 $\pm$ 36.8 156.5 $\pm$ 129.6
[…]
Tasks Hypothesis Half-Cheetah-Fwd-Back(1e7) Half-Cheetah-Vel(1e7) Humanoid-Direc-2D(1e6)
Train&\&&Test Train Test Train Test
PlaNet $\Pi_{2}$ 30.5 $\pm$ 42.9 -198.1 $\pm$ 1.9 -202.1 $\pm$ 1.8 215.9 $\pm$ 72.3 220.6 $\pm$ 75.3
Dreamer $\Pi_{2}$ 127.4 $\pm$ 181.8 -151.4 $\pm$ 0.4 -169.4 $\pm$ 1.2 260.5 $\pm$ 48.9 263.5 $\pm$ 52.3
RL2(zero-shot) $\Pi_{3}$ 1070.7 $\pm$ 109.7 -70.3 $\pm$ 6.7 191.9 $\pm$ 50.8
RL2(few-shot) $\Pi_{3}$ 1006.9 $\pm$ 26.4 -146.9 $\pm$ 0.4 268.8 $\pm$ 30.2
MAML(few-shot) 429.3 $\pm$ 81.4 -121.0 $\pm$ 37.1 205.3 $\pm$ 34.7
VariBAD(zero-shot) $\Pi_{3}$ 1177.5 $\pm$ 94.9 -58.4 $\pm$ 20.6 260.3 $\pm$ 61.6
TAD $\Pi_{3}$ 1455.8 $\pm$ 78.3 -49.3 $\pm$ 1.9 -47.1 $\pm$ 0.3 339.5 $\pm$ 78.7 335.5 $\pm$ 70.5
[…]
5 Experiments
We now present empirical results to answer the questions:
  • Can we verify the analyses about TDR, i.e., the expressive abilities of $\Pi_{1},\Pi_{2}$ are severely restricted in task distributions with high TDR?
  • What about the generalization ability of TAD in both image-based and state-based environments?
  • Can TAD be extended to more general settings like dynamic generalization?
5.1 Experimental Setup
Image-based Control.
To answer the first question and verify analyses about TDR, we first consider lots of different task combinations in DeepMind control suite (DMC) Tassa et al. (2018). These task combinations include (1) two task combinations of Walker-stand&\&&walk Mu et al. (2022) and Acrobot-swingup&\&&sparse, which share the same optimal actions with TDR being 0; (2) two task combinations with non-zero TDR of Reacher-easy&\&&hard Mu et al. (2022) and Walker-walk&\&&prostrate. For example, Walker-walk and Walker-prostrate hope a two-leg robot to maintain its height above a certain threshold for walking and below another certain threshold respectively, yielding almost opposite optimal Q functions with huge TDR. In Appendix C.1, we introduce more details about them and four other task combinations.
To evaluate generalization, we extend tasks in DMC and design three task distributions: (1) Cheetah_normal-_\__speed($\alpha,\ \beta$), which extends Cheetah-run and hopes the agent to run within the target speed interval $(\alpha-\beta,\ \alpha+\beta)$; (2) Pendulum_normal-_\__angle($\alpha,\ \beta$), which extends Pendulum-swingup and hopes an inverted pendulum to stay its pole within the target angle interval $(\arccos\alpha,\ \arccos\beta)$; and (3) Walker_normal-_\__speed($\alpha,\ \beta$), which is based on Walker-run and requires the planar walker to run within the target speed interval $(\alpha-\beta,\ \alpha+\beta)$. For each task distribution, we sample $4$ tasks for training and $2$ additional tasks for testing. More details are in Appendix C.2.
State-based Control.
To demonstrate the scalability of TAD, we also consider some state-based continuous control task distributions about robotic locomotion, simulated via the MuJoCo simulator Todorov et al. (2012). Following previous work Finn et al. (2017), our chosen tasks distributions include (1) Half-Cheetah-Fwd-Back, which owns two opposite tasks; and (2) Half-Cheetah-Vel and Humanoid-Direc-2D, which are task distributions with 100 training tasks 30 testing tasks. More details of these task distributions are in Appendix D.
Baselines.
In DMC, we choose two model-free methods: CURL Laskin et al. (2020) and SAC+AE Yarats et al. (2021), which employ Markovian policies and belong to $\Pi_{1}$. In addition, we choose two classic world models, PlaNet Hafner et al. (2019b) and Dreamer Hafner et al. (2019a), which utilize historical state-actions in policies and belong to $\Pi_{2}$. In state-based control, besides PlaNet and Dreamer, we also take some meta RL methods like MAML Finn et al. (2017), RL2 Duan et al. (2016), and VariBAD Zintgraf et al. (2019), including zero-shot and few-shot evaluation, as reference.
Metrics.
For the task combinations, we evaluate the average return of all tasks to verify TDR. For the task generalization setting, we train agents in training tasks and evaluate their generalization abilities in testing tasks. For all experiments, we train 3 policies with different seeds and calculate the mean and std to mitigate the effects of randomness.
Tasks Acrobot-Cartpole-Pendulum Walker-Cheetah-Hopper Cheetah-run__\__mass Walker-walk__\__mass
Train&\&&Test Train&\&&Test Train Test Train Test
Dreamer 541.3 $\pm$ 4.0 299.4 $\pm$ 17.2 717.8 $\pm$ 27.9 711.7 $\pm$ 38.8 889.9 $\pm$ 114.1 903.3 $\pm$ 102.8
TAD 667.7 $\pm$ 6.4 554.7 $\pm$ 23.6 754.3 $\pm$ 22.9 738.2 $\pm$ 41.1 957.8 $\pm$ 35.1 963.1 $\pm$ 32.6
[…]
Figure 2: Visualization of Dreamer (upper) and TAD (below) for handling Walker-walk&\&&Prostrate. For each algorithm in each task, we show 4 frame observations of a trajectory in the right part. In the left part, we draw t-SNE results of all states in trajectories of Walker-walk (blue points) and Walker-prostrate (red points).
5.2 Experimental Results for TDR
To answer the first question and validate our analyses of TDR, we report the optimal mean $\pm$ std of the average cumulative return of different task combinations in Table 5. Comprehensive results of all 8 task combinations and training curves are in Appendix C.3. The results show that TAD outperforms baselines in all environments, especially in those with high TDR. In task combinations where TDR is 0, like Walker-stand&\&&walk or Acrobot-swingup&\&&sparse, different tasks share the same optimal action and $\Pi_{1},\Pi_{2}$ own the optimal policy, as derived in Theorem 3. In these cases, $\Pi_{1},\Pi_{2}$ performs well and TAD can surpass baselines in Acrobot-swingup&\&&sparse since it can better distinguish different tasks. In the latter two environments with non-zero TDR, TAD achieves a notable improvement compared with baselines. Especially, in Walker-walk&\&&prostrate, where two tasks own almost opposite optimal Q functions and TDR is significantly huge, policies belong to $\Pi_{1},\Pi_{2}$ and can not differentiate different tasks, aligning with our Theorem 3. In contrast, TAD can distinguish between tasks by utilizing $\Pi_{3}$ and improves performance conspicuously.
Moreover, we present some visualization to better understand how TAD works in Fig. 2. We sample trajectories from the trained agent via Dreamer as well as TAD in Walker-walk&\&&prostrate and visualize some observations (right part of Fig. 2) as well as all states (left part of Fig. 2), of which the dimensions are reduced for visualization by t-SNE Van der Maaten and Hinton (2008). As shown here, states in TAD of different tasks are clearly distinguished, while Dreamer can not differentiate them and perform the same in different tasks. Consequently, TAD can successfully distinguish different tasks and handle them at the same time effectively. Videos of these trajectories are in supplementary materials.
5.3 Experimental Results for Task Generalization
To answer the second question, we report the results under the task generalization setting of image-based and state-based environments in Table 2 and Table 3, respectively. In Table 2, TAD shows powerful generalization ability in image-based environments as it can significantly enhance the average cumulative return in both training tasks and testing tasks compared with baselines. This demonstrates that TAD can both handle multiple training tasks simultaneously and generalize to unseen testing tasks effectively. We also provide some visualization results in Appendix C.4 with corresponding videos in supplementary materials.
Additionally, in Table 3, we compare TAD with some meta RL methods in state-based environments, of which the training timesteps are 1e7, 1e7, and 1e6, respectively. Many context-based meta-RL methods, including RL2, PEARL and VariBAD, utilize reward into policies and belong to $\Pi_{3}$, thus they can distinguish different tasks, which also verify our analyses. As shown in Table 3, TAD can achieve a significant improvement in the average cumulative return for both training tasks and testing tasks, since TAD is aware of task information and can generalize to unseen tasks.
5.4 Ablation Study
Reward Signal and Task ID.
In Fig. 3, we do ablation studies with two baselines: Dreamer(w/ r), which only integrates $\Pi_{3}$ into Dreamer, i.e., including reward signals into policies, and Dreamer(w/ task ID), which includes one-hot task ID into policies in both training and testing. Results show that just providing reward signals can help Dreamer distinguish different tasks, which verifies our analyses in TDR. Also, TAD shows superior performance even compared with Dreamer(w/ task ID), which demonstrates that our task model and corresponding ELBO are effective for task generalization.
Figure 3: Ablation study on Reward Signal and Task ID.
Extension to Dynamic Generalization.
To answer the third question, we further evaluate TAD in more general settings with different observations, dynamics, and/or actions. As TAD utilizes all historical information to infer the environment, it can be directly applied to these settings. We design task distribution with different embodiments (Acrobot-Cartpole-Pendulum, Walker-Cheetah-Hopper), as well as different dynamics (Cheetah-run__\__mass, Walker-walk__\__mass). Table 4 shows that TAD can achieve much greater performance and better convergence compared to Dreamer, which indicates TAD’s potential in further handling dynamic generalization and even cross-embodiment tasks. More details about the environments and results are in Appendix E.1.
5.5 Limitations and Ethical Statement
In terms of limitations, TAD assumes that the task context is continuously related to historical information for generalizing to unseen tasks. Thus TAD might be difficult to generalize in sparse-reward settings. We further demonstrate that without extra knowledge or finetuning, zero-shot generalization to unseen tasks with extremely sparse rewards is impossible since there is no way to distinguish different tasks (Appendix B.7). Fortunately, in relatively sparse reward settings, we conduct studies in Appendix E.2 to show that TAD can infer the current task and generalize to unseen tasks well. There are no serious ethical issues as this is basic research.
6 Conclusion
In this work, we propose a novel framework of TAD that distinguishes different tasks via all historical information and utilizes novel reward-informed world models to capture invariant latent features. In TAD, we calculate the corresponding variational lower bound of the data log-likelihood, which includes a novel loss term to distinguish different tasks via states. To explain components in TAD, we introduce a novel metric TDR to capture the relevance of the task distribution and show that Markovian policies perform poorly in tasks with high TDR. Experiments in image-based and state-based settings demonstrate that TAD can remarkably improve the performance of handling different tasks meanwhile, especially for high TDR ones, and successfully generalize to unseen tasks.
Appendix A Pseudo Code of TAD
We provide the pseudo code of TAD in Algorithm 2.
Algorithm 2 Task Aware Dreamer (TAD)
0:  $M$ training tasks $\{\mathcal{M}_{m}\}_{m=1}^{M}$, $M$ replay buffers $\{\mathcal{D}_{m}\}_{m=1}^{M}$, $N$ testing tasks $\{\mathcal{M}_{M+n}\}_{n=1}^{N}$, initialize neural network parameters of world models, the policy, and the critic
1:  while not converge do
2:     //$Model\ Training$
3:     for update $\text{update step}=1,2,...,U$ do
4:        Sample observation-action-reward pairs form each replay buffer $\{(o_{t}^{i},a_{t}^{i},r_{t}^{i})_{t=1}^{T}\}\sim\mathcal{D}_{i},i=1,2,...,M$
5:        Calculate the deterministic state $h$ and further calculate model states $s$.
6:        Update the world models via optimizing Eq. (6).
7:        Collect imagined trajectories from each $s_{t}$ via the policy and the world models.
8:        Use these imagined trajectories to update the policy and the critic.
9:     end for
10:     //$Data\ Collection$
11:     for $m=1,2,...,M$ do
12:        $o_{1}\leftarrow\mathcal{M}_{m}.reset()$
13:        for sample $\text{sample step}=1,2,...,S$ do
14:           Compute $h_{t},s_{t}$ and sample action $a_{t}$ via the policy.
15:           $r_{t},o_{t+1}\leftarrow\mathcal{M}_{m}.step(a_{t})$
16:        end for
17:        Add these data to the replay buffer $\mathcal{D}_{m}$.
18:     end for
19:  end while
20:  //$Model\ Evaluation$
21:  for $n=1,2,...,N$ do
22:     $o_{1}\leftarrow\mathcal{M}_{M+n}.reset()$
23:     while the environment not done do
24:        Compute $h_{t},s_{t}$ and sample action $a_{t}$ via the policy.
25:        $r_{t},o_{t+1}\leftarrow\mathcal{M}_{M+n}.step(a_{t})$
26:     end while
27:  end for
Appendix B Proof of Theorems
In this section, we will provide detailed proofs of theorems.
B.1 The Proof of Theorem 1
In this part, we give an informed version of Theorem 1 with detailed proof as
Theorem 5.
Given $M$ tasks $\{\mathcal{M}_{m}\}_{m=1}^{M}$, we consider the product space $\mathcal{H}=\mathcal{Q}^{M}$ composed of M state-Q function spaces, $\forall\{q_{m}\}_{m=1}^{M}\in\mathcal{H}$, $q_{m}:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ is some Q function. Given $M$ corresponding dataset $D^{m}=\{(s_{t}^{m},a_{t}^{m},r_{t}^{m},s_{t+1}^{m})\}$, we can consider three hypothesis classes of optimal Q function $\mathcal{H}_{1},\mathcal{H}_{2},\mathcal{H}_{3}$, which are subsets of $\mathcal{H}$:
$\begin{split}\mathcal{H}_{1}=&\{(q_{m})_{m=1}^{M}|q_{m}(s_{t}^{m},a_{t}^{m})=r_{t}^{m}+\gamma\max_{a^{\prime}}q_{m}(s_{t+1}^{m},a^{\prime})\}\\ \mathcal{H}_{2}=&\{(q_{m})_{m=1}^{M}|\exists(p_{m})_{m=1}^{M}:p_{m}(s_{t}^{m},a_{t}^{m})=s_{t+1}^{m},\\ &q_{m}(s,a)=r^{m}(s,a)+\gamma\max_{a^{\prime}}q_{m}(p_{m}(s,a),a^{\prime}),\forall s,a\}\\ \mathcal{H}_{3}=&\{(q_{m})_{m=1}^{M}|\exists p:p(s_{t}^{m},a_{t}^{m})=s_{t+1}^{m},\\ &q_{m}(s,a)=r^{m}(s,a)+\gamma\max_{a^{\prime}}q_{m}(p(s,a),a^{\prime}),\forall s,a\}.\end{split}$ (8)
Then we have $\mathcal{H}_{3}\subseteq\mathcal{H}_{2}\subseteq\mathcal{H}_{1}$.
Proof.
First, we will show that $\mathcal{H}_{2}\subseteq\mathcal{H}_{1}$, following the proof in Young et al. (2022) that considers the setting that the rewards of all environments are the same.
$\begin{split}&\{q_{m}\}\in\mathcal{H}_{2}\\ \Leftrightarrow&\exists\{p_{m}\}:p_{m}(s_{t}^{m},a_{t}^{m})=s_{t+1}^{m},\forall m,t\\ &q_{m}(s,a)=r^{m}(s,a)+\gamma\max_{a^{\prime}}q_{m}(p_{m}(s,a),a^{\prime}),\forall m,s,a\\ \Rightarrow&\exists\{p_{m}\}:p_{m}(s_{t}^{m},a_{t}^{m})=s_{t+1}^{m},\forall m,t\\ &q_{m}(s_{t}^{m},a_{t}^{m})=r^{m}_{t}+\gamma\max_{a^{\prime}}q_{m}(p_{m}(s_{t}^{m},a_{t}^{m}),a^{\prime}),\forall m,t\\ \Rightarrow&q_{m}(s_{t}^{m},a_{t}^{m})=r^{m}_{t}+\gamma\max_{a^{\prime}}q_{m}(p_{m}(s_{t}^{m},a_{t}^{m}),a^{\prime}),\forall m,t\\ \Leftrightarrow&\{q_{m}\}\in\mathcal{H}_{1},\end{split}$ (9)
thus we have $\mathcal{H}_{2}\subseteq\mathcal{H}_{1}$. Next, we prove that $\mathcal{H}_{3}\subseteq\mathcal{H}_{2}$, which is mainly because we can utilize similar dynamic structures from different tasks to narrow down the hypothesis spaces of the dynamic model.
$\begin{split}&\{q_{m}\}\in\mathcal{H}_{3}\\ \Leftrightarrow&\exists p:p(s_{t}^{m},a_{t}^{m})=s_{t+1}^{m},\forall m,t\\ &q_{m}(s,a)=r^{m}(s,a)+\gamma\max_{a^{\prime}}q_{m}(p(s,a),a^{\prime}),\forall m,s,a\\ \Rightarrow&\exists p_{m}:p_{m}(s_{t}^{m},a_{t}^{m})=s_{t+1}^{m},\forall m,t\\ &q_{m}(s,a)=r^{m}(s,a)+\gamma\max_{a^{\prime}}q_{m}(p_{m}(s,a),a^{\prime}),\forall m,s,a\\ \Leftrightarrow&\{q_{m}\}\in\mathcal{H}_{2},\end{split}$ (10)
thus we have $\mathcal{H}_{3}\subseteq\mathcal{H}_{2}$. ∎
B.2 The Derivation of the ELBO
We use ${\color[rgb]{0,139,69}q}$ to represent $q(s_{1:T}|o_{1:T},a_{1:T},r_{1:T})$, ${\color[rgb]{0,0,1}\hat{q}}$ to present $q(s_{t}|o_{\leq t},a_{<t},r_{<t})$, ${\color[rgb]{1,0,0}\tilde{q}}$ to represent $q(s_{t-1}|o_{\leq(t-1)},r_{<(t-1)},a_{<(t-1)})$, and we have
$\begin{split}&\ln p(o_{1:T},r_{1:T},\mathcal{M}|a_{1:T})\\ =&\ln\mathbb{E}_{p(s_{1:T}|a_{1:T})}\left[p(o_{1:T},r_{1:T},\mathcal{M}|s_{1:T})\right]\\ =&\ln\mathbb{E}_{p(s_{1:T}|a_{1:T})}\left[p(o_{1:T},r_{1:T}|\mathcal{M},s_{1:T})p(\mathcal{M}|s_{1:T})\right]\\ =&\ln\mathbb{E}_{p(s_{1:T}|a_{1:T})}\left[p(\mathcal{M}|s_{1:T})\prod_{t=1}^{T}p(o_{t},r_{t}|\mathcal{M},s_{t})\right]\\ =&\ln\mathbb{E}_{{\color[rgb]{0,139,69}q}}\left[p(\mathcal{M}|s_{1:T})\prod_{t=1}^{T}p(o_{t},r_{t}|\mathcal{M},s_{t})\frac{p(s_{t}|s_{t-1},a_{t-1})}{q(s_{t}|o_{\leq t},r_{<t},a_{<t})}\right]\\ \geq&\mathbb{E}_{{\color[rgb]{0,139,69}q}}[\ln p(\mathcal{M}|s_{1:T})+\sum_{t=1}^{T}\ln p(o_{t},r_{t}|\mathcal{M},s_{t})]\\ +&\mathbb{E}_{{\color[rgb]{0,139,69}q}}\sum_{t=1}^{T}\left[\ln p(s_{t}|s_{t-1},a_{t-1})-\ln q(s_{t}|o_{\leq t},r_{<t},a_{<t})\right]\\ =&\mathbb{E}_{{\color[rgb]{0,139,69}q}}[\ln p(\mathcal{M}|s_{1:T})]\\ +&\sum_{t=1}^{T}\left[\mathbb{E}_{{\color[rgb]{0,139,69}q}}\ln p(o_{t},r_{t}|\mathcal{M},s_{t})-\mathbb{E}_{{\color[rgb]{0,139,69}q}}\ln\frac{q(s_{t}|o_{\leq t},r_{<t},a_{<t})}{p(s_{t}|s_{t-1},a_{t-1})}\right]\\ =&\mathbb{E}_{{\color[rgb]{0,139,69}q}}[\ln p(\mathcal{M}|s_{1:T})]+\sum_{t=1}^{T}\mathbb{E}_{{\color[rgb]{0,0,1}\hat{q}}}[\ln p(o_{t},r_{t}|\mathcal{M},s_{t})]\\ -&\sum_{t=1}^{T}\mathbb{E}_{{\color[rgb]{0,0,1}\hat{q}}{\color[rgb]{1,0,0}\tilde{q}}}\left[\ln\frac{q(s_{t}|o_{\leq t},r_{<t},a_{<t})}{p(s_{t}|s_{t-1},a_{t-1})}\right]\\ =&\mathbb{E}_{{\color[rgb]{0,139,69}q}}[\ln p(\mathcal{M}|s_{1:T})]+\sum_{t=1}^{T}\mathbb{E}_{{\color[rgb]{0,0,1}\hat{q}}}[\ln p(o_{t},r_{t}|\mathcal{M},s_{t})]\\ -&\sum_{t=1}^{T}\mathbb{E}_{{\color[rgb]{1,0,0}\tilde{q}}}[\mathrm{KL}\left(q(s_{t}|o_{\leq t},r_{<t},a_{<t})\|p(s_{t}|s_{t-1},a_{t-1})\right)].\end{split}$ (11)
Thus we have proven the ELBO.
B.3 The Proof of Theorem 2
Proof.
We first prove that for $\forall\pi\in\Pi_{2}$, we have $\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}[J_{\mathcal{M}}(\pi)]=J_{\bar{\mathcal{M}}}(\pi)$.
For any $\mathcal{M}=(\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R}_{\mathcal{M}},\gamma)\sim\mathcal{T}$, we can use the policy $\pi$ to interact with $\mathcal{M}$ and get the trajectory $\tau=(s_{0}^{\mathcal{M}},a_{0}^{\mathcal{M}},r_{1}^{\mathcal{M}},s_{1}^{\mathcal{M}},a_{1}^{\mathcal{M}},r_{2}^{\mathcal{M}},...)$. Since the dynamic transition $\mathcal{P}$ is the same for all $\mathcal{M}$ and the policy $\pi\in\Pi_{2}$ only depends on historical states and actions, we naturally have that the distribution of all states and actions $(s_{0}^{\mathcal{M}},a_{0}^{\mathcal{M}},s_{1}^{\mathcal{M}},a_{1}^{\mathcal{M}},...)$ are the same for all $\mathcal{M}\sim\mathcal{T}$ as well as $\bar{\mathcal{M}}$. Consequently, we have
$\begin{split}&\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}[J_{\mathcal{M}}(\pi)]\\ =&\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}\mathbb{E}_{\tau\sim\mathcal{P},\pi}[R_{\mathcal{M}}(\tau)]=\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}\mathbb{E}_{\tau\sim\mathcal{P},\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}r_{t}^{\mathcal{M}}\right]\\ =&\mathbb{E}_{\tau\sim\mathcal{P},\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}[r_{t}^{\mathcal{M}}]\right]\\ =&\mathbb{E}_{\tau\sim\mathcal{P},\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}[\mathcal{R}_{\mathcal{M}}(s_{t}^{\mathcal{M}},a_{t}^{\mathcal{M}})]\right]\\ =&\mathbb{E}_{\tau\sim\mathcal{P},\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}[\bar{\mathcal{R}}(s_{t}^{\mathcal{M}},a_{t}^{\mathcal{M}})]\right]\\ =&\mathbb{E}_{\tau\sim\mathcal{P},\pi}[R_{\bar{\mathcal{M}}}(\tau)]=J_{\bar{\mathcal{M}}}(\pi).\end{split}$ (12)
It is well known that the optimal policy in single MDP is memory-less, i.e., $\max_{\pi\in\Pi_{2}}J_{\bar{\mathcal{M}}}(\pi)=\max_{\pi\in\Pi_{1}}J_{\bar{\mathcal{M}}}(\pi)$. Consequently, we have
$\begin{split}J_{\mathcal{T}}^{2}&=\max_{\pi\in\Pi_{2}}\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}[J_{\mathcal{M}}(\pi)]=\max_{\pi\in\Pi_{2}}J_{\bar{\mathcal{M}}}(\pi)\\ &=\max_{\pi\in\Pi_{1}}J_{\bar{\mathcal{M}}}(\pi)=\max_{\pi\in\Pi_{1}}\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}[J_{\mathcal{M}}(\pi)]=J_{\mathcal{T}}^{1}\\ &\leq\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}\left[\max_{\pi\in\Pi_{1}}J_{\mathcal{M}}(\pi)\right]=J_{\mathcal{T}}^{*}.\end{split}$ (13)
Thus we have proven this result. ∎
B.4 The Proof of Theorem 3
Proof.
Our proof follows some previous work Kakade and Langford (2002); Ying et al. (2022). First, we consider the bellman equation of value function of $\pi,\pi_{\mathcal{M}}^{*}\in\Pi_{1}$ in $\mathcal{M}$ as
$\begin{split}V_{\mathcal{M},\pi}(s)=&\sum_{a}\pi(a|s)\left[\mathcal{R}(s,a)+\gamma\sum_{s^{\prime}}\mathcal{P}(s^{\prime}|s,a)V_{\mathcal{M},\pi}(s^{\prime})\right],\\ V_{\mathcal{M},\pi_{\mathcal{M}}^{*}}(s)=&\sum_{a}\pi_{\mathcal{M}}^{*}(a|s)\left[\mathcal{R}(s,a)+\gamma\sum_{s^{\prime}}\mathcal{P}(s^{\prime}|s,a)V_{\mathcal{M},\pi_{\mathcal{M}}^{*}}(s^{\prime})\right].\end{split}$
Defining $\Delta V(s)\triangleq V_{\mathcal{M},\pi}(s)-V_{\mathcal{M},\pi_{\mathcal{M}}^{*}}(s)$ as the difference of these two value functions, we can further deduce that
$\begin{split}&V_{\mathcal{M},\pi}(s)-V_{\mathcal{M},\pi_{\mathcal{M}}^{*}}(s)\\ =&\gamma\sum_{a}\Delta\pi(a|s)\sum_{s^{\prime}}\mathcal{P}(s^{\prime}|s,a)V_{\mathcal{M},\pi_{\mathcal{M}}^{*}}(s^{\prime})\\ +&\gamma\sum_{a}\pi(a|s)\sum_{s^{\prime}}\mathcal{P}(s^{\prime}|s,a)\Delta V(s^{\prime})+\sum_{a}\Delta\pi(a|s)\mathcal{R}(s,a)\\ =&\sum_{a}\Delta\pi(a|s)Q_{\mathcal{M},\pi_{\mathcal{M}}^{*}}(s,a)\\ +&\gamma\sum_{a}\pi(a|s)\sum_{s^{\prime}}\mathcal{P}(s^{\prime}|s,a)\Delta V(s^{\prime}),\end{split}$ (14)
here $\Delta\pi(a|s)=\pi(a|s)-\pi_{\mathcal{M}}^{*}(a|s)$. Since Eq. (14) holds for any $s$, thus we calculate its expectation for $s\sim d_{\mathcal{M}}^{\pi_{\mathcal{M}}^{*}}$:
$\begin{split}&\sum_{s}d_{\mathcal{M}}^{\pi}(s)\Delta V(s)\\ =&\sum_{s}d_{\mathcal{M}}^{\pi}(s)[V_{\mathcal{M},\pi}(s)-V_{\mathcal{M},\pi_{\mathcal{M}}^{*}}(s)]\\ =&\sum_{s}d_{\mathcal{M}}^{\pi}(s)\sum_{a}\Delta\pi(a|s)Q_{\mathcal{M},\pi_{\mathcal{M}}^{*}}(s,a)\\ +&\gamma\sum_{s}d_{\mathcal{M}}^{\pi}(s)\sum_{a}\pi(a|s)\sum_{s^{\prime}}\mathcal{P}(s^{\prime}|s,a)\Delta V(s^{\prime})\\ =&\sum_{s}d_{\mathcal{M}}^{\pi}(s)\sum_{a}\Delta\pi(a|s)Q_{\mathcal{M},\pi_{\mathcal{M}}^{*}}(s,a)\\ +&\sum_{s^{\prime}}\Delta V(s^{\prime})\left[\gamma\sum_{s}d_{\mathcal{M}}^{\pi}(s)\sum_{a}\pi(a|s)\mathcal{P}(s^{\prime}|s,a)\right].\end{split}$ (15)
Since $\gamma\sum_{s^{\prime}}d_{\mathcal{M}}^{\pi}(s^{\prime})\sum_{a}\pi(a|s^{\prime})\mathcal{P}(s|s^{\prime},a)=d_{\mathcal{M}}^{\pi}(s)-(1-\gamma)\mathcal{P}(s_{0}=s)$, we have
$\begin{split}\sum_{s}d_{\mathcal{M}}^{\pi}(s)&\Delta V(s)=\sum_{s}d_{\mathcal{M}}^{\pi}(s)\sum_{a}\Delta\pi(a|s)Q_{\mathcal{M},\pi_{\mathcal{M}}^{*}}(s,a)\\ +&\sum_{s^{\prime}}\Delta V(s^{\prime})\left[d_{\mathcal{M}}^{\pi}(s^{\prime})-(1-\gamma)\mathcal{P}(s_{0}=s^{\prime})\right].\end{split}$ (16)
By moving the second term of the right part in Eq. (16) to the left part, we can deduce that
$\begin{split}&(1-\gamma)\sum_{s^{\prime}}\Delta V(s^{\prime})\mathcal{P}(s_{0}=s^{\prime})\\ =&\sum_{s}d_{\mathcal{M}}^{\pi}(s)\sum_{a}\Delta\pi(a|s)Q_{\mathcal{M},\pi_{\mathcal{M}}^{*}}(s,a),\end{split}$ (17)
thus we can calculate that
[…]
Consequently, we have
$\begin{split}&\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}\left[J_{\mathcal{M}}(\pi_{\mathcal{M}}^{*})-J_{\mathcal{M}}(\pi)\right]\\ =&\frac{1}{1-\gamma}\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}\mathbb{E}_{s\sim d_{\mathcal{M},\pi}(\cdot)}\left[\max_{a}Q_{\mathcal{M}}^{*}(s,a)\right.\\ &\left.-\int_{a}\pi(a|s)Q_{\mathcal{M}}^{*}(s,a)da\right]\\ \geq&\frac{1}{1-\gamma}\mathbb{E}_{s\sim d_{\mathcal{M},\pi}(\cdot)}\left[\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}\max_{a}Q_{\mathcal{M}}^{*}(s,a)\right.\\ &\left.-\max_{a}\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}Q_{\mathcal{M}}^{*}(s,a)\right]\\ =&\frac{1}{1-\gamma}\mathbb{E}_{s\sim d_{\mathcal{M},\pi}}\left[D_{\text{TDR}}(\mathcal{T},s)\right].\end{split}$ (19)
Since $J_{\mathcal{M}}(\pi_{\mathcal{M}}^{*})=\max_{\pi\in\Pi_{1}}J_{\mathcal{M}}(\pi)$, we have
$\begin{split}&\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}\max_{\pi\in\Pi_{1}}J_{\mathcal{M}}(\pi)-\max_{\pi\in\Pi_{1}}\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}J_{\mathcal{M}}(\pi)\\ \geq&\frac{1}{1-\gamma}\mathbb{E}_{s\sim d_{\mathcal{M},\pi^{*}}}\left[D_{\text{TDR}}(\mathcal{T},s)\right],\end{split}$ (20)
Thus we have proven this result. ∎
B.5 Expressive Ability of $\Pi_{3}$
Proposition 1.
For $\forall\epsilon_{1},\epsilon_{2}$ satisfying $0<\epsilon_{1}\leq 1,0<\epsilon_{2}\leq 1$, there exists a task distribution $\mathcal{T}$ satisfying that
$\begin{split}J_{\mathcal{T}}^{1}=J_{\mathcal{T}}^{2}\leq\epsilon_{1},~{}~{}J_{\mathcal{T}}^{3}\geq 1-\epsilon_{2},~{}~{}J_{\mathcal{T}}^{*}=1.\end{split}$ (21)
Proof.
Given fixed discount factor $\gamma\in(0,1)$, we first take $n\in\mathbb{N}$ satisfying $n\geq\frac{1}{\epsilon_{1}}$, $A=\epsilon_{2}\frac{n}{n-1}\frac{1-\gamma}{1-\gamma^{n}}$, and $B=\frac{1-\gamma}{\gamma^{n+1}}\left(1-\epsilon_{2}\frac{n}{n-1}\right)$. We construct state sets $\mathcal{S}=\{s_{t}^{l}\}(t=0,1,...,\infty,l=1,...,n)$ and action sets $\mathcal{A}=\{a_{j}\}_{j=1}^{n}$. Then we construct $n$ tasks $\mathcal{M}_{i}=(\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R}_{i},\gamma),i=1,...,n$, which share the same dynamic $\mathcal{P}$ and different reward functions $\mathcal{R}_{i}$. The initial state of each task is $s_{0}^{1}$, and the dynamic as well as reward functions are as below
$\begin{split}&\mathcal{P}(s_{t}^{l},a_{j})=\mathbb{I}(s=s_{t+1}^{j}),\quad j,l=1,...,n;\ t=0,1,...,\infty\\ &\mathcal{R}_{i}(s_{t}^{l},a_{j})=f(t)\mathbb{I}(i=j),\quad i,j,l=1,...,n;\ t=0,1,...\infty\\ &f(t)=\left\{\begin{array}[]{ll}A,&t\leq n-1\\ B,&t\geq n\\ \end{array}\right.\end{split}$ (22)
Take $\mathcal{T}$ as the uniform distribution over $\mathcal{M}_{1},...,\mathcal{M}_{n}$, thus we have
$\begin{split}J_{\mathcal{M}_{i}}^{*}=&A+A\gamma+...+A\gamma^{n-1}+B\gamma^{n}+B\gamma^{n+1}+...\\ =&A\frac{1-\gamma^{n}}{1-\gamma}+B\frac{\gamma^{n+1}}{1-\gamma}=1\\ J_{\mathcal{T}}^{*}=&\frac{1}{n}\sum_{i=1}^{n}J_{\mathcal{M}_{i}}^{*}=1.\\ \end{split}$ (23)
Since our construction satisfies $\mathbb{E}_{\mathcal{T}}[\mathcal{R}_{i}(s_{k}^{l},a_{j})]=\frac{f(k)}{n}$, for $\forall\pi\in\Pi_{2}$, we have
$\begin{split}J_{\mathcal{T}}(\pi)=&\frac{1}{n}(A+A\gamma+...+A\gamma^{n-1}+B\gamma^{n}+B\gamma^{n+1}+...)\\ =&\frac{1}{n},\\ J_{\mathcal{T}}^{1}=&J_{\mathcal{T}}^{2}=\max_{\pi\in\Pi_{2}}J_{\mathcal{T}}(\pi)=\frac{1}{n}\leq\epsilon_{1}.\end{split}$ (24)
Moreover, we can construction an agent $\hat{\pi}\in\Pi_{3}$ that takes action via the historical trajectory $\hat{\tau}_{t}=(\hat{s}_{0},\hat{a}_{0},\hat{r}_{0},...\hat{s}_{t})$:
$\begin{split}\hat{\pi}(a_{j}|\hat{\tau}_{t})=&\mathbb{I}(j=t+1),\quad t=0,1,...,n-1\\ \hat{\pi}(a_{j}|\hat{\tau}_{t})=&\mathbb{I}(j=i),\quad t=n,...,\infty\\ \end{split}$ (25)
here $i=\arg\max\{\hat{r}_{0},\hat{r}_{1},...,\hat{r}_{n-1}\}+1$, thus we have
$\begin{split}&J_{\mathcal{T}}^{3}\geq J_{\mathcal{T}}(\hat{\pi})\\ =&\frac{1}{n}(A+A\gamma+...+A\gamma^{n-1})+B\gamma^{n}+B\gamma^{n+1}+...\\ =&\frac{A}{n}\frac{1-\gamma^{n}}{1-\gamma}+B\frac{\gamma^{n+1}}{1-\gamma}=1-\frac{(n-1)A}{n}\frac{1-\gamma^{n}}{1-\gamma}\\ \geq&1-\epsilon_{2}.\end{split}$ (26)
Thus we have proven this result. ∎
B.6 Proof and Discussion of Theorem 4
In this part, we introduce an informed version of Theorem 4, about why optimizing $p(\mathcal{M}|l)\forall l\in\mathcal{L}$ is beneficial for task generalization, with detailed proofs. Recall that we consider the policy hypothesis $\mathcal{H}_{3}$ here, that each policy $\pi:\mathcal{L}\rightarrow\Delta(\mathcal{A}),\mathcal{L}=\cup_{t=1}^{\infty}\mathcal{L}_{t},\mathcal{L}_{t}=(\mathcal{S}\times\mathcal{A}\times\mathbb{R})^{t-1}\times\mathcal{S}$. As directly such $\mathcal{S}-\mathcal{A}-\mathcal{R}$ memorized policy is difficult, we consider an alternative MDP as ~$\tilde{\mathcal{M}}=(\mathcal{L},\mathcal{A},\mathcal{P}_{\mathcal{M}},\mathcal{R}_{\mathcal{M}},\gamma)$. For $\forall l=(s_{1},a_{1},r_{1},...,s_{t})\in\mathcal{L}_{t}\subseteq\mathcal{L}$, we can sample the action $a_{t}$ from the distribution $\pi(\cdot|l)$. Then the environment will feedback the reward signal $r_{t}=\mathcal{R}_{\mathcal{M}}(l,a_{t})=\mathcal{R}_{\mathcal{M}}(s_{t},a_{t})$, we can sample $s_{t+1}$ from the distribution $\mathcal{P}(\cdot|s_{t},a_{t})$, and the next environment state will be $l^{\prime}=(s_{1},a_{1},r_{1},...,s_{t},a_{t},r_{t},s_{t+1})\in\mathcal{L}_{t+1}$. In summary, we set the distribution $p(l^{\prime}|l,a_{t})$ as the corresponding dynamic $\mathcal{P}_{\mathcal{M}}$. Notice that although all $\mathcal{M}\in\mathcal{T}$ shares the same dynamic $\mathcal{P}$, their new dynamic $\mathcal{P}_{\mathcal{M}}$ are different since the dynamic is related to the given reward. An obvious advantage of introducing $\tilde{\mathcal{M}}$ is that $\pi\in\Pi_{3}$ is now Markovian in $\tilde{\mathcal{M}}$ and it is much easier to analyze its performance.
Obviously, we can set $J_{\mathcal{M}}(\pi)=J_{\tilde{\mathcal{M}}}(\pi),Q_{\mathcal{M}}^{*}(s,a)=Q_{\tilde{\mathcal{M}}}^{*}(s,a)$ to simplify the notation, and we can prove that
Theorem 6.
For any policy $\pi\in\Pi_{3}$, we have
$\begin{split}&J_{\mathcal{T}}^{*}-\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}\left[J_{\mathcal{M}}(\pi)\right]\\ =&\frac{1}{1-\gamma}\int_{\mathcal{L}}p(l)\left[\int p(\mathcal{M}|l)\max_{a}Q_{\mathcal{M}}^{*}(l,a)d\mathcal{M}\right.\\ &\qquad\ -\left.\int_{a,\mathcal{M}}\pi(a|l)p(\mathcal{M}|l)Q_{\mathcal{M}}^{*}(l,a)dad\mathcal{M}\right]dl\\ \geq&\frac{1}{1-\gamma}\int_{\mathcal{L}}p(l)\left[\int p(\mathcal{M}|l)\max_{a}Q_{\mathcal{M}}^{*}(l,a)d\mathcal{M}\right.\\ &\qquad\qquad\quad-\left.\max_{a}\int_{\mathcal{M}}p(\mathcal{M}|l)Q_{\mathcal{M}}^{*}(l,a)d\mathcal{M}\right]dl,\end{split}$ (27)
here $p(l)$ is a distribution of $\mathcal{L}$ related to $\mathcal{T},\pi$ and $p(\mathcal{M}|l)$ is the task posterior related to $\pi$.
Proof.
$\forall\pi\in\mathcal{H}_{3}$, as $\pi$ is Markovian in $\tilde{\mathcal{M}}$, we can directly utilize the proof of Theorem 3 from the beginning to Eq. (18), and the only difference is that the dynamics in $\tilde{\mathcal{M}}$ are different but the dynamics in $\mathcal{M}$ are the same. Thus we need to change the Eq. (19) as
$\begin{split}&\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}\left[J_{\mathcal{M}}(\pi_{\mathcal{M}}^{*})-J_{\mathcal{M}}(\pi)\right]\\ =&\frac{1}{1-\gamma}\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}\mathbb{E}_{l\sim d_{\tilde{\mathcal{M}},\pi}(\cdot)}\left[\max_{a}Q_{\mathcal{M}}^{*}(l,a)\right.\\ &\left.-\int_{a}\pi(a|l)Q_{\mathcal{M}}^{*}(l,a)da\right]\\ =&\frac{1}{1-\gamma}\int p(\mathcal{M})\int_{\mathcal{L}}d_{\tilde{\mathcal{M}},\pi}(l)\left[\max_{a}Q_{\mathcal{M}}^{*}(l,a)\right.\\ &\left.-\int_{a}\pi(a|l)Q_{\mathcal{M}}^{*}(l,a)da\right]\\ =&\frac{1}{1-\gamma}\int_{\mathcal{L}}p(l)\int p(\mathcal{M}|l)\left[\max_{a}Q_{\mathcal{M}}^{*}(l,a)\right.\\ &\left.-\int_{a}\pi(a|l)Q_{\mathcal{M}}^{*}(l,a)da\right]\\ \geq&\frac{1}{1-\gamma}\int_{\mathcal{L}}p(l)\left[\int p(\mathcal{M}|l)\max_{a}Q_{\mathcal{M}}^{*}(l,a)d\mathcal{M}\right.\\ &\qquad\qquad\quad-\left.\max_{a}\int_{\mathcal{M}}Q_{\mathcal{M}}^{*}(l,a)d\mathcal{M}\right]dl,\end{split}$ (28)
[…]
Finally, we will show that maximizing $p(\mathcal{M}|l)$ is helpful for task generalization. In the training stage, we will sample a task $\mathcal{M}$ and corresponding state-action-reward pairs $l$, thus optimizing $p(\mathcal{M}|l)$ will make it to be closer to some Dirac distribution. In such a situation, for each $l$, we can infer a “most possible” posterior task $\mathcal{M}_{l}$ with high $p(\mathcal{M}_{l}|l)$, thus we can approximately take $\pi(l)=\arg\max_{a}Q_{\mathcal{M}_{l}}^{*}(l,a)$ and the optimal gap calculated by E.q. (27) will be controlled.
B.7 Sparse Reward Task
In this part, we show that generalizing to tasks with the same dynamics and sparse rewards without extra knowledge (like context) is extremely difficult and sometimes impossible. It is because we cannot distinguish them via historical information. Here we construct an example.
Assume there are $n$ MDPs, $\mathcal{M}_{i}(i=1,...,n)$, each MDP share the same state set $\mathcal{S}=\{s_{1},...,s_{T}\}$ and action set $\mathcal{A}=\{a_{1},...,a_{n}\}$. The initial state is $s_{1}$ and the dynamic is that $\mathcal{P}(s_{t+1}|s_{t},a_{i})=1(t=1,...,T-1,i=1,...,n),\mathcal{P}(s_{T}|s_{T},a_{i})=1(i=1,...,n)$. As for the reward function, we define that
$\mathcal{R}_{\mathcal{M}_{i}}(s_{T-1},a_{i})=1,i=1,...,n.$ (29)
and the reward function is 0 otherwise. In this case, any policies (including markovian, state-action memorized, and state-action-reward memorized) in this task distribution can only handle one task since they can not distinguish them.
Appendix C Experimental Details for DMControl
C.1 Details of All Task Combinations
[…]
  • Walker-stand&\&&walk. This task combination includes two tasks: Walker-stand and Walker-walk, which both require a two-leg robot to maintain its height above a certain threshold, and Walker-walk also hopes it to move faster than a certain speed, i.e., their reward functions can be roughly described as
    $\begin{split}\mathcal{R}_{\text{stand}}&=\text{tolerance}(\text{height},(1.2,\infty)),\\ \mathcal{R}_{\text{walk}}&=\text{tolerance}(\text{height},(1.2,\infty))\\ &*\text{tolerance}(\text{speed},(1,\infty)).\end{split}$ (30)
    Therefore, for all states, the optimal action of Walker-walk is also optimal in Walker-stand and TDR here is 0.
  • Cartpole-balance&\&&sparse. This task combination includes two tasks: Cartpole-balance and Cartpole-balance__\__sparse, which both hope to balance an unactuated pole with dense and sparse rewards respectively. The optimal actions of these tasks are both hope to balance the agent and thus TDR here is 0.
  • Acrobot-swingup&\&&sparse. This task combination includes two tasks: Acrobot-swingup and Acrobot-swingup__\__sparse, which both hope to swing up a double pendulum with dense and sparse rewards respectively. Also, the optimal actions of these tasks are both hoped to swingup the agent and thus TDR here is 0.
  • Quadruped-walk&\&&run. This task combination includes two tasks: Quadruped-walk and Quadruped-run, which both require a quadruped agent to move faster than a target speed. The target speed of the running task is larger than the walking task, i.e., their reward functions can be roughly described as
    $\begin{split}&\mathcal{R}_{\text{walk}}=\text{tolerance}(\text{speed},(0.5,\infty)),\\ &\mathcal{R}_{\text{run}}=\text{tolerance}(\text{speed},(5.0,\infty)).\end{split}$ (31)
    Therefore, TDR here is still 0 since optimal actions of Quadruped-run are also optimal in Quadruped-walk.
  • Reacher-easy&\&&hard. This task combination includes two tasks: Reacher-easy and Reacher-hard, which both hope to control a two-link planar reacher to prostrate the target sphere. The sizes of target spheres in these two tasks are different, i.e., their reward functions can be roughly described as
    $\begin{split}&\mathcal{R}_{\text{easy}}=\text{tolerance}(\text{distance},(0.0,0.05)),\\ &\mathcal{R}_{\text{hard}}=\text{tolerance}(\text{distance},(0.0,0.015)).\end{split}$ (32)
    Thus TDR here is not 0 but still small.
  • Hopper-stand&\&&hop. This task combination includes two tasks: Hopper-stand and Hopper-hop, which both hope the height of a one-legged hopper is larger than the target height. While Hopper-stand hopes the control force is 0, Hopper-hop hopes to control the robot to hop forward and its speed is larger than the target speed. Their reward functions can be roughly described as
    $\begin{split}\mathcal{R}_{\text{stand}}&=\text{tolerance}(\text{height},(0.6,2))\\ &*\text{tolerance}(\text{speed},(2.0,\infty)),\\ \mathcal{R}_{\text{hop}}&=\text{tolerance}(\text{height},(0.6,2))\\ &*\text{tolerance}(\text{control},(0.0,0.0)).\end{split}$ (33)
    Since one agent can not hop when controlling the force as 0, TDR here is not 0 but not huge.
  • Walker-stand&\&&prostrate. This task combination includes two tasks: Walker-stand and Walker-prostrate. Walker-stand hopes the height of a two-leg robot to be larger than the target height. And we design the task Walker-prostrate, which hopes its height to be lower than another target height. Their reward functions can be roughly described as
    $\begin{split}&\mathcal{R}_{\text{stand}}=\text{tolerance}(\text{height},(1.2,\infty)),\\ &\mathcal{R}_{\text{prostrate}}=\text{tolerance}(\text{height},(0.0,0.2)).\end{split}$ (34)
    Thus TDR here is huge since the optimal Q functions of these two tasks are almost the opposite.
  • Walker-walk&\&&prostrate. This task combination includes two tasks: Walker-walk and Walker-prostrate. Walker-walk hopes the height of the improved planar walker to be larger than a target height, while Walker-prostrate hopes it to be lower than a target height. Moreover, Walker-walk hopes the speed of the robot is larger than another target speed, i.e., their reward functions can be roughly described as
    $\begin{split}\mathcal{R}_{\text{walk}}&=\text{tolerance}(\text{height},(1.2,\infty))\\ &*\text{tolerance}(\text{speed},(1.0,\infty)),\\ \mathcal{R}_{\text{prostrate}}&=\text{tolerance}(\text{height},(0.0,0.2)).\end{split}$ (35)
    Thus TDR here is also huge since the optimal Q functions of these two tasks are almost the opposite.
Figure 4: Average cumulative reward curves over different task combinations in DMC for CURL, SAC+AE, PlaNet, Dreamer, and TAD. The x-axes indicate the number of episodes interacting with the environment, and the y-axes indicate the performance, including average rewards with std.
Tasks Hypothesis Walker-stand&\&&walk Cartpole-balance&\&&sparse Acrobot-swingup&\&&sparse Quadruped-walk&\&&run
CURL $\Pi_{1}$ 563.4 $\pm$ 16.0 996.8 $\pm$ 0.8 11.0 $\pm$ 1.8 203.3 $\pm$ 69.9
SAC+AE $\Pi_{1}$ 683.8 $\pm$ 164.3 993.4 $\pm$ 0.6 11.7 $\pm$ 5.4 203.2 $\pm$ 70.2
PlaNet $\Pi_{2}$ 990.6 $\pm$ 3.8 412.8 $\pm$ 65.9 12.2 $\pm$ 9.8 305.3 $\pm$ 206.6
Dreamer $\Pi_{2}$ 992.3 $\pm$ 1.7 997.1 $\pm$ 1.1 159.9 $\pm$ 108.4 345.4 $\pm$ 218.3
TAD $\Pi_{3}$ 993.9 $\pm$ 1.0 998.9 $\pm$ 0.3 217.3 $\pm$ 107.1 505.0 $\pm$ 119.1
Tasks Hypothesis Reacher-easy&\&&hard Hopper-stand&\&&hop Walker-stand&\&&prostrate Walker-walk&\&&prostrate
CURL $\Pi_{1}$ 552.5 $\pm$ 88.9 137.1 $\pm$ 9.2 496.3 $\pm$ 8.7 441.0 $\pm$ 7.7
SAC+AE $\Pi_{1}$ 669.5 $\pm$ 99.4 118.2 $\pm$ 83.1 504.4 $\pm$ 6.7 445.8 $\pm$ 2.6
PlaNet $\Pi_{2}$ 341.0 $\pm$ 228.3 25.9 $\pm$ 25.7 544.3 $\pm$ 68.6 445.4 $\pm$ 5.1
Dreamer $\Pi_{2}$ 506.1 $\pm$ 382.9 436.9 $\pm$ 18.9 510.0 $\pm$ 6.7 449.7 $\pm$ 12.5
TAD $\Pi_{3}$ 823.3 $\pm$ 237.4 503.9 $\pm$ 22.2 941.1 $\pm$ 4.0 913.7 $\pm$ 11.5
[…]
C.2 Details of All Task Distributions
Now we introduce the three task distributions in our experiments, which are designed based on existing tasks in DeepMind control suite for testing the generalization of trained agents.
  • Cheetah_normal-_\__speed($\alpha,\beta$). This task distribution is designed in this paper with parameter $0\leq\beta\leq\alpha$, based on the task Cheetah__\__run in DeepMind control suite, and hopes the Cheetah robot can run with the target speed interval $(\alpha-\beta,\alpha+\beta)$.
    $\begin{split}&\mathcal{R}_{\text{Cheetah}\_\text{speed}}(\alpha,\beta)=\text{tolerance}(\text{speed},(\alpha-\beta,\alpha+\beta)).\end{split}$ (36)
    We train the agents in tasks with parameters ($0.5,0.2$), ($1.5,0.2$), ($2.0,0.2$), ($3.0,0.2$) and test them in tasks with parameters ($1.0,0.2$), ($2,5,0.2$).
  • Pendulum_normal-_\__angle($\alpha,\beta$). This task distribution is designed in this paper with parameter $-1\leq\alpha\leq\beta\leq 1$, based on the task Pendulumh__\__swingup in DeepMind control suite, and hopes the Pendulum robot can swing up within the target angle interval $(\arccos\alpha,\arccos\beta)$.
    $\begin{split}\mathcal{R}_{\text{Pendulum}\_\text{angle}}(\alpha,\beta)=\text{tolerance}(\text{angle},(&\arccos\alpha,\\ &\arccos\beta)).\end{split}$ (37)
    Training tasks are with parameters ($-0.95,-0.9$), ($-0.85,-0.8$), ($-0.8,-0.75$), ($-0.7,-0.65$) and testing tasks are with parameters ($-0.9,-0.85$), ($-0,75,-0.7$).
  • Walker_normal-_\__speed($\alpha,\beta$). This task distribution is designed in this paper with parameter $0\leq\beta\leq\alpha$, based on the task Walker__\__run in DeepMind control suite, and hopes the Walker robot can run within the target speed interval $(\alpha-\beta,\alpha+\beta)$.
    $\begin{split}&\mathcal{R}_{\text{Walker}\_\text{speed}}(\alpha,\beta)=\text{tolerance}(\text{speed},(\alpha-\beta,\alpha+\beta)).\end{split}$ (38)
    We train the agents in tasks with parameters ($0.5,0.2$), ($1.5,0.2$), ($2.0,0.2$), ($3.0,0.2$) and test in tasks with parameters ($1.0,0.2$), ($2,5,0.2$).
Moreover, we introduce some details about our experiments. Our codes are based on Python and the deep learning library PyTorch. All algorithms are trained on one NVIDIA GeForce RTX 2080 Ti. As for the hyper-parameters, we follow previous works Ha and Schmidhuber (2018); Hafner et al. (2019b, a) and select 2 as the action repeat for all experiments following Hafner et al. (2019a).
C.3 Full Experimental Results for TDR
In tasks where TDR is 0, like Walker-stand&\&&walk and Cartpole-balance&\&&sparse, the optimal performance of TAD and baselines are similar since these tasks share the same optimal action and $\Pi_{1},\Pi_{2}$ own the optimal policy, as derived in Theorem 2. In these cases, TAD still converges faster and shows more robust performance. In Acrobot-swingup&\&&sparse and Quadruped-walk&\&&run, although TDR of these tasks is 0, TAD can surpass baselines since it can better distinguish different tasks. Finally, in the last four environments with non-zero TDR, TAD achieves a notable improvement compared with baselines. Especially, in Walker-stand&\&&prostrate and Walker-walk&\&&prostrate, where TDR is significantly huge, baselines perform poorly as these policies belong to $\Pi_{1},\Pi_{2}$ and can not differentiate different tasks, which also aligns with our Theorem 2. In contrast, TAD can distinguish between tasks by utilizing reward signals and improves performance conspicuously.
C.4 Visualization Results for Task Generalization
Figure 5: Visualization of the trained TAD agent in the task distribution of Cheetah__\__speed. We plot the speed of the agent as a function of timesteps in all tasks.
Moreover, for each task sampled from the task distribution Cheetah__\__speed (here parameters $(3.0,0.2)$, $(2.0,0.2)$, $(1.5,0.2)$, $(0.5,0.2)$ are for training tasks and parameters $(2.5,0.2),(0.5,0.2)$ are for testing tasks), we plot the speed of the agent as a function of the timestep in Fig. 5. As depicted, for each task, the agent trained by TAD will quickly improve its speed until reaching the target speed and then keep its speed since the speed determines whether it has met the task requirements via utilizing historical information. Consequently, TAD is aware of different tasks and can successfully generalize to unseen test tasks. We also provide videos of these trajectories in supplementary materials.
Appendix D Experimental Details for MuJoCo
We here introduce state-based control tasks (Half-CheetahFwd-Back, Half-Cheetah-Vel, and Humanoid-Direc-2D) in detail, following the setting of previous meta RL works Finn et al. (2017); Rakelly et al. (2019).
  • Half-Cheetah-Fwd-Back. This task distribution includes two tasks: moving forward and moving backward.
  • Half-Cheetah-Vel. This task distribution hopes the agent to move forward and achieve the target velocity. There are 100 training tasks and 30 testing tasks for experiments.
  • Humanoid-Direc-2D. This task distribution hopes the agent to move in the target direction. There are 100 training tasks and 30 testing tasks for experiments.
Moreover, we introduce some details about our experiments. Our codes are based on Python and the deep learning library PyTorch. All algorithms are trained on one NVIDIA GeForce RTX 2080 Ti. We select 1 as the action repeat for all experiments following.
Appendix E Ablation Study
E.1 Dynamic Generalization
As TAD utilizes all historical information to infer the environment, it can be directly applied to more general settings with different observations, dynamics, and/or actions. To evaluate the performance of TAD in these settings, we have conducted the following four experiments based on DMControl:
  • Acrobot-Cartpole-Pendulum: includes 7 tasks of artpole-balance, cartpole-balance__\__sparse, cartpole-swingup, cartpole-swingup__\__sparse, acrobot-swingup, acrobot-swingup__\__sparse, and pendulum-swingup. All these tasks aim to control a rod-shaped robot, while they own different embodiments, dynamics, and observations.
  • Walker-Cheetah-Hopper: includes 6 tasks of walker-prostrate, walker-stand, walker-walk, cheetah-run, hopper-stand, and hopper-hop. The tasks own different embodiments, dynamics, actions, and observations.
  • Cheetah-run-mass($m$): This task distribution is based on the task Cheetah-run and the mass of the robot is $m$ times that of the standard task. Thus different tasks own different dynamics. We train the agents in tasks with $m=0.6,1.0,1.2,1.6$ and test them in tasks with parameters $m=0.8,1.4$.
  • Walker-walk-mass($m$): This task distribution is based on the task Walker-walk and the mass of the robot is $m$ times that of the standard task. Thus different tasks own different dynamics. We train the agents in tasks with $m=0.6,1.0,1.2,1.6$ and test them in tasks with parameters $m=0.8,1.4$.
Then we test Dreamer and TAD in these four settings and report the results. TAD can achieve much greater performance and better convergence compared to Dreamer, as it can better infer the current task. This experiment indicates TAD’s potential in further handling dynamic generalization and even cross-embodiment tasks.
E.2 Sparse Reward
$\beta$ 0.2 0.15 0.1
Train Test Train Test Train Test
Dreamer 250.2 $\pm$ 9.6 3.0 $\pm$ 2.2 247.5 $\pm$ 1.0 0.0 $\pm$ 0.0 175.8 $\pm$ 55.8 13.5 $\pm$ 13.5
TAD 951.9 $\pm$ 3.3 876.9 $\pm$ 51.1 927.6 $\pm$ 2.6 800.1 $\pm$ 121.4 608.5 $\pm$ 321.4 491.6 $\pm$ 389.4
Table 6: Average cumulative reward (mean $\pm$ one std) over different target region (smaller $\beta$ represents smaller target region and more sparse return) of the best policy trained by Dreamer and TAD in Cheetah__\__Speed. For each $\beta$, we train agents in the train tasks and evaluate them in both train and test environments. Numbers greater than 95 percent of the best performance for each environment are bold.
SR 0.0 0.8 0.9
Train Test Train Test Train Test
Dreamer 250.2 $\pm$ 9.6 3.0 $\pm$ 2.2 237.4 $\pm$ 14.9 28.9 $\pm$ 14.6 168.7 $\pm$ 96.2 157.2 $\pm$ 188.4
TAD 951.9 $\pm$ 3.3 876.9 $\pm$ 51.1 841.9 $\pm$ 154.9 546.1 $\pm$ 291.6 $7$$\pm$ 170.1 716.8 $\pm$ 136.5
Table 7: Average cumulative reward (mean $\pm$ one std) over different sparse rates of the best policy trained by Dreamer and TAD in Cheetah__\__Speed. For each sparse rate, we train agents in the train tasks and evaluate them in both train and test environments. Numbers greater than 95 percent of the best performance for each environment are bold.
In this part, we will evaluate TAD in more challenging settings with sparse rewards. First, we evaluate Dreamer and TAD in Cheetah__\__speed with different $\beta$, which identifies the region of target speeds. With smaller $\beta$, the reward signals are more sparse since the target intervals are smaller. In the main experiment, we take $\beta=0.2$, and here we evaluate in $\beta=0.2,0.15,0.1$, of which the result is reported in Table 6. For each $\beta$, we take the training parameters ($0.5,\beta$), ($1.5,\beta$), ($2.0,\beta$), ($3.0,\beta$) and test them in tasks with parameters ($1.0,\beta$), ($2.5,\beta$). As shown in Table 6, with the decreasing of $\beta$, the performance of Dreamer and TAD decreases since reward signals are sparse so exploration here is much more difficult. However, our TAD still significantly outperforms Dreamer and shows strong generalization abilities, which shows that TAD can effectively utilize historical information, even sparse rewards.
Moreover, we design Cheetah__\__speed__\__sparse based on Cheetah__\__speed. In Cheetah__\__speed__\__sparse($n$), we make the reward function sparse, i.e., the output reward is the same as Cheetah__\__speed every $n$ timesteps (in step $n-1,2n-1,...$) and 0 otherwise, of which the sparse rate (SR) is $(n-1)/n$. We supplement experiments to evaluate the performance of Dreamer and TAD with $n=5$ (SR=0.8) and $n=10$ (SR=0.9). As shown in Table 7, with the increasing of SR, although the performance of TAD decrease since inferring task context from sparse reward is extremely difficult, TAD still shows strong performance in train tasks and generalizes well to unseen test tasks.
Appendix F Ethics Issues and Border Impact
Designing agents that can generalize to unseen tasks is a major concern in reinforcement learning. This work focuses on task generalization in reinforcement learning and proposes a novel algorithm Reward Informed Dreamer. One of the potential negative impacts is that algorithms using deep neural networks, which lack interoperability and theoretical guarantee. If we hope to apply them in real-world applications, they may face security and robustness issues, and a possible way is to develop more explainable methods. There are no serious ethical issues as this is basic research. We hope our work can inspire more research on designing agents with stronger generalization abilities.

### v4

@@
 Abstract
-A long-standing goal of reinforcement learning is to acquire agents that can learn on training tasks and generalize well on unseen tasks that may share a similar dynamic but with different reward functions.
-The ability to generalize across tasks is important as it determines an agent’s adaptability to real-world scenarios where reward mechanisms might vary.
-In this work, we first show that training a general world model can utilize similar structures in these tasks and help train more generalizable agents.
+A long-standing goal of reinforcement learning is to acquire agents that can learn on various training tasks and generalize well on unseen tasks that may share a similar dynamic but with different reward functions.
+The ability to generalize across tasks is significant for real-world applications where the robot needs to adapt to varying reward mechanisms with the same embodiment.
+In this work, we first indicate that training general world models can utilize similar structures in these tasks and benefit training more generalizable agents.
 Extending world models into the task generalization setting, we introduce a novel method named Task Aware Dreamer (TAD), which integrates reward-informed features to identify consistent latent characteristics across tasks.
 Within TAD, we compute the variational lower bound of sample data log-likelihood, which introduces a new term designed to differentiate tasks using their states, as the optimization objective of our reward-informed world models.
-To demonstrate the advantages of the reward-informed policy in TAD, we introduce a new metric called Task Distribution Relevance (TDR) which quantitatively measures the relevance of different tasks.
-For tasks exhibiting a high TDR, i.e., the tasks differ significantly, we illustrate that Markovian policies struggle to distinguish them, thus it is necessary to utilize reward-informed policies in TAD.
-Extensive experiments in both image-based and state-based tasks show that TAD can significantly improve the performance of handling different tasks simultaneously, especially for those with high TDR, and display a strong generalization ability to unseen tasks.
-(a) Overview of this work
-(b) Probabilistic graphical model designs
-Figure 1: (a) An overview.
-Given a task distribution, we train the agent in training tasks and hope it to zero-shot generalize to testing tasks.
+To demonstrate the advantages of the reward-informed policy utilized in TAD for handling the task distribution, we introduce a novel metric named Task Distribution Relevance (TDR) which quantitatively measures the relevance of different tasks.
+For tasks exhibiting a high TDR, i.e., these tasks differ significantly, we illustrate that Markovian policies struggle to distinguish them, thus it is necessary to utilize reward-informed policies in TAD.
+Extensive experiments in both image-based and state-based settings show that TAD can significantly improve the performance of handling different tasks meanwhile, especially for those with high TDR, and display a strong generalization ability to unseen tasks.
+Figure 1: An overview.
+Given a task distribution, we train the agent in training tasks and hope it to zero-shot generalize to test tasks.
 For improving the generalization, we propose TAD, which utilizes $\Pi_{3}$ to encode all historical information for inferring the current task and novel reward-informed world models for capturing invariant latent features.
-(b) Probabilistic graphical model designs for the single-task setting (left) and the task-distribution setting (right).
+1 Introduction
+Deep Reinforcement Learning (DRL) has demonstrated significant advancements in diverse fields (Mnih et al.
+2016; Silver et al.
+2016), and a key factor in these achievements is an agent’s proficiency in assimilating lessons from special training tasks.
+This tendency towards specialization harms the wider real-world application of DRL, where broad generalization across various tasks is necessary.
+One primary challenge in generalization is the correspondence in the underlying dynamics of different tasks, which exhibit distinct reward structures.
+For example, we may need to control the same robot to handle different tasks, which can be rewarded for speed or caution in a navigating robot.
+Developing generalizable agents that recognize and handle these subtle variations is still an area of keen interest and exploration in DRL.
+For developing such generalizable agents, a promising pathway is to train general world models (Ha and Schmidhuber 2018; Hafner et al.
+2019b, a) that help the agent understand the world and make decisions.
+In this work, we first provide insight that general world models benefit improving the sample efficiency in handling the task distribution, especially by utilizing similar dynamic structures of trajectories sampled from different tasks (Theorem 1).
+Consequently, we propose a novel framework named Task Aware Dreamer (TAD) to improve the generalization ability of agents via general world models.
+As current world models are primarily designed for the single-task setting (left of Fig.
+2), we consider the corresponding probabilistic graphical model for the task-distribution setting (right of Fig.
+2) and propose novel reward-informed world models that can capture invariant structures across tasks in the distribution.
+Then we compute the variational lower bound of the data log-likelihood as the primary training objective for optimizing the reward-informed world models, which incorporate a novel task context term that categorizes tasks based on their states.
+In practice, we implement this optimization objective via two alternative methods: cross-entropy (TAD-CE) and supervised-contrastive (TAD-SC).
+Figure 2: Probabilistic graphical model designs for the single-task setting (left) and the task-distribution setting (right).
 The latter inspires the design of reward-informed world models.
 Solid and dashed lines represent the generative process and the inference model, respectively.
-1 Introduction
-Deep Reinforcement Learning (DRL) has demonstrated significant advancements in diverse fields Mnih et al.
-(2016); Hafner et al.
-(2019a) , and a key factor in these accomplishments is an agent’s proficiency in assimilating lessons from dedicated training tasks.
-However, when the objective transitions from optimizing performance in specific tasks to achieving broad generalization across related tasks, new challenges emerge.
-One primary source of these challenges is the congruence in the underlying dynamics of different tasks, which exhibit distinct reward structures.
-To illustrate, two tasks might involve navigating a robot, but in one, the robot is rewarded for speed, while in the other, it is rewarded for caution.
-This tendency towards specialization harms the wider applicability of DRL in real-world scenarios.
-Developing generalizable agents that recognize and handle these subtle variations continues to be an area of keen interest and exploration in DRL.
-For designing such generalizable agents, a promising pathway is to train a general world model Ha and Schmidhuber (2018); Hafner et al.
-(2019b) that helps the agent understand the world and make the decision.
-In this research, we first provide insight into how a unified dynamic model can help improve the sample efficiency in handling the distribution of tasks, especially by utilizing similar structures of trajectories sampled from different tasks (Theorem 1).
-Consequently, we propose a novel framework named Task Aware Dreamer (TAD), which extends existing world models into the task generalization setting.
-As current world models are primarily designed for the single-task setting (left of Fig.
-1(b)), we consider the corresponding probabilistic graphical model of the task distribution setting (right of Fig.
-1(b)) and extend single-task world model into novel reward-informed world models, which can capture learnable structures in the task distribution.
-To optimize these models, we compute the variational lower bound of the data log-likelihood while incorporating a novel term that categorizes tasks based on their states, serving as the primary training objective for the reward-informed world models.
-Furthermore, we conduct detailed analyses of the components of TAD, which are simple but effective for handling task distribution.
-First, we point out that the reward-informed policy used in TAD is necessary for task generalization.
-While prior studies Ghosh et al.
-(2021) have highlighted that generalizing to unseen environments results in partial observability, rendering all Markovian policies suboptimal, they often overlook the relationship between the task distribution and the expressiveness of the chosen policy hypothesis set.
-To bridge this gap, we propose a novel metric of Task Distribution Relevance (TDR), which encapsulates the relevance of different tasks within the distribution through their optimal Q functions.
-Furthermore, we prove that both the commonly adopted hypothesis set $\Pi_{1}$ of Markovian policies and the set $\Pi_{2}$ of policies that encode historical states and actions, are sub-optimal under the task-distribution setting and struggle to generalize.
-This sub-optimality is proven to be related to TDR (Theorem 3), i.e., for task distributions with high TDR, the performance of these two policy hypotheses might degenerate significantly as they fail to identify the current task, a phenomenon we also demonstrate by experiments (Sec.
+Moreover, we theoretically explain TAD’s components, like reward-informed policies and task optimization terms, are effective for handling task distribution.
+About the policy hypothesis, although several meta RL methods (Zintgraf et al.
+2019; Rimon et al.
+2024) have chosen the reward-informed policy hypothesis $\Pi_{3}$, the relationship between the task distribution and the expressiveness of the policy hypothesis set is still unclear.
+In other words, why the commonly adopted hypothesis set $\Pi_{1}$ of Markovian policies and the set $\Pi_{2}$ of policies encoding historical states and actions are not suitable for task distribution?
+To answer this question, we propose a novel metric of Task Distribution Relevance (TDR), encapsulating the relevance of different tasks within the distribution through their optimal Q functions.
+We then prove that both $\Pi_{1}$ and $\Pi_{2}$ are sub-optimal under the task-distribution setting.
+This sub-optimality is related to TDR (Theorem 3), i.e., for task distributions with high TDR, the performance of these two policy hypotheses might degenerate significantly, a phenomenon we also demonstrate in experiments (Sec.
 5.2).
-Consequently, the reward-informed policy hypothesis $\Pi_{3}$ utilized in TAD is necessary for handling task generalization.
-Besides the policy chosen, we also discuss the novel optimization term, i.e., categorizing tasks based on hidden state, in TAD and theoretically show that it can effectively reduce the gap between the policy return and the optimal return (Theorem 4).
-Extensive experiments are performed on task distributions based on DeepMind control suite Tassa et al.
-(2018) and MuJoCo simulator Todorov et al.
-(2012), which are image-based and state-based respectively.
-Agents are trained on specific tasks and evaluated for their generalization capabilities on unseen evaluation tasks.
-Our findings corroborate our analyses, indicating that $\Pi_{1}$ and $\Pi_{2}$ falter in managing task distributions characterized by a high TDR.
-Contrastingly, TAD excels in discerning and simultaneously managing varied tasks, significantly surpassing benchmark performance.
-This underscores TAD’s impressive sample efficiency and adaptability in both image-based and state-based contexts.
-Additionally, our ablation studies spotlight TAD’s versatility, showcasing its prowess in dynamic generalization and handling cross-embodiment tasks.
-In summary, our contributions are:
-  • We present a theoretical observation that a general world model can utilize similar structures of different tasks and improve the sample efficiency for training a generalizable agent for the task distribution.
-  • Extending existing single-task world models into the task distribution setting, we then formulate a new algorithm, termed TAD, leveraging the variational lower bound for calculations and showcasing superior generalization capabilities in contrast to baselines.
-  • To analyze TAD’s components, we analyze policy generalization across task distributions and introduce a new metric TDR, which quantifies the distribution relevance.
-Our findings indicate that Markovian policies tend to be sub-optimal, with performance gaps correlating with TDR.
-Further, we demonstrate that the new optimization term in TAD is effective for task generalization.
+This result explains why TAD and previous meta RL methods choose $\Pi_{3}$ for handling the task distribution.
+Besides the policy chosen, we also discuss our task optimization term in TAD and prove that it can effectively reduce the gap between the policy return and the optimal return (Theorem 4).
+We evaluate the task generalization ability of TAD in extensive experiments, including DeepMind control suite (Tassa et al.
+2018) and MuJoCo (Todorov, Erez, and Tassa 2012), which are image-based and state-based respectively.
+Agents are trained on various tasks and evaluated on unseen tasks.
+Results corroborate our analyses, indicating that $\Pi_{1}$ and $\Pi_{2}$ falter in managing task distributions characterized by a high TDR.
+Contrastingly, TAD excels in simultaneously managing varied tasks and outperforms all baselines, including SOTA model-based meta RL method MAMBA (Rimon et al.
+2024).
+Additionally, our ablation studies highlight TAD’s versatility, showing its prowess in dynamic generalization and handling cross-embodiment tasks.
+Overall, our contributions include:
+  • We present theoretical insights that general world models can utilize similar structures across tasks and improve the sample efficiency for task generalization (Sec. 4.1).
+  • Our TAD extends world models for the task distribution with the corresponding variational lower bound (Sec. 4.2).
+  • We theoretically analyze TAD’s components are effective for task generalization, with a novel metric TDR to quantify the distribution relevance (Sec. 4.3).
+  • Extensive experiments show that TAD can outperform various SOTA baselines and exhibit better generalization flexibility over image-based and state-based settings (Sec. 5).
 2 Related Work
-2.1 Generalization in Reinforcement Learning
-Contemporary RL methodologies consistently struggle when tasked with generalizing to new environments Song et al.
-(2019).
-Prior studies have crafted an array of training methodologies, such as loss regularization Cobbe et al.
-(2019); Wang et al.
-(2020), successor representation Touati and Ollivier (2021); Touati et al.
-(2022), network architecture design Lee et al.
-(2019); Raileanu and Fergus (2021), and data augmentation Raileanu et al.
-(2021); Hansen and Wang (2021).
-Despite this, there are few investigations into the mechanism behind policy generalization in unseen environments and its relationship with the distribution of those environments.
+Generalization in RL.
+Current RL methods always struggle to generalize to new tasks (Song et al.
+2019).
+Prior works have studied an array of training strategies, like loss regularization (Cobbe et al.
+2019; Wang et al.
+2020), successor representation (Touati and Ollivier 2021; Touati, Rapin, and Ollivier 2022), network architecture design (Lee et al.
+2019; Raileanu and Fergus 2021), lifelong learning (Chen and Liu 2018; Mendez, van Seijen, and Eaton 2022), data augmentation (Raileanu et al.
+2021; Hansen and Wang 2021), etc.
+Besides these, some works investigate the connection between policy generalization and the distribution of those environments.
 Ghosh et al.
-took the initial step in demonstrating that generalizing to unseen environments introduces partial observability, thereby rendering deterministic Markovian policies sub-optimal.
-Also, some studies Lee et al.
-(2020); Ghosh et al.
-(2021) have experimentally indicated that stochastic or non-Markovian policies can yield superior generalization compared to deterministic Markovian policies.
-Nevertheless, expressive capabilities of differing hypothesis sets and their connection to the environment distribution remain understudied areas.
-This knowledge gap is significant for understanding current algorithms and the design of more robust algorithms with enhanced generalization capacities.
-There are several topics closely related to generalization, like multi-task RL and meta RL.
-Multi-task RL Yang et al.
-(2020); Sodhani et al.
-(2021); Lee et al.
-(2022) primarily aims for the trained agent to excel across all training tasks.
-For boosting the generalization, Meta RL seeks to enable the trained agents to adapt to new tasks with as little data as possible, including gradient-based Finn et al.
-(2017) and context-based methods Duan et al.
-(2016); Rakelly et al.
-(2019).
-There are also some model-based methods Nagabandi et al.
-(2018); Clavera et al.
-(2018); Zhao et al.
-(2021) utilize learned models to boost the sample efficiency.
-Though some context-based methods like VariBAD Zintgraf et al.
-(2019) show zero-shot generalization ability, it is still significant to analyze the generalization in RL and design corresponding algorithms.
-2.2 World Models
-The concept of world models Ha and Schmidhuber (2018) aims to better learn environmental representations, which has potential advantages for generalization as it can capture invariant features across tasks.
-Classical methods utilize the Recurrent State Space Model (RSSM) Hafner et al.
-(2019b) for planning like PlaNet Hafner et al.
-(2019b) and policy learning like Dreamer Hafner et al.
-(2019a, 2020, 2023).
-Subsequent research explored reconstruction-free world models Deng et al.
-(2022), preserving representational capacity by avoiding reconstructing task-irrelevant elements.
-Also, techniques like temporal predictive coding Nguyen et al.
-(2021), cooperative reconstruction Fu et al.
-(2021), and Denoised MDP Wang et al.
-(2022) have been proposed for more effective task-relevant information encoding.
-Recent works Chen et al.
-(2021); Robine et al.
-(2023) integrate Transformer Vaswani et al.
-(2017) into world models for capturing long-term dependencies.
-Other studies propose using world models to extract environmental invariant features by learning from videos Seo et al.
-(2022) or exploration Sekar et al.
-(2020); Xu et al.
-(2022), and then fine-tuning to new tasks.
+demonstrates that generalizing to unseen environments introduces partial observability, thereby rendering deterministic Markovian policies sub-optimal.
+Also, some studies (Lee et al.
+2020; Ghosh et al.
+2021) experimentally indicate that stochastic or non-Markovian policies can improve the generalization ability.
+However, the expressive abilities of differing hypothesis sets and their connection to the environment distribution, which are significant for developing more generalizable agents, remain understudied.
+Multi-task RL and Meta RL.
+These two topics are closely related to generalization in RL.
+Multi-task RL (Yang et al.
+2020; Sodhani, Zhang, and Pineau 2021; Lee et al.
+2022; Xu et al.
+2022a) primarily aims to excel across all training tasks but is difficult to zero-shot generalize to unseen tasks.
+For boosting the generalization, Meta RL seeks to enable the trained agents to adapt to new tasks with few episodes, including gradient-based (Finn, Abbeel, and Levine 2017) and context-based methods (Duan et al.
+2016; Rakelly et al.
+2019).
+There are also some model-based methods (Nagabandi et al.
+2018; Rimon et al.
+2024) utilize learned models to boost the sample efficiency.
+Though some context-based methods like VariBAD (Zintgraf et al.
+2019) show zero-shot generalization ability, it is still significant to directly analyze the generalization in RL and design corresponding algorithms.
+World Models.
+World models (Ha and Schmidhuber 2018) aims to better learn environmental representations, which has potential advantages for generalization as it can capture invariant features across tasks.
+Classical methods utilize the Recurrent State Space Model (RSSM) (Hafner et al.
+2019b) for planning (Hafner et al.
+2019b) and policy learning (Hafner et al.
+2019a, 2020, 2023).
+Subsequent research explored reconstruction-free world models (Deng, Jang, and Ahn 2022), temporal predictive coding (Nguyen et al.
+2021), cooperative reconstruction (Fu et al.
+2021), and Denoised MDP (Wang et al.
+2022) for more effective task-relevant information encoding.
+World models are also utilized to extract environmental invariant features by learning from videos (Seo et al.
+2022) or exploration (Sekar et al.
+2020; Xu et al.
+2022b), and then fine-tuning to new tasks.
 However, most world models are designed for the single-task setting and struggle to manage multiple tasks without fine-tuning, limiting their effectiveness for zero-shot generalization to unseen tasks.
 3 Preliminary
-We consider the setting with a task distribution $\mathcal{T}$ of MDPs, where different tasks own the same dynamic and different rewards.
-Formally, each MDP $\mathcal{M}\sim\mathcal{T}$ is represented as $\mathcal{M}=(\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R}_{\mathcal{M}},\gamma)$.
+We consider the setting with a task distribution $\mathcal{T}$ of Partially observable Markov decision processes (POMDPs), where different tasks own the same dynamic and different rewards.
+Formally, each POMDP $\mathcal{M}\sim\mathcal{T}$ can be represented as $\mathcal{M}=(\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R}_{\mathcal{M}},\Omega,\mathcal{O})$.
 Here $\mathcal{S}$ and $\mathcal{A}$ denote the state and action spaces, respectively.
-For $\forall(s,a)\in\mathcal{S}\times\mathcal{A}$, $\mathcal{P}(\cdot|s,a)$ is a distribution over $\mathcal{S}$, representing the dynamic, $\mathcal{R}_{\mathcal{M}}(s,a)$ is the reward function of $\mathcal{M}$, and $\gamma\in(0,1)$ is a discount factor.
-Although deterministic Markovian policy is well known to be optimal for handling the single task setting, some recent work Ghosh et al.
-(2021) has demonstrated that generalizing to unseen environments introduces partial observability, thus deterministic Markovian policy will be sub-optimal and we need to encode more historical information.
-Also, previous multi-task methods Rakelly et al.
-(2019); Zintgraf et al.
-(2019) will utilize all historical information to choose the action.
-Formally, given policy $\pi$, at each timestep $t$, the agent will use the whole history trajectory $(s_{0},a_{0},r_{0},s_{1},a_{1},...,s_{t})$ to sample action $a_{t}$, arrive at the next state $s_{t+1}\sim\mathcal{P}(\cdot|s_{t},a_{t})$, and get the current reward $r_{t}=\mathcal{R}_{\mathcal{M}}(s_{t},a_{t})$.
-The performance of policy $\pi$ in $\mathcal{M}$ is defined as the expected discounted return of the trajectory $\tau$: $J_{\mathcal{M}}(\pi)=\mathbb{E}_{\tau\sim\pi}\left[R(\tau)\triangleq\sum_{t=0}^{\infty}\gamma^{t}r_{t}\right]$, where $R(\tau)$ represents the discounted return of $\tau$.
-And we hope to find the optimal policy over $\mathcal{T}$, i.e., $\max_{\pi}\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}\left[J_{\mathcal{M}}(\pi)\right]$.
+For $\forall(s,a)\in\mathcal{S}\times\mathcal{A}$, $\mathcal{P}(\cdot|s,a)$ is the state transition probability that is Markovian, $\mathcal{R}_{\mathcal{M}}(s,a)$ is its reward function.
+The state is unobserved for the agent, which can only access the observation from the observation space $\Omega$ calculated by the observation function $\mathcal{O}(\cdot|s)$.
+Following previous meta RL and generalization methods, we consider policies that encoder all historical information (we prove its necessity in Sec.
+4.3).
+Formally, at each timestep $t$, the agent with the policy $\pi$ will use the whole history trajectory $(o_{0},a_{0},r_{0},o_{1},...,o_{t})$ to sample action $a_{t}$, arrive at the next state $s_{t+1}\sim\mathcal{P}(\cdot|s_{t},a_{t})$, and get the current reward $r_{t}=\mathcal{R}_{\mathcal{M}}(s_{t},a_{t})$.
+The performance of policy $\pi$ in $\mathcal{M}$ is defined as the expected discounted return: $J_{\mathcal{M}}(\pi)=\mathbb{E}_{\tau\sim\pi}\left[R(\tau)\triangleq\sum_{t=0}^{\infty}\gamma^{t}r_{t}\right]$.
+Our objective is to maximize the return over $\mathcal{T}$, i.e., $\max_{\pi}\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}\left[J_{\mathcal{M}}(\pi)\right]$.
 In practice, given the task distribution $\mathcal{T}$, we sample $M$ training tasks $\{\mathcal{M}_{m}\}_{m=1}^{M}$ for optimizing the agent, i.e., maximizing $\frac{1}{M}\sum_{m=1}^{M}J_{\mathcal{M}_{m}}(\pi)$.
-In the testing stage, we will sample $N$ unseen testing tasks $\{\mathcal{M}_{M+n}\}_{n=1}^{N}$ to evaluate its generalization ability, i.e., evaluating $\frac{1}{N}\sum_{n=1}^{N}J_{\mathcal{M}_{M+n}}(\pi)$.
+In the testing stage, we will sample $N$ unseen test tasks $\{\mathcal{M}_{M+n}\}_{n=1}^{N}$ to evaluate its generalization ability, i.e., evaluating $\frac{1}{N}\sum_{n=1}^{N}J_{\mathcal{M}_{M+n}}(\pi)$.
 4 Task Aware Dreamer
-In this section, we propose a novel framework of Task Aware Dreamer (TAD) for handling task generalization, including the reward-informed world models and the optimization.
-Finally, we will provide theoretical analyses of designs in TAD.
+In this section, we first demonstrate that general world models benefit task generalization.
+Then we introduce our reward-informed world models and propose a novel framework of Task Aware Dreamer (TAD) for handling task generalization.
+Finally, we provide theoretical analyses of designs in TAD.
 4.1 Reward-Informed World Models
-Our first observation is that model-based methods, in addition to proving to improve sample efficiency in the single-task setting Janner et al.
-(2019); Hafner et al.
-(2019a), are also effective in narrowing down the hypothesis space of the optimal Q function when considering task generalization:
-Theorem 1 (Informally, a detailed version and proof are in Appendix B.1).
-Given $M$ tasks $\{\mathcal{M}_{m}\}_{m=1}^{M}$ and corresponding dataset $D^{m}$ sampled from $\mathcal{M}_{m}$.
-We can consider three hypothesis classes of optimal Q function $\mathcal{H}_{1},\mathcal{H}_{2},\mathcal{H}_{3}$ as
-  • $\mathcal{H}_{1}$ includes all optimal Q functions satisfying the Bellman optimality equation for all transitions in $\mathcal{D}_{m}$.
-  • For any dynamics $p_{m}$ satisfying the dynamic transition in $\mathcal{D}_{m}$, $\mathcal{H}_{2}$ includes all optimal Q functions satisfying the Bellman optimality equation for all transitions consistent with $p_{m}$.
-  • For any dynamics $p$ satisfying all dynamic transition in $\{\mathcal{D}_{m}\}_{m=1}^{M}$, $\mathcal{H}_{3}$ includes all optimal Q functions satisfying the Bellman optimality equation for all transitions consistent with $p$.
-We have that $\mathcal{H}_{3}\subseteq\mathcal{H}_{2}\subseteq\mathcal{H}_{1}$.
-Here $\mathcal{H}_{2}\subseteq\mathcal{H}_{1}$ utilizes the generalization ability of the trained world model, which is similar to the single-task setting.
-Moreover, $\mathcal{H}_{3}\subseteq\mathcal{H}_{2}$ is based on similar dynamic structures across different tasks, which benefits learning the general world model.
-In summary, this result inherits the previous result Young et al.
-(2022) where the reward function is unchanged and inspires us that training a general world model can narrow down the hypothesis of possible Q function as well as benefit training the agent.
-As traditional world models Hafner et al.
-(2019b, a) mainly designed for the single-task setting, we first analyze the probabilistic graphical model of the task-distribution setting in Fig.
-1(b) and further build reward-informed world models for handling the task distribution.
-As shown in Fig.
-1(b), the reward does not only rely on the previous states and actions but also on the current task.
-Thus the joint distribution can be calculated as
+Our first observation is that general world models are effective in narrowing down the hypothesis space of the optimal Q function when handling task generalization:
+Theorem 1 (Proof in Appendix A.1).
+Set $\mathcal{Q}$ as the space of observation-action Q functions.
+Given $M$ tasks $\{\mathcal{M}_{m}\}_{m=1}^{M}$ and corresponding dataset $\mathcal{D}_{m}=\{(o_{t}^{m},a_{t}^{m},r_{t}^{m},o_{t+1}^{m})\}$, we set the product space $\mathcal{H}=\mathcal{Q}^{M}$ composed of M spaces, i.e., $\forall\{q_{m}\}_{m=1}^{M}\in\mathcal{H}$, $q_{m}:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ belongs to $\mathcal{Q}$.
+Considering the following three hypothesis classes $\mathcal{H}_{1},\mathcal{H}_{2},\mathcal{H}_{3}\subseteq\mathcal{H}$:
+$\begin{split}\mathcal{H}_{1}=&\{(q_{m})_{m=1}^{M}|q_{m}(o_{t}^{m},a_{t}^{m})=r_{t}^{m}+\gamma\max_{a^{\prime}}q_{m}(o_{t+1}^{m},a^{\prime})\}\\ \mathcal{H}_{2}=&\{(q_{m})_{m=1}^{M}|\exists(p_{m})_{m=1}^{M}:p_{m}(o_{t}^{m},a_{t}^{m})=o_{t+1}^{m},\\ &\exists(r_{m})_{m=1}^{M}:r_{m}(o_{t}^{m},a_{t}^{m})=r_{t}^{m},\\ &q_{m}(o,a)=r_{m}(o,a)+\gamma\max_{a^{\prime}}q_{m}(p_{m}(o,a),a^{\prime}),\forall o,a\}\\ \mathcal{H}_{3}=&\{(q_{m})_{m=1}^{M}|\exists p:p(o_{t}^{m},a_{t}^{m})=o_{t+1}^{m},\\ &\exists(r_{m})_{m=1}^{M}:r_{m}(o_{t}^{m},a_{t}^{m})=r_{t}^{m},\\ &q_{m}(o,a)=r_{m}(o,a)+\gamma\max_{a^{\prime}}q_{m}(p(o,a),a^{\prime}),\forall o,a\}.\end{split}$
+Then we have $\mathcal{H}_{3}\subseteq\mathcal{H}_{2}\subseteq\mathcal{H}_{1}$.
+Here $\mathcal{H}_{1}$ own Q functions satisfying optimal Bellman equation with data from $\mathcal{D}_{m}$, while $\mathcal{H}_{2}$ and $\mathcal{H}_{3}$ own Q functions satisfying optimal Bellman function with data from world models, which are trained with the single dataset and all datasets, respectively.
+Consequently, $\mathcal{H}_{2}\subseteq\mathcal{H}_{1}$ utilizes the generalization ability of world models, which extends the results in the fixed-task setting (Young et al.
+2023).
+Moreover, $\mathcal{H}_{3}\subseteq\mathcal{H}_{2}$ illustrate similar dynamic structures across different tasks benefit learning general world models.
+Consequently, training general world models can narrow down the hypothesis of possible Q function and benefit agent training.
+As existing world models (Hafner et al.
+2019b, a) are mainly designed for the single-task setting, we propose reward-informed world models for the task-distribution setting.
+We first analyze the probabilistic graphical model of the task-distribution setting shown in Fig.
+2, where rewards do not only rely on the previous states and actions but also the current task.
+Thus we can calculate the joint distribution as
 $p(s_{1:T},o_{1:T},r_{1:T},a_{1:T-1},\mathcal{M})$ (1)
 $=$ $p(\mathcal{M})\prod_{t=1}p(s_{t+1}|s_{t},a_{t})p(o_{t}|s_{t})p(r_{t}|s_{t},\mathcal{M}),$
-and we choose the reward-informed inference model (analyses about the necessity of utilizing reward-informed policy in task generalization is in Sec. 4.3) to approximate state posteriors as
+and we choose the reward-informed inference model (the necessity of reward-informed policy in task generalization is discussed in Sec. 4.3) to approximate state posteriors as
 $q(s_{1:T}|o_{1:T},a_{1:T},r_{1:T})=\prod_{t=1}^{T}q(s_{t}|s_{t-1},a_{t-1},r_{t-1},o_{t}).$ (2)
-Further, we can construct the variational lower bound of the log-likelihood on the data as
+Based on this inference model, we can construct the variational lower bound of the log-likelihood on the data as
 $\begin{split}&\ln p(o_{1:T},r_{1:T},\mathcal{M}|a_{1:T})\\ \geq&\sum_{t=1}^{T}\mathbb{E}_{q(s_{t}|o_{\leq t},a_{<t},r_{<t})}\left[\ln p(o_{t},r_{t}|\mathcal{M},s_{t})\right.\\ -&\left.\text{KL}\left(q(s_{t}|o_{\leq t},r_{<t},a_{<t})\|p(s_{t}|s_{t-1},a_{t-1})\right)\right]\\ +&\mathbb{E}_{q(s_{1:T}|o_{1:T},a_{1:T},r_{1:T})}\left[\ln p(\mathcal{M}|s_{1:T})\right].\end{split}$ (3)
-It is a general form of the single task setting Hafner et al.
-(2019b) and the detailed derivation is in Appendix B.2.
-Here the first two terms in Eq.(3) are for reconstructing observations, predicting rewards, and inferring states, which are similar to the single-task setting.
-The last novel term is dedicated to predicting the current task from historical information, which is beneficial for improving the generalization as it encourages inferring the current task context, and we use one-hot task vectors to present $\mathcal{M}$ during training.
-Extending RSSM via Eq.(3), our reward-informed world models consist of:
+This result is a general form of the single-task setting (Hafner et al.
+2019b) and its derivation is in Appendix A.2.
+The first two terms in Eq.(LABEL:eq_elbo) are for reconstructing observations, predicting rewards, and inferring states, which are similar to the single-task setting.
+The last novel term is dedicated to predicting the current task from historical information, which is beneficial for improving the generalization as it encourages inferring the current task context.
+Extending RSSM via Eq.(LABEL:eq_elbo), our reward-informed world models consist of:
 Deterministic state model: $h_{t}=f(h_{t-1},s_{t-1},a_{t-1},r_{t-1}),$
@@
 Task model: $p_{\theta}(\mathcal{M}|h_{t},s_{t}).$
-Here hidden state $h_{t}$ encodes historical states, actions, and rewards, by using gated recurrent unit (GRU) Chung et al.
-(2014) as $f(\cdot)=\text{GRU}(\cdot)$.
-Then the transition model, observation model, and reward model further predict state, observation, and reward respectively.
-Finally, the task model infers the current task to distinguish different tasks’ trajectories.
+Here hidden state $h_{t}$ encodes historical states, actions, and rewards, by using gated recurrent unit (GRU) (Chung et al.
+2014) as $f(\cdot)=\text{GRU}(\cdot)$.
+Then the transition model, observation model, reward model, and task model further predict state, observation, reward, and task context respectively.
 4.2 Optimization
 Based on the above analyses, we now introduce the training of TAD in detail.
-Following previous work Hafner et al.
-(2019a), we adopt an alternating training approach between the reward-informed world models and the policy.
-To balance different tasks, in the stage of collecting data, TAD utilizes different replay buffers $\{\mathcal{D}_{m}\}_{m=1}^{M}$ to store trajectories sampled from $\{\mathcal{M}_{m}\}_{m=1}^{M}$, respectively.
-Then in the training stage, TAD will sample data from each replay buffer and train the reward-informed world models via the objective, which follows Eq.
-(3), as
-$\begin{split}&L_{\text{TAD}}=\sum_{i=1}^{M}\mathbb{E}_{q}\left[\sum_{t=1}^{T}\ln p_{\theta}(o_{t}^{i}|h_{t}^{i},s_{t}^{i})+\sum_{t=1}^{T}\ln p_{\theta}(r_{t}^{i}|h_{t}^{i},s_{t}^{i})\right.\\ &-\left.\sum_{t=1}^{T}D_{\text{KL}}(q(s_{t}^{i}|h_{t}^{i},o_{t}^{i})\|p_{\theta}(s_{t}^{i}|h_{t}^{i}))+\sum_{t=1}^{T}\ln p_{\theta}(\mathcal{M}_{i}|h_{t}^{i},s_{t}^{i})\right].\end{split}$ (4)
+Following previous world models (Hafner et al.
+2019a), we adopt an alternating training approach between the reward-informed world models and the policy.
+To balance different tasks, when collecting data, TAD utilizes $M$ replay buffers $\{\mathcal{D}_{m}\}_{m=1}^{M}$ to store trajectories sampled from $\{\mathcal{M}_{m}\}_{m=1}^{M}$, respectively.
+Then, TAD samples data from each replay buffer and trains the reward-informed world models via the optimization objective following Eq.
+(LABEL:eq_elbo) as
+$\begin{split}&L_{\text{TAD}}=\sum_{i=1}^{M}\mathbb{E}_{q}\left[\sum_{t=1}^{T}\ln p_{\theta}(o_{t}^{i}|h_{t}^{i},s_{t}^{i})+\sum_{t=1}^{T}\ln p_{\theta}(r_{t}^{i}|h_{t}^{i},s_{t}^{i})\right.\\ &-\left.\sum_{t=1}^{T}D_{\text{KL}}(q(s_{t}^{i}|h_{t}^{i},o_{t}^{i})\|p_{\theta}(s_{t}^{i}|h_{t}^{i}))+L_{\text{task}}\right].\end{split}$ (4)
 In Eq.
-(4), the first three items are similar to Dreamer for reconstructing observations, predicting rewards, and inferring states.
-The novel last term is calculated in Eq.
-(3) for encouraging the agent to encode task information into hidden states and infer the current task based on these hidden states, which benefits generalizing to unseen tasks (We demonstrate this in Sec.
+(LABEL:eq_loss), the first three items are similar to Dreamer for reconstructing observations, predicting rewards, and inferring states.
+Besides them, TAD includes an additional task term $L_{\text{task}}$ for predicting different tasks and learning task-aware embedding.
+In detail, we provide two alternatives: cross-entropy and self-contrastive, which are introduced below (We theoretically demonstrate their effectiveness in Sec.
 4.3).
-In terms of training the actor-critic, both of which are parameterized neural networks, we follow the actor-critic learning approach in Dreamer with a more powerful hypothesis set $\Pi_{3}$ (The necessity is shown in Sec.
-4.3).
-When optimizing the policy, TAD samples a series of states from the replay buffer and starts from them to imagine trajectories via interacting with reward-informed world models, which is fixed at this time.
-Here reward-informed world models can capture invariant features, which are beneficial for the agent to train more effectively and gain better generalization.
-After obtaining imagined trajectories, the actor-critic can be optimized via maximizing the $\lambda$-return Schulman et al.
-(2015) and regressing the targets calculated by TD Sutton and Barto (2018), respectively.
+Cross-Entropy.
+The last term in Eq.
+LABEL:eq_elbo indicates that we need to maximize the log probability of the task context to distinguish different tasks via historical information.
+Thus TAD-CE directly maximizes the log probability over different tasks represented by one-hot vectors following previous works (Yang et al.
+2020) and $L_{\text{task}}$ is set as below.
+$\begin{split}L_{\text{task}}=\sum_{t=1}^{T}\ln p_{\theta}(m^{i}|h_{t}^{i},s_{t}^{i}).\end{split}$ (5)
+Supervised-Contrastive.
+Besides directly maximizing the log probability, we further propose TAD-SC, borrowing the idea of supervised contractive learning (Khosla et al.
+2020).
+In detail, TAD-SC keeps the task embeddings of the same task closer and the task embeddings of different tasks far apart, which benefits reward-informed world models to better distinguish different tasks.
+Formally, we assume that the task model maps all sampled data as $\{m_{j}\}_{j=1}^{M\times T}$ and set $L_{\text{task}}$ as
+$\begin{split}L_{\text{task}}=\sum_{j=1}^{M\times T}\sum_{a\in A(j)}\ln\frac{\exp(m_{j}\cdot m_{a}/\tau)}{\sum_{b\neq j}\exp(m_{j}\cdot m_{b}/\tau)},\end{split}$ (6)
+here $A(j)$ is the set of indices that are sampled with the same task of $m_{j}$, and $\tau$ is the temperature parameter, which is set as 0.1 following previous works (Khosla et al. 2020).
+In terms of training the actor-critic that are parameterized neural networks, we extend the actor-critic learning in Dreamer to the task distribution.
+We first sample a series of states from the replay buffer and start from them to imagine trajectories via our reward-informed world models (results are in Fig.
+3), which can capture invariant features and are beneficial for the agent to gain better generalization.
+After obtaining imagined trajectories, the actor-critic is optimized via maximizing the $\lambda$-return (Schulman et al.
+2015) and regressing the TD targets (Sutton and Barto 2018), respectively.
 Algorithm 1 Task Aware Dreamer (TAD)
-0:  $M$ training tasks $\{\mathcal{M}_{m}\}_{m=1}^{M}$, $M$ replay buffers $\{\mathcal{D}_{m}\}_{m=1}^{M}$, $N$ testing tasks $\{\mathcal{M}_{M+n}\}_{n=1}^{N}$, initialize parameters of world models, the policy, and the critic.
-1:  while not converge do
+0:  $M$ training tasks $\{\mathcal{M}_{m}\}_{m=1}^{M}$, $M$ replay buffers $\{\mathcal{D}_{m}\}_{m=1}^{M}$, $N$ test tasks $\{\mathcal{M}_{M+n}\}_{n=1}^{N}$, initialize parameters of world models, the policy, and the critic.
+[…]
 2:     for update $\text{update step}=1,2,...,U$ do
@@
 4:        Calculate the deterministic state $h$ and further calculate model states $s$.
-5:        Update the world models via optimizing Eq. (4).
+5:        Update the world models via optimizing Eq. (LABEL:eq_loss).
 6:        Collect imagined trajectories from each $s$ via the policy and the world models and use these imagined trajectories to update the policy and the critic.
@@
 8:     Collect trajectories from $\mathcal{M}_{m}(m=1,2,...,M)$ and store them into the replay buffer $\mathcal{D}_{m}$.
-9:  end while
+9:  end for
 10:  Evaluate the agent in testing environments $\{\mathcal{M}_{M+n}\}$.
 4.3 Theoretical Analyses
-In this part, we will provide some analyses to show that components of TAD are simple but effective for task generalization.
-Is it necessary to use policies that utilize all historical information for task generalization?
-Below we will introduce $3$ types of widely used policy hypotheses and show that $\Pi_{3}$, used in TAD, is necessary for handling the task generalization.
+Below, we provide theoretical analyses to show that TAD’s components are simple but effective for task generalization.
+Are policies that utilize all historical information in TAD necessary for task generalization?
+Below we introduce $3$ types of widely used policy hypotheses and show that $\Pi_{3}$, used in TAD, is necessary for handling task generalization.
   1. 1.
-    Markovian policy set $\Pi_{1}$ Sutton and Barto (2018); Yarats et al.
-(2021), i.e., $\Pi_{1}=\{\pi|\pi:\mathcal{S}\rightarrow\Delta(\mathcal{A})\}$, here $\Delta(\mathcal{A})$ represents a distribution over $\mathcal{A}$, which is widely used and optimal for the single-task setting;
+    Markovian policy set $\Pi_{1}$ (Sutton and Barto 2018; Yarats et al.
+2021), i.e., $\Pi_{1}=\{\pi|\pi:\mathcal{S}\rightarrow\Delta(\mathcal{A})\}$, here $\Delta(\mathcal{A})$ represents a distribution over $\mathcal{A}$, which is widely used and optimal for the single-task setting;
   2. 2.
-    $\mathcal{S}$-$\mathcal{A}$ memorized policy set $\Pi_{2}$ Hafner et al.
-(2019a, 2020); Lee et al.
-(2020), i.e., $\Pi_{2}=\{\pi|\pi:\mathcal{H}\rightarrow\Delta(\mathcal{A})\}$, here $\mathcal{H}=\cup_{t=1}^{\infty}\mathcal{H}_{t},\mathcal{H}_{t}=(\mathcal{S}\times\mathcal{A})^{t-1}\times\mathcal{S}$;
+    $\mathcal{S}$-$\mathcal{A}$ memorized policy set $\Pi_{2}$ (Hafner et al.
+2019a, 2020; Lee et al.
+2020), i.e., $\Pi_{2}=\{\pi|\pi:\mathcal{H}\rightarrow\Delta(\mathcal{A})\}$, here $\mathcal{H}=\cup_{t=1}^{\infty}\mathcal{H}_{t},\mathcal{H}_{t}=(\mathcal{S}\times\mathcal{A})^{t-1}\times\mathcal{S}$;
   3. 3.
-    $\mathcal{S}$-$\mathcal{A}$$-\mathcal{R}$ memorized policy set $\Pi_{3}$ Rakelly et al.
-(2019); Zintgraf et al.
-(2019), i.e., $\Pi_{3}=\{\pi|\pi:\mathcal{L}\rightarrow\Delta(\mathcal{A})\}$, here $\mathcal{L}=\cup_{t=1}^{\infty}\mathcal{L}_{t},\mathcal{L}_{t}=(\mathcal{S}\times\mathcal{A}\times\mathbb{R})^{t-1}\times\mathcal{S}$.
+    $\mathcal{S}$-$\mathcal{A}$$-\mathcal{R}$ memorized policy set $\Pi_{3}$ (Zintgraf et al.
+2019; Rimon et al.
+2024), i.e., $\Pi_{3}=\{\pi|\pi:\mathcal{L}\rightarrow\Delta(\mathcal{A})\}$, here $\mathcal{L}=\cup_{t=1}^{\infty}\mathcal{L}_{t},\mathcal{L}_{t}=(\mathcal{S}\times\mathcal{A}\times\mathbb{R})^{t-1}\times\mathcal{S}$.
 As illustrated in Fig.
-1(a), naturally $\Pi_{1}\subseteq\Pi_{2}\subseteq\Pi_{3}$.
-Next, we will analyze their expressive ability under task distribution $\mathcal{T}$.
+1, naturally $\Pi_{1}\subseteq\Pi_{2}\subseteq\Pi_{3}$.
+Now we will analyze their expressive ability under task distribution $\mathcal{T}$.
 Denote $J_{\mathcal{T}}^{*}\triangleq\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}\left[\max_{\pi}J_{\mathcal{M}}(\pi)\right]$ as the optimal return under $\mathcal{T}$, and $J_{\mathcal{T}}^{i}\triangleq\max_{\pi\in\Pi_{i}}\left[\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}J_{\mathcal{M}}(\pi)\right]$ as the optimal return for $\Pi_{i}(i=1,2,3)$ under $\mathcal{T}$.
-Our first main result shows that, although $\Pi_{1}\subseteq\Pi_{2}$, they own the same expressive ability, i.e., $J_{\mathcal{T}}^{1}=J_{\mathcal{T}}^{2}$, and are both sub-optimal, summarized as:
-Theorem 2 (Sub-Optimality of $\Pi_{1},\Pi_{2}$. Proof in Appendix B.3).
+Our first result shows that, although $\Pi_{1}\subseteq\Pi_{2}$, they own the same expressive ability, i.e., $J_{\mathcal{T}}^{1}=J_{\mathcal{T}}^{2}$, and are both sub-optimal:
+Theorem 2 (Sub-Optimality of $\Pi_{1},\Pi_{2}$. Proof in Appendix A.3).
 We set ¯$\bar{\mathcal{M}}=(\mathcal{S},\mathcal{A},\mathcal{P},\bar{\mathcal{R}},\gamma)$, here ¯$\bar{\mathcal{R}}=\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}[\mathcal{R}_{\mathcal{M}}]$.
@@
 Also, as $\Pi_{1}$ and $\Pi_{2}$ only choose actions via current state or historical state-action pairs, they cannot distinguish different tasks and are both sub-optimal.
-To quantitatively analyze the characteristic of $\mathcal{T}$ and the gap between $J_{\mathcal{T}}^{1},J_{\mathcal{T}}^{2}$ and $J_{\mathcal{T}}^{*}$, we propose a novel metric named Task Distribution Relevance (TDR) of the distribution $\mathcal{T}$ as
+To quantitatively analyze the characteristic of $\mathcal{T}$ and the gap between $J_{\mathcal{T}}^{1},J_{\mathcal{T}}^{2}$ and $J_{\mathcal{T}}^{*}$, we propose a novel metric Task Distribution Relevance (TDR) of the distribution $\mathcal{T}$ as
 Definition 1 (Task Distribution Relevance).
 For any task distribution $\mathcal{T}$ and state $s$, the Task Distribution Relevance of $\mathcal{T}$ and $s$ is defined as
-[…]
+$\begin{split}D_{\text{TDR}}(\mathcal{T},s)&=\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}[\max_{a}Q_{\mathcal{M}}^{*}(s,a)]\\ &-\max_{a}\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}\left[Q_{\mathcal{M}}^{*}(s,a)\right].\end{split}$ (7)
 Intuitively, TDR describes the relevance of $\mathcal{T}$ via optimal $Q$ functions, which determine the distribution of optimal actions in corresponding tasks. Based on TDR, we can bound the gap:
-Theorem 3 (Proof in Appendix B.4).
+Theorem 3 (Proof in Appendix A.4).
 Assume $\pi_{\mathcal{M}}^{*}=\mathop{\arg\max}_{\pi}J_{\mathcal{M}}(\pi)$, for $\forall\pi\in\Pi_{1}$, we have
-[…]
-[…]
+$\begin{split}J_{\mathcal{T}}^{*}-\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}\left[J_{\mathcal{M}}(\pi)\right]\geq&\frac{1}{1-\gamma}\mathbb{E}_{s\sim d_{\mathcal{M},\pi}}[D_{\text{TDR}}(\mathcal{T},s)]\end{split}$ (8)
+Thus $J_{\mathcal{T}}^{*}-J_{\mathcal{T}}^{2}=J_{\mathcal{T}}^{*}-J_{\mathcal{T}}^{1}\geq\frac{1}{1-\gamma}\mathbb{E}_{s\sim d_{\mathcal{M},\pi^{*}}}\left[D_{\text{TDR}}(\mathcal{T},s)\right]$, here $\pi^{*}=\arg\max_{\pi\in\Pi_{1}}J_{\bar{\mathcal{M}}}(\pi)$.
 Theorem 3 demonstrates that the gap between $J_{\mathcal{T}}^{1},J_{\mathcal{T}}^{2}$ and $J_{\mathcal{T}}^{*}$ is related to the TDR of $\mathcal{T}$.
@@
 This conclusion is further verified empirically in the experimental section.
-Moreover, we can show that $J_{\mathcal{T}}^{1},J_{\mathcal{T}}^{2}$ might be arbitrarily small when $J_{\mathcal{T}}^{3}$ might be arbitrarily close to $J_{\mathcal{T}}^{*}$ (More details are in Appendix B.5), which demonstrates that $\Pi_{3}$ owns stronger expressive ability.
+Moreover, we can show that $J_{\mathcal{T}}^{1},J_{\mathcal{T}}^{2}$ might be arbitrarily small when $J_{\mathcal{T}}^{3}$ might be arbitrarily close to $J_{\mathcal{T}}^{*}$ (More details are in Appendix A.5), which demonstrates that $\Pi_{3}$ owns stronger expressive ability.
 Consequently, it is necessary to utilize $\Pi_{3}$ to distinguish different tasks for enhancing expression ability and generalization over the task distribution.
@@
 As the input space of $\Pi_{3}$ is $\mathcal{L}$, i.e., all partial trajectories, our major result analyzes the relation between the policy $\pi\in\Pi_{3}$ and the task posterior $p(\mathcal{M}|l),l\in\mathcal{L}$.
-Notice that $p_{\theta}(\mathcal{M}|h_{t},s_{t})$ can be regard a deformation of $p(\mathcal{M}|l),l\in\mathcal{L}$ as $h_{t}$ encodes all historical information.
-Theorem 4 (Informally, detailed analyses and proof are in Appendix B.6).
+As $h_{t}$ encodes all historical information, $p_{\theta}(\mathcal{M}|h_{t},s_{t})$ can be rewritten as $p(\mathcal{M}|l),l\in\mathcal{L}$.
+Theorem 4 (Informally, detailed analyses and proof are in Appendix A.6).
 For any policy $\pi\in\Pi_{3}$, we have
-$\begin{split}&J_{\mathcal{T}}^{*}-\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}\left[J_{\mathcal{M}}(\pi)\right]\\ =&\frac{1}{1-\gamma}\int_{\mathcal{L}}p(l)dl\left[\int p(\mathcal{M}|l)\max_{a}Q_{\mathcal{M}}^{*}(l,a)d\mathcal{M}\right.\\ &\qquad\qquad\qquad-\left.\max_{a}\int p(\mathcal{M}|l)Q_{\mathcal{M}}^{*}(l,a)d\mathcal{M}\right],\end{split}$ (7)
+$\begin{split}&J_{\mathcal{T}}^{*}-\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}\left[J_{\mathcal{M}}(\pi)\right]\\ =&\frac{1}{1-\gamma}\int_{\mathcal{L}}p(l)\left[\int p(\mathcal{M}|l)\max_{a}Q_{\mathcal{M}}^{*}(l,a)d\mathcal{M}\right.\\ &\qquad\ -\left.\int_{a,\mathcal{M}}\pi(a|l)p(\mathcal{M}|l)Q_{\mathcal{M}}^{*}(l,a)dad\mathcal{M}\right]dl,\end{split}$ (9)
 here $p(l)$ is a distribution of $\mathcal{L}$ related to $\mathcal{T},\pi$ and $p(\mathcal{M}|l)$ is the task posterior related to $\pi$.
 Consequently, maximizing $p(\mathcal{M}|l)$, i.e., making the distribution of $p(\mathcal{M}|l)$ to be closer to some Dirac distribution, can significantly reduce the right part of Eq.
-(7), thus is effective for improving the generalization ability of $\pi$.
-More details and discussion are also provided in Appendix B.6.
-Tasks Hypothesis Walker-stand&\&&walk Acrobot-swingup&\&&sparse Reacher-easy&\&&hard Walker-walk&\&&prostrate
-CURL $\Pi_{1}$ 563.4 $\pm$ 16.0 11.0 $\pm$ 1.8 552.5 $\pm$ 88.9 441.0 $\pm$ 7.7
-SAC+AE $\Pi_{1}$ 683.8 $\pm$ 164.3 11.7 $\pm$ 5.4 669.5 $\pm$ 99.4 445.8 $\pm$ 2.6
-PlaNet $\Pi_{2}$ 990.6 $\pm$ 3.8 12.2 $\pm$ 9.8 341.0 $\pm$ 228.3 445.4 $\pm$ 5.1
-Dreamer $\Pi_{2}$ 992.3 $\pm$ 1.7 159.9 $\pm$ 108.4 506.1 $\pm$ 382.9 449.7 $\pm$ 12.5
-TAD $\Pi_{3}$ 993.9 $\pm$ 1.0 217.3 $\pm$ 107.1 823.3 $\pm$ 237.4 913.7 $\pm$ 11.5
+(LABEL:eq_pi3), thus is effective for improving the generalization ability of $\pi$.
+More details and discussion are also provided in Appendix A.6.
+Algorithm Hypothesis Cartpole-balance Walker-stand&walk Cheetah-run&run__\__back
+&balance__\__sparse &prostrate&flip &flip&flip__\__back
+CURL $\Pi_{1}$ 994.5 $\pm$ 3.6 254.1 $\pm$ 9.2 229.7 $\pm$ 10.9
+SAC+AE $\Pi_{1}$ 992.5 $\pm$ 2.6 256.9 $\pm$ 5.9 225.8 $\pm$ 10.1
+PlaNet $\Pi_{2}$ 309.5 $\pm$ 59.9 606.7 $\pm$ 152.9 244.8 $\pm$ 17.8
+Dreamer $\Pi_{2}$ 974.2 $\pm$ 5.8 722.2 $\pm$ 12.6 241.1 $\pm$ 19.5
+MAMBA $\Pi_{3}$ 994.7 $\pm$ 3.2 436.8 $\pm$ 116.1 375.6 $\pm$ 44.0
+TAD-CE (Ours) $\Pi_{3}$ 998.9 $\pm$ 0.4 778.9 $\pm$ 63.1 549.6 $\pm$ 28.6
+TAD-SC (Ours) $\Pi_{3}$ 982.6 $\pm$ 2.0 807.8 $\pm$ 85.2 588.8 $\pm$ 20.2
 […]
-Tasks Hypothesis Cheetah__\__speed Pendulum__\__angle Walker__\__speed
+Algorithms Hypothesis Cheetah__\__speed Pendulum__\__angle Walker__\__speed
 Train Test Train Test Train Test
-CURL $\Pi_{1}$ 211.7 $\pm$ 13.7 57.4 $\pm$ 26.6 137.1 $\pm$ 8.6 57.4 $\pm$ 21.8 127.0 $\pm$ 33.7 77.5 $\pm$ 11.5
-SAC+AE $\Pi_{1}$ 182.2 $\pm$ 7.6 115.2 $\pm$ 10.1 144.0 $\pm$ 4.5 68.6 $\pm$ 3.7 136.8 $\pm$ 34.4 27.5 $\pm$ 10.5
-PlaNet $\Pi_{2}$ 176.6 $\pm$ 25.9 83.0 $\pm$ 52.2 99.8 $\pm$ 28.3 80.5 $\pm$ 24.2 173.9 $\pm$ 19.3 58.4 $\pm$ 23.7
-Dreamer $\Pi_{2}$ 250.2 $\pm$ 9.6 3.0 $\pm$ 2.2 92.2 $\pm$ 19.9 79.6 $\pm$ 14.8 197.6 $\pm$ 24.6 10.0 $\pm$ 6.5
-TAD $\Pi_{3}$ 951.9 $\pm$ 3.3 876.9 $\pm$ 51.1 344.8 $\pm$ 2.8 132.6 $\pm$ 14.2 241.2 $\pm$ 36.8 156.5 $\pm$ 129.6
+CURL $\Pi_{1}$ 211.7 $\pm$ 13.7 57.4 $\pm$ 26.6 140.2 $\pm$ 1.7 46.1 $\pm$ 29.8 127.0 $\pm$ 33.7 77.5 $\pm$ 11.5
+SAC+AE $\Pi_{1}$ 182.2 $\pm$ 7.6 115.2 $\pm$ 10.1 130.6 $\pm$ 12.2 89.0 $\pm$ 25.7 136.8 $\pm$ 34.4 27.5 $\pm$ 10.5
+PlaNet $\Pi_{2}$ 176.6 $\pm$ 25.9 83.0 $\pm$ 52.2 92.5 $\pm$ 31.3 70.6 $\pm$ 18.4 173.9 $\pm$ 19.3 58.4 $\pm$ 23.7
+Dreamer $\Pi_{2}$ 250.2 $\pm$ 9.6 3.0 $\pm$ 2.2 87.8 $\pm$ 16.1 87.3 $\pm$ 20.5 197.6 $\pm$ 24.6 10.0 $\pm$ 6.5
+MAMBA $\Pi_{3}$ 568.0 $\pm$ 229.1 475.7 $\pm$ 316.6 153.8 $\pm$ 34.7 121.1 $\pm$ 29.2 104.5 $\pm$ 25.7 99.9 $\pm$ 43.1
+TAD-CE (Ours) $\Pi_{3}$ 937.4 $\pm$ 9.8 909.8 $\pm$ 21.9 283.9 $\pm$ 16.2 163.8 $\pm$ 53.0 241.2 $\pm$ 36.8 156.5 $\pm$ 129.6
+TAD-SC (Ours) $\Pi_{3}$ 919.3 $\pm$ 21.9 906.7 $\pm$ 21.7 204.4 $\pm$ 62.2 143.9 $\pm$ 70.7 159.0 $\pm$ 36.9 104.3 $\pm$ 43.7
 […]
-Tasks Hypothesis Half-Cheetah-Fwd-Back(1e7) Half-Cheetah-Vel(1e7) Humanoid-Direc-2D(1e6)
-Train&\&&Test Train Test Train Test
+Algorithms Hypothesis Half-Cheetah-Fwd-Back(1e7) Half-Cheetah-Vel(1e7) Humanoid-Direc-2D(1e6)
+Train&Test Train Test Train Test
 PlaNet $\Pi_{2}$ 30.5 $\pm$ 42.9 -198.1 $\pm$ 1.9 -202.1 $\pm$ 1.8 215.9 $\pm$ 72.3 220.6 $\pm$ 75.3
@@
 VariBAD(zero-shot) $\Pi_{3}$ 1177.5 $\pm$ 94.9 -58.4 $\pm$ 20.6 260.3 $\pm$ 61.6
-TAD $\Pi_{3}$ 1455.8 $\pm$ 78.3 -49.3 $\pm$ 1.9 -47.1 $\pm$ 0.3 339.5 $\pm$ 78.7 335.5 $\pm$ 70.5
+TAD-CE (Ours) $\Pi_{3}$ 1455.8 $\pm$ 78.3 -49.3 $\pm$ 1.9 -47.1 $\pm$ 0.3 339.5 $\pm$ 78.7 335.5 $\pm$ 70.5
+TAD-SC (Ours) $\Pi_{3}$ 1541.5 $\pm$ 114.8 -50.5 $\pm$ 1.6 -49.6 $\pm$ 1.6 260.2 $\pm$ 185.0 249.0 $\pm$ 168.9
 […]
 5 Experiments
-We now present empirical results to answer the questions:
-  • Can we verify the analyses about TDR, i.e., the expressive abilities of $\Pi_{1},\Pi_{2}$ are severely restricted in task distributions with high TDR?
-  • What about the generalization ability of TAD in both image-based and state-based environments?
-  • Can TAD be extended to more general settings like dynamic generalization?
+We now present empirical results to answer the following questions:
+  • Can we verify the analyses about TDR, i.e., the expressive abilities of $\Pi_{1},\Pi_{2}$ are severely restricted in task distributions with high TDR? (Sec. 5.2)
+  • How about TAD’s generalization ability when handling image-based and state-based observations? (Sec. 5.3)
+  • Can TAD be extended to more general settings like dynamic generalization? (Sec. 5.4)
 5.1 Experimental Setup
 Image-based Control.
-To answer the first question and verify analyses about TDR, we first consider lots of different task combinations in DeepMind control suite (DMC) Tassa et al.
-(2018).
-These task combinations include (1) two task combinations of Walker-stand&\&&walk Mu et al.
-(2022) and Acrobot-swingup&\&&sparse, which share the same optimal actions with TDR being 0; (2) two task combinations with non-zero TDR of Reacher-easy&\&&hard Mu et al.
-(2022) and Walker-walk&\&&prostrate.
-For example, Walker-walk and Walker-prostrate hope a two-leg robot to maintain its height above a certain threshold for walking and below another certain threshold respectively, yielding almost opposite optimal Q functions with huge TDR.
-In Appendix C.1, we introduce more details about them and four other task combinations.
-To evaluate generalization, we extend tasks in DMC and design three task distributions: (1) Cheetah_normal-_\__speed($\alpha,\ \beta$), which extends Cheetah-run and hopes the agent to run within the target speed interval $(\alpha-\beta,\ \alpha+\beta)$; (2) Pendulum_normal-_\__angle($\alpha,\ \beta$), which extends Pendulum-swingup and hopes an inverted pendulum to stay its pole within the target angle interval $(\arccos\alpha,\ \arccos\beta)$; and (3) Walker_normal-_\__speed($\alpha,\ \beta$), which is based on Walker-run and requires the planar walker to run within the target speed interval $(\alpha-\beta,\ \alpha+\beta)$.
-For each task distribution, we sample $4$ tasks for training and $2$ additional tasks for testing.
+To verify our analyses of TDR in Sec.
+4.3, we consider several task combinations in DeepMind Control suite (DMC) (Tassa et al.
+2018): (1) Cartpole-balance&balance__\__sparse, which shares the same optimal actions with TDR 0; (2) two task combinations with non-zero TDR of Walker-stand&walk&prostrat&flip, and Cheetah-run&run__\__back&flip&flip__\__flip__\__back, which are widely used in multi-task unsupervised RL (Sekar et al.
+2020; Laskin et al.
+2021).
+For example, Cheetah-run and Cheetah-flip hope a two-leg robot to move forward by running and flip around the torso, respectively (see Fig.
+3), yielding almost opposite optimal Q functions with huge TDR.
+In Appendix C.1, we introduce more details about these task combinations.
+To evaluate generalization with image-based observations, we extend tasks in DMC and design three task distributions: (1) Cheetah__\__speed($\alpha,\ \beta$), which extends Cheetah-run and hopes the agent to run within the target speed interval $(\alpha-\beta,\ \alpha+\beta)$; (2) Pendulum__\__angle($\alpha,\ \beta$), extending Pendulum-swingup to keep pendulum’s pole within the target angle interval $(\arccos\alpha,\ \arccos\beta)$; and (3) Walker__\__speed($\alpha,\ \beta$), which is based on Walker-run and requires the planar walker to run within the target speed interval $(\alpha-\beta,\ \alpha+\beta)$.
+For each task distribution, we sample $4$ training tasks and $2$ additional test tasks.
 More details are in Appendix C.2.
 State-based Control.
-To demonstrate the scalability of TAD, we also consider some state-based continuous control task distributions about robotic locomotion, simulated via the MuJoCo simulator Todorov et al.
-(2012).
-Following previous work Finn et al.
-(2017), our chosen tasks distributions include (1) Half-Cheetah-Fwd-Back, which owns two opposite tasks; and (2) Half-Cheetah-Vel and Humanoid-Direc-2D, which are task distributions with 100 training tasks 30 testing tasks.
+To demonstrate the scalability of TAD, we also consider some state-based continuous robotic control task distributions simulated via MuJoCo (Todorov, Erez, and Tassa 2012).
+Following previous work (Finn, Abbeel, and Levine 2017), we choose tasks distributions: (1) Half-Cheetah-Fwd-Back, which owns two opposite tasks; and (2) Half-Cheetah-Vel and Humanoid-Direc-2D, which are task distributions with 100 training tasks 30 test tasks.
 More details of these task distributions are in Appendix D.
 Baselines.
-In DMC, we choose two model-free methods: CURL Laskin et al.
-(2020) and SAC+AE Yarats et al.
-(2021), which employ Markovian policies and belong to $\Pi_{1}$.
-In addition, we choose two classic world models, PlaNet Hafner et al.
-(2019b) and Dreamer Hafner et al.
-(2019a), which utilize historical state-actions in policies and belong to $\Pi_{2}$.
-In state-based control, besides PlaNet and Dreamer, we also take some meta RL methods like MAML Finn et al.
-(2017), RL2 Duan et al.
-(2016), and VariBAD Zintgraf et al.
-(2019), including zero-shot and few-shot evaluation, as reference.
+In DMC, we choose two model-free methods employing Markovian policies $\Pi_{1}$: CURL (Laskin, Srinivas, and Abbeel 2020) and SAC+AE (Yarats et al.
+2021).
+In addition, we choose two classic world models, PlaNet (Hafner et al.
+2019b) and Dreamer (Hafner et al.
+2019a), which utilize historical state-actions in policies and belong to $\Pi_{2}$.
+Moreover, we take a SOTA model-based meta RL method belonging to $\Pi_{3}$: MAMBA (Rimon et al.
+2024).
+In state-based control, besides PlaNet and Dreamer, we take some meta RL methods like MAML (Finn, Abbeel, and Levine 2017), RL2 (Duan et al.
+2016), and VariBAD (Zintgraf et al.
+2019), including zero-shot and few-shot evaluation, as reference.
 Metrics.
 For the task combinations, we evaluate the average return of all tasks to verify TDR.
-For the task generalization setting, we train agents in training tasks and evaluate their generalization abilities in testing tasks.
-For all experiments, we train 3 policies with different seeds and calculate the mean and std to mitigate the effects of randomness.
-Tasks Acrobot-Cartpole-Pendulum Walker-Cheetah-Hopper Cheetah-run__\__mass Walker-walk__\__mass
-Train&\&&Test Train&\&&Test Train Test Train Test
-Dreamer 541.3 $\pm$ 4.0 299.4 $\pm$ 17.2 717.8 $\pm$ 27.9 711.7 $\pm$ 38.8 889.9 $\pm$ 114.1 903.3 $\pm$ 102.8
-TAD 667.7 $\pm$ 6.4 554.7 $\pm$ 23.6 754.3 $\pm$ 22.9 738.2 $\pm$ 41.1 957.8 $\pm$ 35.1 963.1 $\pm$ 32.6
-[…]
-Figure 2: Visualization of Dreamer (upper) and TAD (below) for handling Walker-walk&\&&Prostrate.
-For each algorithm in each task, we show 4 frame observations of a trajectory in the right part.
-In the left part, we draw t-SNE results of all states in trajectories of Walker-walk (blue points) and Walker-prostrate (red points).
+For the task generalization settings, we train agents in training tasks and evaluate their generalization abilities in test tasks.
+For all experiments, we repeat 5 different random seeds and report the mean $\pm$ std to mitigate the effects of randomness following previous works (Hafner et al.
+2019a; Rimon et al.
+2024).
+Figure 3: Sampled trajectories and imaginary trajectories of TAD for different tasks (Cheetah-run and Cheetah-flip).
+Figure 4: The t-SNE clustering of state embeddings for different tasks sampled via Dreamer, TAD-CE, and TAD-SC.
 5.2 Experimental Results for TDR
-To answer the first question and validate our analyses of TDR, we report the optimal mean $\pm$ std of the average cumulative return of different task combinations in Table 5.
-Comprehensive results of all 8 task combinations and training curves are in Appendix C.3.
-The results show that TAD outperforms baselines in all environments, especially in those with high TDR.
-In task combinations where TDR is 0, like Walker-stand&\&&walk or Acrobot-swingup&\&&sparse, different tasks share the same optimal action and $\Pi_{1},\Pi_{2}$ own the optimal policy, as derived in Theorem 3.
-In these cases, $\Pi_{1},\Pi_{2}$ performs well and TAD can surpass baselines in Acrobot-swingup&\&&sparse since it can better distinguish different tasks.
-In the latter two environments with non-zero TDR, TAD achieves a notable improvement compared with baselines.
-Especially, in Walker-walk&\&&prostrate, where two tasks own almost opposite optimal Q functions and TDR is significantly huge, policies belong to $\Pi_{1},\Pi_{2}$ and can not differentiate different tasks, aligning with our Theorem 3.
-In contrast, TAD can distinguish between tasks by utilizing $\Pi_{3}$ and improves performance conspicuously.
-Moreover, we present some visualization to better understand how TAD works in Fig.
-2.
-We sample trajectories from the trained agent via Dreamer as well as TAD in Walker-walk&\&&prostrate and visualize some observations (right part of Fig.
-2) as well as all states (left part of Fig.
-2), of which the dimensions are reduced for visualization by t-SNE Van der Maaten and Hinton (2008).
-As shown here, states in TAD of different tasks are clearly distinguished, while Dreamer can not differentiate them and perform the same in different tasks.
-Consequently, TAD can successfully distinguish different tasks and handle them at the same time effectively.
-Videos of these trajectories are in supplementary materials.
+To validate our analyses of TDR, we report different algorithms’ performance of different task combinations in Table 1.
+Results show that TAD-CE and TAD-SC outperform all baselines, especially in those environments with high TDR.
+As derived in Theorem 3, in task combinations where TDR is 0 like Cartpole-balance&balance__\__sparse, different tasks share the same optimal action and $\Pi_{1},\Pi_{2}$ own the optimal policy, thus baselines like CURL and Dreamer perform well.
+Conversely, in other task combinations, different tasks’ optimal Q functions may differ a lot (like Chetah-run and Cheetah-run__\__back) and TDR is significantly huge.
+Thus methods with policies in $\Pi_{1},\Pi_{2}$ like CURL and Dreamer can not differentiate different tasks and perform poorly.
+Aligning with Theorem 3, baselines utilizing $\Pi_{3}$ like MAMBA outperform other baselines and TAD improves performance conspicuously.
+Moreover, we present some visualization to better understand how TAD works.
+Fig, 3 shows video predictions of TAD for different tasks (Cheetah-run and Cheetah-flip).
+We use the same agent trained by TAD to sample trajectories for these two tasks and show their trajectories in lines 1 and 3 respectively.
+Given the first 7 steps as the context (their observations are similar but the rewards are different), we directly imagine the future 55 steps by our trained reward-informed world model, of which the results are in lines 2 and 4 of Fig.
+3 respectively.
+As shown here, TAD performs differently to handle different tasks only by receiving different rewards.
+Moreover, our reward-informed world models demonstrate long-term prediction and high-quality reconstruction capabilities, which reveals the potential for training future large-scale world models.
+In Fig.
+4, we sample trajectories from agents trained by Dreamer, TAD-CE, as well as TAD-SC, and visualize the states of different tasks, of which the dimensions are reduced for visualization by t-SNE (Van der Maaten and Hinton 2008).
+As shown here, states in TAD-CE and TAD-SC of different tasks are clearly distinguished, while Dreamer can not differentiate them and perform the same in different tasks.
+This result demonstrates TAD can effectively learn task-aware information and distinguish different tasks.
+Videos of different tasks are provided in supplementary materials.
 5.3 Experimental Results for Task Generalization
-To answer the second question, we report the results under the task generalization setting of image-based and state-based environments in Table 2 and Table 3, respectively.
-In Table 2, TAD shows powerful generalization ability in image-based environments as it can significantly enhance the average cumulative return in both training tasks and testing tasks compared with baselines.
-This demonstrates that TAD can both handle multiple training tasks simultaneously and generalize to unseen testing tasks effectively.
-We also provide some visualization results in Appendix C.4 with corresponding videos in supplementary materials.
-Additionally, in Table 3, we compare TAD with some meta RL methods in state-based environments, of which the training timesteps are 1e7, 1e7, and 1e6, respectively.
-Many context-based meta-RL methods, including RL2, PEARL and VariBAD, utilize reward into policies and belong to $\Pi_{3}$, thus they can distinguish different tasks, which also verify our analyses.
-As shown in Table 3, TAD can achieve a significant improvement in the average cumulative return for both training tasks and testing tasks, since TAD is aware of task information and can generalize to unseen tasks.
+To answer the second question, we report the task generalization results of image-based and state-based environments in Table 2-3.
+In Table 2, TAD shows more powerful generalization abilities than all baselines.
+This demonstrates that TAD can both handle multiple training tasks simultaneously and generalize to unseen test tasks effectively.
+We also provide some visualization results in Appendix C.3 with videos in supplementary materials.
+Additionally, in Table 3, we compare TAD with existing meta RL methods in state-based environments, of which the training timesteps are 1e7, 1e7, and 1e6, respectively.
+Many context-based meta RL methods, including RL2, PEARL, and VariBAD, utilize historical rewards and belong to $\Pi_{3}$, thus they can distinguish different tasks, which also verify our analyses.
+As shown in Table 3, TAD achieves a significant improvement for both training tasks and test tasks, since TAD is aware of task information for generalizing to unseen tasks.
 5.4 Ablation Study
-Reward Signal and Task ID.
+Reward Signals.
 In Fig.
-3, we do ablation studies with two baselines: Dreamer(w/ r), which only integrates $\Pi_{3}$ into Dreamer, i.e., including reward signals into policies, and Dreamer(w/ task ID), which includes one-hot task ID into policies in both training and testing.
-Results show that just providing reward signals can help Dreamer distinguish different tasks, which verifies our analyses in TDR.
-Also, TAD shows superior performance even compared with Dreamer(w/ task ID), which demonstrates that our task model and corresponding ELBO are effective for task generalization.
-Figure 3: Ablation study on Reward Signal and Task ID.
+5, we do ablation studies about the task term in TAD, i.e., we consider Dreamer(w/ r) that only integrates $\Pi_{3}$ into Dreamer.
+Results show that just providing rewards can help Dreamer distinguish different tasks, which verifies our analyses in TDR.
+Also, TAD-CE and TAD-SC show superior performance, demonstrating that our task model and corresponding ELBO are effective for task generalization.
+Also, we do ablation studies on reward signals for model-free methods like CURL in Appendix E.1.
+Figure 5: Ablation study on Reward Signal.
 Extension to Dynamic Generalization.
-To answer the third question, we further evaluate TAD in more general settings with different observations, dynamics, and/or actions.
+To answer the third question, we evaluate TAD in more general settings with different observations, dynamics, and/or actions.
 As TAD utilizes all historical information to infer the environment, it can be directly applied to these settings.
 We design task distribution with different embodiments (Acrobot-Cartpole-Pendulum, Walker-Cheetah-Hopper), as well as different dynamics (Cheetah-run__\__mass, Walker-walk__\__mass).
-Table 4 shows that TAD can achieve much greater performance and better convergence compared to Dreamer, which indicates TAD’s potential in further handling dynamic generalization and even cross-embodiment tasks.
-More details about the environments and results are in Appendix E.1.
-5.5 Limitations and Ethical Statement
+More details about environments and results are in Appendix E.2, where TAD achieves much greater performance compared to baselines and show potential in further handling dynamic generalization and even cross-embodiment tasks.
+5.5 Limitations and Discussion
 In terms of limitations, TAD assumes that the task context is continuously related to historical information for generalizing to unseen tasks.
 Thus TAD might be difficult to generalize in sparse-reward settings.
-We further demonstrate that without extra knowledge or finetuning, zero-shot generalization to unseen tasks with extremely sparse rewards is impossible since there is no way to distinguish different tasks (Appendix B.7).
-Fortunately, in relatively sparse reward settings, we conduct studies in Appendix E.2 to show that TAD can infer the current task and generalize to unseen tasks well.
-There are no serious ethical issues as this is basic research.
+We further demonstrate that without extra knowledge or finetuning, zero-shot generalization to unseen tasks with extremely sparse rewards is impossible since there is no way to distinguish different tasks (Appendix A.7).
+Fortunately, in relatively sparse reward settings, we conduct experiments in Appendix E.3 to show that TAD can infer the current task and generalize to unseen tasks well.
 6 Conclusion
-In this work, we propose a novel framework of TAD that distinguishes different tasks via all historical information and utilizes novel reward-informed world models to capture invariant latent features.
+In this work, we propose a novel framework of TAD that handles different tasks via all historical information and utilizes novel reward-informed world models to capture invariant latent features.
 In TAD, we calculate the corresponding variational lower bound of the data log-likelihood, which includes a novel loss term to distinguish different tasks via states.
@@
 Experiments in image-based and state-based settings demonstrate that TAD can remarkably improve the performance of handling different tasks meanwhile, especially for high TDR ones, and successfully generalize to unseen tasks.
-Appendix A Pseudo Code of TAD
-We provide the pseudo code of TAD in Algorithm 2.
+Appendix A Proof of Theorems
+In this section, we will provide detailed proofs of theorems.
+A.1 The Proof of Theorem 1
+Proof.
+First, we will show that $\mathcal{H}_{2}\subseteq\mathcal{H}_{1}$, following the proof in (Young et al. 2023) that considers the setting that the rewards of all environments are the same.
+$\begin{split}&\{q_{m}\}\in\mathcal{H}_{2}\\ \Leftrightarrow&\exists(p_{m})_{m=1}^{M}:p_{m}(o_{t}^{m},a_{t}^{m})=o_{t+1}^{m},\\ &\exists(r_{m})_{m=1}^{M}:r_{m}(o_{t}^{m},a_{t}^{m})=r_{t}^{m},\\ &q_{m}(o,a)=r_{m}(o,a)+\gamma\max_{a^{\prime}}q_{m}(p_{m}(o,a),a^{\prime}),\forall o,a\\ \Rightarrow&\exists(r_{m})_{m=1}^{M}::p_{m}(o_{t}^{m},a_{t}^{m})=o_{t+1}^{m},\\ &\exists(r_{m})_{m=1}^{M}:r_{m}(o_{t}^{m},a_{t}^{m})=r_{t}^{m},\\ &q_{m}(o_{t}^{m},a_{t}^{m})=r^{m}_{t}+\gamma\max_{a^{\prime}}q_{m}(p_{m}(o_{t}^{m},a_{t}^{m}),a^{\prime}),\forall m,t\\ \Rightarrow&q_{m}(o_{t}^{m},a_{t}^{m})=r^{m}_{t}+\gamma\max_{a^{\prime}}q_{m}(o_{t+1}^{m},a^{\prime}),\forall m,t\\ \Leftrightarrow&\{q_{m}\}\in\mathcal{H}_{1},\end{split}$ (10)
+thus we have $\mathcal{H}_{2}\subseteq\mathcal{H}_{1}$.
+Next, we prove that $\mathcal{H}_{3}\subseteq\mathcal{H}_{2}$, which is mainly because we can utilize similar dynamic structures from different tasks to narrow down the hypothesis spaces of the dynamic model.
+$\begin{split}&\{q_{m}\}\in\mathcal{H}_{3}\\ \Leftrightarrow&\exists p:p(o_{t}^{m},a_{t}^{m})=o_{t+1}^{m},\forall m,t\\ &\exists(r_{m})_{m=1}^{M}:r_{m}(o_{t}^{m},a_{t}^{m})=r_{t}^{m},\\ &q_{m}(o,a)=r_{m}(o,a)+\gamma\max_{a^{\prime}}q_{m}(p(o,a),a^{\prime}),\forall m,o,a\\ \Rightarrow&\exists(p_{m})_{m=1}^{M}:p_{m}(o_{t}^{m},a_{t}^{m})=o_{t+1}^{m},\forall m,t\\ &\exists(r_{m})_{m=1}^{M}:r_{m}(o_{t}^{m},a_{t}^{m})=r_{t}^{m},\\ &q_{m}(o,a)=r_{m}(o,a)+\gamma\max_{a^{\prime}}q_{m}(p_{m}(o,a),a^{\prime}),\forall m,o,a\\ \Leftrightarrow&\{q_{m}\}\in\mathcal{H}_{2},\end{split}$ (11)
+thus we have $\mathcal{H}_{3}\subseteq\mathcal{H}_{2}$. ∎
+A.2 The Derivation of the ELBO
+We use ${\color[rgb]{0,139,69}q}$ to represent $q(s_{1:T}|o_{1:T},a_{1:T},r_{1:T})$, ${\color[rgb]{0,0,1}\hat{q}}$ to present $q(s_{t}|o_{\leq t},a_{<t},r_{<t})$, ${\color[rgb]{1,0,0}\tilde{q}}$ to represent $q(s_{t-1}|o_{\leq(t-1)},r_{<(t-1)},a_{<(t-1)})$, and we have
+$\begin{split}&\ln p(o_{1:T},r_{1:T},\mathcal{M}|a_{1:T})\\ =&\ln\mathbb{E}_{p(s_{1:T}|a_{1:T})}\left[p(o_{1:T},r_{1:T},\mathcal{M}|s_{1:T})\right]\\ =&\ln\mathbb{E}_{p(s_{1:T}|a_{1:T})}\left[p(o_{1:T},r_{1:T}|\mathcal{M},s_{1:T})p(\mathcal{M}|s_{1:T})\right]\\ =&\ln\mathbb{E}_{p(s_{1:T}|a_{1:T})}\left[p(\mathcal{M}|s_{1:T})\prod_{t=1}^{T}p(o_{t},r_{t}|\mathcal{M},s_{t})\right]\\ =&\ln\mathbb{E}_{{\color[rgb]{0,139,69}q}}\left[p(\mathcal{M}|s_{1:T})\prod_{t=1}^{T}p(o_{t},r_{t}|\mathcal{M},s_{t})\frac{p(s_{t}|s_{t-1},a_{t-1})}{q(s_{t}|o_{\leq t},r_{<t},a_{<t})}\right]\\ \geq&\mathbb{E}_{{\color[rgb]{0,139,69}q}}[\ln p(\mathcal{M}|s_{1:T})+\sum_{t=1}^{T}\ln p(o_{t},r_{t}|\mathcal{M},s_{t})]\\ +&\mathbb{E}_{{\color[rgb]{0,139,69}q}}\sum_{t=1}^{T}\left[\ln p(s_{t}|s_{t-1},a_{t-1})-\ln q(s_{t}|o_{\leq t},r_{<t},a_{<t})\right]\\ =&\mathbb{E}_{{\color[rgb]{0,139,69}q}}[\ln p(\mathcal{M}|s_{1:T})]\\ +&\sum_{t=1}^{T}\left[\mathbb{E}_{{\color[rgb]{0,139,69}q}}\ln p(o_{t},r_{t}|\mathcal{M},s_{t})-\mathbb{E}_{{\color[rgb]{0,139,69}q}}\ln\frac{q(s_{t}|o_{\leq t},r_{<t},a_{<t})}{p(s_{t}|s_{t-1},a_{t-1})}\right]\\ =&\mathbb{E}_{{\color[rgb]{0,139,69}q}}[\ln p(\mathcal{M}|s_{1:T})]+\sum_{t=1}^{T}\mathbb{E}_{{\color[rgb]{0,0,1}\hat{q}}}[\ln p(o_{t},r_{t}|\mathcal{M},s_{t})]\\ -&\sum_{t=1}^{T}\mathbb{E}_{{\color[rgb]{0,0,1}\hat{q}}{\color[rgb]{1,0,0}\tilde{q}}}\left[\ln\frac{q(s_{t}|o_{\leq t},r_{<t},a_{<t})}{p(s_{t}|s_{t-1},a_{t-1})}\right]\\ =&\mathbb{E}_{{\color[rgb]{0,139,69}q}}[\ln p(\mathcal{M}|s_{1:T})]+\sum_{t=1}^{T}\mathbb{E}_{{\color[rgb]{0,0,1}\hat{q}}}[\ln p(o_{t},r_{t}|\mathcal{M},s_{t})]\\ -&\sum_{t=1}^{T}\mathbb{E}_{{\color[rgb]{1,0,0}\tilde{q}}}[\mathrm{KL}\left(q(s_{t}|o_{\leq t},r_{<t},a_{<t})\|p(s_{t}|s_{t-1},a_{t-1})\right)].\end{split}$ (12)
+Thus we have proven the ELBO.
+A.3 The Proof of Theorem 2
+Proof.
+We first prove that for $\forall\pi\in\Pi_{2}$, we have $\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}[J_{\mathcal{M}}(\pi)]=J_{\bar{\mathcal{M}}}(\pi)$.
+For any $\mathcal{M}=(\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R}_{\mathcal{M}},\gamma)\sim\mathcal{T}$, we can use the policy $\pi$ to interact with $\mathcal{M}$ and get the trajectory $\tau=(s_{0}^{\mathcal{M}},a_{0}^{\mathcal{M}},r_{1}^{\mathcal{M}},s_{1}^{\mathcal{M}},a_{1}^{\mathcal{M}},r_{2}^{\mathcal{M}},...)$.
+Since the dynamic transition $\mathcal{P}$ is the same for all $\mathcal{M}$ and the policy $\pi\in\Pi_{2}$ only depends on historical states and actions, we naturally have that the distribution of all states and actions $(s_{0}^{\mathcal{M}},a_{0}^{\mathcal{M}},s_{1}^{\mathcal{M}},a_{1}^{\mathcal{M}},...)$ are the same for all $\mathcal{M}\sim\mathcal{T}$ as well as $\bar{\mathcal{M}}$.
+Consequently, we have
+$\begin{split}&\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}[J_{\mathcal{M}}(\pi)]\\ =&\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}\mathbb{E}_{\tau\sim\mathcal{P},\pi}[R_{\mathcal{M}}(\tau)]=\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}\mathbb{E}_{\tau\sim\mathcal{P},\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}r_{t}^{\mathcal{M}}\right]\\ =&\mathbb{E}_{\tau\sim\mathcal{P},\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}[r_{t}^{\mathcal{M}}]\right]\\ =&\mathbb{E}_{\tau\sim\mathcal{P},\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}[\mathcal{R}_{\mathcal{M}}(s_{t}^{\mathcal{M}},a_{t}^{\mathcal{M}})]\right]\\ =&\mathbb{E}_{\tau\sim\mathcal{P},\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}[\bar{\mathcal{R}}(s_{t}^{\mathcal{M}},a_{t}^{\mathcal{M}})]\right]\\ =&\mathbb{E}_{\tau\sim\mathcal{P},\pi}[R_{\bar{\mathcal{M}}}(\tau)]=J_{\bar{\mathcal{M}}}(\pi).\end{split}$ (13)
+It is well known that the optimal policy in single MDP is memory-less, i.e., $\max_{\pi\in\Pi_{2}}J_{\bar{\mathcal{M}}}(\pi)=\max_{\pi\in\Pi_{1}}J_{\bar{\mathcal{M}}}(\pi)$. Consequently, we have
+$\begin{split}J_{\mathcal{T}}^{2}&=\max_{\pi\in\Pi_{2}}\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}[J_{\mathcal{M}}(\pi)]=\max_{\pi\in\Pi_{2}}J_{\bar{\mathcal{M}}}(\pi)\\ &=\max_{\pi\in\Pi_{1}}J_{\bar{\mathcal{M}}}(\pi)=\max_{\pi\in\Pi_{1}}\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}[J_{\mathcal{M}}(\pi)]=J_{\mathcal{T}}^{1}\\ &\leq\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}\left[\max_{\pi\in\Pi_{1}}J_{\mathcal{M}}(\pi)\right]=J_{\mathcal{T}}^{*}.\end{split}$ (14)
+Thus we have proven this result. ∎
+A.4 The Proof of Theorem 3
+Proof.
+Our proof follows some previous work (Kakade and Langford 2002; Ying et al.
+2022).
+First, we consider the bellman equation of value function of $\pi,\pi_{\mathcal{M}}^{*}\in\Pi_{1}$ in $\mathcal{M}$ as
+$\begin{split}V_{\mathcal{M},\pi}(s)=&\sum_{a}\pi(a|s)\left[\mathcal{R}(s,a)+\gamma\sum_{s^{\prime}}\mathcal{P}(s^{\prime}|s,a)V_{\mathcal{M},\pi}(s^{\prime})\right],\\ V_{\mathcal{M},\pi_{\mathcal{M}}^{*}}(s)=&\sum_{a}\pi_{\mathcal{M}}^{*}(a|s)\left[\mathcal{R}(s,a)+\gamma\sum_{s^{\prime}}\mathcal{P}(s^{\prime}|s,a)V_{\mathcal{M},\pi_{\mathcal{M}}^{*}}(s^{\prime})\right].\end{split}$
+Defining $\Delta V(s)\triangleq V_{\mathcal{M},\pi}(s)-V_{\mathcal{M},\pi_{\mathcal{M}}^{*}}(s)$ as the difference of these two value functions, we can further deduce that
+$\begin{split}&V_{\mathcal{M},\pi}(s)-V_{\mathcal{M},\pi_{\mathcal{M}}^{*}}(s)\\ =&\gamma\sum_{a}\Delta\pi(a|s)\sum_{s^{\prime}}\mathcal{P}(s^{\prime}|s,a)V_{\mathcal{M},\pi_{\mathcal{M}}^{*}}(s^{\prime})\\ +&\gamma\sum_{a}\pi(a|s)\sum_{s^{\prime}}\mathcal{P}(s^{\prime}|s,a)\Delta V(s^{\prime})+\sum_{a}\Delta\pi(a|s)\mathcal{R}(s,a)\\ =&\sum_{a}\Delta\pi(a|s)Q_{\mathcal{M},\pi_{\mathcal{M}}^{*}}(s,a)\\ +&\gamma\sum_{a}\pi(a|s)\sum_{s^{\prime}}\mathcal{P}(s^{\prime}|s,a)\Delta V(s^{\prime}),\end{split}$ (15)
+here $\Delta\pi(a|s)=\pi(a|s)-\pi_{\mathcal{M}}^{*}(a|s)$. Since Eq. (LABEL:15) holds for any $s$, thus we calculate its expectation for $s\sim d_{\mathcal{M}}^{\pi_{\mathcal{M}}^{*}}$:
+$\begin{split}&\sum_{s}d_{\mathcal{M}}^{\pi}(s)\Delta V(s)\\ =&\sum_{s}d_{\mathcal{M}}^{\pi}(s)[V_{\mathcal{M},\pi}(s)-V_{\mathcal{M},\pi_{\mathcal{M}}^{*}}(s)]\\ =&\sum_{s}d_{\mathcal{M}}^{\pi}(s)\sum_{a}\Delta\pi(a|s)Q_{\mathcal{M},\pi_{\mathcal{M}}^{*}}(s,a)\\ +&\gamma\sum_{s}d_{\mathcal{M}}^{\pi}(s)\sum_{a}\pi(a|s)\sum_{s^{\prime}}\mathcal{P}(s^{\prime}|s,a)\Delta V(s^{\prime})\\ =&\sum_{s}d_{\mathcal{M}}^{\pi}(s)\sum_{a}\Delta\pi(a|s)Q_{\mathcal{M},\pi_{\mathcal{M}}^{*}}(s,a)\\ +&\sum_{s^{\prime}}\Delta V(s^{\prime})\left[\gamma\sum_{s}d_{\mathcal{M}}^{\pi}(s)\sum_{a}\pi(a|s)\mathcal{P}(s^{\prime}|s,a)\right].\end{split}$ (16)
+Since $\gamma\sum_{s^{\prime}}d_{\mathcal{M}}^{\pi}(s^{\prime})\sum_{a}\pi(a|s^{\prime})\mathcal{P}(s|s^{\prime},a)=d_{\mathcal{M}}^{\pi}(s)-(1-\gamma)\mathcal{P}(s_{0}=s)$, we have
+$\begin{split}\sum_{s}d_{\mathcal{M}}^{\pi}(s)&\Delta V(s)=\sum_{s}d_{\mathcal{M}}^{\pi}(s)\sum_{a}\Delta\pi(a|s)Q_{\mathcal{M},\pi_{\mathcal{M}}^{*}}(s,a)\\ +&\sum_{s^{\prime}}\Delta V(s^{\prime})\left[d_{\mathcal{M}}^{\pi}(s^{\prime})-(1-\gamma)\mathcal{P}(s_{0}=s^{\prime})\right].\end{split}$ (17)
+By moving the second term of the right part in Eq. (17) to the left part, we can deduce that
+$\begin{split}&(1-\gamma)\sum_{s^{\prime}}\Delta V(s^{\prime})\mathcal{P}(s_{0}=s^{\prime})\\ =&\sum_{s}d_{\mathcal{M}}^{\pi}(s)\sum_{a}\Delta\pi(a|s)Q_{\mathcal{M},\pi_{\mathcal{M}}^{*}}(s,a),\end{split}$ (18)
+thus we can calculate that
+[…]
+Consequently, we have
+$\begin{split}&\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}\left[J_{\mathcal{M}}(\pi_{\mathcal{M}}^{*})-J_{\mathcal{M}}(\pi)\right]\\ =&\frac{1}{1-\gamma}\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}\mathbb{E}_{s\sim d_{\mathcal{M},\pi}(\cdot)}\left[\max_{a}Q_{\mathcal{M}}^{*}(s,a)\right.\\ &\left.-\int_{a}\pi(a|s)Q_{\mathcal{M}}^{*}(s,a)da\right]\\ \geq&\frac{1}{1-\gamma}\mathbb{E}_{s\sim d_{\mathcal{M},\pi}(\cdot)}\left[\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}\max_{a}Q_{\mathcal{M}}^{*}(s,a)\right.\\ &\left.-\max_{a}\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}Q_{\mathcal{M}}^{*}(s,a)\right]\\ =&\frac{1}{1-\gamma}\mathbb{E}_{s\sim d_{\mathcal{M},\pi}}\left[D_{\text{TDR}}(\mathcal{T},s)\right].\end{split}$ (20)
+Since $J_{\mathcal{M}}(\pi_{\mathcal{M}}^{*})=\max_{\pi\in\Pi_{1}}J_{\mathcal{M}}(\pi)$, we have
+$\begin{split}&\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}\max_{\pi\in\Pi_{1}}J_{\mathcal{M}}(\pi)-\max_{\pi\in\Pi_{1}}\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}J_{\mathcal{M}}(\pi)\\ \geq&\frac{1}{1-\gamma}\mathbb{E}_{s\sim d_{\mathcal{M},\pi^{*}}}\left[D_{\text{TDR}}(\mathcal{T},s)\right],\end{split}$ (21)
+Thus we have proven this result. ∎
+A.5 Expressive Ability of $\Pi_{3}$
+Proposition 1.
+For $\forall\epsilon_{1},\epsilon_{2}$ satisfying $0<\epsilon_{1}\leq 1,0<\epsilon_{2}\leq 1$, there exists a task distribution $\mathcal{T}$ satisfying that
+$\begin{split}J_{\mathcal{T}}^{1}=J_{\mathcal{T}}^{2}\leq\epsilon_{1},~{}~{}J_{\mathcal{T}}^{3}\geq 1-\epsilon_{2},~{}~{}J_{\mathcal{T}}^{*}=1.\end{split}$ (22)
+Proof.
+Given fixed discount factor $\gamma\in(0,1)$, we first take $n\in\mathbb{N}$ satisfying $n\geq\frac{1}{\epsilon_{1}}$, $A=\epsilon_{2}\frac{n}{n-1}\frac{1-\gamma}{1-\gamma^{n}}$, and $B=\frac{1-\gamma}{\gamma^{n+1}}\left(1-\epsilon_{2}\frac{n}{n-1}\right)$.
+We construct state sets $\mathcal{S}=\{s_{t}^{l}\}(t=0,1,...,\infty,l=1,...,n)$ and action sets $\mathcal{A}=\{a_{j}\}_{j=1}^{n}$.
+Then we construct $n$ tasks $\mathcal{M}_{i}=(\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R}_{i},\gamma),i=1,...,n$, which share the same dynamic $\mathcal{P}$ and different reward functions $\mathcal{R}_{i}$.
+The initial state of each task is $s_{0}^{1}$, and the dynamic as well as reward functions are as below
+$\begin{split}&\mathcal{P}(s_{t}^{l},a_{j})=\mathbb{I}(s=s_{t+1}^{j}),\quad j,l=1,...,n;\ t=0,1,...,\infty\\ &\mathcal{R}_{i}(s_{t}^{l},a_{j})=f(t)\mathbb{I}(i=j),\quad i,j,l=1,...,n;\ t=0,1,...\infty\\ &f(t)=\left\{\begin{array}[]{ll}A,&t\leq n-1\\ B,&t\geq n\\ \end{array}\right.\end{split}$ (23)
+Take $\mathcal{T}$ as the uniform distribution over $\mathcal{M}_{1},...,\mathcal{M}_{n}$, thus we have
+$\begin{split}J_{\mathcal{M}_{i}}^{*}=&A+A\gamma+...+A\gamma^{n-1}+B\gamma^{n}+B\gamma^{n+1}+...\\ =&A\frac{1-\gamma^{n}}{1-\gamma}+B\frac{\gamma^{n+1}}{1-\gamma}=1\\ J_{\mathcal{T}}^{*}=&\frac{1}{n}\sum_{i=1}^{n}J_{\mathcal{M}_{i}}^{*}=1.\\ \end{split}$ (24)
+Since our construction satisfies $\mathbb{E}_{\mathcal{T}}[\mathcal{R}_{i}(s_{k}^{l},a_{j})]=\frac{f(k)}{n}$, for $\forall\pi\in\Pi_{2}$, we have
+$\begin{split}J_{\mathcal{T}}(\pi)=&\frac{1}{n}(A+A\gamma+...+A\gamma^{n-1}+B\gamma^{n}+B\gamma^{n+1}+...)\\ =&\frac{1}{n},\\ J_{\mathcal{T}}^{1}=&J_{\mathcal{T}}^{2}=\max_{\pi\in\Pi_{2}}J_{\mathcal{T}}(\pi)=\frac{1}{n}\leq\epsilon_{1}.\end{split}$ (25)
+Moreover, we can construction an agent $\hat{\pi}\in\Pi_{3}$ that takes action via the historical trajectory $\hat{\tau}_{t}=(\hat{s}_{0},\hat{a}_{0},\hat{r}_{0},...\hat{s}_{t})$:
+$\begin{split}\hat{\pi}(a_{j}|\hat{\tau}_{t})=&\mathbb{I}(j=t+1),\quad t=0,1,...,n-1\\ \hat{\pi}(a_{j}|\hat{\tau}_{t})=&\mathbb{I}(j=i),\quad t=n,...,\infty\\ \end{split}$ (26)
+here $i=\arg\max\{\hat{r}_{0},\hat{r}_{1},...,\hat{r}_{n-1}\}+1$, thus we have
+$\begin{split}&J_{\mathcal{T}}^{3}\geq J_{\mathcal{T}}(\hat{\pi})\\ =&\frac{1}{n}(A+A\gamma+...+A\gamma^{n-1})+B\gamma^{n}+B\gamma^{n+1}+...\\ =&\frac{A}{n}\frac{1-\gamma^{n}}{1-\gamma}+B\frac{\gamma^{n+1}}{1-\gamma}=1-\frac{(n-1)A}{n}\frac{1-\gamma^{n}}{1-\gamma}\\ \geq&1-\epsilon_{2}.\end{split}$ (27)
+Thus we have proven this result. ∎
+A.6 Proof and Discussion of Theorem 4
+In this part, we introduce an informed version of Theorem 4, about why optimizing $p(\mathcal{M}|l)\forall l\in\mathcal{L}$ is beneficial for task generalization, with detailed proofs.
+Recall that we consider the policy hypothesis $\mathcal{H}_{3}$ here, that each policy $\pi:\mathcal{L}\rightarrow\Delta(\mathcal{A}),\mathcal{L}=\cup_{t=1}^{\infty}\mathcal{L}_{t},\mathcal{L}_{t}=(\mathcal{S}\times\mathcal{A}\times\mathbb{R})^{t-1}\times\mathcal{S}$.
+As directly such $\mathcal{S}-\mathcal{A}-\mathcal{R}$ memorized policy is difficult, we consider an alternative MDP as ~$\tilde{\mathcal{M}}=(\mathcal{L},\mathcal{A},\mathcal{P}_{\mathcal{M}},\mathcal{R}_{\mathcal{M}},\gamma)$.
+For $\forall l=(s_{1},a_{1},r_{1},...,s_{t})\in\mathcal{L}_{t}\subseteq\mathcal{L}$, we can sample the action $a_{t}$ from the distribution $\pi(\cdot|l)$.
+Then the environment will feedback the reward signal $r_{t}=\mathcal{R}_{\mathcal{M}}(l,a_{t})=\mathcal{R}_{\mathcal{M}}(s_{t},a_{t})$, we can sample $s_{t+1}$ from the distribution $\mathcal{P}(\cdot|s_{t},a_{t})$, and the next environment state will be $l^{\prime}=(s_{1},a_{1},r_{1},...,s_{t},a_{t},r_{t},s_{t+1})\in\mathcal{L}_{t+1}$.
+In summary, we set the distribution $p(l^{\prime}|l,a_{t})$ as the corresponding dynamic $\mathcal{P}_{\mathcal{M}}$.
+Notice that although all $\mathcal{M}\in\mathcal{T}$ shares the same dynamic $\mathcal{P}$, their new dynamic $\mathcal{P}_{\mathcal{M}}$ are different since the dynamic is related to the given reward.
+An obvious advantage of introducing $\tilde{\mathcal{M}}$ is that $\pi\in\Pi_{3}$ is now Markovian in $\tilde{\mathcal{M}}$ and it is much easier to analyze its performance.
+Obviously, we can set $J_{\mathcal{M}}(\pi)=J_{\tilde{\mathcal{M}}}(\pi),Q_{\mathcal{M}}^{*}(s,a)=Q_{\tilde{\mathcal{M}}}^{*}(s,a)$ to simplify the notation, and we can prove that
+Theorem 5.
+For any policy $\pi\in\Pi_{3}$, we have
+$\begin{split}&J_{\mathcal{T}}^{*}-\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}\left[J_{\mathcal{M}}(\pi)\right]\\ =&\frac{1}{1-\gamma}\int_{\mathcal{L}}p(l)\left[\int p(\mathcal{M}|l)\max_{a}Q_{\mathcal{M}}^{*}(l,a)d\mathcal{M}\right.\\ &\qquad\ -\left.\int_{a,\mathcal{M}}\pi(a|l)p(\mathcal{M}|l)Q_{\mathcal{M}}^{*}(l,a)dad\mathcal{M}\right]dl\\ \geq&\frac{1}{1-\gamma}\int_{\mathcal{L}}p(l)\left[\int p(\mathcal{M}|l)\max_{a}Q_{\mathcal{M}}^{*}(l,a)d\mathcal{M}\right.\\ &\qquad\qquad\quad-\left.\max_{a}\int_{\mathcal{M}}p(\mathcal{M}|l)Q_{\mathcal{M}}^{*}(l,a)d\mathcal{M}\right]dl,\end{split}$ (28)
+here $p(l)$ is a distribution of $\mathcal{L}$ related to $\mathcal{T},\pi$ and $p(\mathcal{M}|l)$ is the task posterior related to $\pi$.
+Proof.
+$\forall\pi\in\mathcal{H}_{3}$, as $\pi$ is Markovian in $\tilde{\mathcal{M}}$, we can directly utilize the proof of Theorem 3 from the beginning to Eq.
+(LABEL:app_eq_11), and the only difference is that the dynamics in $\tilde{\mathcal{M}}$ are different but the dynamics in $\mathcal{M}$ are the same.
+Thus we need to change the Eq.
+(LABEL:app_eq_12) as
+$\begin{split}&\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}\left[J_{\mathcal{M}}(\pi_{\mathcal{M}}^{*})-J_{\mathcal{M}}(\pi)\right]\\ =&\frac{1}{1-\gamma}\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}\mathbb{E}_{l\sim d_{\tilde{\mathcal{M}},\pi}(\cdot)}\left[\max_{a}Q_{\mathcal{M}}^{*}(l,a)\right.\\ &\left.-\int_{a}\pi(a|l)Q_{\mathcal{M}}^{*}(l,a)da\right]\\ =&\frac{1}{1-\gamma}\int p(\mathcal{M})\int_{\mathcal{L}}d_{\tilde{\mathcal{M}},\pi}(l)\left[\max_{a}Q_{\mathcal{M}}^{*}(l,a)\right.\\ &\left.-\int_{a}\pi(a|l)Q_{\mathcal{M}}^{*}(l,a)da\right]\\ =&\frac{1}{1-\gamma}\int_{\mathcal{L}}p(l)\int p(\mathcal{M}|l)\left[\max_{a}Q_{\mathcal{M}}^{*}(l,a)\right.\\ &\left.-\int_{a}\pi(a|l)Q_{\mathcal{M}}^{*}(l,a)da\right]\\ \geq&\frac{1}{1-\gamma}\int_{\mathcal{L}}p(l)\left[\int p(\mathcal{M}|l)\max_{a}Q_{\mathcal{M}}^{*}(l,a)d\mathcal{M}\right.\\ &\qquad\qquad\quad-\left.\max_{a}\int_{\mathcal{M}}Q_{\mathcal{M}}^{*}(l,a)d\mathcal{M}\right]dl,\end{split}$ (29)
+[…]
+Finally, we will show that maximizing $p(\mathcal{M}|l)$ is helpful for task generalization.
+In the training stage, we will sample a task $\mathcal{M}$ and corresponding state-action-reward pairs $l$, thus optimizing $p(\mathcal{M}|l)$ will make it to be closer to some Dirac distribution.
+In such a situation, for each $l$, we can infer a “most possible” posterior task $\mathcal{M}_{l}$ with high $p(\mathcal{M}_{l}|l)$, thus we can approximately take $\pi(l)=\arg\max_{a}Q_{\mathcal{M}_{l}}^{*}(l,a)$ and the optimal gap calculated by Eq.
+(LABEL:app_ep_20) will be controlled.
+A.7 Sparse Reward Task
+In this part, we show that generalizing to tasks with the same dynamics and sparse rewards without extra knowledge (like context) is extremely difficult and sometimes impossible.
+It is because we cannot distinguish them via historical information.
+Here we construct an example.
+Assume there are $n$ MDPs, $\mathcal{M}_{i}(i=1,...,n)$, each MDP share the same state set $\mathcal{S}=\{s_{1},...,s_{T}\}$ and action set $\mathcal{A}=\{a_{1},...,a_{n}\}$.
+The initial state is $s_{1}$ and the dynamic is that $\mathcal{P}(s_{t+1}|s_{t},a_{i})=1(t=1,...,T-1,i=1,...,n),\mathcal{P}(s_{T}|s_{T},a_{i})=1(i=1,...,n)$.
+As for the reward function, we define that
+$\mathcal{R}_{\mathcal{M}_{i}}(s_{T-1},a_{i})=1,i=1,...,n.$ (30)
+and the reward function is 0 otherwise.
+In this case, any policies (including markovian, state-action memorized, and state-action-reward memorized) in this task distribution can only handle one task since they can not distinguish them.
+Appendix B Pseudo Code of TAD
+The detailed pseudo code of TAD is provided in Algorithm 2.
 Algorithm 2 Task Aware Dreamer (TAD)
-0:  $M$ training tasks $\{\mathcal{M}_{m}\}_{m=1}^{M}$, $M$ replay buffers $\{\mathcal{D}_{m}\}_{m=1}^{M}$, $N$ testing tasks $\{\mathcal{M}_{M+n}\}_{n=1}^{N}$, initialize neural network parameters of world models, the policy, and the critic
+0:  $M$ training tasks $\{\mathcal{M}_{m}\}_{m=1}^{M}$, $M$ replay buffers $\{\mathcal{D}_{m}\}_{m=1}^{M}$, $N$ test tasks $\{\mathcal{M}_{M+n}\}_{n=1}^{N}$, initialize neural network parameters of world models, the policy, and the critic
 1:  while not converge do
@@
 27:  end for
-Appendix B Proof of Theorems
-In this section, we will provide detailed proofs of theorems.
-B.1 The Proof of Theorem 1
-In this part, we give an informed version of Theorem 1 with detailed proof as
-Theorem 5.
-Given $M$ tasks $\{\mathcal{M}_{m}\}_{m=1}^{M}$, we consider the product space $\mathcal{H}=\mathcal{Q}^{M}$ composed of M state-Q function spaces, $\forall\{q_{m}\}_{m=1}^{M}\in\mathcal{H}$, $q_{m}:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ is some Q function.
-Given $M$ corresponding dataset $D^{m}=\{(s_{t}^{m},a_{t}^{m},r_{t}^{m},s_{t+1}^{m})\}$, we can consider three hypothesis classes of optimal Q function $\mathcal{H}_{1},\mathcal{H}_{2},\mathcal{H}_{3}$, which are subsets of $\mathcal{H}$:
-$\begin{split}\mathcal{H}_{1}=&\{(q_{m})_{m=1}^{M}|q_{m}(s_{t}^{m},a_{t}^{m})=r_{t}^{m}+\gamma\max_{a^{\prime}}q_{m}(s_{t+1}^{m},a^{\prime})\}\\ \mathcal{H}_{2}=&\{(q_{m})_{m=1}^{M}|\exists(p_{m})_{m=1}^{M}:p_{m}(s_{t}^{m},a_{t}^{m})=s_{t+1}^{m},\\ &q_{m}(s,a)=r^{m}(s,a)+\gamma\max_{a^{\prime}}q_{m}(p_{m}(s,a),a^{\prime}),\forall s,a\}\\ \mathcal{H}_{3}=&\{(q_{m})_{m=1}^{M}|\exists p:p(s_{t}^{m},a_{t}^{m})=s_{t+1}^{m},\\ &q_{m}(s,a)=r^{m}(s,a)+\gamma\max_{a^{\prime}}q_{m}(p(s,a),a^{\prime}),\forall s,a\}.\end{split}$ (8)
-Then we have $\mathcal{H}_{3}\subseteq\mathcal{H}_{2}\subseteq\mathcal{H}_{1}$.
-Proof.
-First, we will show that $\mathcal{H}_{2}\subseteq\mathcal{H}_{1}$, following the proof in Young et al. (2022) that considers the setting that the rewards of all environments are the same.
-$\begin{split}&\{q_{m}\}\in\mathcal{H}_{2}\\ \Leftrightarrow&\exists\{p_{m}\}:p_{m}(s_{t}^{m},a_{t}^{m})=s_{t+1}^{m},\forall m,t\\ &q_{m}(s,a)=r^{m}(s,a)+\gamma\max_{a^{\prime}}q_{m}(p_{m}(s,a),a^{\prime}),\forall m,s,a\\ \Rightarrow&\exists\{p_{m}\}:p_{m}(s_{t}^{m},a_{t}^{m})=s_{t+1}^{m},\forall m,t\\ &q_{m}(s_{t}^{m},a_{t}^{m})=r^{m}_{t}+\gamma\max_{a^{\prime}}q_{m}(p_{m}(s_{t}^{m},a_{t}^{m}),a^{\prime}),\forall m,t\\ \Rightarrow&q_{m}(s_{t}^{m},a_{t}^{m})=r^{m}_{t}+\gamma\max_{a^{\prime}}q_{m}(p_{m}(s_{t}^{m},a_{t}^{m}),a^{\prime}),\forall m,t\\ \Leftrightarrow&\{q_{m}\}\in\mathcal{H}_{1},\end{split}$ (9)
-thus we have $\mathcal{H}_{2}\subseteq\mathcal{H}_{1}$.
-Next, we prove that $\mathcal{H}_{3}\subseteq\mathcal{H}_{2}$, which is mainly because we can utilize similar dynamic structures from different tasks to narrow down the hypothesis spaces of the dynamic model.
-$\begin{split}&\{q_{m}\}\in\mathcal{H}_{3}\\ \Leftrightarrow&\exists p:p(s_{t}^{m},a_{t}^{m})=s_{t+1}^{m},\forall m,t\\ &q_{m}(s,a)=r^{m}(s,a)+\gamma\max_{a^{\prime}}q_{m}(p(s,a),a^{\prime}),\forall m,s,a\\ \Rightarrow&\exists p_{m}:p_{m}(s_{t}^{m},a_{t}^{m})=s_{t+1}^{m},\forall m,t\\ &q_{m}(s,a)=r^{m}(s,a)+\gamma\max_{a^{\prime}}q_{m}(p_{m}(s,a),a^{\prime}),\forall m,s,a\\ \Leftrightarrow&\{q_{m}\}\in\mathcal{H}_{2},\end{split}$ (10)
-thus we have $\mathcal{H}_{3}\subseteq\mathcal{H}_{2}$. ∎
-B.2 The Derivation of the ELBO
-We use ${\color[rgb]{0,139,69}q}$ to represent $q(s_{1:T}|o_{1:T},a_{1:T},r_{1:T})$, ${\color[rgb]{0,0,1}\hat{q}}$ to present $q(s_{t}|o_{\leq t},a_{<t},r_{<t})$, ${\color[rgb]{1,0,0}\tilde{q}}$ to represent $q(s_{t-1}|o_{\leq(t-1)},r_{<(t-1)},a_{<(t-1)})$, and we have
-$\begin{split}&\ln p(o_{1:T},r_{1:T},\mathcal{M}|a_{1:T})\\ =&\ln\mathbb{E}_{p(s_{1:T}|a_{1:T})}\left[p(o_{1:T},r_{1:T},\mathcal{M}|s_{1:T})\right]\\ =&\ln\mathbb{E}_{p(s_{1:T}|a_{1:T})}\left[p(o_{1:T},r_{1:T}|\mathcal{M},s_{1:T})p(\mathcal{M}|s_{1:T})\right]\\ =&\ln\mathbb{E}_{p(s_{1:T}|a_{1:T})}\left[p(\mathcal{M}|s_{1:T})\prod_{t=1}^{T}p(o_{t},r_{t}|\mathcal{M},s_{t})\right]\\ =&\ln\mathbb{E}_{{\color[rgb]{0,139,69}q}}\left[p(\mathcal{M}|s_{1:T})\prod_{t=1}^{T}p(o_{t},r_{t}|\mathcal{M},s_{t})\frac{p(s_{t}|s_{t-1},a_{t-1})}{q(s_{t}|o_{\leq t},r_{<t},a_{<t})}\right]\\ \geq&\mathbb{E}_{{\color[rgb]{0,139,69}q}}[\ln p(\mathcal{M}|s_{1:T})+\sum_{t=1}^{T}\ln p(o_{t},r_{t}|\mathcal{M},s_{t})]\\ +&\mathbb{E}_{{\color[rgb]{0,139,69}q}}\sum_{t=1}^{T}\left[\ln p(s_{t}|s_{t-1},a_{t-1})-\ln q(s_{t}|o_{\leq t},r_{<t},a_{<t})\right]\\ =&\mathbb{E}_{{\color[rgb]{0,139,69}q}}[\ln p(\mathcal{M}|s_{1:T})]\\ +&\sum_{t=1}^{T}\left[\mathbb{E}_{{\color[rgb]{0,139,69}q}}\ln p(o_{t},r_{t}|\mathcal{M},s_{t})-\mathbb{E}_{{\color[rgb]{0,139,69}q}}\ln\frac{q(s_{t}|o_{\leq t},r_{<t},a_{<t})}{p(s_{t}|s_{t-1},a_{t-1})}\right]\\ =&\mathbb{E}_{{\color[rgb]{0,139,69}q}}[\ln p(\mathcal{M}|s_{1:T})]+\sum_{t=1}^{T}\mathbb{E}_{{\color[rgb]{0,0,1}\hat{q}}}[\ln p(o_{t},r_{t}|\mathcal{M},s_{t})]\\ -&\sum_{t=1}^{T}\mathbb{E}_{{\color[rgb]{0,0,1}\hat{q}}{\color[rgb]{1,0,0}\tilde{q}}}\left[\ln\frac{q(s_{t}|o_{\leq t},r_{<t},a_{<t})}{p(s_{t}|s_{t-1},a_{t-1})}\right]\\ =&\mathbb{E}_{{\color[rgb]{0,139,69}q}}[\ln p(\mathcal{M}|s_{1:T})]+\sum_{t=1}^{T}\mathbb{E}_{{\color[rgb]{0,0,1}\hat{q}}}[\ln p(o_{t},r_{t}|\mathcal{M},s_{t})]\\ -&\sum_{t=1}^{T}\mathbb{E}_{{\color[rgb]{1,0,0}\tilde{q}}}[\mathrm{KL}\left(q(s_{t}|o_{\leq t},r_{<t},a_{<t})\|p(s_{t}|s_{t-1},a_{t-1})\right)].\end{split}$ (11)
-Thus we have proven the ELBO.
-B.3 The Proof of Theorem 2
-Proof.
-We first prove that for $\forall\pi\in\Pi_{2}$, we have $\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}[J_{\mathcal{M}}(\pi)]=J_{\bar{\mathcal{M}}}(\pi)$.
-For any $\mathcal{M}=(\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R}_{\mathcal{M}},\gamma)\sim\mathcal{T}$, we can use the policy $\pi$ to interact with $\mathcal{M}$ and get the trajectory $\tau=(s_{0}^{\mathcal{M}},a_{0}^{\mathcal{M}},r_{1}^{\mathcal{M}},s_{1}^{\mathcal{M}},a_{1}^{\mathcal{M}},r_{2}^{\mathcal{M}},...)$.
-Since the dynamic transition $\mathcal{P}$ is the same for all $\mathcal{M}$ and the policy $\pi\in\Pi_{2}$ only depends on historical states and actions, we naturally have that the distribution of all states and actions $(s_{0}^{\mathcal{M}},a_{0}^{\mathcal{M}},s_{1}^{\mathcal{M}},a_{1}^{\mathcal{M}},...)$ are the same for all $\mathcal{M}\sim\mathcal{T}$ as well as $\bar{\mathcal{M}}$.
-Consequently, we have
-$\begin{split}&\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}[J_{\mathcal{M}}(\pi)]\\ =&\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}\mathbb{E}_{\tau\sim\mathcal{P},\pi}[R_{\mathcal{M}}(\tau)]=\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}\mathbb{E}_{\tau\sim\mathcal{P},\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}r_{t}^{\mathcal{M}}\right]\\ =&\mathbb{E}_{\tau\sim\mathcal{P},\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}[r_{t}^{\mathcal{M}}]\right]\\ =&\mathbb{E}_{\tau\sim\mathcal{P},\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}[\mathcal{R}_{\mathcal{M}}(s_{t}^{\mathcal{M}},a_{t}^{\mathcal{M}})]\right]\\ =&\mathbb{E}_{\tau\sim\mathcal{P},\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}[\bar{\mathcal{R}}(s_{t}^{\mathcal{M}},a_{t}^{\mathcal{M}})]\right]\\ =&\mathbb{E}_{\tau\sim\mathcal{P},\pi}[R_{\bar{\mathcal{M}}}(\tau)]=J_{\bar{\mathcal{M}}}(\pi).\end{split}$ (12)
-It is well known that the optimal policy in single MDP is memory-less, i.e., $\max_{\pi\in\Pi_{2}}J_{\bar{\mathcal{M}}}(\pi)=\max_{\pi\in\Pi_{1}}J_{\bar{\mathcal{M}}}(\pi)$. Consequently, we have
-$\begin{split}J_{\mathcal{T}}^{2}&=\max_{\pi\in\Pi_{2}}\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}[J_{\mathcal{M}}(\pi)]=\max_{\pi\in\Pi_{2}}J_{\bar{\mathcal{M}}}(\pi)\\ &=\max_{\pi\in\Pi_{1}}J_{\bar{\mathcal{M}}}(\pi)=\max_{\pi\in\Pi_{1}}\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}[J_{\mathcal{M}}(\pi)]=J_{\mathcal{T}}^{1}\\ &\leq\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}\left[\max_{\pi\in\Pi_{1}}J_{\mathcal{M}}(\pi)\right]=J_{\mathcal{T}}^{*}.\end{split}$ (13)
-Thus we have proven this result. ∎
-B.4 The Proof of Theorem 3
-Proof.
-Our proof follows some previous work Kakade and Langford (2002); Ying et al.
-(2022).
-First, we consider the bellman equation of value function of $\pi,\pi_{\mathcal{M}}^{*}\in\Pi_{1}$ in $\mathcal{M}$ as
-$\begin{split}V_{\mathcal{M},\pi}(s)=&\sum_{a}\pi(a|s)\left[\mathcal{R}(s,a)+\gamma\sum_{s^{\prime}}\mathcal{P}(s^{\prime}|s,a)V_{\mathcal{M},\pi}(s^{\prime})\right],\\ V_{\mathcal{M},\pi_{\mathcal{M}}^{*}}(s)=&\sum_{a}\pi_{\mathcal{M}}^{*}(a|s)\left[\mathcal{R}(s,a)+\gamma\sum_{s^{\prime}}\mathcal{P}(s^{\prime}|s,a)V_{\mathcal{M},\pi_{\mathcal{M}}^{*}}(s^{\prime})\right].\end{split}$
-Defining $\Delta V(s)\triangleq V_{\mathcal{M},\pi}(s)-V_{\mathcal{M},\pi_{\mathcal{M}}^{*}}(s)$ as the difference of these two value functions, we can further deduce that
-$\begin{split}&V_{\mathcal{M},\pi}(s)-V_{\mathcal{M},\pi_{\mathcal{M}}^{*}}(s)\\ =&\gamma\sum_{a}\Delta\pi(a|s)\sum_{s^{\prime}}\mathcal{P}(s^{\prime}|s,a)V_{\mathcal{M},\pi_{\mathcal{M}}^{*}}(s^{\prime})\\ +&\gamma\sum_{a}\pi(a|s)\sum_{s^{\prime}}\mathcal{P}(s^{\prime}|s,a)\Delta V(s^{\prime})+\sum_{a}\Delta\pi(a|s)\mathcal{R}(s,a)\\ =&\sum_{a}\Delta\pi(a|s)Q_{\mathcal{M},\pi_{\mathcal{M}}^{*}}(s,a)\\ +&\gamma\sum_{a}\pi(a|s)\sum_{s^{\prime}}\mathcal{P}(s^{\prime}|s,a)\Delta V(s^{\prime}),\end{split}$ (14)
-here $\Delta\pi(a|s)=\pi(a|s)-\pi_{\mathcal{M}}^{*}(a|s)$. Since Eq. (14) holds for any $s$, thus we calculate its expectation for $s\sim d_{\mathcal{M}}^{\pi_{\mathcal{M}}^{*}}$:
-$\begin{split}&\sum_{s}d_{\mathcal{M}}^{\pi}(s)\Delta V(s)\\ =&\sum_{s}d_{\mathcal{M}}^{\pi}(s)[V_{\mathcal{M},\pi}(s)-V_{\mathcal{M},\pi_{\mathcal{M}}^{*}}(s)]\\ =&\sum_{s}d_{\mathcal{M}}^{\pi}(s)\sum_{a}\Delta\pi(a|s)Q_{\mathcal{M},\pi_{\mathcal{M}}^{*}}(s,a)\\ +&\gamma\sum_{s}d_{\mathcal{M}}^{\pi}(s)\sum_{a}\pi(a|s)\sum_{s^{\prime}}\mathcal{P}(s^{\prime}|s,a)\Delta V(s^{\prime})\\ =&\sum_{s}d_{\mathcal{M}}^{\pi}(s)\sum_{a}\Delta\pi(a|s)Q_{\mathcal{M},\pi_{\mathcal{M}}^{*}}(s,a)\\ +&\sum_{s^{\prime}}\Delta V(s^{\prime})\left[\gamma\sum_{s}d_{\mathcal{M}}^{\pi}(s)\sum_{a}\pi(a|s)\mathcal{P}(s^{\prime}|s,a)\right].\end{split}$ (15)
-Since $\gamma\sum_{s^{\prime}}d_{\mathcal{M}}^{\pi}(s^{\prime})\sum_{a}\pi(a|s^{\prime})\mathcal{P}(s|s^{\prime},a)=d_{\mathcal{M}}^{\pi}(s)-(1-\gamma)\mathcal{P}(s_{0}=s)$, we have
-$\begin{split}\sum_{s}d_{\mathcal{M}}^{\pi}(s)&\Delta V(s)=\sum_{s}d_{\mathcal{M}}^{\pi}(s)\sum_{a}\Delta\pi(a|s)Q_{\mathcal{M},\pi_{\mathcal{M}}^{*}}(s,a)\\ +&\sum_{s^{\prime}}\Delta V(s^{\prime})\left[d_{\mathcal{M}}^{\pi}(s^{\prime})-(1-\gamma)\mathcal{P}(s_{0}=s^{\prime})\right].\end{split}$ (16)
-By moving the second term of the right part in Eq. (16) to the left part, we can deduce that
-$\begin{split}&(1-\gamma)\sum_{s^{\prime}}\Delta V(s^{\prime})\mathcal{P}(s_{0}=s^{\prime})\\ =&\sum_{s}d_{\mathcal{M}}^{\pi}(s)\sum_{a}\Delta\pi(a|s)Q_{\mathcal{M},\pi_{\mathcal{M}}^{*}}(s,a),\end{split}$ (17)
-thus we can calculate that
-[…]
-Consequently, we have
-$\begin{split}&\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}\left[J_{\mathcal{M}}(\pi_{\mathcal{M}}^{*})-J_{\mathcal{M}}(\pi)\right]\\ =&\frac{1}{1-\gamma}\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}\mathbb{E}_{s\sim d_{\mathcal{M},\pi}(\cdot)}\left[\max_{a}Q_{\mathcal{M}}^{*}(s,a)\right.\\ &\left.-\int_{a}\pi(a|s)Q_{\mathcal{M}}^{*}(s,a)da\right]\\ \geq&\frac{1}{1-\gamma}\mathbb{E}_{s\sim d_{\mathcal{M},\pi}(\cdot)}\left[\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}\max_{a}Q_{\mathcal{M}}^{*}(s,a)\right.\\ &\left.-\max_{a}\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}Q_{\mathcal{M}}^{*}(s,a)\right]\\ =&\frac{1}{1-\gamma}\mathbb{E}_{s\sim d_{\mathcal{M},\pi}}\left[D_{\text{TDR}}(\mathcal{T},s)\right].\end{split}$ (19)
-Since $J_{\mathcal{M}}(\pi_{\mathcal{M}}^{*})=\max_{\pi\in\Pi_{1}}J_{\mathcal{M}}(\pi)$, we have
-$\begin{split}&\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}\max_{\pi\in\Pi_{1}}J_{\mathcal{M}}(\pi)-\max_{\pi\in\Pi_{1}}\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}J_{\mathcal{M}}(\pi)\\ \geq&\frac{1}{1-\gamma}\mathbb{E}_{s\sim d_{\mathcal{M},\pi^{*}}}\left[D_{\text{TDR}}(\mathcal{T},s)\right],\end{split}$ (20)
-Thus we have proven this result. ∎
-B.5 Expressive Ability of $\Pi_{3}$
-Proposition 1.
-For $\forall\epsilon_{1},\epsilon_{2}$ satisfying $0<\epsilon_{1}\leq 1,0<\epsilon_{2}\leq 1$, there exists a task distribution $\mathcal{T}$ satisfying that
-$\begin{split}J_{\mathcal{T}}^{1}=J_{\mathcal{T}}^{2}\leq\epsilon_{1},~{}~{}J_{\mathcal{T}}^{3}\geq 1-\epsilon_{2},~{}~{}J_{\mathcal{T}}^{*}=1.\end{split}$ (21)
-Proof.
-Given fixed discount factor $\gamma\in(0,1)$, we first take $n\in\mathbb{N}$ satisfying $n\geq\frac{1}{\epsilon_{1}}$, $A=\epsilon_{2}\frac{n}{n-1}\frac{1-\gamma}{1-\gamma^{n}}$, and $B=\frac{1-\gamma}{\gamma^{n+1}}\left(1-\epsilon_{2}\frac{n}{n-1}\right)$.
-We construct state sets $\mathcal{S}=\{s_{t}^{l}\}(t=0,1,...,\infty,l=1,...,n)$ and action sets $\mathcal{A}=\{a_{j}\}_{j=1}^{n}$.
-Then we construct $n$ tasks $\mathcal{M}_{i}=(\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R}_{i},\gamma),i=1,...,n$, which share the same dynamic $\mathcal{P}$ and different reward functions $\mathcal{R}_{i}$.
-The initial state of each task is $s_{0}^{1}$, and the dynamic as well as reward functions are as below
-$\begin{split}&\mathcal{P}(s_{t}^{l},a_{j})=\mathbb{I}(s=s_{t+1}^{j}),\quad j,l=1,...,n;\ t=0,1,...,\infty\\ &\mathcal{R}_{i}(s_{t}^{l},a_{j})=f(t)\mathbb{I}(i=j),\quad i,j,l=1,...,n;\ t=0,1,...\infty\\ &f(t)=\left\{\begin{array}[]{ll}A,&t\leq n-1\\ B,&t\geq n\\ \end{array}\right.\end{split}$ (22)
-Take $\mathcal{T}$ as the uniform distribution over $\mathcal{M}_{1},...,\mathcal{M}_{n}$, thus we have
-$\begin{split}J_{\mathcal{M}_{i}}^{*}=&A+A\gamma+...+A\gamma^{n-1}+B\gamma^{n}+B\gamma^{n+1}+...\\ =&A\frac{1-\gamma^{n}}{1-\gamma}+B\frac{\gamma^{n+1}}{1-\gamma}=1\\ J_{\mathcal{T}}^{*}=&\frac{1}{n}\sum_{i=1}^{n}J_{\mathcal{M}_{i}}^{*}=1.\\ \end{split}$ (23)
-Since our construction satisfies $\mathbb{E}_{\mathcal{T}}[\mathcal{R}_{i}(s_{k}^{l},a_{j})]=\frac{f(k)}{n}$, for $\forall\pi\in\Pi_{2}$, we have
-$\begin{split}J_{\mathcal{T}}(\pi)=&\frac{1}{n}(A+A\gamma+...+A\gamma^{n-1}+B\gamma^{n}+B\gamma^{n+1}+...)\\ =&\frac{1}{n},\\ J_{\mathcal{T}}^{1}=&J_{\mathcal{T}}^{2}=\max_{\pi\in\Pi_{2}}J_{\mathcal{T}}(\pi)=\frac{1}{n}\leq\epsilon_{1}.\end{split}$ (24)
-Moreover, we can construction an agent $\hat{\pi}\in\Pi_{3}$ that takes action via the historical trajectory $\hat{\tau}_{t}=(\hat{s}_{0},\hat{a}_{0},\hat{r}_{0},...\hat{s}_{t})$:
-$\begin{split}\hat{\pi}(a_{j}|\hat{\tau}_{t})=&\mathbb{I}(j=t+1),\quad t=0,1,...,n-1\\ \hat{\pi}(a_{j}|\hat{\tau}_{t})=&\mathbb{I}(j=i),\quad t=n,...,\infty\\ \end{split}$ (25)
-here $i=\arg\max\{\hat{r}_{0},\hat{r}_{1},...,\hat{r}_{n-1}\}+1$, thus we have
-$\begin{split}&J_{\mathcal{T}}^{3}\geq J_{\mathcal{T}}(\hat{\pi})\\ =&\frac{1}{n}(A+A\gamma+...+A\gamma^{n-1})+B\gamma^{n}+B\gamma^{n+1}+...\\ =&\frac{A}{n}\frac{1-\gamma^{n}}{1-\gamma}+B\frac{\gamma^{n+1}}{1-\gamma}=1-\frac{(n-1)A}{n}\frac{1-\gamma^{n}}{1-\gamma}\\ \geq&1-\epsilon_{2}.\end{split}$ (26)
-Thus we have proven this result. ∎
-B.6 Proof and Discussion of Theorem 4
-In this part, we introduce an informed version of Theorem 4, about why optimizing $p(\mathcal{M}|l)\forall l\in\mathcal{L}$ is beneficial for task generalization, with detailed proofs.
-Recall that we consider the policy hypothesis $\mathcal{H}_{3}$ here, that each policy $\pi:\mathcal{L}\rightarrow\Delta(\mathcal{A}),\mathcal{L}=\cup_{t=1}^{\infty}\mathcal{L}_{t},\mathcal{L}_{t}=(\mathcal{S}\times\mathcal{A}\times\mathbb{R})^{t-1}\times\mathcal{S}$.
-As directly such $\mathcal{S}-\mathcal{A}-\mathcal{R}$ memorized policy is difficult, we consider an alternative MDP as ~$\tilde{\mathcal{M}}=(\mathcal{L},\mathcal{A},\mathcal{P}_{\mathcal{M}},\mathcal{R}_{\mathcal{M}},\gamma)$.
-For $\forall l=(s_{1},a_{1},r_{1},...,s_{t})\in\mathcal{L}_{t}\subseteq\mathcal{L}$, we can sample the action $a_{t}$ from the distribution $\pi(\cdot|l)$.
-Then the environment will feedback the reward signal $r_{t}=\mathcal{R}_{\mathcal{M}}(l,a_{t})=\mathcal{R}_{\mathcal{M}}(s_{t},a_{t})$, we can sample $s_{t+1}$ from the distribution $\mathcal{P}(\cdot|s_{t},a_{t})$, and the next environment state will be $l^{\prime}=(s_{1},a_{1},r_{1},...,s_{t},a_{t},r_{t},s_{t+1})\in\mathcal{L}_{t+1}$.
-In summary, we set the distribution $p(l^{\prime}|l,a_{t})$ as the corresponding dynamic $\mathcal{P}_{\mathcal{M}}$.
-Notice that although all $\mathcal{M}\in\mathcal{T}$ shares the same dynamic $\mathcal{P}$, their new dynamic $\mathcal{P}_{\mathcal{M}}$ are different since the dynamic is related to the given reward.
-An obvious advantage of introducing $\tilde{\mathcal{M}}$ is that $\pi\in\Pi_{3}$ is now Markovian in $\tilde{\mathcal{M}}$ and it is much easier to analyze its performance.
-Obviously, we can set $J_{\mathcal{M}}(\pi)=J_{\tilde{\mathcal{M}}}(\pi),Q_{\mathcal{M}}^{*}(s,a)=Q_{\tilde{\mathcal{M}}}^{*}(s,a)$ to simplify the notation, and we can prove that
-Theorem 6.
-For any policy $\pi\in\Pi_{3}$, we have
-$\begin{split}&J_{\mathcal{T}}^{*}-\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}\left[J_{\mathcal{M}}(\pi)\right]\\ =&\frac{1}{1-\gamma}\int_{\mathcal{L}}p(l)\left[\int p(\mathcal{M}|l)\max_{a}Q_{\mathcal{M}}^{*}(l,a)d\mathcal{M}\right.\\ &\qquad\ -\left.\int_{a,\mathcal{M}}\pi(a|l)p(\mathcal{M}|l)Q_{\mathcal{M}}^{*}(l,a)dad\mathcal{M}\right]dl\\ \geq&\frac{1}{1-\gamma}\int_{\mathcal{L}}p(l)\left[\int p(\mathcal{M}|l)\max_{a}Q_{\mathcal{M}}^{*}(l,a)d\mathcal{M}\right.\\ &\qquad\qquad\quad-\left.\max_{a}\int_{\mathcal{M}}p(\mathcal{M}|l)Q_{\mathcal{M}}^{*}(l,a)d\mathcal{M}\right]dl,\end{split}$ (27)
-here $p(l)$ is a distribution of $\mathcal{L}$ related to $\mathcal{T},\pi$ and $p(\mathcal{M}|l)$ is the task posterior related to $\pi$.
-Proof.
-$\forall\pi\in\mathcal{H}_{3}$, as $\pi$ is Markovian in $\tilde{\mathcal{M}}$, we can directly utilize the proof of Theorem 3 from the beginning to Eq.
-(18), and the only difference is that the dynamics in $\tilde{\mathcal{M}}$ are different but the dynamics in $\mathcal{M}$ are the same.
-Thus we need to change the Eq.
-(19) as
-$\begin{split}&\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}\left[J_{\mathcal{M}}(\pi_{\mathcal{M}}^{*})-J_{\mathcal{M}}(\pi)\right]\\ =&\frac{1}{1-\gamma}\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}\mathbb{E}_{l\sim d_{\tilde{\mathcal{M}},\pi}(\cdot)}\left[\max_{a}Q_{\mathcal{M}}^{*}(l,a)\right.\\ &\left.-\int_{a}\pi(a|l)Q_{\mathcal{M}}^{*}(l,a)da\right]\\ =&\frac{1}{1-\gamma}\int p(\mathcal{M})\int_{\mathcal{L}}d_{\tilde{\mathcal{M}},\pi}(l)\left[\max_{a}Q_{\mathcal{M}}^{*}(l,a)\right.\\ &\left.-\int_{a}\pi(a|l)Q_{\mathcal{M}}^{*}(l,a)da\right]\\ =&\frac{1}{1-\gamma}\int_{\mathcal{L}}p(l)\int p(\mathcal{M}|l)\left[\max_{a}Q_{\mathcal{M}}^{*}(l,a)\right.\\ &\left.-\int_{a}\pi(a|l)Q_{\mathcal{M}}^{*}(l,a)da\right]\\ \geq&\frac{1}{1-\gamma}\int_{\mathcal{L}}p(l)\left[\int p(\mathcal{M}|l)\max_{a}Q_{\mathcal{M}}^{*}(l,a)d\mathcal{M}\right.\\ &\qquad\qquad\quad-\left.\max_{a}\int_{\mathcal{M}}Q_{\mathcal{M}}^{*}(l,a)d\mathcal{M}\right]dl,\end{split}$ (28)
-[…]
-Finally, we will show that maximizing $p(\mathcal{M}|l)$ is helpful for task generalization.
-In the training stage, we will sample a task $\mathcal{M}$ and corresponding state-action-reward pairs $l$, thus optimizing $p(\mathcal{M}|l)$ will make it to be closer to some Dirac distribution.
-In such a situation, for each $l$, we can infer a “most possible” posterior task $\mathcal{M}_{l}$ with high $p(\mathcal{M}_{l}|l)$, thus we can approximately take $\pi(l)=\arg\max_{a}Q_{\mathcal{M}_{l}}^{*}(l,a)$ and the optimal gap calculated by E.q.
-(27) will be controlled.
-B.7 Sparse Reward Task
-In this part, we show that generalizing to tasks with the same dynamics and sparse rewards without extra knowledge (like context) is extremely difficult and sometimes impossible.
-It is because we cannot distinguish them via historical information.
-Here we construct an example.
-Assume there are $n$ MDPs, $\mathcal{M}_{i}(i=1,...,n)$, each MDP share the same state set $\mathcal{S}=\{s_{1},...,s_{T}\}$ and action set $\mathcal{A}=\{a_{1},...,a_{n}\}$.
-The initial state is $s_{1}$ and the dynamic is that $\mathcal{P}(s_{t+1}|s_{t},a_{i})=1(t=1,...,T-1,i=1,...,n),\mathcal{P}(s_{T}|s_{T},a_{i})=1(i=1,...,n)$.
-As for the reward function, we define that
-$\mathcal{R}_{\mathcal{M}_{i}}(s_{T-1},a_{i})=1,i=1,...,n.$ (29)
-and the reward function is 0 otherwise.
-In this case, any policies (including markovian, state-action memorized, and state-action-reward memorized) in this task distribution can only handle one task since they can not distinguish them.
 Appendix C Experimental Details for DMControl
@@
 […]
-  • Walker-stand&\&&walk.
-This task combination includes two tasks: Walker-stand and Walker-walk, which both require a two-leg robot to maintain its height above a certain threshold, and Walker-walk also hopes it to move faster than a certain speed, i.e., their reward functions can be roughly described as
-    $\begin{split}\mathcal{R}_{\text{stand}}&=\text{tolerance}(\text{height},(1.2,\infty)),\\ \mathcal{R}_{\text{walk}}&=\text{tolerance}(\text{height},(1.2,\infty))\\ &*\text{tolerance}(\text{speed},(1,\infty)).\end{split}$ (30)
-    Therefore, for all states, the optimal action of Walker-walk is also optimal in Walker-stand and TDR here is 0.
-  • Cartpole-balance&\&&sparse.
+  • Cartpole-balance&sparse.
 This task combination includes two tasks: Cartpole-balance and Cartpole-balance__\__sparse, which both hope to balance an unactuated pole with dense and sparse rewards respectively.
-The optimal actions of these tasks are both hope to balance the agent and thus TDR here is 0.
-  • Acrobot-swingup&\&&sparse.
-This task combination includes two tasks: Acrobot-swingup and Acrobot-swingup__\__sparse, which both hope to swing up a double pendulum with dense and sparse rewards respectively.
-Also, the optimal actions of these tasks are both hoped to swingup the agent and thus TDR here is 0.
-  • Quadruped-walk&\&&run.
-This task combination includes two tasks: Quadruped-walk and Quadruped-run, which both require a quadruped agent to move faster than a target speed.
-The target speed of the running task is larger than the walking task, i.e., their reward functions can be roughly described as
-    $\begin{split}&\mathcal{R}_{\text{walk}}=\text{tolerance}(\text{speed},(0.5,\infty)),\\ &\mathcal{R}_{\text{run}}=\text{tolerance}(\text{speed},(5.0,\infty)).\end{split}$ (31)
-    Therefore, TDR here is still 0 since optimal actions of Quadruped-run are also optimal in Quadruped-walk.
-  • Reacher-easy&\&&hard.
-This task combination includes two tasks: Reacher-easy and Reacher-hard, which both hope to control a two-link planar reacher to prostrate the target sphere.
-The sizes of target spheres in these two tasks are different, i.e., their reward functions can be roughly described as
-    $\begin{split}&\mathcal{R}_{\text{easy}}=\text{tolerance}(\text{distance},(0.0,0.05)),\\ &\mathcal{R}_{\text{hard}}=\text{tolerance}(\text{distance},(0.0,0.015)).\end{split}$ (32)
-    Thus TDR here is not 0 but still small.
-  • Hopper-stand&\&&hop.
-This task combination includes two tasks: Hopper-stand and Hopper-hop, which both hope the height of a one-legged hopper is larger than the target height.
-While Hopper-stand hopes the control force is 0, Hopper-hop hopes to control the robot to hop forward and its speed is larger than the target speed.
-Their reward functions can be roughly described as
-    $\begin{split}\mathcal{R}_{\text{stand}}&=\text{tolerance}(\text{height},(0.6,2))\\ &*\text{tolerance}(\text{speed},(2.0,\infty)),\\ \mathcal{R}_{\text{hop}}&=\text{tolerance}(\text{height},(0.6,2))\\ &*\text{tolerance}(\text{control},(0.0,0.0)).\end{split}$ (33)
-    Since one agent can not hop when controlling the force as 0, TDR here is not 0 but not huge.
-  • Walker-stand&\&&prostrate.
-This task combination includes two tasks: Walker-stand and Walker-prostrate.
+The optimal actions of these tasks are both hoped to balance the agent and thus TDR here is 0.
+  • Walker-stand&walk&prostrate&flip.
+This task combination includes four tasks: Walker-stand, Walker-walk, Walker-prostrate, and Walker-flip.
 Walker-stand hopes the height of a two-leg robot to be larger than the target height.
-And we design the task Walker-prostrate, which hopes its height to be lower than another target height.
-Their reward functions can be roughly described as
-    $\begin{split}&\mathcal{R}_{\text{stand}}=\text{tolerance}(\text{height},(1.2,\infty)),\\ &\mathcal{R}_{\text{prostrate}}=\text{tolerance}(\text{height},(0.0,0.2)).\end{split}$ (34)
-    Thus TDR here is huge since the optimal Q functions of these two tasks are almost the opposite.
-  • Walker-walk&\&&prostrate.
-This task combination includes two tasks: Walker-walk and Walker-prostrate.
-Walker-walk hopes the height of the improved planar walker to be larger than a target height, while Walker-prostrate hopes it to be lower than a target height.
-Moreover, Walker-walk hopes the speed of the robot is larger than another target speed, i.e., their reward functions can be roughly described as
-    $\begin{split}\mathcal{R}_{\text{walk}}&=\text{tolerance}(\text{height},(1.2,\infty))\\ &*\text{tolerance}(\text{speed},(1.0,\infty)),\\ \mathcal{R}_{\text{prostrate}}&=\text{tolerance}(\text{height},(0.0,0.2)).\end{split}$ (35)
-    Thus TDR here is also huge since the optimal Q functions of these two tasks are almost the opposite.
-Figure 4: Average cumulative reward curves over different task combinations in DMC for CURL, SAC+AE, PlaNet, Dreamer, and TAD.
-The x-axes indicate the number of episodes interacting with the environment, and the y-axes indicate the performance, including average rewards with std.
-Tasks Hypothesis Walker-stand&\&&walk Cartpole-balance&\&&sparse Acrobot-swingup&\&&sparse Quadruped-walk&\&&run
-CURL $\Pi_{1}$ 563.4 $\pm$ 16.0 996.8 $\pm$ 0.8 11.0 $\pm$ 1.8 203.3 $\pm$ 69.9
-SAC+AE $\Pi_{1}$ 683.8 $\pm$ 164.3 993.4 $\pm$ 0.6 11.7 $\pm$ 5.4 203.2 $\pm$ 70.2
-PlaNet $\Pi_{2}$ 990.6 $\pm$ 3.8 412.8 $\pm$ 65.9 12.2 $\pm$ 9.8 305.3 $\pm$ 206.6
-Dreamer $\Pi_{2}$ 992.3 $\pm$ 1.7 997.1 $\pm$ 1.1 159.9 $\pm$ 108.4 345.4 $\pm$ 218.3
-TAD $\Pi_{3}$ 993.9 $\pm$ 1.0 998.9 $\pm$ 0.3 217.3 $\pm$ 107.1 505.0 $\pm$ 119.1
-Tasks Hypothesis Reacher-easy&\&&hard Hopper-stand&\&&hop Walker-stand&\&&prostrate Walker-walk&\&&prostrate
-CURL $\Pi_{1}$ 552.5 $\pm$ 88.9 137.1 $\pm$ 9.2 496.3 $\pm$ 8.7 441.0 $\pm$ 7.7
-SAC+AE $\Pi_{1}$ 669.5 $\pm$ 99.4 118.2 $\pm$ 83.1 504.4 $\pm$ 6.7 445.8 $\pm$ 2.6
-PlaNet $\Pi_{2}$ 341.0 $\pm$ 228.3 25.9 $\pm$ 25.7 544.3 $\pm$ 68.6 445.4 $\pm$ 5.1
-Dreamer $\Pi_{2}$ 506.1 $\pm$ 382.9 436.9 $\pm$ 18.9 510.0 $\pm$ 6.7 449.7 $\pm$ 12.5
-TAD $\Pi_{3}$ 823.3 $\pm$ 237.4 503.9 $\pm$ 22.2 941.1 $\pm$ 4.0 913.7 $\pm$ 11.5
+Walker-walk hopes the height of the improved planar walker to be larger than a target height and the speed of the robot to be larger than another target speed.
+Walker-prostrate hopes the height of the robot to be lower than the target height.
+Finally, Walker-flip hopes the robot to stand and move forward to the target speed by executing a rapid twist and jump.
+In detail, their reward functions can be roughly described as
+    $\begin{split}\mathcal{R}_{\text{stand}}=&\text{tolerance}(\text{height},(1.2,\infty)),\\ \mathcal{R}_{\text{walk}}=&\text{tolerance}(\text{height},(1.2,\infty))\\ &*\text{tolerance}(\text{speed},(1,\infty)).\\ \mathcal{R}_{\text{prostrate}}=&\text{tolerance}(\text{height},(0.0,0.2)).\\ \mathcal{R}_{\text{flip}}=&\text{tolerance}(\text{height},(1.2,\infty))\\ &*\text{tolerance}(\text{angmomentum},5,\infty).\end{split}$ (31)
+    In this situation, for all states, the optimal actions of Walker-walk/stand and Walker-prostrate are opposite, and TDR here is huge.
+  • Cheetah-run&run__\__back&flip&flip__\__back.
+This task combination includes four tasks: Cheetah-run, Cheetah-run__\__back, Cheetah-flip, and Cheetah-flip__\__back.
+Cheetah-run hopes to control a running planar biped to run forward within a target speed.
+Cheetah-run__\__back, differently, hopes to control the Cheetah robot to run backward within a target speed.
+Cheetah-flip hopes the robot to move forward to the target speed by executing a rapid twist and jump.
+Similarly, Cheetah-flip__\__back hopes to control the robot to move backward by flipping.
+    $\begin{split}\mathcal{R}_{\text{Cheetah}\_\text{run}}=&\text{tolerance}(\text{speed},(10,\infty)).\\ \mathcal{R}_{\text{Cheetah}\_\text{run$\_$back}}=&\text{tolerance}(-\text{speed},(10,\infty)).\\ \mathcal{R}_{\text{Cheetah}\_\text{flip}}=&\text{tolerance}(\text{angmomentum},(5,\infty)).\\ \mathcal{R}_{\text{Cheetah}\_\text{flip$\_$back}}=&\text{tolerance}(-\text{angmomentum},(5,\infty)).\end{split}$ (32)
+    Obviously, in this situation, for all states, the optimal actions of Cheetah-run and Cheetah-run__\__back are opposite, and TDR here is also huge.
+Algorithm Hypothesis Cartpole-balance Walker-stand&walk Cheetah-run&run__\__back
+&balance__\__sparse &prostrate&flip &flip&flip__\__back
+CURL $\Pi_{1}$ 994.5 $\pm$ 3.6 254.1 $\pm$ 9.2 229.7 $\pm$ 10.9
+CURL (w/ r) $\Pi_{3}$ 987.3 $\pm$ 12.9 265.0 $\pm$ 4.7 236.1 $\pm$ 5.8
+SAC+AE $\Pi_{1}$ 992.5 $\pm$ 2.6 256.9 $\pm$ 5.9 225.8 $\pm$ 10.1
+SAC+AE (w/ r) $\Pi_{3}$ 988.5 $\pm$ 7.0 257.2 $\pm$ 8.5 239.9 $\pm$ 12.4
+Table 4: Performance (mean $\pm$ std) in DMC of CURL, CURL (w/ r), SAC+AE, and SAC+AE (w/ r).
+Tasks Acrobot-Cartpole-Pendulum Walker-Cheetah-Hopper Cheetah-run__\__mass Walker-walk__\__mass
+Train&Test Train&Test Train Test Train Test
+Dreamer 541.3 $\pm$ 4.0 299.4 $\pm$ 17.2 717.8 $\pm$ 27.9 711.7 $\pm$ 38.8 889.9 $\pm$ 114.1 903.3 $\pm$ 102.8
+TAD 667.7 $\pm$ 6.4 554.7 $\pm$ 23.6 754.3 $\pm$ 22.9 738.2 $\pm$ 41.1 957.8 $\pm$ 35.1 963.1 $\pm$ 32.6
 […]
@@
 Now we introduce the three task distributions in our experiments, which are designed based on existing tasks in DeepMind control suite for testing the generalization of trained agents.
-  • Cheetah_normal-_\__speed($\alpha,\beta$).
+  • Cheetah__\__speed($\alpha,\beta$).
 This task distribution is designed in this paper with parameter $0\leq\beta\leq\alpha$, based on the task Cheetah__\__run in DeepMind control suite, and hopes the Cheetah robot can run with the target speed interval $(\alpha-\beta,\alpha+\beta)$.
@@
     We train the agents in tasks with parameters ($0.5,0.2$), ($1.5,0.2$), ($2.0,0.2$), ($3.0,0.2$) and test them in tasks with parameters ($1.0,0.2$), ($2,5,0.2$).
-  • Pendulum_normal-_\__angle($\alpha,\beta$).
+  • Pendulum__\__angle($\alpha,\beta$).
 This task distribution is designed in this paper with parameter $-1\leq\alpha\leq\beta\leq 1$, based on the task Pendulumh__\__swingup in DeepMind control suite, and hopes the Pendulum robot can swing up within the target angle interval $(\arccos\alpha,\arccos\beta)$.
     $\begin{split}\mathcal{R}_{\text{Pendulum}\_\text{angle}}(\alpha,\beta)=\text{tolerance}(\text{angle},(&\arccos\alpha,\\ &\arccos\beta)).\end{split}$ (34)
-    Training tasks are with parameters ($-0.95,-0.9$), ($-0.85,-0.8$), ($-0.8,-0.75$), ($-0.7,-0.65$) and testing tasks are with parameters ($-0.9,-0.85$), ($-0,75,-0.7$).
-  • Walker_normal-_\__speed($\alpha,\beta$).
+    Training tasks are with parameters ($-0.95,-0.9$), ($-0.85,-0.8$), ($-0.8,-0.75$), ($-0.7,-0.65$) and test tasks are with parameters ($-0.9,-0.85$), ($-0,75,-0.7$).
+  • Walker__\__speed($\alpha,\beta$).
 This task distribution is designed in this paper with parameter $0\leq\beta\leq\alpha$, based on the task Walker__\__run in DeepMind control suite, and hopes the Walker robot can run within the target speed interval $(\alpha-\beta,\alpha+\beta)$.
@@
 All algorithms are trained on one NVIDIA GeForce RTX 2080 Ti.
-As for the hyper-parameters, we follow previous works Ha and Schmidhuber (2018); Hafner et al.
-(2019b, a) and select 2 as the action repeat for all experiments following Hafner et al.
-(2019a).
-C.3 Full Experimental Results for TDR
-In tasks where TDR is 0, like Walker-stand&\&&walk and Cartpole-balance&\&&sparse, the optimal performance of TAD and baselines are similar since these tasks share the same optimal action and $\Pi_{1},\Pi_{2}$ own the optimal policy, as derived in Theorem 2.
-In these cases, TAD still converges faster and shows more robust performance.
-In Acrobot-swingup&\&&sparse and Quadruped-walk&\&&run, although TDR of these tasks is 0, TAD can surpass baselines since it can better distinguish different tasks.
-Finally, in the last four environments with non-zero TDR, TAD achieves a notable improvement compared with baselines.
-Especially, in Walker-stand&\&&prostrate and Walker-walk&\&&prostrate, where TDR is significantly huge, baselines perform poorly as these policies belong to $\Pi_{1},\Pi_{2}$ and can not differentiate different tasks, which also aligns with our Theorem 2.
-In contrast, TAD can distinguish between tasks by utilizing reward signals and improves performance conspicuously.
-C.4 Visualization Results for Task Generalization
-Figure 5: Visualization of the trained TAD agent in the task distribution of Cheetah__\__speed. We plot the speed of the agent as a function of timesteps in all tasks.
-Moreover, for each task sampled from the task distribution Cheetah__\__speed (here parameters $(3.0,0.2)$, $(2.0,0.2)$, $(1.5,0.2)$, $(0.5,0.2)$ are for training tasks and parameters $(2.5,0.2),(0.5,0.2)$ are for testing tasks), we plot the speed of the agent as a function of the timestep in Fig.
-5.
+For each seed and each task setting, it will take around 3 days.
+As for the hyper-parameters, we follow previous works (Ha and Schmidhuber 2018; Hafner et al.
+2019b, a) and select 2 as the action repeat for all experiments following (Hafner et al.
+2019a).
+C.3 Visualization Results for Task Generalization
+Moreover, for each task sampled from the task distribution Cheetah__\__speed (here parameters $(3.0,0.2)$, $(2.0,0.2)$, $(1.5,0.2)$, $(0.5,0.2)$ are for training tasks and parameters $(2.5,0.2),(0.5,0.2)$ are for test tasks), we plot the speed of the agent as a function of the timestep in Fig.
+6.
 As depicted, for each task, the agent trained by TAD will quickly improve its speed until reaching the target speed and then keep its speed since the speed determines whether it has met the task requirements via utilizing historical information.
@@
 We also provide videos of these trajectories in supplementary materials.
+Figure 6: Visualization of the trained TAD agent in the task distribution of Cheetah__\__speed. We plot the speed of the agent as a function of timesteps in all tasks.
 Appendix D Experimental Details for MuJoCo
-We here introduce state-based control tasks (Half-CheetahFwd-Back, Half-Cheetah-Vel, and Humanoid-Direc-2D) in detail, following the setting of previous meta RL works Finn et al.
-(2017); Rakelly et al.
-(2019).
+We here introduce state-based control tasks, including Half-CheetahFwd-Back, Half-Cheetah-Vel, and Humanoid-Direc-2D, in detail, following the setting of previous meta RL works (Finn, Abbeel, and Levine 2017; Rakelly et al.
+2019).
   • Half-Cheetah-Fwd-Back. This task distribution includes two tasks: moving forward and moving backward.
-  • Half-Cheetah-Vel. This task distribution hopes the agent to move forward and achieve the target velocity. There are 100 training tasks and 30 testing tasks for experiments.
-  • Humanoid-Direc-2D. This task distribution hopes the agent to move in the target direction. There are 100 training tasks and 30 testing tasks for experiments.
+  • Half-Cheetah-Vel. This task distribution hopes the agent to move forward and achieve the target velocity. There are 100 training tasks and 30 test tasks for experiments.
+  • Humanoid-Direc-2D. This task distribution hopes the agent to move in the target direction. There are 100 training tasks and 30 test tasks for experiments.
 Moreover, we introduce some details about our experiments.
@@
 All algorithms are trained on one NVIDIA GeForce RTX 2080 Ti.
+For each seed and each task setting, it will take around 1 day.
 We select 1 as the action repeat for all experiments following.
+$\beta$ 0.2 0.15 0.1
+Train Test Train Test Train Test
+Dreamer 250.2 $\pm$ 9.6 3.0 $\pm$ 2.2 247.5 $\pm$ 1.0 0.0 $\pm$ 0.0 175.8 $\pm$ 55.8 13.5 $\pm$ 13.5
+TAD 951.9 $\pm$ 3.3 876.9 $\pm$ 51.1 927.6 $\pm$ 2.6 800.1 $\pm$ 121.4 608.5 $\pm$ 321.4 491.6 $\pm$ 389.4
+Table 6: Average cumulative reward (mean $\pm$ one std) over different target region (smaller $\beta$ represents smaller target region and more sparse return) of the best policy trained by Dreamer and TAD in Cheetah__\__Speed.
+For each $\beta$, we train agents in the train tasks and evaluate them in both train and test environments.
+Numbers greater than 95 percent of the best performance for each environment are bold.
+SR 0.0 0.8 0.9
+Train Test Train Test Train Test
+Dreamer 250.2 $\pm$ 9.6 3.0 $\pm$ 2.2 237.4 $\pm$ 14.9 28.9 $\pm$ 14.6 168.7 $\pm$ 96.2 157.2 $\pm$ 188.4
+TAD 951.9 $\pm$ 3.3 876.9 $\pm$ 51.1 841.9 $\pm$ 154.9 546.1 $\pm$ 291.6 $7$$\pm$ 170.1 716.8 $\pm$ 136.5
+Table 7: Average cumulative reward (mean $\pm$ one std) over different sparse rates of the best policy trained by Dreamer and TAD in Cheetah__\__Speed.
+For each sparse rate, we train agents in the train tasks and evaluate them in both train and test environments.
+Numbers greater than 95 percent of the best performance for each environment are bold.
 Appendix E Ablation Study
-E.1 Dynamic Generalization
+E.1 Ablation Study on Reward Signals
+In this part, we do ablation studies on reward signals for model-free methods including CURL and SAC+AE, which both belong to $\Pi_{1}$.
+We design CURL (w/ r) and SAC+AE (w/ r) by directly adding reward signals into the observation based on CURL and SAC+AE, respectively.
+As shown in Table.
+4, the performance of CURL (w/ r) and SAC+AE (w/ r) is similar to CURL and SAC+AE.
+The major reason is that CURL (w/ r) and SAC+AE (w/ r)only utilize the observation and reward of the current timestep without historical information.
+Thus it is still difficult for them to distinguish different tasks.
+E.2 Dynamic Generalization
 As TAD utilizes all historical information to infer the environment, it can be directly applied to more general settings with different observations, dynamics, and/or actions.
@@
 This experiment indicates TAD’s potential in further handling dynamic generalization and even cross-embodiment tasks.
-E.2 Sparse Reward
-$\beta$ 0.2 0.15 0.1
-Train Test Train Test Train Test
-Dreamer 250.2 $\pm$ 9.6 3.0 $\pm$ 2.2 247.5 $\pm$ 1.0 0.0 $\pm$ 0.0 175.8 $\pm$ 55.8 13.5 $\pm$ 13.5
-TAD 951.9 $\pm$ 3.3 876.9 $\pm$ 51.1 927.6 $\pm$ 2.6 800.1 $\pm$ 121.4 608.5 $\pm$ 321.4 491.6 $\pm$ 389.4
-Table 6: Average cumulative reward (mean $\pm$ one std) over different target region (smaller $\beta$ represents smaller target region and more sparse return) of the best policy trained by Dreamer and TAD in Cheetah__\__Speed.
-For each $\beta$, we train agents in the train tasks and evaluate them in both train and test environments.
-Numbers greater than 95 percent of the best performance for each environment are bold.
-SR 0.0 0.8 0.9
-Train Test Train Test Train Test
-Dreamer 250.2 $\pm$ 9.6 3.0 $\pm$ 2.2 237.4 $\pm$ 14.9 28.9 $\pm$ 14.6 168.7 $\pm$ 96.2 157.2 $\pm$ 188.4
-TAD 951.9 $\pm$ 3.3 876.9 $\pm$ 51.1 841.9 $\pm$ 154.9 546.1 $\pm$ 291.6 $7$$\pm$ 170.1 716.8 $\pm$ 136.5
-Table 7: Average cumulative reward (mean $\pm$ one std) over different sparse rates of the best policy trained by Dreamer and TAD in Cheetah__\__Speed.
-For each sparse rate, we train agents in the train tasks and evaluate them in both train and test environments.
-Numbers greater than 95 percent of the best performance for each environment are bold.
+E.3 Sparse Reward
 In this part, we will evaluate TAD in more challenging settings with sparse rewards.
@@
 We supplement experiments to evaluate the performance of Dreamer and TAD with $n=5$ (SR=0.8) and $n=10$ (SR=0.9).
-As shown in Table 7, with the increasing of SR, although the performance of TAD decrease since inferring task context from sparse reward is extremely difficult, TAD still shows strong performance in train tasks and generalizes well to unseen test tasks.
-Appendix F Ethics Issues and Border Impact
+As shown in Table 7, with the increasing of SR, although the performance of TAD decreases since inferring task context from sparse reward is extremely difficult, TAD still shows strong performance in train tasks and generalizes well to unseen test tasks.
+Appendix F Ethics Issues and Broader Impact
 Designing agents that can generalize to unseen tasks is a major concern in reinforcement learning.

### v5

@@
 Abstract
-A long-standing goal of reinforcement learning is to acquire agents that can learn on various training tasks and generalize well on unseen tasks that may share a similar dynamic but with different reward functions.
-The ability to generalize across tasks is significant for real-world applications where the robot needs to adapt to varying reward mechanisms with the same embodiment.
-In this work, we first indicate that training general world models can utilize similar structures in these tasks and benefit training more generalizable agents.
-Extending world models into the task generalization setting, we introduce a novel method named Task Aware Dreamer (TAD), which integrates reward-informed features to identify consistent latent characteristics across tasks.
-Within TAD, we compute the variational lower bound of sample data log-likelihood, which introduces a new term designed to differentiate tasks using their states, as the optimization objective of our reward-informed world models.
-To demonstrate the advantages of the reward-informed policy utilized in TAD for handling the task distribution, we introduce a novel metric named Task Distribution Relevance (TDR) which quantitatively measures the relevance of different tasks.
-For tasks exhibiting a high TDR, i.e., these tasks differ significantly, we illustrate that Markovian policies struggle to distinguish them, thus it is necessary to utilize reward-informed policies in TAD.
-Extensive experiments in both image-based and state-based settings show that TAD can significantly improve the performance of handling different tasks meanwhile, especially for those with high TDR, and display a strong generalization ability to unseen tasks.
-Figure 1: An overview.
-Given a task distribution, we train the agent in training tasks and hope it to zero-shot generalize to test tasks.
-For improving the generalization, we propose TAD, which utilizes $\Pi_{3}$ to encode all historical information for inferring the current task and novel reward-informed world models for capturing invariant latent features.
+A long-standing goal of reinforcement learning is to achieve agents that can learn on various training tasks and generalize well on unseen tasks with similar transition dynamics but different reward functions.
+For instance, a household robot needs to use the same embodiment to complete various tasks characterized by distinct reward functions.
+To develop agents capable of handling this challenge, we construct Reward-Informed World Models (RIWM) that learns task-specific features while utilizing similar transition dynamics of trajectories collected from different tasks.
+The world model is optimized with a novel algorithm named Task Aware Dreamer (TAD), which employs a task context term to guide the world model to differentiate tasks and thus enhance task generalization.
+We further provide theoretical analyses to show that TAD’s key components are essential for handling task generalization.
+To establish the necessity of the policy hypothesis set utilized in TAD, we introduce a novel metric named Task Distribution Relevance (TDR), which quantitatively measures the relevance of different tasks.
+Tasks with high TAD often have vastly different optimal polices and thus are impossible to solve with traditional Markovian policy hypothesis sets, thereby necessitating the reward-informed policy used in TAD.
+Extensive experiments in both image-based and state-based benchmarks show that TAD can significantly improve the performance of handling various training tasks, especially for those with high TDR, and exhibits strong zero-shot generalization ability to unseen tasks.
+We also conduct ablation studies that demonstrate TAD’s potential to handle transition variations in cross-embodiment scenarios.
+keywords:
+reinforcement learning, task generalization, world models, model-based reinforcement learning, deep learning
 1 Introduction
-Deep Reinforcement Learning (DRL) has demonstrated significant advancements in diverse fields (Mnih et al.
-2016; Silver et al.
-2016), and a key factor in these achievements is an agent’s proficiency in assimilating lessons from special training tasks.
-This tendency towards specialization harms the wider real-world application of DRL, where broad generalization across various tasks is necessary.
-One primary challenge in generalization is the correspondence in the underlying dynamics of different tasks, which exhibit distinct reward structures.
-For example, we may need to control the same robot to handle different tasks, which can be rewarded for speed or caution in a navigating robot.
-Developing generalizable agents that recognize and handle these subtle variations is still an area of keen interest and exploration in DRL.
-For developing such generalizable agents, a promising pathway is to train general world models (Ha and Schmidhuber 2018; Hafner et al.
-2019b, a) that help the agent understand the world and make decisions.
-In this work, we first provide insight that general world models benefit improving the sample efficiency in handling the task distribution, especially by utilizing similar dynamic structures of trajectories sampled from different tasks (Theorem 1).
-Consequently, we propose a novel framework named Task Aware Dreamer (TAD) to improve the generalization ability of agents via general world models.
-As current world models are primarily designed for the single-task setting (left of Fig.
-2), we consider the corresponding probabilistic graphical model for the task-distribution setting (right of Fig.
-2) and propose novel reward-informed world models that can capture invariant structures across tasks in the distribution.
-Then we compute the variational lower bound of the data log-likelihood as the primary training objective for optimizing the reward-informed world models, which incorporate a novel task context term that categorizes tasks based on their states.
-In practice, we implement this optimization objective via two alternative methods: cross-entropy (TAD-CE) and supervised-contrastive (TAD-SC).
-Figure 2: Probabilistic graphical model designs for the single-task setting (left) and the task-distribution setting (right).
-The latter inspires the design of reward-informed world models.
-Solid and dashed lines represent the generative process and the inference model, respectively.
-Moreover, we theoretically explain TAD’s components, like reward-informed policies and task optimization terms, are effective for handling task distribution.
-About the policy hypothesis, although several meta RL methods (Zintgraf et al.
-2019; Rimon et al.
-2024) have chosen the reward-informed policy hypothesis $\Pi_{3}$, the relationship between the task distribution and the expressiveness of the policy hypothesis set is still unclear.
-In other words, why the commonly adopted hypothesis set $\Pi_{1}$ of Markovian policies and the set $\Pi_{2}$ of policies encoding historical states and actions are not suitable for task distribution?
-To answer this question, we propose a novel metric of Task Distribution Relevance (TDR), encapsulating the relevance of different tasks within the distribution through their optimal Q functions.
-We then prove that both $\Pi_{1}$ and $\Pi_{2}$ are sub-optimal under the task-distribution setting.
-This sub-optimality is related to TDR (Theorem 3), i.e., for task distributions with high TDR, the performance of these two policy hypotheses might degenerate significantly, a phenomenon we also demonstrate in experiments (Sec.
+Although Deep Reinforcement Learning (DRL) agents have demonstrated significant advancements in various fields [1, 2], their achievements are usually limited to specific training tasks.
+This tendency toward specialization harms the real-world applicability of DRL agents, where robust generalization across diverse tasks is necessary.
+For example, a household robot needs to handle different tasks that share the same underlying transition dynamics (i.e., the same embodiment and physical environment) but are characterized by distinct reward functions.
+Developing generalizable agents that can recognize and handle these subtle variations remains an active and important area of research in DRL.
+Compared to the traditional single-task setting, we formalize the problem as the task-distribution setting (Sec.
+3), where each training/testing task is sampled from a task distribution $\mathcal{T}$.
+To develop agents capable of handling this situation, we show that the world model [3, 4, 5] is a promising pathway to address the challenges posed by the variability of tasks.
+In this work, we first demonstrate that world models can enhance the sample efficiency when handling the task-distribution setting, particularly by exploiting shared dynamic structures of trajectories sampled from different tasks (Theorem 4.1).
+While current world models are primarily tailored for the single-task setting (left of Fig.
+1b), we extend the probabilistic graphical model to handle the task-distribution setting (right of Fig.
+1b) by introducing the task random variable $\mathcal{M}$.
+Building upon this framework, we extend the traditional world [3, 4, 5] models to our Reward-Informed World Models (RIWM, Eq.
+(LABEL:eq_world_model)).
+RIWM incorporates historical reward signals to optimize latent states and integrates a task model to infer the task context of the trajectory.
+We further propose a novel algorithm named Task Aware Dreamer (TAD) to optimize the aforementioned RIWM by maximizing the variational evidence lower bound.
+The derived task context term encourages the world model to learn compact representations that effectively distinguish among tasks, thereby enhancing the agents’ ability to generalize across tasks.
+In practice, we implement this objective using two alternative approaches: a cross-entropy–based method (TAD-CE) and a supervised contrastive method (TAD-SC).
+(a) Overview of this work.
+(b) Probabilistic graphical model designs.
+Figure 1: (a) An overview.
+Given a task distribution, the agent is trained on the training tasks and is expected to generalize to the test tasks in a zero-shot manner.
+Here $\Pi_{1}$ represents the set of Markovian policies, while $\Pi_{2}$ denotes all policies that include historical state and action information.
+To improve generalization ability, we propose TAD, which utilizes $\Pi_{3}$ to encode all historical information for inferring the current task and introduces novel Reward-Informed World Models to capture task-specific latent features.
+(b) Probabilistic graphical model designed for the single-task setting (left) and the task-distribution setting (right).
+The latter illustrates the design of Reward-Informed World Models, where solid and dashed lines denote the generative process and the inference model, respectively.
+To validate the effectiveness of the proposed method, we theoretically demonstrate that the key innovations of TAD are essential for addressing task generalization.
+Specifically, we first establish the necessity of adopting the reward-informed policy hypothesis $\Pi_{3}$ in the task-distribution setting.
+Although several meta RL methods [6, 7] practically choose $\Pi_{3}$, the relationship between the task distribution and the expressiveness of the policy hypothesis set remains insufficiently understood.
+In particular, we address the following question: Why are the commonly adopted hypothesis set $\Pi_{1}$ of Markovian policies and set $\Pi_{2}$ of policies that encode historical states and actions inadequate for the task-distribution setting?
+To answer this question, we propose a novel metric named Task Distribution Relevance (TDR), which encapsulates the relevance among tasks within the distribution through their optimal Q-functions.
+We then prove that both $\Pi_{1}$ and $\Pi_{2}$ are sub-optimal in the task-distribution setting.
+This sub-optimality is characterized by TDR (Theorem 4.4).
+Specifically, for task distributions with high TDR, the performance of these two policy hypothesis sets can degrade significantly, a phenomenon we also empirically demonstrate in experiments (Sec .
 5.2).
-This result explains why TAD and previous meta RL methods choose $\Pi_{3}$ for handling the task distribution.
-Besides the policy chosen, we also discuss our task optimization term in TAD and prove that it can effectively reduce the gap between the policy return and the optimal return (Theorem 4).
-We evaluate the task generalization ability of TAD in extensive experiments, including DeepMind control suite (Tassa et al.
-2018) and MuJoCo (Todorov, Erez, and Tassa 2012), which are image-based and state-based respectively.
-Agents are trained on various tasks and evaluated on unseen tasks.
-Results corroborate our analyses, indicating that $\Pi_{1}$ and $\Pi_{2}$ falter in managing task distributions characterized by a high TDR.
-Contrastingly, TAD excels in simultaneously managing varied tasks and outperforms all baselines, including SOTA model-based meta RL method MAMBA (Rimon et al.
-2024).
-Additionally, our ablation studies highlight TAD’s versatility, showing its prowess in dynamic generalization and handling cross-embodiment tasks.
-Overall, our contributions include:
-  • We present theoretical insights that general world models can utilize similar structures across tasks and improve the sample efficiency for task generalization (Sec. 4.1).
-  • Our TAD extends world models for the task distribution with the corresponding variational lower bound (Sec. 4.2).
-  • We theoretically analyze TAD’s components are effective for task generalization, with a novel metric TDR to quantify the distribution relevance (Sec. 4.3).
-  • Extensive experiments show that TAD can outperform various SOTA baselines and exhibit better generalization flexibility over image-based and state-based settings (Sec. 5).
+This result provides a theoretical justification for the adoption of $\Pi_{3}$ in TAD as well as in prior meta-RL methods when addressing task-distribution settings.
+Beyond the choice of policy hypothesis set, we further analyze the task context term in TAD and prove that it effectively reduces the gap between the policy return and the optimal return (Theorem 4.5).
+We conduct extensive experiments to evaluate TAD’s ability to enhance task generalization, using commonly adopted benchmarks from the DeepMind Control Suite [8] and MuJoCo [9].
+Both training and testing tasks are drawn from a task distribution $\mathcal{T}$, which requires the agents to simultaneously learn from multiple training tasks and generalize to potentially unseen test tasks.
+The experimental results corroborate our theoretical analysis, showing that policies from $\Pi_{1}$ and $\Pi_{2}$ struggle to handle tasks from distributions with a high TDR.
+Conversely, TAD not only effectively handles diverse training tasks but also exhibits better zero-shot generalization performance than the few-shot adaptation performance of the SOTA meta-RL method MAMBA [7].
+Moreover, our ablation studies demonstrate the versatility of TAD, highlighting its potential for dynamics generalization and handling cross-embodiment task settings.
+Overall, our contributions are summarized as follows:
+  • We provide theoretical insights showing that world models can exploit shared transition dynamics across tasks to improve sample efficiency for task generalization (Sec. 4.1).
+  • We extend world models to the task-distribution setting by proposing TAD, which introduces a task context term that encourages representations to capture task-relevant information (Sec. 4.2).
+  • We theoretically demonstrate the effectiveness of TAD’s components for task generalization and introduce a novel metric named TDR to quantify task distribution relevance (Sec. 4.3).
+  • Extensive experiments show that TAD outperforms a range of SOTA baselines and exhibits strong generalization flexibility across both image-based and state-based benchmarks (Sec. 5).
 2 Related work
 Generalization in RL.
-Current RL methods always struggle to generalize to new tasks (Song et al.
-2019).
-Prior works have studied an array of training strategies, like loss regularization (Cobbe et al.
-2019; Wang et al.
-2020), successor representation (Touati and Ollivier 2021; Touati, Rapin, and Ollivier 2022), network architecture design (Lee et al.
-2019; Raileanu and Fergus 2021), lifelong learning (Chen and Liu 2018; Mendez, van Seijen, and Eaton 2022), data augmentation (Raileanu et al.
-2021; Hansen and Wang 2021), etc.
-Besides these, some works investigate the connection between policy generalization and the distribution of those environments.
-Ghosh et al.
-demonstrates that generalizing to unseen environments introduces partial observability, thereby rendering deterministic Markovian policies sub-optimal.
-Also, some studies (Lee et al.
-2020; Ghosh et al.
-2021) experimentally indicate that stochastic or non-Markovian policies can improve the generalization ability.
-However, the expressive abilities of differing hypothesis sets and their connection to the environment distribution, which are significant for developing more generalizable agents, remain understudied.
+Current RL agents often struggle to generalize to new tasks [10].
+Prior work has explored a wide range of training strategies to address this challenge, including loss regularization [11, 12], successor representations [13, 14], network architecture design [15, 16], data augmentation [17, 18], leveraging trajectory similarities [19], and contrastive objectives [20].
+Beyond these approaches, several studies have investigated the relationship between policy generalization and the distribution over environments.
+In particular, LEEP [21] demonstrates that generalizing to unseen environments introduces partial observability, which makes deterministic Markovian policies suboptimal.
+Moreover, empirical evidence from prior works [22, 21] suggests that stochastic or non-Markovian policies can enhance generalization ability.
+Despite these insights, a theoretical gap remains regarding the expressiveness of different hypothesis sets and their relationship to the environment distribution, both of which are crucial for developing more generalizable agents.
 Multi-task RL and Meta RL.
 These two topics are closely related to generalization in RL.
-Multi-task RL (Yang et al.
-2020; Sodhani, Zhang, and Pineau 2021; Lee et al.
-2022; Xu et al.
-2022a) primarily aims to excel across all training tasks but is difficult to zero-shot generalize to unseen tasks.
-For boosting the generalization, Meta RL seeks to enable the trained agents to adapt to new tasks with few episodes, including gradient-based (Finn, Abbeel, and Levine 2017) and context-based methods (Duan et al.
-2016; Rakelly et al.
-2019).
-There are also some model-based methods (Nagabandi et al.
-2018; Rimon et al.
-2024) utilize learned models to boost the sample efficiency.
-Though some context-based methods like VariBAD (Zintgraf et al.
-2019) show zero-shot generalization ability, it is still significant to directly analyze the generalization in RL and design corresponding algorithms.
+Multi-task RL [23, 24, 25, 26] primarily focuses on excelling across all training tasks, but often struggles to generalize to previously unseen tasks.
+In contrast, Meta RL aims to equip agents with the ability to rapidly adapt to new tasks with a few episodes, through approaches such as gradient-based [27] and context-based methods [28, 29].
+In addition, several model-based approaches [30, 7] leverage learned environment models to improve sample efficiency.
+Although certain context-based methods, such as VariBAD [31], demonstrate zero-shot generalization capabilities, it remains important to directly study the mechanisms of generalization in RL and to develop algorithms specifically tailored to this objective.
 World models.
-World models (Ha and Schmidhuber 2018) aims to better learn environmental representations, which has potential advantages for generalization as it can capture invariant features across tasks.
-Classical methods utilize the Recurrent State Space Model (RSSM) (Hafner et al.
-2019b) for planning (Hafner et al.
-2019b) and policy learning (Hafner et al.
-2019a, 2020, 2023).
-Subsequent research explored reconstruction-free world models (Deng, Jang, and Ahn 2022), temporal predictive coding (Nguyen et al.
-2021), cooperative reconstruction (Fu et al.
-2021), and Denoised MDP (Wang et al.
-2022) for more effective task-relevant information encoding.
-World models are also utilized to extract environmental invariant features by learning from videos (Seo et al.
-2022) or exploration (Sekar et al.
-2020; Xu et al.
-2022b), and then fine-tuning to new tasks.
-However, most world models are designed for the single-task setting and struggle to manage multiple tasks without fine-tuning, limiting their effectiveness for zero-shot generalization to unseen tasks.
+World models [3] aim to learn rich representations of the environment, offering potential advantages for generalization by capturing task-invariant features.
+Classical approaches typically adopt the Recurrent State Space Model (RSSM) [4] for planning [4] and policy learning [5, 32, 33].
+Subsequent work has proposed reconstruction-free world models [34], temporal predictive coding [35], cooperative reconstruction [36], and Denoised MDPs [37] to more effectively encode task-relevant information.
+World models have also been leveraged to extract environment-invariant features through videos [38] or unsupervised exploration [39, 40], followed by fine-tuning on downstream tasks.
+Despite these advances, most existing world models are designed for single-task settings and require task-specific fine-tuning to handle multiple tasks, which limits their ability to achieve zero-shot generalization to unseen tasks.
 3 Preliminary
-We consider the setting with a task distribution $\mathcal{T}$ of Partially observable Markov decision processes (POMDPs), where different tasks own the same dynamic and different rewards.
-Formally, each POMDP $\mathcal{M}\sim\mathcal{T}$ can be represented as $\mathcal{M}=(\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R}_{\mathcal{M}},\Omega,\mathcal{O})$.
-Here $\mathcal{S}$ and $\mathcal{A}$ denote the state and action spaces, respectively.
-For $\forall(s,a)\in\mathcal{S}\times\mathcal{A}$, $\mathcal{P}(\cdot|s,a)$ is the state transition probability that is Markovian, $\mathcal{R}_{\mathcal{M}}(s,a)$ is its reward function.
-The state is unobserved for the agent, which can only access the observation from the observation space $\Omega$ calculated by the observation function $\mathcal{O}(\cdot|s)$.
-Following previous meta RL and generalization methods, we consider policies that encoder all historical information (we prove its necessity in Sec.
+We consider a setting with a task distribution $\mathcal{T}$ over Partially Observable Markov Decision Processes (POMDPs), where different tasks share identical dynamics but differ in their reward functions.
+Formally, each POMDP $\mathcal{M}\sim\mathcal{T}$ is defined as $\mathcal{M}=(\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R}_{\mathcal{M}},\Omega,\mathcal{O})$.
+Here, $\mathcal{S}$ and $\mathcal{A}$ denote the state and action spaces, respectively.
+For $\forall(s,a)\in\mathcal{S}\times\mathcal{A}$, $\mathcal{P}(\cdot|s,a)$ specifies the Markovian state transition dynamics, while $\mathcal{R}_{\mathcal{M}}(s,a)$ denotes the task-specific reward function.
+The underlying state $s$ is not directly observable by the agent.
+Instead, the agent receives observations from the observation space $\Omega$ according to the observation function $\mathcal{O}(\cdot\mid s)$.
+Following prior work on meta-RL and generalization, we consider policies that encode all historical information (we prove the necessity of this assumption in Sec.
 4.3).
-Formally, at each timestep $t$, the agent with the policy $\pi$ will use the whole history trajectory $(o_{0},a_{0},r_{0},o_{1},...,o_{t})$ to sample action $a_{t}$, arrive at the next state $s_{t+1}\sim\mathcal{P}(\cdot|s_{t},a_{t})$, and get the current reward $r_{t}=\mathcal{R}_{\mathcal{M}}(s_{t},a_{t})$.
-The performance of policy $\pi$ in $\mathcal{M}$ is defined as the expected discounted return: $J_{\mathcal{M}}(\pi)=\mathbb{E}_{\tau\sim\pi}\left[R(\tau)\triangleq\sum_{t=0}^{\infty}\gamma^{t}r_{t}\right]$.
-Our objective is to maximize the return over $\mathcal{T}$, i.e., $\max_{\pi}\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}\left[J_{\mathcal{M}}(\pi)\right]$.
-In practice, given the task distribution $\mathcal{T}$, we sample $M$ training tasks $\{\mathcal{M}_{m}\}_{m=1}^{M}$ for optimizing the agent, i.e., maximizing $\frac{1}{M}\sum_{m=1}^{M}J_{\mathcal{M}_{m}}(\pi)$.
-In the testing stage, we will sample $N$ unseen test tasks $\{\mathcal{M}_{M+n}\}_{n=1}^{N}$ to evaluate its generalization ability, i.e., evaluating $\frac{1}{N}\sum_{n=1}^{N}J_{\mathcal{M}_{M+n}}(\pi)$.
+Formally, at each time step $t$, the agent with the policy $\pi$ will utilize the whole historical trajectory $(o_{0},a_{0},r_{0},o_{1},...,o_{t})$ to sample action $a_{t}$, arrive at the next state $s_{t+1}\sim\mathcal{P}(\cdot|s_{t},a_{t})$, and receive current reward $r_{t}=\mathcal{R}_{\mathcal{M}}(s_{t},a_{t})$.
+The performance of policy $\pi$ on task $\mathcal{M}$ is measured by the expected discounted return $J_{\mathcal{M}}(\pi)=\mathbb{E}_{\tau\sim\pi}\left[R(\tau)\triangleq\sum_{t=0}^{\infty}\gamma^{t}r_{t}\right]$.
+Our objective is to learn a policy that maximizes the expected return over the task distribution $\mathcal{T}$: $\max_{\pi}\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}\left[J_{\mathcal{M}}(\pi)\right]$.
+In practice, given the task distribution $\mathcal{T}$, we sample $M$ training tasks $\{\mathcal{M}_{m}\}_{m=1}^{M}$ to optimize the agent by maximizing the empirical objective $\frac{1}{M}\sum_{m=1}^{M}J_{\mathcal{M}_{m}}(\pi)$.
+At test time, we sample $N$ previously unseen tasks $\{\mathcal{M}_{M+n}\}_{n=1}^{N}$ to evaluate the agent’s generalization performance, measured by $\frac{1}{N}\sum_{n=1}^{N}J_{\mathcal{M}_{M+n}}(\pi)$.
 4 Task aware dreamer
-In this section, we first demonstrate that general world models benefit task generalization.
-Then we introduce our reward-informed world models and propose a novel framework of Task Aware Dreamer (TAD) for handling task generalization.
-Finally, we provide theoretical analyses of designs in TAD.
+In this section, we begin by establishing that world models are beneficial for task generalization.
+We then introduce Reward-Informed World Models (RIWM) and propose Task-Aware Dreamer (TAD) as a novel approach to promote generalization.
+Finally, we provide theoretical analyses demonstrating that TAD can effectively enhance performance across various tasks.
 4.1 Reward-informed world models
-Our first observation is that general world models are effective in narrowing down the hypothesis space of the optimal Q function when handling task generalization:
-Theorem 1 (Proof in Appendix A.1).
-Set $\mathcal{Q}$ as the space of observation-action Q functions.
-Given $M$ tasks $\{\mathcal{M}_{m}\}_{m=1}^{M}$ and corresponding dataset $\mathcal{D}_{m}=\{(o_{t}^{m},a_{t}^{m},r_{t}^{m},o_{t+1}^{m})\}$, we set the product space $\mathcal{H}=\mathcal{Q}^{M}$ composed of M spaces, i.e., $\forall\{q_{m}\}_{m=1}^{M}\in\mathcal{H}$, $q_{m}:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ belongs to $\mathcal{Q}$.
-Considering the following three hypothesis classes $\mathcal{H}_{1},\mathcal{H}_{2},\mathcal{H}_{3}\subseteq\mathcal{H}$:
-$\begin{split}\mathcal{H}_{1}=&\{(q_{m})_{m=1}^{M}|q_{m}(o_{t}^{m},a_{t}^{m})=r_{t}^{m}+\gamma\max_{a^{\prime}}q_{m}(o_{t+1}^{m},a^{\prime})\}\\ \mathcal{H}_{2}=&\{(q_{m})_{m=1}^{M}|\exists(p_{m})_{m=1}^{M}:p_{m}(o_{t}^{m},a_{t}^{m})=o_{t+1}^{m},\\ &\exists(r_{m})_{m=1}^{M}:r_{m}(o_{t}^{m},a_{t}^{m})=r_{t}^{m},\\ &q_{m}(o,a)=r_{m}(o,a)+\gamma\max_{a^{\prime}}q_{m}(p_{m}(o,a),a^{\prime}),\forall o,a\}\\ \mathcal{H}_{3}=&\{(q_{m})_{m=1}^{M}|\exists p:p(o_{t}^{m},a_{t}^{m})=o_{t+1}^{m},\\ &\exists(r_{m})_{m=1}^{M}:r_{m}(o_{t}^{m},a_{t}^{m})=r_{t}^{m},\\ &q_{m}(o,a)=r_{m}(o,a)+\gamma\max_{a^{\prime}}q_{m}(p(o,a),a^{\prime}),\forall o,a\}.\end{split}$
-Then we have $\mathcal{H}_{3}\subseteq\mathcal{H}_{2}\subseteq\mathcal{H}_{1}$.
-Here $\mathcal{H}_{1}$ own Q functions satisfying optimal Bellman equation with data from $\mathcal{D}_{m}$, while $\mathcal{H}_{2}$ and $\mathcal{H}_{3}$ own Q functions satisfying optimal Bellman function with data from world models, which are trained with the single dataset and all datasets, respectively.
-Consequently, $\mathcal{H}_{2}\subseteq\mathcal{H}_{1}$ utilizes the generalization ability of world models, which extends the results in the fixed-task setting (Young et al.
-2023).
-Moreover, $\mathcal{H}_{3}\subseteq\mathcal{H}_{2}$ illustrate similar dynamic structures across different tasks benefit learning general world models.
-Consequently, training general world models can narrow down the hypothesis of possible Q function and benefit agent training.
-As existing world models (Hafner et al.
-2019b, a) are mainly designed for the single-task setting, we propose reward-informed world models for the task-distribution setting.
-We first analyze the probabilistic graphical model of the task-distribution setting shown in Fig.
-2, where rewards do not only rely on the previous states and actions but also the current task.
-Thus we can calculate the joint distribution as
-$p(s_{1:T},o_{1:T},r_{1:T},a_{1:T-1},\mathcal{M})$ (1)
-$=$ $p(\mathcal{M})\prod_{t=1}p(s_{t+1}|s_{t},a_{t})p(o_{t}|s_{t})p(r_{t}|s_{t},\mathcal{M}),$
-and we choose the reward-informed inference model (the necessity of reward-informed policy in task generalization is discussed in Sec. 4.3) to approximate state posteriors as
+Our first observation is that world models are effective in narrowing the hypothesis space of the optimal Q-function when handling task generalization:
+Theorem 4.1 (Proof in Appendix A.1).
+Set $\mathcal{Q}$ denotes the space of observation-action Q-functions.
+Given $M$ tasks $\{\mathcal{M}_{m}\}_{m=1}^{M}$ and corresponding dataset $\mathcal{D}_{m}=\{(o_{t}^{m},a_{t}^{m},r_{t}^{m},o_{t+1}^{m})\}$, we use $\mathcal{H}=\mathcal{Q}^{M}$ to represent the product space of $M$ Q-function spaces, i.e., $\forall\{q_{m}\}_{m=1}^{M}\in\mathcal{H}$, $q_{m}:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ belongs to $\mathcal{Q}$.
+Considering the following three hypothesis sets $\mathcal{H}_{1},\mathcal{H}_{2},\mathcal{H}_{3}\subseteq\mathcal{H}$:
+$\begin{split}\mathcal{H}_{1}=&\{(q_{m})_{m=1}^{M}|q_{m}(o_{t}^{m},a_{t}^{m})=r_{t}^{m}+\gamma\max_{a^{\prime}}q_{m}(o_{t+1}^{m},a^{\prime})\}\\ \mathcal{H}_{2}=&\{(q_{m})_{m=1}^{M}|\exists(p_{m})_{m=1}^{M}:p_{m}(o_{t}^{m},a_{t}^{m})=o_{t+1}^{m},\\ &\exists(r_{m})_{m=1}^{M}:r_{m}(o_{t}^{m},a_{t}^{m})=r_{t}^{m},q_{m}(o,a)=r_{m}(o,a)+\gamma\max_{a^{\prime}}q_{m}(p_{m}(o,a),a^{\prime}),\forall o,a\}\\ \mathcal{H}_{3}=&\{(q_{m})_{m=1}^{M}|\exists p:p(o_{t}^{m},a_{t}^{m})=o_{t+1}^{m},\\ &\exists(r_{m})_{m=1}^{M}:r_{m}(o_{t}^{m},a_{t}^{m})=r_{t}^{m},q_{m}(o,a)=r_{m}(o,a)+\gamma\max_{a^{\prime}}q_{m}(p(o,a),a^{\prime}),\forall o,a\}.\end{split}$
+The hypothesis sets satisfy $\mathcal{H}_{3}\subseteq\mathcal{H}_{2}\subseteq\mathcal{H}_{1}$.
+Here, $\mathcal{H}_{1}$ contains Q-functions that satisfy the optimal Bellman equation with data from $\mathcal{D}_{m}$.
+In contrast, $\mathcal{H}_{2}$ and $\mathcal{H}_{3}$ contain Q-functions that satisfy the optimal Bellman equation using data generated by world models trained on a single dataset and on all datasets, respectively.
+Consequently, $\mathcal{H}_{2}\subseteq\mathcal{H}_{1}$ utilizes the generalization ability of world models, extending prior results in the fixed-task setting [41].
+Moreover, $\mathcal{H}_{3}\subseteq\mathcal{H}_{2}$ suggests that shared dynamic structures across tasks benefit learning world models.
+Overall, Theorem 4.1 indicates that training world models narrows the hypothesis space of plausible Q-functions, thereby facilitating agent learning and improving generalization.
+As existing world models [4, 5] are primarily designed for the single-task setting, we propose Reward-Informed World Models (RIWM) for the task-distribution setting.
+We begin by analyzing the probabilistic graphical model of this setting, illustrated in Fig.
+1b, where the rewards depend not only on previous states and actions but also on the current task.
+Accordingly, the joint distribution can be expressed as:
+$p(s_{1:T},o_{1:T},r_{1:T},a_{1:T-1},\mathcal{M})=p(\mathcal{M})\prod_{t=1}p(s_{t+1}|s_{t},a_{t})p(o_{t}|s_{t})p(r_{t}|s_{t},\mathcal{M}),$ (1)
+and we choose to adopt the reward-informed inference model (the necessity of reward-informed policy in task generalization is discussed in Sec. 4.3) to approximate state posteriors as:
 $q(s_{1:T}|o_{1:T},a_{1:T},r_{1:T})=\prod_{t=1}^{T}q(s_{t}|s_{t-1},a_{t-1},r_{t-1},o_{t}).$ (2)
-Based on this inference model, we can construct the variational lower bound of the log-likelihood on the data as
-$\begin{split}&\ln p(o_{1:T},r_{1:T},\mathcal{M}|a_{1:T})\\ \geq&\sum_{t=1}^{T}\mathbb{E}_{q(s_{t}|o_{\leq t},a_{<t},r_{<t})}\left[\ln p(o_{t},r_{t}|\mathcal{M},s_{t})\right.\\ -&\left.\text{KL}\left(q(s_{t}|o_{\leq t},r_{<t},a_{<t})\|p(s_{t}|s_{t-1},a_{t-1})\right)\right]\\ +&\mathbb{E}_{q(s_{1:T}|o_{1:T},a_{1:T},r_{1:T})}\left[\ln p(\mathcal{M}|s_{1:T})\right].\end{split}$ (3)
-This result is a general form of the single-task setting (Hafner et al.
-2019b) and its derivation is in Appendix A.2.
-The first two terms in Eq.(LABEL:eq_elbo) are for reconstructing observations, predicting rewards, and inferring states, which are similar to the single-task setting.
-The last novel term is dedicated to predicting the current task from historical information, which is beneficial for improving the generalization as it encourages inferring the current task context.
-Extending RSSM via Eq.(LABEL:eq_elbo), our reward-informed world models consist of:
-Deterministic state model: $h_{t}=f(h_{t-1},s_{t-1},a_{t-1},r_{t-1}),$
+Based on this inference model, we can construct the variational lower bound of the action-conditioned probability as:
+$\begin{split}\ln p(o_{1:T},r_{1:T},\mathcal{M}|a_{1:T})\geq&\sum_{t=1}^{T}\mathbb{E}_{q(s_{t}|o_{\leq t},a_{<t},r_{<t})}\left[\ln p(o_{t},r_{t}|\mathcal{M},s_{t})-\mathrm{KL}\left(q(s_{t}|o_{\leq t},r_{<t},a_{<t})\|p(s_{t}|s_{t-1},a_{t-1})\right)\right]\\ &+\mathbb{E}_{q(s_{1:T}|o_{1:T},a_{1:T},r_{1:T})}\left[\ln p(\mathcal{M}|s_{1:T})\right].\end{split}$ (3)
+This result is a general form of the single-task setting [4], with the full derivation in Appendix A.2.
+The first two terms in Eq.
+(3) correspond to reconstructing observations and predicting rewards, which are similar to the single-task setting.
+The final novel term is dedicated to predicting the current task from historical information, which enhances generalization by encouraging the model to infer the current task context.
+Building on RSSM via Eq.
+(3), our RIWM consists of:
+Deterministic state model: $h_{t}=f(h_{t-1},s_{t-1},a_{t-1},r_{t-1}),$ (4)
 Transition model: $p_{\theta}(s_{t}|h_{t}),$
@@
 Task model: $p_{\theta}(\mathcal{M}|h_{t},s_{t}).$
-Here hidden state $h_{t}$ encodes historical states, actions, and rewards, by using gated recurrent unit (GRU) (Chung et al.
-2014) as $f(\cdot)=\text{GRU}(\cdot)$.
-Then the transition model, observation model, reward model, and task model further predict state, observation, reward, and task context respectively.
+Here, the hidden state $h_{t}$ encodes historical states, actions, and rewards using a gated recurrent unit (GRU) [42], i.e., $f(\cdot)=\text{GRU}(\cdot)$.
+The transition, observation, reward, and task models then predict the next state, observation, reward, and task context, respectively.
 4.2 Optimization
-Based on the above analyses, we now introduce the training of TAD in detail.
-Following previous world models (Hafner et al.
-2019a), we adopt an alternating training approach between the reward-informed world models and the policy.
-To balance different tasks, when collecting data, TAD utilizes $M$ replay buffers $\{\mathcal{D}_{m}\}_{m=1}^{M}$ to store trajectories sampled from $\{\mathcal{M}_{m}\}_{m=1}^{M}$, respectively.
-Then, TAD samples data from each replay buffer and trains the reward-informed world models via the optimization objective following Eq.
-(LABEL:eq_elbo) as
-$\begin{split}&L_{\text{TAD}}=\sum_{i=1}^{M}\mathbb{E}_{q}\left[\sum_{t=1}^{T}\ln p_{\theta}(o_{t}^{i}|h_{t}^{i},s_{t}^{i})+\sum_{t=1}^{T}\ln p_{\theta}(r_{t}^{i}|h_{t}^{i},s_{t}^{i})\right.\\ &-\left.\sum_{t=1}^{T}D_{\text{KL}}(q(s_{t}^{i}|h_{t}^{i},o_{t}^{i})\|p_{\theta}(s_{t}^{i}|h_{t}^{i}))+L_{\text{task}}\right].\end{split}$ (4)
+Based on the above analyses, we now describe the training of TAD in detail. Following prior world model approaches [5], TAD adopts an alternating training scheme between the RIWM and the policy.
+To ensure balanced learning across tasks, TAD maintains $M$ replay buffers $\{\mathcal{D}_{m}\}_{m=1}^{M}$ to store trajectories collected from the corresponding task $\{\mathcal{M}_{m}\}_{m=1}^{M}$.
+During training, TAD samples data from each replay buffer and updates the RIWM by optimizing the objective in Eq.
+(3) as:
+$\begin{split}&L_{\text{TAD}}=\sum_{i=1}^{M}\mathbb{E}_{q}\left[\sum_{t=1}^{T}\ln p_{\theta}(o_{t}^{i}|h_{t}^{i},s_{t}^{i})+\sum_{t=1}^{T}\ln p_{\theta}(r_{t}^{i}|h_{t}^{i},s_{t}^{i})\right.-\left.\sum_{t=1}^{T}D_{\text{KL}}(q(s_{t}^{i}|h_{t}^{i},o_{t}^{i})\|p_{\theta}(s_{t}^{i}|h_{t}^{i}))+L_{\text{task}}\right].\end{split}$ (5)
 In Eq.
-(LABEL:eq_loss), the first three items are similar to Dreamer for reconstructing observations, predicting rewards, and inferring states.
-Besides them, TAD includes an additional task term $L_{\text{task}}$ for predicting different tasks and learning task-aware embedding.
-In detail, we provide two alternatives: cross-entropy and self-contrastive, which are introduced below (We theoretically demonstrate their effectiveness in Sec.
-4.3).
+(LABEL:eq_loss), the first three terms are similar to Dreamer for reconstructing observations, predicting rewards, and inferring states.
+In addition, TAD introduces a novel task context term $L_{\text{task}}$ to distinguish tasks and learn task-aware embeddings.
+Specifically, we propose two alternative formulations for this term: cross-entropy and self-contrastive, which are described in detail below.
+A theoretical justification for their effectiveness is provided in Sec.
+4.3.
 Cross-entropy.
-The last term in Eq.
-LABEL:eq_elbo indicates that we need to maximize the log probability of the task context to distinguish different tasks via historical information.
-Thus TAD-CE directly maximizes the log probability over different tasks represented by one-hot vectors following previous works (Yang et al.
-2020) and $L_{\text{task}}$ is set as below.
+The final term in Eq.
+(3) emphasizes the need to maximize the log-probability of the task context, enabling the model to distinguish tasks based on historical information.
+Following prior work [23], TAD-CE directly maximizes this log-probability over different tasks represented by one-hot vectors.
+The task loss $L_{\text{task}}$ is thus defined as:
 $\begin{split}L_{\text{task}}=\sum_{t=1}^{T}\ln p_{\theta}(m^{i}|h_{t}^{i},s_{t}^{i}).\end{split}$ (6)
 Supervised-contrastive.
-Besides directly maximizing the log probability, we further propose TAD-SC, borrowing the idea of supervised contractive learning (Khosla et al.
-2020).
-In detail, TAD-SC keeps the task embeddings of the same task closer and the task embeddings of different tasks far apart, which benefits reward-informed world models to better distinguish different tasks.
-Formally, we assume that the task model maps all sampled data as $\{m_{j}\}_{j=1}^{M\times T}$ and set $L_{\text{task}}$ as
+In addition to directly maximizing the log-probability, we further propose TAD-SC, inspired by supervised contrastive learning [43].
+Specifically, TAD-SC encourages task embeddings from the same task to be close while pushing embeddings from different tasks apart, enabling RIWM to distinguish tasks more effectively.
+Formally, we assume that the task model maps all sampled data as $\{m_{j}\}_{j=1}^{M\times T}$ and define $L_{\text{task}}$ as:
 $\begin{split}L_{\text{task}}=\sum_{j=1}^{M\times T}\sum_{a\in A(j)}\ln\frac{\exp(m_{j}\cdot m_{a}/\tau)}{\sum_{b\neq j}\exp(m_{j}\cdot m_{b}/\tau)},\end{split}$ (7)
-here $A(j)$ is the set of indices that are sampled with the same task of $m_{j}$, and $\tau$ is the temperature parameter, which is set as 0.1 following previous works (Khosla et al. 2020).
-In terms of training the actor-critic that are parameterized neural networks, we extend the actor-critic learning in Dreamer to the task distribution.
-We first sample a series of states from the replay buffer and start from them to imagine trajectories via our reward-informed world models (results are in Fig.
-3), which can capture invariant features and are beneficial for the agent to gain better generalization.
-After obtaining imagined trajectories, the actor-critic is optimized via maximizing the $\lambda$-return (Schulman et al.
-2015) and regressing the TD targets (Sutton and Barto 2018), respectively.
+where $A(j)$ is the set of data indices that are collected from task $m_{j}$, and $\tau$ is the temperature parameter, which is set as 0.1 following previous works [43].
+For training the actor-critic networks, we extend the actor-critic learning in Dreamer to the task distribution.
+Specifically, we first sample a set of states from the replay buffers and use our RIWM to imagine trajectories starting from these states (illustrated in Fig.
+2).
+These imagined trajectories can capture task-specific features and are thus beneficial for the agent to gain better generalization.
+Once the imagined trajectories are obtained, the actor and critic are optimized by maximizing the $\lambda$-return [44] and regressing the TD targets [45], respectively.
 Algorithm 1 Task Aware Dreamer (TAD)
-0:  $M$ training tasks $\{\mathcal{M}_{m}\}_{m=1}^{M}$, $M$ replay buffers $\{\mathcal{D}_{m}\}_{m=1}^{M}$, $N$ test tasks $\{\mathcal{M}_{M+n}\}_{n=1}^{N}$, initialize parameters of world models, the policy, and the critic.
-[…]
-2:     for update $\text{update step}=1,2,...,U$ do
+0:$M$ training tasks $\{\mathcal{M}_{m}\}_{m=1}^{M}$, $M$ replay buffers $\{\mathcal{D}_{m}\}_{m=1}^{M}$, NN test tasks $\{\mathcal{M}_{M+n}\}_{n=1}^{N}$, initialize parameters of world models, the policy, and the critic.
+1:for $\text{iteration step}=1,2,$... do
+2:  for $\text{update step}=1,2,$... do
 3:   Sample $o$-$a$-$r$ pairs $\{(o_{t}^{i},a_{t}^{i},r_{t}^{i})_{t=1}^{T}\}$ form each replay buffer $\mathcal{D}_{i},i=1,2,...,M$
@@
 Are policies that utilize all historical information in TAD necessary for task generalization?
-Below we introduce $3$ types of widely used policy hypotheses and show that $\Pi_{3}$, used in TAD, is necessary for handling task generalization.
+Below, we introduce $3$ types of widely used policy hypothesis sets and show that $\Pi_{3}$, used in TAD, is necessary for handling task generalization.
   1. 1.
-    Markovian policy set $\Pi_{1}$ (Sutton and Barto 2018; Yarats et al.
-2021), i.e., $\Pi_{1}=\{\pi|\pi:\mathcal{S}\rightarrow\Delta(\mathcal{A})\}$, here $\Delta(\mathcal{A})$ represents a distribution over $\mathcal{A}$, which is widely used and optimal for the single-task setting;
+    Markovian policy set $\Pi_{1}$ [45, 46], i.e., $\Pi_{1}=\{\pi|\pi:\mathcal{S}\rightarrow\Delta(\mathcal{A})\}$, where $\Delta(\mathcal{A})$ represents a distribution over $\mathcal{A}$, which is widely used and optimal for the single-task setting;
   2. 2.
-    $\mathcal{S}$-$\mathcal{A}$ memorized policy set $\Pi_{2}$ (Hafner et al.
-2019a, 2020; Lee et al.
-2020), i.e., $\Pi_{2}=\{\pi|\pi:\mathcal{H}\rightarrow\Delta(\mathcal{A})\}$, here $\mathcal{H}=\cup_{t=1}^{\infty}\mathcal{H}_{t},\mathcal{H}_{t}=(\mathcal{S}\times\mathcal{A})^{t-1}\times\mathcal{S}$;
+    $\mathcal{S}$-$\mathcal{A}$ memorized policy set $\Pi_{2}$ [5, 32, 22], i.e., $\Pi_{2}=\{\pi|\pi:\mathcal{H}\rightarrow\Delta(\mathcal{A})\}$, where $\mathcal{H}=\cup_{t=1}^{\infty}\mathcal{H}_{t},\mathcal{H}_{t}=(\mathcal{S}\times\mathcal{A})^{t-1}\times\mathcal{S}$;
   3. 3.
-    $\mathcal{S}$-$\mathcal{A}$$-\mathcal{R}$ memorized policy set $\Pi_{3}$ (Zintgraf et al.
-2019; Rimon et al.
-2024), i.e., $\Pi_{3}=\{\pi|\pi:\mathcal{L}\rightarrow\Delta(\mathcal{A})\}$, here $\mathcal{L}=\cup_{t=1}^{\infty}\mathcal{L}_{t},\mathcal{L}_{t}=(\mathcal{S}\times\mathcal{A}\times\mathbb{R})^{t-1}\times\mathcal{S}$.
+    $\mathcal{S}$-$\mathcal{A}$-$\mathcal{R}$ memorized policy set $\Pi_{3}$ [31, 7], i.e., $\Pi_{3}=\{\pi|\pi:\mathcal{L}\rightarrow\Delta(\mathcal{A})\}$, where $\mathcal{L}=\cup_{t=1}^{\infty}\mathcal{L}_{t},\mathcal{L}_{t}=(\mathcal{S}\times\mathcal{A}\times\mathbb{R})^{t-1}\times\mathcal{S}$.
 As illustrated in Fig.
-1, naturally $\Pi_{1}\subseteq\Pi_{2}\subseteq\Pi_{3}$.
-Now we will analyze their expressive ability under task distribution $\mathcal{T}$.
-Denote $J_{\mathcal{T}}^{*}\triangleq\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}\left[\max_{\pi}J_{\mathcal{M}}(\pi)\right]$ as the optimal return under $\mathcal{T}$, and $J_{\mathcal{T}}^{i}\triangleq\max_{\pi\in\Pi_{i}}\left[\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}J_{\mathcal{M}}(\pi)\right]$ as the optimal return for $\Pi_{i}(i=1,2,3)$ under $\mathcal{T}$.
+1a, these hypothesis sets naturally satisfy $\Pi_{1}\subseteq\Pi_{2}\subseteq\Pi_{3}$.
+We then analyze their expressive ability under task distribution $\mathcal{T}$.
+Let $J_{\mathcal{T}}^{*}\triangleq\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}\left[\max_{\pi}J_{\mathcal{M}}(\pi)\right]$ denote the optimal return under $\mathcal{T}$, and $J_{\mathcal{T}}^{i}\triangleq\max_{\pi\in\Pi_{i}}\left[\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}J_{\mathcal{M}}(\pi)\right]$ the optimal return achievable by policies in $\Pi_{i}(i=1,2,3)$ under $\mathcal{T}$.
 Our first result shows that, although $\Pi_{1}\subseteq\Pi_{2}$, they own the same expressive ability, i.e., $J_{\mathcal{T}}^{1}=J_{\mathcal{T}}^{2}$, and are both sub-optimal:
-Theorem 2 (Sub-Optimality of $\Pi_{1},\Pi_{2}$. Proof in Appendix A.3).
-We set ¯$\bar{\mathcal{M}}=(\mathcal{S},\mathcal{A},\mathcal{P},\bar{\mathcal{R}},\gamma)$, here ¯$\bar{\mathcal{R}}=\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}[\mathcal{R}_{\mathcal{M}}]$.
+Theorem 4.2 (Sub-Optimality of $\Pi_{1},\Pi_{2}$. Proof in Appendix A.3).
+We set $\bar{\mathcal{M}}=(\mathcal{S},\mathcal{A},\mathcal{P},\bar{\mathcal{R}},\gamma)$, here $\bar{\mathcal{R}}=\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}[\mathcal{R}_{\mathcal{M}}]$.
 For $\forall\pi\in\Pi_{2}$, we have $\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}[J_{\mathcal{M}}(\pi)]=J_{\bar{\mathcal{M}}}(\pi)$ and further $J_{\mathcal{T}}^{1}=J_{\mathcal{T}}^{2}\leq J_{\mathcal{T}}^{*}$.
-Theorem 2 reveals that the cumulative returns of policies in $\Pi_{1}$ and $\Pi_{2}$ are the same as their returns in the “average” MDP $\bar{\mathcal{M}}$, where the reward function is the average of reward functions in different tasks.
-Also, as $\Pi_{1}$ and $\Pi_{2}$ only choose actions via current state or historical state-action pairs, they cannot distinguish different tasks and are both sub-optimal.
-To quantitatively analyze the characteristic of $\mathcal{T}$ and the gap between $J_{\mathcal{T}}^{1},J_{\mathcal{T}}^{2}$ and $J_{\mathcal{T}}^{*}$, we propose a novel metric Task Distribution Relevance (TDR) of the distribution $\mathcal{T}$ as
-Definition 1 (Task Distribution Relevance).
-For any task distribution $\mathcal{T}$ and state $s$, the Task Distribution Relevance of $\mathcal{T}$ and $s$ is defined as
-$\begin{split}D_{\text{TDR}}(\mathcal{T},s)&=\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}[\max_{a}Q_{\mathcal{M}}^{*}(s,a)]\\ &-\max_{a}\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}\left[Q_{\mathcal{M}}^{*}(s,a)\right].\end{split}$ (7)
-Intuitively, TDR describes the relevance of $\mathcal{T}$ via optimal $Q$ functions, which determine the distribution of optimal actions in corresponding tasks. Based on TDR, we can bound the gap:
-Theorem 3 (Proof in Appendix A.4).
+Theorem 4.2 reveals that the cumulative returns of policies in $\Pi_{1}$ and $\Pi_{2}$ are equivalent as their returns in the “average” MDP $\bar{\mathcal{M}}$, where the reward function is the average of reward functions in different tasks.
+Moreover, since policies in $\Pi_{1}$ and $\Pi_{2}$ only select actions based on current state or historical state-action pairs, they are unable to distinguish different tasks and are thus both sub-optimal.
+To quantitatively analyze the relationship between $\mathcal{T}$ and the gap between $J_{\mathcal{T}}^{1},J_{\mathcal{T}}^{2}$ and $J_{\mathcal{T}}^{*}$, we propose a novel metric named Task Distribution Relevance (TDR), which characterize the distribution $\mathcal{T}$ as:
+Definition 4.3 (Task Distribution Relevance).
+For any task distribution $\mathcal{T}$ and state $s$, the Task Distribution Relevance of $\mathcal{T}$ and $s$ is defined as:
+$\begin{split}D_{\text{TDR}}(\mathcal{T},s)&=\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}[\max_{a}Q_{\mathcal{M}}^{*}(s,a)]-\max_{a}\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}\left[Q_{\mathcal{M}}^{*}(s,a)\right].\end{split}$ (8)
+Intuitively, TDR quantifies the relevance of tasks in $\mathcal{T}$ via optimal $Q$ functions, which determine the choice of optimal actions in corresponding tasks.
+Based on TDR, we can derive a bound on the performance gap:
+Theorem 4.4 (Regret of Markovian Policies).
 Assume $\pi_{\mathcal{M}}^{*}=\mathop{\arg\max}_{\pi}J_{\mathcal{M}}(\pi)$, for $\forall\pi\in\Pi_{1}$, we have
-$\begin{split}J_{\mathcal{T}}^{*}-\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}\left[J_{\mathcal{M}}(\pi)\right]\geq&\frac{1}{1-\gamma}\mathbb{E}_{s\sim d_{\mathcal{M},\pi}}[D_{\text{TDR}}(\mathcal{T},s)]\end{split}$ (8)
+$\begin{split}J_{\mathcal{T}}^{*}-\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}\left[J_{\mathcal{M}}(\pi)\right]\geq&\frac{1}{1-\gamma}\mathbb{E}_{s\sim d_{\mathcal{M},\pi}}[D_{\text{TDR}}(\mathcal{T},s)].\end{split}$ (9)
 Thus $J_{\mathcal{T}}^{*}-J_{\mathcal{T}}^{2}=J_{\mathcal{T}}^{*}-J_{\mathcal{T}}^{1}\geq\frac{1}{1-\gamma}\mathbb{E}_{s\sim d_{\mathcal{M},\pi^{*}}}\left[D_{\text{TDR}}(\mathcal{T},s)\right]$, here $\pi^{*}=\arg\max_{\pi\in\Pi_{1}}J_{\bar{\mathcal{M}}}(\pi)$.
-Theorem 3 demonstrates that the gap between $J_{\mathcal{T}}^{1},J_{\mathcal{T}}^{2}$ and $J_{\mathcal{T}}^{*}$ is related to the TDR of $\mathcal{T}$.
-When considering $\mathcal{T}$ with high TDR, i.e., the optimal $Q$ values in different tasks differ greatly, the performance of $\Pi_{1}$ and $\Pi_{2}$ will be extremely poor since they cannot differentiate different tasks and their expressive abilities are significantly limited.
-This conclusion is further verified empirically in the experimental section.
-Moreover, we can show that $J_{\mathcal{T}}^{1},J_{\mathcal{T}}^{2}$ might be arbitrarily small when $J_{\mathcal{T}}^{3}$ might be arbitrarily close to $J_{\mathcal{T}}^{*}$ (More details are in Appendix A.5), which demonstrates that $\Pi_{3}$ owns stronger expressive ability.
-Consequently, it is necessary to utilize $\Pi_{3}$ to distinguish different tasks for enhancing expression ability and generalization over the task distribution.
+Theorem 4.4 demonstrates that the gap between $J_{\mathcal{T}}^{1},J_{\mathcal{T}}^{2}$ and $J_{\mathcal{T}}^{*}$ is closely related to the TDR of $\mathcal{T}$.
+For task distributions $\mathcal{T}$ with high TDR, i.e., the optimal $Q$ values vary greatly across tasks, the performance of $\Pi_{1}$ and $\Pi_{2}$ can be extremely poor since they cannot differentiate between tasks and their expressive abilities are significantly limited.
+This conclusion is further confirmed empirically in the experimental section.
+Moreover, we show that $J_{\mathcal{T}}^{1},J_{\mathcal{T}}^{2}$ can be arbitrarily small when $J_{\mathcal{T}}^{3}$ can be arbitrarily close to $J_{\mathcal{T}}^{*}$ (see Appendix A.5 for details), demonstrating that $\Pi_{3}$ possesses stronger expressive ability.
+Consequently, it is necessary to utilize policies in $\Pi_{3}$ to distinguish different tasks and thus improve generalization across the task distribution.
 Is optimizing $p_{\theta}(\mathcal{M}|h_{t},s_{t})$ helpful for task generalization?
-Now we will show that, although $p_{\theta}(\mathcal{M}|h_{t},s_{t})$ is a simple term, optimizing it can be effective for reaching the generalizable agent in the task distribution.
-As the input space of $\Pi_{3}$ is $\mathcal{L}$, i.e., all partial trajectories, our major result analyzes the relation between the policy $\pi\in\Pi_{3}$ and the task posterior $p(\mathcal{M}|l),l\in\mathcal{L}$.
-As $h_{t}$ encodes all historical information, $p_{\theta}(\mathcal{M}|h_{t},s_{t})$ can be rewritten as $p(\mathcal{M}|l),l\in\mathcal{L}$.
-Theorem 4 (Informally, detailed analyses and proof are in Appendix A.6).
-For any policy $\pi\in\Pi_{3}$, we have
-$\begin{split}&J_{\mathcal{T}}^{*}-\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}\left[J_{\mathcal{M}}(\pi)\right]\\ =&\frac{1}{1-\gamma}\int_{\mathcal{L}}p(l)\left[\int p(\mathcal{M}|l)\max_{a}Q_{\mathcal{M}}^{*}(l,a)d\mathcal{M}\right.\\ &\qquad\ -\left.\int_{a,\mathcal{M}}\pi(a|l)p(\mathcal{M}|l)Q_{\mathcal{M}}^{*}(l,a)dad\mathcal{M}\right]dl,\end{split}$ (9)
-here $p(l)$ is a distribution of $\mathcal{L}$ related to $\mathcal{T},\pi$ and $p(\mathcal{M}|l)$ is the task posterior related to $\pi$.
-Consequently, maximizing $p(\mathcal{M}|l)$, i.e., making the distribution of $p(\mathcal{M}|l)$ to be closer to some Dirac distribution, can significantly reduce the right part of Eq.
-(LABEL:eq_pi3), thus is effective for improving the generalization ability of $\pi$.
-More details and discussion are also provided in Appendix A.6.
-Algorithm Hypothesis Cartpole-balance Walker-stand&walk Cheetah-run&run__\__back
-&balance__\__sparse &prostrate&flip &flip&flip__\__back
+Now we will show that, although $p_{\theta}(\mathcal{M}|h_{t},s_{t})$ is a simple term, optimizing it can be effective for obtaining a generalizable agent in the task distribution setting.
+As the input space of $\Pi_{3}$ is $\mathcal{L}$, i.e., all historical trajectories, our main result analyzes the relationship between a policy $\pi\in\Pi_{3}$ and the task posterior $p(\mathcal{M}|l),l\in\mathcal{L}$.
+As $h_{t}$ encodes all historical information, $p_{\theta}(\mathcal{M}|h_{t},s_{t})$ can equivalently be expressed as $p(\mathcal{M}|l),l\in\mathcal{L}$.
+Theorem 4.5 (Informal; detailed analyses and proof in Appendix A.6).
+For $\forall\pi\in\Pi_{3}$, we have
+$\begin{split}J_{\mathcal{T}}^{*}-\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}\left[J_{\mathcal{M}}(\pi)\right]=&\frac{1}{1-\gamma}\int_{\mathcal{L}}p(l)\left[\int p(\mathcal{M}|l)\max_{a}Q_{\mathcal{M}}^{*}(l,a)d\mathcal{M}\right.-\left.\int_{a,\mathcal{M}}\pi(a|l)p(\mathcal{M}|l)Q_{\mathcal{M}}^{*}(l,a)dad\mathcal{M}\right]dl,\end{split}$ (10)
+where $p(l)$ is a distribution of $\mathcal{L}$ related to $\mathcal{T},\pi$ and $p(\mathcal{M}|l)$ is the task posterior related to $\pi$.
+Consequently, maximizing $p(\mathcal{M}|l)$, i.e., making the distribution of $p(\mathcal{M}|l)$ closer to a certain Dirac distribution, can significantly reduce the right part of Eq.
+(10), and thus is effective for improving the generalization ability of $\pi$.
+Additional details and discussion are provided in Appendix A.6.
+5 Experiments
+We now present empirical results to answer the following questions:
+  • Can we verify the analyses about TDR, i.e., the expressive abilities of $\Pi_{1},\Pi_{2}$ are severely restricted in task distributions with high TDR? (Sec. 5.2)
+  • How well does TAD generalize across tasks when handling image-based and state-based observations? (Sec. 5.3)
+  • Can TAD be extended to more general settings, such as dynamics generalization? (Sec. 5.4)
+5.1 Experimental setup
+Image-based control.
+To verify our analyses of TDR in Sec.
+4.3, we consider several task combinations in the DeepMind Control suite (DMC) [8]: (1) Cartpole-balance&balance_\_sparse, which shares the same optimal actions with TDR=0=0; (2) Walker-stand&walk&prostrat&flip, and Cheetah-run&run_\_back&flip&flip_\_flip_\_back, which are widely used in multi-task unsupervised RL [39, 47] and include task combinations with non-zero TDR.
+For example, Cheetah-run and Cheetah-flip hope a two-legged robot to move forward by running and flip around the torso, respectively (see Fig.
+2), yielding distinct optimal Q-functions and thus leading to huge TDR.
+Additional details about these task combinations are provided in Appendix C.1.
+To evaluate generalization with image-based observations, we extend classic tasks in DMC and design three task distributions: (1) Cheetah_\_speed($\alpha,\ \beta$), which extends Cheetah-run and hopes the agent to run within the target speed interval $(\alpha-\beta,\ \alpha+\beta)$; (2) Pendulum_\_angle($\alpha,\ \beta$), extending Pendulum-swingup to keep pendulum’s pole within the target angle interval $(\arccos\alpha,\ \arccos\beta)$; and (3) Walker_\_speed($\alpha,\ \beta$), which is based on Walker-run and requires the planar walker to run within the target speed interval $(\alpha-\beta,\ \alpha+\beta)$.
+For each task distribution, we sample $4$ training tasks and $2$ additional test tasks.
+More details are provided in Appendix C.2.
+State-based control.
+To demonstrate the scalability of TAD, we also consider several state-based continuous robotic control task distributions simulated via MuJoCo [48].
+Following previous work [27, 29], we select the following task distributions: (1) Half-Cheetah-Fwd-Back, consisting of two opposite tasks; (2) Half-Cheetah-Vel and Humanoid-Direc-2D, each comprising 100 training tasks and 30 test tasks.
+More details of these task distributions are provided in Appendix D.
+Baselines.
+In DMC, we choose two model-free methods using Markovian policies $\Pi_{1}$: CURL [49] and SAC+AE [46].
+In addition, we include two classic world models, PlaNet [4] and Dreamer [5], and two advanced contrastive learning-based methods, TACO [50] and TD-MPC [51], which utilize historical state-actions in policies and thus belong to $\Pi_{2}$.
+Moreover, we evaluate a SOTA model-based meta RL method belonging to $\Pi_{3}$: MAMBA [7].
+For state-based control, besides PlaNet and Dreamer, we include meta-RL baselines such as MAML [27], RL2 [28], and VariBAD [31], including zero-shot and few-shot evaluation as references.
+Metrics.
+For the task combinations, we evaluate the average return of all tasks to verify TDR.
+For the task generalization setting, agents are trained on the training tasks and evaluated on the test tasks to assess their generalization performance.
+All experiments are repeated with 5 different random seeds, and results are reported as the mean $\pm$ std to mitigate the effects of randomness following previous works [5, 7].
+Algorithm Hypothesis Cartpole-balance Walker-stand&walk Cheetah-run&run_\_back
+&balance_\_sparse &prostrate&flip &flip&flip_\_back
 CURL $\Pi_{1}$ 994.5 $\pm$ 3.6 254.1 $\pm$ 9.2 229.7 $\pm$ 10.9
+CURL (w/ r) $\Pi_{3}$ 987.3 $\pm$ 12.9 265.0 $\pm$ 4.7 236.1 $\pm$ 5.8
 SAC+AE $\Pi_{1}$ 992.5 $\pm$ 2.6 256.9 $\pm$ 5.9 225.8 $\pm$ 10.1
+SAC+AE (w/ r) $\Pi_{3}$ 988.5 $\pm$ 7.0 257.2 $\pm$ 8.5 239.9 $\pm$ 12.4
 PlaNet $\Pi_{2}$ 309.5 $\pm$ 59.9 606.7 $\pm$ 152.9 244.8 $\pm$ 17.8
 Dreamer $\Pi_{2}$ 974.2 $\pm$ 5.8 722.2 $\pm$ 12.6 241.1 $\pm$ 19.5
+TD-MPC $\Pi_{2}$ 614.0 $\pm$ 8.6 609.5 $\pm$ 49.6 467.4 $\pm$ 52.6
+TACO $\Pi_{2}$ 740.0 $\pm$ 69.4 540.2 $\pm$ 139.2 200.6 $\pm$ 49.8
 MAMBA $\Pi_{3}$ 994.7 $\pm$ 3.2 436.8 $\pm$ 116.1 375.6 $\pm$ 44.0
@@
 […]
-Algorithms Hypothesis Cheetah__\__speed Pendulum__\__angle Walker__\__speed
+Algorithms Hypothesis Cheetah_\_speed Pendulum_\_angle Walker_\_speed
 Train Test Train Test Train Test
@@
 […]
-5 Experiments
-We now present empirical results to answer the following questions:
-  • Can we verify the analyses about TDR, i.e., the expressive abilities of $\Pi_{1},\Pi_{2}$ are severely restricted in task distributions with high TDR? (Sec. 5.2)
-  • How about TAD’s generalization ability when handling image-based and state-based observations? (Sec. 5.3)
-  • Can TAD be extended to more general settings like dynamic generalization? (Sec. 5.4)
-5.1 Experimental Setup
-Image-based Control.
-To verify our analyses of TDR in Sec.
-4.3, we consider several task combinations in DeepMind Control suite (DMC) (Tassa et al.
-2018): (1) Cartpole-balance&balance__\__sparse, which shares the same optimal actions with TDR 0; (2) two task combinations with non-zero TDR of Walker-stand&walk&prostrat&flip, and Cheetah-run&run__\__back&flip&flip__\__flip__\__back, which are widely used in multi-task unsupervised RL (Sekar et al.
-2020; Laskin et al.
-2021).
-For example, Cheetah-run and Cheetah-flip hope a two-leg robot to move forward by running and flip around the torso, respectively (see Fig.
-3), yielding almost opposite optimal Q functions with huge TDR.
-In Appendix C.1, we introduce more details about these task combinations.
-To evaluate generalization with image-based observations, we extend tasks in DMC and design three task distributions: (1) Cheetah__\__speed($\alpha,\ \beta$), which extends Cheetah-run and hopes the agent to run within the target speed interval $(\alpha-\beta,\ \alpha+\beta)$; (2) Pendulum__\__angle($\alpha,\ \beta$), extending Pendulum-swingup to keep pendulum’s pole within the target angle interval $(\arccos\alpha,\ \arccos\beta)$; and (3) Walker__\__speed($\alpha,\ \beta$), which is based on Walker-run and requires the planar walker to run within the target speed interval $(\alpha-\beta,\ \alpha+\beta)$.
-For each task distribution, we sample $4$ training tasks and $2$ additional test tasks.
-More details are in Appendix C.2.
-State-based Control.
-To demonstrate the scalability of TAD, we also consider some state-based continuous robotic control task distributions simulated via MuJoCo (Todorov, Erez, and Tassa 2012).
-Following previous work (Finn, Abbeel, and Levine 2017), we choose tasks distributions: (1) Half-Cheetah-Fwd-Back, which owns two opposite tasks; and (2) Half-Cheetah-Vel and Humanoid-Direc-2D, which are task distributions with 100 training tasks 30 test tasks.
-More details of these task distributions are in Appendix D.
-Baselines.
-In DMC, we choose two model-free methods employing Markovian policies $\Pi_{1}$: CURL (Laskin, Srinivas, and Abbeel 2020) and SAC+AE (Yarats et al.
-2021).
-In addition, we choose two classic world models, PlaNet (Hafner et al.
-2019b) and Dreamer (Hafner et al.
-2019a), which utilize historical state-actions in policies and belong to $\Pi_{2}$.
-Moreover, we take a SOTA model-based meta RL method belonging to $\Pi_{3}$: MAMBA (Rimon et al.
-2024).
-In state-based control, besides PlaNet and Dreamer, we take some meta RL methods like MAML (Finn, Abbeel, and Levine 2017), RL2 (Duan et al.
-2016), and VariBAD (Zintgraf et al.
-2019), including zero-shot and few-shot evaluation, as reference.
-Metrics.
-For the task combinations, we evaluate the average return of all tasks to verify TDR.
-For the task generalization settings, we train agents in training tasks and evaluate their generalization abilities in test tasks.
-For all experiments, we repeat 5 different random seeds and report the mean $\pm$ std to mitigate the effects of randomness following previous works (Hafner et al.
-2019a; Rimon et al.
-2024).
+5.2 Experimental results for TDR
+To validate our analysis of TDR, we report the performance of different algorithms across different task combinations in Table 1.
+The results show that TAD-CE and TAD-SC outperform all baselines, particularly in environments with high TDR.
+As illustrated by Theorem 4.4, in task combinations with zero TDR, since different tasks share the same optimal action, policy hypothesis sets $\Pi_{1},\Pi_{2}$ are sufficient to obtain the optimal policy.
+Therefore, baselines like CURL and Dreamer also perform well in such scenarios, e.g., Cartpole-balance&balance_\_sparse.
+Conversely, in task combinations where the optimal Q-functions vary significantly across tasks, such as Cheetah-run and Cheetah-run_\_back, the TDR is significantly huge.
+In these cases, methods with policies in $\Pi_{1},\Pi_{2}$ (e.g., CURL and Dreamer) cannot differentiate between tasks and perform poorly.
+Consistent with Theorem 4.4, baselines leveraging $\Pi_{3}$ (e.g., MAMBA) outperform other baselines, and TAD further improves performance substantially.
+Moreover, we provide qualitative visualizations to further illustrate how TAD operates.
+Fig.
+2 presents future predictions generated by TAD for two tasks (Cheetah-run and Cheetah-flip).
+Using the same agent trained with TAD, we collect trajectories for both tasks and display them in the first and third rows, respectively.
+Given the first 7 steps as context, where the observations are visually similar while the rewards differ, we roll out the next 55 steps purely in imagination using the trained RIWM.
+The corresponding predicted trajectories are shown in the second and fourth rows of Figure 2.
+As observed, TAD produces distinct predictions across tasks solely based on the reward signals, indicating that the agent successfully infers task context from historical information.
+In addition, the accurate long-horizon predictions and high-quality reconstructions produced by RIWM highlight its potential as a foundation for training larger-scale world models.
+In Figure 3a, we further visualize the latent states of world models trained with Dreamer, TAD-CE, and TAD-SC.
+Trajectories from different tasks are embedded into a two-dimensional space using t-SNE [52].
+The latent representations from TAD-CE and TAD-SC form clearly separated clusters across tasks, whereas Dreamer produces overlapping embeddings and fails to distinguish task context.
+This visualization confirms that TAD effectively learns task-aware representations and differentiates between tasks at the representation level.
 Figure 2: Sampled trajectories and imaginary trajectories of TAD for different tasks (Cheetah-run and Cheetah-flip).
-Figure 4: The t-SNE clustering of state embeddings for different tasks sampled via Dreamer, TAD-CE, and TAD-SC.
-5.2 Experimental Results for TDR
-To validate our analyses of TDR, we report different algorithms’ performance of different task combinations in Table 1.
-Results show that TAD-CE and TAD-SC outperform all baselines, especially in those environments with high TDR.
-As derived in Theorem 3, in task combinations where TDR is 0 like Cartpole-balance&balance__\__sparse, different tasks share the same optimal action and $\Pi_{1},\Pi_{2}$ own the optimal policy, thus baselines like CURL and Dreamer perform well.
-Conversely, in other task combinations, different tasks’ optimal Q functions may differ a lot (like Chetah-run and Cheetah-run__\__back) and TDR is significantly huge.
-Thus methods with policies in $\Pi_{1},\Pi_{2}$ like CURL and Dreamer can not differentiate different tasks and perform poorly.
-Aligning with Theorem 3, baselines utilizing $\Pi_{3}$ like MAMBA outperform other baselines and TAD improves performance conspicuously.
-Moreover, we present some visualization to better understand how TAD works.
-Fig, 3 shows video predictions of TAD for different tasks (Cheetah-run and Cheetah-flip).
-We use the same agent trained by TAD to sample trajectories for these two tasks and show their trajectories in lines 1 and 3 respectively.
-Given the first 7 steps as the context (their observations are similar but the rewards are different), we directly imagine the future 55 steps by our trained reward-informed world model, of which the results are in lines 2 and 4 of Fig.
-3 respectively.
-As shown here, TAD performs differently to handle different tasks only by receiving different rewards.
-Moreover, our reward-informed world models demonstrate long-term prediction and high-quality reconstruction capabilities, which reveals the potential for training future large-scale world models.
-In Fig.
-4, we sample trajectories from agents trained by Dreamer, TAD-CE, as well as TAD-SC, and visualize the states of different tasks, of which the dimensions are reduced for visualization by t-SNE (Van der Maaten and Hinton 2008).
-As shown here, states in TAD-CE and TAD-SC of different tasks are clearly distinguished, while Dreamer can not differentiate them and perform the same in different tasks.
-This result demonstrates TAD can effectively learn task-aware information and distinguish different tasks.
-Videos of different tasks are provided in supplementary materials.
 5.3 Experimental results for task generalization
-To answer the second question, we report the task generalization results of image-based and state-based environments in Table 2-3.
-In Table 2, TAD shows more powerful generalization abilities than all baselines.
-This demonstrates that TAD can both handle multiple training tasks simultaneously and generalize to unseen test tasks effectively.
-We also provide some visualization results in Appendix C.3 with videos in supplementary materials.
-Additionally, in Table 3, we compare TAD with existing meta RL methods in state-based environments, of which the training timesteps are 1e7, 1e7, and 1e6, respectively.
-Many context-based meta RL methods, including RL2, PEARL, and VariBAD, utilize historical rewards and belong to $\Pi_{3}$, thus they can distinguish different tasks, which also verify our analyses.
-As shown in Table 3, TAD achieves a significant improvement for both training tasks and test tasks, since TAD is aware of task information for generalizing to unseen tasks.
+To address the second question, we report task generalization results on both image-based and state-based environments in Table 23.
+In Table 2, TAD exhibits stronger generalization capability than all baselines, indicating that it can simultaneously handle multiple training tasks and effectively generalize to unseen test tasks.
+Additional qualitative visualizations are provided in Appendix C.3.
+Furthermore, Table 3 compares TAD with existing meta-RL methods in state-based environments, where the training horizons are $1e7$, $1e7$, and $1e6$, respectively.
+Many context-based meta-RL methods, such as RL2, PEARL, and VariBAD, utilize historical rewards and thus belong to $\Pi_{3}$.
+As a result, they are capable of distinguishing different tasks, which is consistent with our theoretical analysis.
+As shown in Table 3, TAD achieves substantial performance gains on both training and test tasks, demonstrating that leveraging task-aware representations enables more effective generalization to unseen tasks.
 5.4 Ablation study
@@
 In Fig.
-5, we do ablation studies about the task term in TAD, i.e., we consider Dreamer(w/ r) that only integrates $\Pi_{3}$ into Dreamer.
-Results show that just providing rewards can help Dreamer distinguish different tasks, which verifies our analyses in TDR.
-Also, TAD-CE and TAD-SC show superior performance, demonstrating that our task model and corresponding ELBO are effective for task generalization.
-Also, we do ablation studies on reward signals for model-free methods like CURL in Appendix E.1.
-Figure 5: Ablation study on Reward Signal.
-Extension to Dynamic Generalization.
-To answer the third question, we evaluate TAD in more general settings with different observations, dynamics, and/or actions.
-As TAD utilizes all historical information to infer the environment, it can be directly applied to these settings.
-We design task distribution with different embodiments (Acrobot-Cartpole-Pendulum, Walker-Cheetah-Hopper), as well as different dynamics (Cheetah-run__\__mass, Walker-walk__\__mass).
-More details about environments and results are in Appendix E.2, where TAD achieves much greater performance compared to baselines and show potential in further handling dynamic generalization and even cross-embodiment tasks.
+3b, we present ablation studies on the task-context term in TAD.
+In particular, we evaluate Dreamer(w/ r), which augments Dreamer to $\Pi_{3}$ by incorporating reward conditioning.
+The results show that simply providing reward signals can help Dreamer distinguish different tasks to some degree, which is consistent with our analysis in Sec.
+4.3.
+Nevertheless, TAD-CE and TAD-SC achieve notably stronger performance, indicating that the proposed RIWM and the corresponding optimization method further enhance task generalization ability.
+We also conduct ablations on reward signals for model-free methods such as CURL and SAC+AE, both of which belong to $\Pi_{1}$.
+Specifically, we design CURL (w/ r) and SAC+AE (w/ r) by directly appending the reward to the observation at each timestep.
+As shown in Table 1, their performance remains comparable to the original CURL and SAC+AE implementation.
+This is because these variants only use the reward and observation from the current timestep, without leveraging historical information, and thus still fail to effectively distinguish between different tasks.
+(a) T-SNE visualization.
+(b) Ablation study on Reward Signal.
+Figure 3: (a) The t-SNE clustering of state embeddings for different tasks via Dreamer, TAD-CE, and TAD-SC. (b) The ablation study on the reward signals.
+Extension to dynamics generalization.
+To address the third question, we evaluate TAD in more general settings involving variations in observations, dynamics, and/or action spaces.
+Since TAD infers task information with all historical information, it can be directly applied to these settings without modification.
+We construct task distributions with different embodiments (Acrobot-Cartpole-Pendulum, Walker-Cheetah-Hopper), as well as different dynamics (Cheetah-run_\_mass, Walker-walk_\_mass).
+More details and results are provided in Appendix E.2, where TAD demonstrates substantially stronger performance than the baselines, highlighting its potential for dynamics generalization and even cross-embodiment adaptation.
 5.5 Limitations and discussion
-In terms of limitations, TAD assumes that the task context is continuously related to historical information for generalizing to unseen tasks.
-Thus TAD might be difficult to generalize in sparse-reward settings.
-We further demonstrate that without extra knowledge or finetuning, zero-shot generalization to unseen tasks with extremely sparse rewards is impossible since there is no way to distinguish different tasks (Appendix A.7).
-Fortunately, in relatively sparse reward settings, we conduct experiments in Appendix E.3 to show that TAD can infer the current task and generalize to unseen tasks well.
+Regarding limitations, TAD assumes that task context can be inferred from historical interaction data in order to generalize to unseen tasks.
+Consequently, its effectiveness may be reduced in environments with extremely sparse rewards.
+In Appendix A.7, we further show that without extra prior knowledge or fine-tuning, zero-shot generalization to unseen tasks with extremely sparse rewards is fundamentally infeasible, as the agent lacks sufficient information to distinguish among tasks.
+Nevertheless, in relatively sparse-reward settings, our experiments in Appendix E.3 demonstrate that TAD can still infer the current task and generalize effectively to unseen tasks.
 6 Conclusion
-In this work, we propose a novel framework of TAD that handles different tasks via all historical information and utilizes novel reward-informed world models to capture invariant latent features.
-In TAD, we calculate the corresponding variational lower bound of the data log-likelihood, which includes a novel loss term to distinguish different tasks via states.
-To explain components in TAD, we introduce a novel metric TDR to capture the relevance of the task distribution and show that Markovian policies perform poorly in tasks with high TDR.
-Experiments in image-based and state-based settings demonstrate that TAD can remarkably improve the performance of handling different tasks meanwhile, especially for high TDR ones, and successfully generalize to unseen tasks.
+In this work, we propose Task-Aware Dreamer (TAD), a novel framework that leverages full historical information to handle multiple tasks and employs Reward-Informed World Models (RIWM) to capture task-specific latent features.
+In TAD, we derive a variational lower bound on the data log-likelihood that incorporates a task context term, encouraging the world model to learn task-specific representations.
+To justify the necessity of TAD’s components, we first introduce the Task Distribution Relevance (TDR) metric to characterize the relevance of different tasks from the task distribution.
+We then theoretically show that Markovian policies perform poorly in task distributions with high TDR.
+Extensive experiments in both image-based and state-based settings demonstrate that TAD substantially improves performance in the task distribution setting, particularly for task distributions with high TDR, and successfully generalizes to unseen tasks in a zero-shot manner.
+[…]
+This work was supported by the NSFC Projects (Nos.
+92370124, 92248303, U25B6003, 62350080), Beijing Natural Science Foundation L247011, Tsinghua Institute for Guo Qiang, and the High Performance Computing Center, Tsinghua University.
 Appendix A Proof of theorems
@@
 A.1 The proof of Theorem 1
-Proof.
-First, we will show that $\mathcal{H}_{2}\subseteq\mathcal{H}_{1}$, following the proof in (Young et al. 2023) that considers the setting that the rewards of all environments are the same.
-$\begin{split}&\{q_{m}\}\in\mathcal{H}_{2}\\ \Leftrightarrow&\exists(p_{m})_{m=1}^{M}:p_{m}(o_{t}^{m},a_{t}^{m})=o_{t+1}^{m},\\ &\exists(r_{m})_{m=1}^{M}:r_{m}(o_{t}^{m},a_{t}^{m})=r_{t}^{m},\\ &q_{m}(o,a)=r_{m}(o,a)+\gamma\max_{a^{\prime}}q_{m}(p_{m}(o,a),a^{\prime}),\forall o,a\\ \Rightarrow&\exists(r_{m})_{m=1}^{M}::p_{m}(o_{t}^{m},a_{t}^{m})=o_{t+1}^{m},\\ &\exists(r_{m})_{m=1}^{M}:r_{m}(o_{t}^{m},a_{t}^{m})=r_{t}^{m},\\ &q_{m}(o_{t}^{m},a_{t}^{m})=r^{m}_{t}+\gamma\max_{a^{\prime}}q_{m}(p_{m}(o_{t}^{m},a_{t}^{m}),a^{\prime}),\forall m,t\\ \Rightarrow&q_{m}(o_{t}^{m},a_{t}^{m})=r^{m}_{t}+\gamma\max_{a^{\prime}}q_{m}(o_{t+1}^{m},a^{\prime}),\forall m,t\\ \Leftrightarrow&\{q_{m}\}\in\mathcal{H}_{1},\end{split}$ (10)
+Proof A.1.
+First, we will show that $\mathcal{H}_{2}\subseteq\mathcal{H}_{1}$, following the proof in [39] that considers the setting where the rewards of all environments are the same.
+$\begin{split}\{q_{m}\}\in\mathcal{H}_{2}\Leftrightarrow&\exists(p_{m})_{m=1}^{M}:p_{m}(o_{t}^{m},a_{t}^{m})=o_{t+1}^{m},\\ &\exists(r_{m})_{m=1}^{M}:r_{m}(o_{t}^{m},a_{t}^{m})=r_{t}^{m},q_{m}(o,a)=r_{m}(o,a)+\gamma\max_{a^{\prime}}q_{m}(p_{m}(o,a),a^{\prime}),\forall o,a\\ \Rightarrow&\exists(r_{m})_{m=1}^{M}::p_{m}(o_{t}^{m},a_{t}^{m})=o_{t+1}^{m},\\ &\exists(r_{m})_{m=1}^{M}:r_{m}(o_{t}^{m},a_{t}^{m})=r_{t}^{m},q_{m}(o_{t}^{m},a_{t}^{m})=r^{m}_{t}+\gamma\max_{a^{\prime}}q_{m}(p_{m}(o_{t}^{m},a_{t}^{m}),a^{\prime}),\forall m,t\\ \Rightarrow&q_{m}(o_{t}^{m},a_{t}^{m})=r^{m}_{t}+\gamma\max_{a^{\prime}}q_{m}(o_{t+1}^{m},a^{\prime}),\forall m,t\\ \Leftrightarrow&\{q_{m}\}\in\mathcal{H}_{1},\end{split}$ (11)
 thus we have $\mathcal{H}_{2}\subseteq\mathcal{H}_{1}$.
 Next, we prove that $\mathcal{H}_{3}\subseteq\mathcal{H}_{2}$, which is mainly because we can utilize similar dynamic structures from different tasks to narrow down the hypothesis spaces of the dynamic model.
-$\begin{split}&\{q_{m}\}\in\mathcal{H}_{3}\\ \Leftrightarrow&\exists p:p(o_{t}^{m},a_{t}^{m})=o_{t+1}^{m},\forall m,t\\ &\exists(r_{m})_{m=1}^{M}:r_{m}(o_{t}^{m},a_{t}^{m})=r_{t}^{m},\\ &q_{m}(o,a)=r_{m}(o,a)+\gamma\max_{a^{\prime}}q_{m}(p(o,a),a^{\prime}),\forall m,o,a\\ \Rightarrow&\exists(p_{m})_{m=1}^{M}:p_{m}(o_{t}^{m},a_{t}^{m})=o_{t+1}^{m},\forall m,t\\ &\exists(r_{m})_{m=1}^{M}:r_{m}(o_{t}^{m},a_{t}^{m})=r_{t}^{m},\\ &q_{m}(o,a)=r_{m}(o,a)+\gamma\max_{a^{\prime}}q_{m}(p_{m}(o,a),a^{\prime}),\forall m,o,a\\ \Leftrightarrow&\{q_{m}\}\in\mathcal{H}_{2},\end{split}$ (11)
-thus we have $\mathcal{H}_{3}\subseteq\mathcal{H}_{2}$. ∎
+$\begin{split}\{q_{m}\}\in\mathcal{H}_{3}\Leftrightarrow&\exists p:p(o_{t}^{m},a_{t}^{m})=o_{t+1}^{m},\forall m,t\\ &\exists(r_{m})_{m=1}^{M}:r_{m}(o_{t}^{m},a_{t}^{m})=r_{t}^{m},q_{m}(o,a)=r_{m}(o,a)+\gamma\max_{a^{\prime}}q_{m}(p(o,a),a^{\prime}),\forall m,o,a\\ \Rightarrow&\exists(p_{m})_{m=1}^{M}:p_{m}(o_{t}^{m},a_{t}^{m})=o_{t+1}^{m},\forall m,t\\ &\exists(r_{m})_{m=1}^{M}:r_{m}(o_{t}^{m},a_{t}^{m})=r_{t}^{m},q_{m}(o,a)=r_{m}(o,a)+\gamma\max_{a^{\prime}}q_{m}(p_{m}(o,a),a^{\prime}),\forall m,o,a\\ \Leftrightarrow&\{q_{m}\}\in\mathcal{H}_{2},\end{split}$ (12)
+thus we have $\mathcal{H}_{3}\subseteq\mathcal{H}_{2}$.
 A.2 The derivation of the ELBO
-We use ${\color[rgb]{0,139,69}q}$ to represent $q(s_{1:T}|o_{1:T},a_{1:T},r_{1:T})$, ${\color[rgb]{0,0,1}\hat{q}}$ to present $q(s_{t}|o_{\leq t},a_{<t},r_{<t})$, ${\color[rgb]{1,0,0}\tilde{q}}$ to represent $q(s_{t-1}|o_{\leq(t-1)},r_{<(t-1)},a_{<(t-1)})$, and we have
-$\begin{split}&\ln p(o_{1:T},r_{1:T},\mathcal{M}|a_{1:T})\\ =&\ln\mathbb{E}_{p(s_{1:T}|a_{1:T})}\left[p(o_{1:T},r_{1:T},\mathcal{M}|s_{1:T})\right]\\ =&\ln\mathbb{E}_{p(s_{1:T}|a_{1:T})}\left[p(o_{1:T},r_{1:T}|\mathcal{M},s_{1:T})p(\mathcal{M}|s_{1:T})\right]\\ =&\ln\mathbb{E}_{p(s_{1:T}|a_{1:T})}\left[p(\mathcal{M}|s_{1:T})\prod_{t=1}^{T}p(o_{t},r_{t}|\mathcal{M},s_{t})\right]\\ =&\ln\mathbb{E}_{{\color[rgb]{0,139,69}q}}\left[p(\mathcal{M}|s_{1:T})\prod_{t=1}^{T}p(o_{t},r_{t}|\mathcal{M},s_{t})\frac{p(s_{t}|s_{t-1},a_{t-1})}{q(s_{t}|o_{\leq t},r_{<t},a_{<t})}\right]\\ \geq&\mathbb{E}_{{\color[rgb]{0,139,69}q}}[\ln p(\mathcal{M}|s_{1:T})+\sum_{t=1}^{T}\ln p(o_{t},r_{t}|\mathcal{M},s_{t})]\\ +&\mathbb{E}_{{\color[rgb]{0,139,69}q}}\sum_{t=1}^{T}\left[\ln p(s_{t}|s_{t-1},a_{t-1})-\ln q(s_{t}|o_{\leq t},r_{<t},a_{<t})\right]\\ =&\mathbb{E}_{{\color[rgb]{0,139,69}q}}[\ln p(\mathcal{M}|s_{1:T})]\\ +&\sum_{t=1}^{T}\left[\mathbb{E}_{{\color[rgb]{0,139,69}q}}\ln p(o_{t},r_{t}|\mathcal{M},s_{t})-\mathbb{E}_{{\color[rgb]{0,139,69}q}}\ln\frac{q(s_{t}|o_{\leq t},r_{<t},a_{<t})}{p(s_{t}|s_{t-1},a_{t-1})}\right]\\ =&\mathbb{E}_{{\color[rgb]{0,139,69}q}}[\ln p(\mathcal{M}|s_{1:T})]+\sum_{t=1}^{T}\mathbb{E}_{{\color[rgb]{0,0,1}\hat{q}}}[\ln p(o_{t},r_{t}|\mathcal{M},s_{t})]\\ -&\sum_{t=1}^{T}\mathbb{E}_{{\color[rgb]{0,0,1}\hat{q}}{\color[rgb]{1,0,0}\tilde{q}}}\left[\ln\frac{q(s_{t}|o_{\leq t},r_{<t},a_{<t})}{p(s_{t}|s_{t-1},a_{t-1})}\right]\\ =&\mathbb{E}_{{\color[rgb]{0,139,69}q}}[\ln p(\mathcal{M}|s_{1:T})]+\sum_{t=1}^{T}\mathbb{E}_{{\color[rgb]{0,0,1}\hat{q}}}[\ln p(o_{t},r_{t}|\mathcal{M},s_{t})]\\ -&\sum_{t=1}^{T}\mathbb{E}_{{\color[rgb]{1,0,0}\tilde{q}}}[\mathrm{KL}\left(q(s_{t}|o_{\leq t},r_{<t},a_{<t})\|p(s_{t}|s_{t-1},a_{t-1})\right)].\end{split}$ (12)
-Thus we have proven the ELBO.
+We use $\color[rgb]{0,139,69}q$} to represent $(s_{1:T}|o_{1:T},a_{1:T},r_{1:T}), q^{\hat{q}}$ to present $(s_{t}|o_{\leq t},a_{<t},r_{<t}), q~{\color[rgb]{1,0,0}\tilde{q}}$ to represent $q(s_{t-1}|o_{\leq(t-1)},r_{<(t-1)},a_{<(t-1)})$, and we have
+$\begin{split}&\ln p(o_{1:T},r_{1:T},\mathcal{M}|a_{1:T})\\ =&\ln\mathbb{E}_{p(s_{1:T}|a_{1:T})}\left[p(o_{1:T},r_{1:T},\mathcal{M}|s_{1:T})\right]\\ =&\ln\mathbb{E}_{p(s_{1:T}|a_{1:T})}\left[p(o_{1:T},r_{1:T}|\mathcal{M},s_{1:T})p(\mathcal{M}|s_{1:T})\right]\\ =&\ln\mathbb{E}_{p(s_{1:T}|a_{1:T})}\left[p(\mathcal{M}|s_{1:T})\prod_{t=1}^{T}p(o_{t},r_{t}|\mathcal{M},s_{t})\right]\\ =&\ln\mathbb{E}_{{\color[rgb]{0,139,69}q}}\left[p(\mathcal{M}|s_{1:T})\prod_{t=1}^{T}p(o_{t},r_{t}|\mathcal{M},s_{t})\frac{p(s_{t}|s_{t-1},a_{t-1})}{q(s_{t}|o_{\leq t},r_{<t},a_{<t})}\right]\\ \geq&\mathbb{E}_{{\color[rgb]{0,139,69}q}}[\ln p(\mathcal{M}|s_{1:T})+\sum_{t=1}^{T}\ln p(o_{t},r_{t}|\mathcal{M},s_{t})]+\mathbb{E}_{{\color[rgb]{0,139,69}q}}\sum_{t=1}^{T}\left[\ln p(s_{t}|s_{t-1},a_{t-1})-\ln q(s_{t}|o_{\leq t},r_{<t},a_{<t})\right]\\ =&\mathbb{E}_{{\color[rgb]{0,139,69}q}}[\ln p(\mathcal{M}|s_{1:T})]+\sum_{t=1}^{T}\left[\mathbb{E}_{{\color[rgb]{0,139,69}q}}\ln p(o_{t},r_{t}|\mathcal{M},s_{t})-\mathbb{E}_{{\color[rgb]{0,139,69}q}}\ln\frac{q(s_{t}|o_{\leq t},r_{<t},a_{<t})}{p(s_{t}|s_{t-1},a_{t-1})}\right]\\ =&\mathbb{E}_{{\color[rgb]{0,139,69}q}}[\ln p(\mathcal{M}|s_{1:T})]+\sum_{t=1}^{T}\mathbb{E}_{{\hat{q}}}[\ln p(o_{t},r_{t}|\mathcal{M},s_{t})]-\sum_{t=1}^{T}\mathbb{E}_{{\hat{q}}{\color[rgb]{1,0,0}\tilde{q}}}\left[\ln\frac{q(s_{t}|o_{\leq t},r_{<t},a_{<t})}{p(s_{t}|s_{t-1},a_{t-1})}\right]\\ =&\mathbb{E}_{{\color[rgb]{0,139,69}q}}[\ln p(\mathcal{M}|s_{1:T})]+\sum_{t=1}^{T}\mathbb{E}_{{\hat{q}}}[\ln p(o_{t},r_{t}|\mathcal{M},s_{t})]-\sum_{t=1}^{T}\mathbb{E}_{{\color[rgb]{1,0,0}\tilde{q}}}[\mathrm{KL}\left(q(s_{t}|o_{\leq t},r_{<t},a_{<t})\|p(s_{t}|s_{t-1},a_{t-1})\right)].\end{split}$ (13)
+Thus, we have proven the ELBO.
 A.3 The proof of Theorem 2
-Proof.
+Proof A.2.
 We first prove that for $\forall\pi\in\Pi_{2}$, we have $\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}[J_{\mathcal{M}}(\pi)]=J_{\bar{\mathcal{M}}}(\pi)$.
-For any $\mathcal{M}=(\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R}_{\mathcal{M}},\gamma)\sim\mathcal{T}$, we can use the policy $\pi$ to interact with $\mathcal{M}$ and get the trajectory $\tau=(s_{0}^{\mathcal{M}},a_{0}^{\mathcal{M}},r_{1}^{\mathcal{M}},s_{1}^{\mathcal{M}},a_{1}^{\mathcal{M}},r_{2}^{\mathcal{M}},...)$.
-Since the dynamic transition $\mathcal{P}$ is the same for all $\mathcal{M}$ and the policy $\pi\in\Pi_{2}$ only depends on historical states and actions, we naturally have that the distribution of all states and actions $(s_{0}^{\mathcal{M}},a_{0}^{\mathcal{M}},s_{1}^{\mathcal{M}},a_{1}^{\mathcal{M}},...)$ are the same for all $\mathcal{M}\sim\mathcal{T}$ as well as $\bar{\mathcal{M}}$.
-Consequently, we have
-$\begin{split}&\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}[J_{\mathcal{M}}(\pi)]\\ =&\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}\mathbb{E}_{\tau\sim\mathcal{P},\pi}[R_{\mathcal{M}}(\tau)]=\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}\mathbb{E}_{\tau\sim\mathcal{P},\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}r_{t}^{\mathcal{M}}\right]\\ =&\mathbb{E}_{\tau\sim\mathcal{P},\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}[r_{t}^{\mathcal{M}}]\right]\\ =&\mathbb{E}_{\tau\sim\mathcal{P},\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}[\mathcal{R}_{\mathcal{M}}(s_{t}^{\mathcal{M}},a_{t}^{\mathcal{M}})]\right]\\ =&\mathbb{E}_{\tau\sim\mathcal{P},\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}[\bar{\mathcal{R}}(s_{t}^{\mathcal{M}},a_{t}^{\mathcal{M}})]\right]\\ =&\mathbb{E}_{\tau\sim\mathcal{P},\pi}[R_{\bar{\mathcal{M}}}(\tau)]=J_{\bar{\mathcal{M}}}(\pi).\end{split}$ (13)
+[…]
+$\begin{split}\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}[J_{\mathcal{M}}(\pi)]=&\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}\mathbb{E}_{\tau\sim\mathcal{P},\pi}[R_{\mathcal{M}}(\tau)]=\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}\mathbb{E}_{\tau\sim\mathcal{P},\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}r_{t}^{\mathcal{M}}\right]=\mathbb{E}_{\tau\sim\mathcal{P},\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}[r_{t}^{\mathcal{M}}]\right]\\ =&\mathbb{E}_{\tau\sim\mathcal{P},\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}[\mathcal{R}_{\mathcal{M}}(s_{t}^{\mathcal{M}},a_{t}^{\mathcal{M}})]\right]=\mathbb{E}_{\tau\sim\mathcal{P},\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}[\bar{\mathcal{R}}(s_{t}^{\mathcal{M}},a_{t}^{\mathcal{M}})]\right]\\ =&\mathbb{E}_{\tau\sim\mathcal{P},\pi}[R_{\bar{\mathcal{M}}}(\tau)]=J_{\bar{\mathcal{M}}}(\pi).\end{split}$ (14)
 It is well known that the optimal policy in single MDP is memory-less, i.e., $\max_{\pi\in\Pi_{2}}J_{\bar{\mathcal{M}}}(\pi)=\max_{\pi\in\Pi_{1}}J_{\bar{\mathcal{M}}}(\pi)$. Consequently, we have
-$\begin{split}J_{\mathcal{T}}^{2}&=\max_{\pi\in\Pi_{2}}\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}[J_{\mathcal{M}}(\pi)]=\max_{\pi\in\Pi_{2}}J_{\bar{\mathcal{M}}}(\pi)\\ &=\max_{\pi\in\Pi_{1}}J_{\bar{\mathcal{M}}}(\pi)=\max_{\pi\in\Pi_{1}}\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}[J_{\mathcal{M}}(\pi)]=J_{\mathcal{T}}^{1}\\ &\leq\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}\left[\max_{\pi\in\Pi_{1}}J_{\mathcal{M}}(\pi)\right]=J_{\mathcal{T}}^{*}.\end{split}$ (14)
-Thus we have proven this result. ∎
+$\begin{split}J_{\mathcal{T}}^{2}&=\max_{\pi\in\Pi_{2}}\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}[J_{\mathcal{M}}(\pi)]=\max_{\pi\in\Pi_{2}}J_{\bar{\mathcal{M}}}(\pi)=\max_{\pi\in\Pi_{1}}J_{\bar{\mathcal{M}}}(\pi)=\max_{\pi\in\Pi_{1}}\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}[J_{\mathcal{M}}(\pi)]=J_{\mathcal{T}}^{1}\\ &\leq\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}\left[\max_{\pi\in\Pi_{1}}J_{\mathcal{M}}(\pi)\right]=J_{\mathcal{T}}^{*}.\end{split}$ (15)
+Thus we have proven this result.
 A.4 The Proof of Theorem 3
-Proof.
-Our proof follows some previous work (Kakade and Langford 2002; Ying et al.
-2022).
-First, we consider the bellman equation of value function of $\pi,\pi_{\mathcal{M}}^{*}\in\Pi_{1}$ in $\mathcal{M}$ as
+Proof A.3.
+Our proof follows some previous work [53, 54]. First, we consider the bellman equation of the value function of $\pi,\pi_{\mathcal{M}}^{*}\in\Pi_{1}$ in $\mathcal{M}$ as
 $\begin{split}V_{\mathcal{M},\pi}(s)=&\sum_{a}\pi(a|s)\left[\mathcal{R}(s,a)+\gamma\sum_{s^{\prime}}\mathcal{P}(s^{\prime}|s,a)V_{\mathcal{M},\pi}(s^{\prime})\right],\\ V_{\mathcal{M},\pi_{\mathcal{M}}^{*}}(s)=&\sum_{a}\pi_{\mathcal{M}}^{*}(a|s)\left[\mathcal{R}(s,a)+\gamma\sum_{s^{\prime}}\mathcal{P}(s^{\prime}|s,a)V_{\mathcal{M},\pi_{\mathcal{M}}^{*}}(s^{\prime})\right].\end{split}$
 Defining $\Delta V(s)\triangleq V_{\mathcal{M},\pi}(s)-V_{\mathcal{M},\pi_{\mathcal{M}}^{*}}(s)$ as the difference of these two value functions, we can further deduce that
-$\begin{split}&V_{\mathcal{M},\pi}(s)-V_{\mathcal{M},\pi_{\mathcal{M}}^{*}}(s)\\ =&\gamma\sum_{a}\Delta\pi(a|s)\sum_{s^{\prime}}\mathcal{P}(s^{\prime}|s,a)V_{\mathcal{M},\pi_{\mathcal{M}}^{*}}(s^{\prime})\\ +&\gamma\sum_{a}\pi(a|s)\sum_{s^{\prime}}\mathcal{P}(s^{\prime}|s,a)\Delta V(s^{\prime})+\sum_{a}\Delta\pi(a|s)\mathcal{R}(s,a)\\ =&\sum_{a}\Delta\pi(a|s)Q_{\mathcal{M},\pi_{\mathcal{M}}^{*}}(s,a)\\ +&\gamma\sum_{a}\pi(a|s)\sum_{s^{\prime}}\mathcal{P}(s^{\prime}|s,a)\Delta V(s^{\prime}),\end{split}$ (15)
+$\begin{split}&V_{\mathcal{M},\pi}(s)-V_{\mathcal{M},\pi_{\mathcal{M}}^{*}}(s)\\ =&\gamma\sum_{a}\Delta\pi(a|s)\sum_{s^{\prime}}\mathcal{P}(s^{\prime}|s,a)V_{\mathcal{M},\pi_{\mathcal{M}}^{*}}(s^{\prime})+\gamma\sum_{a}\pi(a|s)\sum_{s^{\prime}}\mathcal{P}(s^{\prime}|s,a)\Delta V(s^{\prime})+\sum_{a}\Delta\pi(a|s)\mathcal{R}(s,a)\\ =&\sum_{a}\Delta\pi(a|s)Q_{\mathcal{M},\pi_{\mathcal{M}}^{*}}(s,a)+\gamma\sum_{a}\pi(a|s)\sum_{s^{\prime}}\mathcal{P}(s^{\prime}|s,a)\Delta V(s^{\prime}),\end{split}$ (16)
 here $\Delta\pi(a|s)=\pi(a|s)-\pi_{\mathcal{M}}^{*}(a|s)$. Since Eq. (LABEL:15) holds for any $s$, thus we calculate its expectation for $s\sim d_{\mathcal{M}}^{\pi_{\mathcal{M}}^{*}}$:
-$\begin{split}&\sum_{s}d_{\mathcal{M}}^{\pi}(s)\Delta V(s)\\ =&\sum_{s}d_{\mathcal{M}}^{\pi}(s)[V_{\mathcal{M},\pi}(s)-V_{\mathcal{M},\pi_{\mathcal{M}}^{*}}(s)]\\ =&\sum_{s}d_{\mathcal{M}}^{\pi}(s)\sum_{a}\Delta\pi(a|s)Q_{\mathcal{M},\pi_{\mathcal{M}}^{*}}(s,a)\\ +&\gamma\sum_{s}d_{\mathcal{M}}^{\pi}(s)\sum_{a}\pi(a|s)\sum_{s^{\prime}}\mathcal{P}(s^{\prime}|s,a)\Delta V(s^{\prime})\\ =&\sum_{s}d_{\mathcal{M}}^{\pi}(s)\sum_{a}\Delta\pi(a|s)Q_{\mathcal{M},\pi_{\mathcal{M}}^{*}}(s,a)\\ +&\sum_{s^{\prime}}\Delta V(s^{\prime})\left[\gamma\sum_{s}d_{\mathcal{M}}^{\pi}(s)\sum_{a}\pi(a|s)\mathcal{P}(s^{\prime}|s,a)\right].\end{split}$ (16)
+$\begin{split}\sum_{s}d_{\mathcal{M}}^{\pi}(s)\Delta V(s)=&\sum_{s}d_{\mathcal{M}}^{\pi}(s)[V_{\mathcal{M},\pi}(s)-V_{\mathcal{M},\pi_{\mathcal{M}}^{*}}(s)]\\ =&\sum_{s}d_{\mathcal{M}}^{\pi}(s)\sum_{a}\Delta\pi(a|s)Q_{\mathcal{M},\pi_{\mathcal{M}}^{*}}(s,a)+\gamma\sum_{s}d_{\mathcal{M}}^{\pi}(s)\sum_{a}\pi(a|s)\sum_{s^{\prime}}\mathcal{P}(s^{\prime}|s,a)\Delta V(s^{\prime})\\ =&\sum_{s}d_{\mathcal{M}}^{\pi}(s)\sum_{a}\Delta\pi(a|s)Q_{\mathcal{M},\pi_{\mathcal{M}}^{*}}(s,a)+\sum_{s^{\prime}}\Delta V(s^{\prime})\left[\gamma\sum_{s}d_{\mathcal{M}}^{\pi}(s)\sum_{a}\pi(a|s)\mathcal{P}(s^{\prime}|s,a)\right].\end{split}$ (17)
 Since $\gamma\sum_{s^{\prime}}d_{\mathcal{M}}^{\pi}(s^{\prime})\sum_{a}\pi(a|s^{\prime})\mathcal{P}(s|s^{\prime},a)=d_{\mathcal{M}}^{\pi}(s)-(1-\gamma)\mathcal{P}(s_{0}=s)$, we have
-$\begin{split}\sum_{s}d_{\mathcal{M}}^{\pi}(s)&\Delta V(s)=\sum_{s}d_{\mathcal{M}}^{\pi}(s)\sum_{a}\Delta\pi(a|s)Q_{\mathcal{M},\pi_{\mathcal{M}}^{*}}(s,a)\\ +&\sum_{s^{\prime}}\Delta V(s^{\prime})\left[d_{\mathcal{M}}^{\pi}(s^{\prime})-(1-\gamma)\mathcal{P}(s_{0}=s^{\prime})\right].\end{split}$ (17)
+$\begin{split}\sum_{s}d_{\mathcal{M}}^{\pi}(s)&\Delta V(s)=\sum_{s}d_{\mathcal{M}}^{\pi}(s)\sum_{a}\Delta\pi(a|s)Q_{\mathcal{M},\pi_{\mathcal{M}}^{*}}(s,a)+\sum_{s^{\prime}}\Delta V(s^{\prime})\left[d_{\mathcal{M}}^{\pi}(s^{\prime})-(1-\gamma)\mathcal{P}(s_{0}=s^{\prime})\right].\end{split}$ (18)
 By moving the second term of the right part in Eq. (18) to the left part, we can deduce that
-$\begin{split}&(1-\gamma)\sum_{s^{\prime}}\Delta V(s^{\prime})\mathcal{P}(s_{0}=s^{\prime})\\ =&\sum_{s}d_{\mathcal{M}}^{\pi}(s)\sum_{a}\Delta\pi(a|s)Q_{\mathcal{M},\pi_{\mathcal{M}}^{*}}(s,a),\end{split}$ (18)
+$\begin{split}(1-\gamma)\sum_{s^{\prime}}\Delta V(s^{\prime})\mathcal{P}(s_{0}=s^{\prime})=&\sum_{s}d_{\mathcal{M}}^{\pi}(s)\sum_{a}\Delta\pi(a|s)Q_{\mathcal{M},\pi_{\mathcal{M}}^{*}}(s,a),\end{split}$ (19)
 thus we can calculate that
-[…]
+$\begin{split}J_{\mathcal{M}}(\pi)-J_{\mathcal{M}}(\pi_{\mathcal{M}}^{*})=&\sum_{s^{\prime}}\Delta V(s^{\prime})\mathcal{P}(s_{0}=s^{\prime})=\frac{1}{1-\gamma}\sum_{s}d_{\mathcal{M}}^{\pi}(s)\sum_{a}\Delta\pi(a|s)Q_{\mathcal{M},\pi_{\mathcal{M}}^{*}}(s,a)\\ =&\frac{1}{1-\gamma}\sum_{s}d_{\mathcal{M}}^{\pi}(s)\sum_{a}[\pi(a|s)-\pi_{\mathcal{M}}^{*}(a|s)]Q_{\mathcal{M},\pi_{\mathcal{M}}^{*}}(s,a)\\ =&\frac{1}{1-\gamma}\mathbb{E}_{s\sim d_{\mathcal{M}}^{\pi}}\mathbb{E}_{a\sim\pi(\cdot|s)}\left(1-\frac{\pi_{\mathcal{M}}^{*}(a|s)}{\pi(a|s)}\right)Q_{\mathcal{M},\pi_{\mathcal{M}}^{*}}(s,a)\\ =&\frac{1}{1-\gamma}\mathbb{E}_{s\sim d_{\mathcal{M}}^{\pi}}\int_{\mathcal{A}}\pi(a|s)\left(1-\frac{\pi_{\mathcal{M}}^{*}(a|s)}{\pi(a|s)}\right)Q_{\mathcal{M}}^{*}(s,a)\mathrm{d}a\\ =&\frac{1}{1-\gamma}\mathbb{E}_{s\sim d_{\mathcal{M}}^{\pi}}\left[\int_{a}\pi(a|s)Q_{\mathcal{M}}^{*}(s,a)da-\max_{a}Q_{\mathcal{M}}^{*}(s,a)\right].\end{split}$ (20)
 Consequently, we have
-$\begin{split}&\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}\left[J_{\mathcal{M}}(\pi_{\mathcal{M}}^{*})-J_{\mathcal{M}}(\pi)\right]\\ =&\frac{1}{1-\gamma}\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}\mathbb{E}_{s\sim d_{\mathcal{M},\pi}(\cdot)}\left[\max_{a}Q_{\mathcal{M}}^{*}(s,a)\right.\\ &\left.-\int_{a}\pi(a|s)Q_{\mathcal{M}}^{*}(s,a)da\right]\\ \geq&\frac{1}{1-\gamma}\mathbb{E}_{s\sim d_{\mathcal{M},\pi}(\cdot)}\left[\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}\max_{a}Q_{\mathcal{M}}^{*}(s,a)\right.\\ &\left.-\max_{a}\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}Q_{\mathcal{M}}^{*}(s,a)\right]\\ =&\frac{1}{1-\gamma}\mathbb{E}_{s\sim d_{\mathcal{M},\pi}}\left[D_{\text{TDR}}(\mathcal{T},s)\right].\end{split}$ (20)
+$\begin{split}\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}\left[J_{\mathcal{M}}(\pi_{\mathcal{M}}^{*})-J_{\mathcal{M}}(\pi)\right]=&\frac{1}{1-\gamma}\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}\mathbb{E}_{s\sim d_{\mathcal{M},\pi}(\cdot)}\left[\max_{a}Q_{\mathcal{M}}^{*}(s,a)-\int_{a}\pi(a|s)Q_{\mathcal{M}}^{*}(s,a)da\right]\\ \geq&\frac{1}{1-\gamma}\mathbb{E}_{s\sim d_{\mathcal{M},\pi}(\cdot)}\left[\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}\max_{a}Q_{\mathcal{M}}^{*}(s,a)-\max_{a}\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}Q_{\mathcal{M}}^{*}(s,a)\right]\\ =&\frac{1}{1-\gamma}\mathbb{E}_{s\sim d_{\mathcal{M},\pi}}\left[D_{\text{TDR}}(\mathcal{T},s)\right].\end{split}$ (21)
 Since $J_{\mathcal{M}}(\pi_{\mathcal{M}}^{*})=\max_{\pi\in\Pi_{1}}J_{\mathcal{M}}(\pi)$, we have
-$\begin{split}&\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}\max_{\pi\in\Pi_{1}}J_{\mathcal{M}}(\pi)-\max_{\pi\in\Pi_{1}}\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}J_{\mathcal{M}}(\pi)\\ \geq&\frac{1}{1-\gamma}\mathbb{E}_{s\sim d_{\mathcal{M},\pi^{*}}}\left[D_{\text{TDR}}(\mathcal{T},s)\right],\end{split}$ (21)
-Thus we have proven this result. ∎
+$\begin{split}\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}\max_{\pi\in\Pi_{1}}J_{\mathcal{M}}(\pi)-\max_{\pi\in\Pi_{1}}\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}J_{\mathcal{M}}(\pi)\geq&\frac{1}{1-\gamma}\mathbb{E}_{s\sim d_{\mathcal{M},\pi^{*}}}\left[D_{\text{TDR}}(\mathcal{T},s)\right],\end{split}$ (22)
+Thus we have proven this result.
 A.5 Expressive ability of $\Pi_{3}$
-Proposition 1.
+Proposition A.4.
 For $\forall\epsilon_{1},\epsilon_{2}$ satisfying $0<\epsilon_{1}\leq 1,0<\epsilon_{2}\leq 1$, there exists a task distribution $\mathcal{T}$ satisfying that
 $\begin{split}J_{\mathcal{T}}^{1}=J_{\mathcal{T}}^{2}\leq\epsilon_{1},~~J_{\mathcal{T}}^{3}\geq 1-\epsilon_{2},~~J_{\mathcal{T}}^{*}=1.\end{split}$ (23)
-Proof.
+Proof A.5.
 Given fixed discount factor $\gamma\in(0,1)$, we first take $n\in\mathbb{N}$ satisfying $n\geq\frac{1}{\epsilon_{1}}$, $A=\epsilon_{2}\frac{n}{n-1}\frac{1-\gamma}{1-\gamma^{n}}$, and $B=\frac{1-\gamma}{\gamma^{n+1}}\left(1-\epsilon_{2}\frac{n}{n-1}\right)$.
@@
 The initial state of each task is $s_{0}^{1}$, and the dynamic as well as reward functions are as below
-$\begin{split}&\mathcal{P}(s_{t}^{l},a_{j})=\mathbb{I}(s=s_{t+1}^{j}),\quad j,l=1,...,n;\ t=0,1,...,\infty\\ &\mathcal{R}_{i}(s_{t}^{l},a_{j})=f(t)\mathbb{I}(i=j),\quad i,j,l=1,...,n;\ t=0,1,...\infty\\ &f(t)=\left\{\begin{array}[]{ll}A,&t\leq n-1\\ B,&t\geq n\\ \end{array}\right.\end{split}$ (23)
+[…]
 Take $\mathcal{T}$ as the uniform distribution over $\mathcal{M}_{1},...,\mathcal{M}_{n}$, thus we have
-$\begin{split}J_{\mathcal{M}_{i}}^{*}=&A+A\gamma+...+A\gamma^{n-1}+B\gamma^{n}+B\gamma^{n+1}+...\\ =&A\frac{1-\gamma^{n}}{1-\gamma}+B\frac{\gamma^{n+1}}{1-\gamma}=1\\ J_{\mathcal{T}}^{*}=&\frac{1}{n}\sum_{i=1}^{n}J_{\mathcal{M}_{i}}^{*}=1.\\ \end{split}$ (24)
+[…]
 Since our construction satisfies $\mathbb{E}_{\mathcal{T}}[\mathcal{R}_{i}(s_{k}^{l},a_{j})]=\frac{f(k)}{n}$, for $\forall\pi\in\Pi_{2}$, we have
-$\begin{split}J_{\mathcal{T}}(\pi)=&\frac{1}{n}(A+A\gamma+...+A\gamma^{n-1}+B\gamma^{n}+B\gamma^{n+1}+...)\\ =&\frac{1}{n},\\ J_{\mathcal{T}}^{1}=&J_{\mathcal{T}}^{2}=\max_{\pi\in\Pi_{2}}J_{\mathcal{T}}(\pi)=\frac{1}{n}\leq\epsilon_{1}.\end{split}$ (25)
+$\begin{split}J_{\mathcal{T}}(\pi)=&\frac{1}{n}(A+A\gamma+...+A\gamma^{n-1}+B\gamma^{n}+B\gamma^{n+1}+...)=\frac{1}{n},\\ J_{\mathcal{T}}^{1}=&J_{\mathcal{T}}^{2}=\max_{\pi\in\Pi_{2}}J_{\mathcal{T}}(\pi)=\frac{1}{n}\leq\epsilon_{1}.\end{split}$ (26)
 Moreover, we can construction an agent $\hat{\pi}\in\Pi_{3}$ that takes action via the historical trajectory $\hat{\tau}_{t}=(\hat{s}_{0},\hat{a}_{0},\hat{r}_{0},...\hat{s}_{t})$:
-$\begin{split}\hat{\pi}(a_{j}|\hat{\tau}_{t})=&\mathbb{I}(j=t+1),\quad t=0,1,...,n-1\\ \hat{\pi}(a_{j}|\hat{\tau}_{t})=&\mathbb{I}(j=i),\quad t=n,...,\infty\\ \end{split}$ (26)
+$\begin{split}\hat{\pi}(a_{j}|\hat{\tau}_{t})=&\mathbb{I}(j=t+1),\quad t=0,1,...,n-1\\ \hat{\pi}(a_{j}|\hat{\tau}_{t})=&\mathbb{I}(j=i),\quad t=n,$$\infty\\ \end{split}$ (27)
 here $i=\arg\max\{\hat{r}_{0},\hat{r}_{1},...,\hat{r}_{n-1}\}+1$, thus we have
-$\begin{split}&J_{\mathcal{T}}^{3}\geq J_{\mathcal{T}}(\hat{\pi})\\ =&\frac{1}{n}(A+A\gamma+...+A\gamma^{n-1})+B\gamma^{n}+B\gamma^{n+1}+...\\ =&\frac{A}{n}\frac{1-\gamma^{n}}{1-\gamma}+B\frac{\gamma^{n+1}}{1-\gamma}=1-\frac{(n-1)A}{n}\frac{1-\gamma^{n}}{1-\gamma}\\ \geq&1-\epsilon_{2}.\end{split}$ (27)
-Thus we have proven this result. ∎
+$\begin{split}J_{\mathcal{T}}^{3}\geq J_{\mathcal{T}}(\hat{\pi})=&\frac{1}{n}(A+A\gamma+...+A\gamma^{n-1})+B\gamma^{n}+B\gamma^{n+1}+...\\ =&\frac{A}{n}\frac{1-\gamma^{n}}{1-\gamma}+B\frac{\gamma^{n+1}}{1-\gamma}=1-\frac{(n-1)A}{n}\frac{1-\gamma^{n}}{1-\gamma}\geq 1-\epsilon_{2}.\end{split}$ (28)
+Thus we have proven this result.
 A.6 Proof and discussion of Theorem 4
@@
 Recall that we consider the policy hypothesis $\mathcal{H}_{3}$ here, that each policy $\pi:\mathcal{L}\rightarrow\Delta(\mathcal{A}),\mathcal{L}=\cup_{t=1}^{\infty}\mathcal{L}_{t},\mathcal{L}_{t}=(\mathcal{S}\times\mathcal{A}\times\mathbb{R})^{t-1}\times\mathcal{S}$.
-As directly such $\mathcal{S}-\mathcal{A}-\mathcal{R}$ memorized policy is difficult, we consider an alternative MDP as ~$\tilde{\mathcal{M}}=(\mathcal{L},\mathcal{A},\mathcal{P}_{\mathcal{M}},\mathcal{R}_{\mathcal{M}},\gamma)$.
+As directly such $\mathcal{S}-\mathcal{A}$-$\mathcal{R}$ memorized policy is difficult, we consider an alternative MDP as $\tilde{\mathcal{M}}=(\mathcal{L},\mathcal{A},\mathcal{P}_{\mathcal{M}},\mathcal{R}_{\mathcal{M}},\gamma)$.
 For $\forall l=(s_{1},a_{1},r_{1},...,s_{t})\in\mathcal{L}_{t}\subseteq\mathcal{L}$, we can sample the action $a_{t}$ from the distribution $\pi(\cdot|l)$.
@@
 In summary, we set the distribution $p(l^{\prime}|l,a_{t})$ as the corresponding dynamic $\mathcal{P}_{\mathcal{M}}$.
-Notice that although all $\mathcal{M}\in\mathcal{T}$ shares the same dynamic $\mathcal{P}$, their new dynamic $\mathcal{P}_{\mathcal{M}}$ are different since the dynamic is related to the given reward.
-An obvious advantage of introducing $\tilde{\mathcal{M}}$ is that $\pi\in\Pi_{3}$ is now Markovian in $\tilde{\mathcal{M}}$ and it is much easier to analyze its performance.
+Notice that although all $\mathcal{M}\in\mathcal{T}$ share the same dynamic $\mathcal{P}$, their new dynamic $\mathcal{P}_{\mathcal{M}}$ are different since the dynamic is related to the given reward.
+An obvious advantage of introducing $\tilde{\mathcal{M}}$ is that $\pi\in\Pi_{3}$ is now Markovian in $\tilde{\mathcal{M}}$, and it is much easier to analyze its performance.
 Obviously, we can set $J_{\mathcal{M}}(\pi)=J_{\tilde{\mathcal{M}}}(\pi),Q_{\mathcal{M}}^{*}(s,a)=Q_{\tilde{\mathcal{M}}}^{*}(s,a)$ to simplify the notation, and we can prove that
-Theorem 5.
+Theorem A.6.
 For any policy $\pi\in\Pi_{3}$, we have
-$\begin{split}&J_{\mathcal{T}}^{*}-\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}\left[J_{\mathcal{M}}(\pi)\right]\\ =&\frac{1}{1-\gamma}\int_{\mathcal{L}}p(l)\left[\int p(\mathcal{M}|l)\max_{a}Q_{\mathcal{M}}^{*}(l,a)d\mathcal{M}\right.\\ &\qquad\ -\left.\int_{a,\mathcal{M}}\pi(a|l)p(\mathcal{M}|l)Q_{\mathcal{M}}^{*}(l,a)dad\mathcal{M}\right]dl\\ \geq&\frac{1}{1-\gamma}\int_{\mathcal{L}}p(l)\left[\int p(\mathcal{M}|l)\max_{a}Q_{\mathcal{M}}^{*}(l,a)d\mathcal{M}\right.\\ &\qquad\qquad\quad-\left.\max_{a}\int_{\mathcal{M}}p(\mathcal{M}|l)Q_{\mathcal{M}}^{*}(l,a)d\mathcal{M}\right]dl,\end{split}$ (28)
+$\begin{split}J_{\mathcal{T}}^{*}-\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}\left[J_{\mathcal{M}}(\pi)\right]=&\frac{1}{1-\gamma}\int_{\mathcal{L}}p(l)\left[\int p(\mathcal{M}|l)\max_{a}Q_{\mathcal{M}}^{*}(l,a)d\mathcal{M}-\int_{a,\mathcal{M}}\pi(a|l)p(\mathcal{M}|l)Q_{\mathcal{M}}^{*}(l,a)dad\mathcal{M}\right]dl\\ \geq&\frac{1}{1-\gamma}\int_{\mathcal{L}}p(l)\left[\int p(\mathcal{M}|l)\max_{a}Q_{\mathcal{M}}^{*}(l,a)d\mathcal{M}-\max_{a}\int_{\mathcal{M}}p(\mathcal{M}|l)Q_{\mathcal{M}}^{*}(l,a)d\mathcal{M}\right]dl,\end{split}$ (29)
 here $p(l)$ is a distribution of $\mathcal{L}$ related to $\mathcal{T},\pi$ and $p(\mathcal{M}|l)$ is the task posterior related to $\pi$.
-Proof.
+Proof A.7.
 $\forall\pi\in\mathcal{H}_{3}$, as $\pi$ is Markovian in $\tilde{\mathcal{M}}$, we can directly utilize the proof of Theorem 3 from the beginning to Eq.
-(LABEL:app_eq_11), and the only difference is that the dynamics in $\tilde{\mathcal{M}}$ are different but the dynamics in $\mathcal{M}$ are the same.
-Thus we need to change the Eq.
-(LABEL:app_eq_12) as
-$\begin{split}&\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}\left[J_{\mathcal{M}}(\pi_{\mathcal{M}}^{*})-J_{\mathcal{M}}(\pi)\right]\\ =&\frac{1}{1-\gamma}\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}\mathbb{E}_{l\sim d_{\tilde{\mathcal{M}},\pi}(\cdot)}\left[\max_{a}Q_{\mathcal{M}}^{*}(l,a)\right.\\ &\left.-\int_{a}\pi(a|l)Q_{\mathcal{M}}^{*}(l,a)da\right]\\ =&\frac{1}{1-\gamma}\int p(\mathcal{M})\int_{\mathcal{L}}d_{\tilde{\mathcal{M}},\pi}(l)\left[\max_{a}Q_{\mathcal{M}}^{*}(l,a)\right.\\ &\left.-\int_{a}\pi(a|l)Q_{\mathcal{M}}^{*}(l,a)da\right]\\ =&\frac{1}{1-\gamma}\int_{\mathcal{L}}p(l)\int p(\mathcal{M}|l)\left[\max_{a}Q_{\mathcal{M}}^{*}(l,a)\right.\\ &\left.-\int_{a}\pi(a|l)Q_{\mathcal{M}}^{*}(l,a)da\right]\\ \geq&\frac{1}{1-\gamma}\int_{\mathcal{L}}p(l)\left[\int p(\mathcal{M}|l)\max_{a}Q_{\mathcal{M}}^{*}(l,a)d\mathcal{M}\right.\\ &\qquad\qquad\quad-\left.\max_{a}\int_{\mathcal{M}}Q_{\mathcal{M}}^{*}(l,a)d\mathcal{M}\right]dl,\end{split}$ (29)
-[…]
+(20), and the only difference is that the dynamics in $\tilde{\mathcal{M}}$ are different but the dynamics in $\mathcal{M}$ are the same.
+Thus, we need to change the Eq.
+(21) as
+$\begin{split}\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}\left[J_{\mathcal{M}}(\pi_{\mathcal{M}}^{*})-J_{\mathcal{M}}(\pi)\right]=&\frac{1}{1-\gamma}\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}\mathbb{E}_{l\sim d_{\tilde{\mathcal{M}},\pi}(\cdot)}\left[\max_{a}Q_{\mathcal{M}}^{*}(l,a)-\int_{a}\pi(a|l)Q_{\mathcal{M}}^{*}(l,a)da\right]\\ =&\frac{1}{1-\gamma}\int p(\mathcal{M})\int_{\mathcal{L}}d_{\tilde{\mathcal{M}},\pi}(l)\left[\max_{a}Q_{\mathcal{M}}^{*}(l,a)-\int_{a}\pi(a|l)Q_{\mathcal{M}}^{*}(l,a)da\right]\\ =&\frac{1}{1-\gamma}\int_{\mathcal{L}}p(l)\int p(\mathcal{M}|l)\left[\max_{a}Q_{\mathcal{M}}^{*}(l,a)-\int_{a}\pi(a|l)Q_{\mathcal{M}}^{*}(l,a)da\right]\\ \geq&\frac{1}{1-\gamma}\int_{\mathcal{L}}p(l)\left[\int p(\mathcal{M}|l)\max_{a}Q_{\mathcal{M}}^{*}(l,a)d\mathcal{M}-\max_{a}\int_{\mathcal{M}}Q_{\mathcal{M}}^{*}(l,a)d\mathcal{M}\right]dl,\end{split}$ (30)
+here $p(l)=\int p(\mathcal{M})d_{\tilde{\mathcal{M}},\pi}(l)d\mathcal{M}$, and $p(\mathcal{M}|l)=p(\mathcal{M})d_{\tilde{\mathcal{M}},\pi}(l)/p(l)$ is the posterior distribution.
 Finally, we will show that maximizing $p(\mathcal{M}|l)$ is helpful for task generalization.
@@
 In such a situation, for each $l$, we can infer a “most possible” posterior task $\mathcal{M}_{l}$ with high $p(\mathcal{M}_{l}|l)$, thus we can approximately take $\pi(l)=\arg\max_{a}Q_{\mathcal{M}_{l}}^{*}(l,a)$ and the optimal gap calculated by Eq.
-(LABEL:app_ep_20) will be controlled.
-A.7 Sparse Reward Task
+(29) will be controlled.
+A.7 Sparse reward tasks
 In this part, we show that generalizing to tasks with the same dynamics and sparse rewards without extra knowledge (like context) is extremely difficult and sometimes impossible.
@@
 Algorithm 2 Task Aware Dreamer (TAD)
-0:  $M$ training tasks $\{\mathcal{M}_{m}\}_{m=1}^{M}$, $M$ replay buffers $\{\mathcal{D}_{m}\}_{m=1}^{M}$, $N$ test tasks $\{\mathcal{M}_{M+n}\}_{n=1}^{N}$, initialize neural network parameters of world models, the policy, and the critic
+0:$M$ training tasks $\{\mathcal{M}_{m}\}_{m=1}^{M}$, $M$ replay buffers $\{\mathcal{D}_{m}\}_{m=1}^{M}$, NN test tasks $\{\mathcal{M}_{M+n}\}_{n=1}^{N}$, initialize neural network parameters of world models, the policy, and the critic
 1:while not converge do
-2:     //$Model\ Training$
-3:     for update $\text{update step}=1,2,...,U$ do
+[…]
+3:  for $\text{update step}=1,2,...,U$ do
 4:   Sample observation-action-reward pairs form each replay buffer $\{(o_{t}^{i},a_{t}^{i},r_{t}^{i})_{t=1}^{T}\}\sim\mathcal{D}_{i},i=1,2,...,M$
@@
 9:  end for
-10:     //$Data\ Collection$
-11:     for $m=1,2,...,M$ do
+[…]
+11:  for m=1,2,,Mm=1,2,...,M do
 12:   $o_{1}\leftarrow\mathcal{M}_{m}.reset()$
-13:        for sample $\text{sample step}=1,2,...,S$ do
+13:   for $\text{sample step}=1,2,...,S$ do
 14:    Compute $h_{t},s_{t}$ and sample action $a_{t}$ via the policy.
@@
 19:end while
-20:  //$Model\ Evaluation$
-21:  for $n=1,2,...,N$ do
+[…]
+21:for n=1,2,,Nn=1,2,...,N do
 22:  $o_{1}\leftarrow\mathcal{M}_{M+n}.reset()$
@@
 C.1 Details of all task combinations
-[…]
+In this part, we first roughly discuss the reward function of tasks in our experiments to better understand their TDR.
+These reward functions always defined by tolerance function in DeepMind control suite [8], which is a smooth function with parameters $\text{tolerance}(x,\text{bounds}=(\text{lower},\text{upper}))$ and hope the value of $x$ is within $(\text{lower},\text{upper})$.
+More details about tolerance function can be found in [8].
   • Cartpole-balance&sparse.
-This task combination includes two tasks: Cartpole-balance and Cartpole-balance__\__sparse, which both hope to balance an unactuated pole with dense and sparse rewards respectively.
-The optimal actions of these tasks are both hoped to balance the agent and thus TDR here is 0.
+This task combination includes two tasks: Cartpole-balance and Cartpole-balance_\_sparse, which both hope to balance an unactuated pole with dense and sparse rewards, respectively.
+The optimal actions of these tasks are both hoped to balance the agent, and thus, TDR here is 0.
   • Walker-stand&walk&prostrate&flip.
 This task combination includes four tasks: Walker-stand, Walker-walk, Walker-prostrate, and Walker-flip.
-Walker-stand hopes the height of a two-leg robot to be larger than the target height.
+Walker-stand hopes the height of a two-legged robot to be larger than the target height.
 Walker-walk hopes the height of the improved planar walker to be larger than a target height and the speed of the robot to be larger than another target speed.
@@
 In detail, their reward functions can be roughly described as
-    $\begin{split}\mathcal{R}_{\text{stand}}=&\text{tolerance}(\text{height},(1.2,\infty)),\\ \mathcal{R}_{\text{walk}}=&\text{tolerance}(\text{height},(1.2,\infty))\\ &*\text{tolerance}(\text{speed},(1,\infty)).\\ \mathcal{R}_{\text{prostrate}}=&\text{tolerance}(\text{height},(0.0,0.2)).\\ \mathcal{R}_{\text{flip}}=&\text{tolerance}(\text{height},(1.2,\infty))\\ &*\text{tolerance}(\text{angmomentum},5,\infty).\end{split}$ (31)
+    $\begin{split}\mathcal{R}_{\text{stand}}=&\text{tolerance}(\text{height},(1.2,\infty)),\\ \mathcal{R}_{\text{walk}}=&\text{tolerance}(\text{height},(1.2,\infty))*\text{tolerance}(\text{speed},(1,\infty)).\\ \mathcal{R}_{\text{prostrate}}=&\text{tolerance}(\text{height},(0.0,0.2)).\\ \mathcal{R}_{\text{flip}}=&\text{tolerance}(\text{height},(1.2,\infty))*\text{tolerance}(\text{angmomentum},5,\infty).\end{split}$ (32)
     In this situation, for all states, the optimal actions of Walker-walk/stand and Walker-prostrate are opposite, and TDR here is huge.
-  • Cheetah-run&run__\__back&flip&flip__\__back.
-This task combination includes four tasks: Cheetah-run, Cheetah-run__\__back, Cheetah-flip, and Cheetah-flip__\__back.
+  • Cheetah-run&run_\_back&flip&flip_\_back.
+This task combination includes four tasks: Cheetah-run, Cheetah-run_\_back, Cheetah-flip, and Cheetah-flip_\_back.
 Cheetah-run hopes to control a running planar biped to run forward within a target speed.
-Cheetah-run__\__back, differently, hopes to control the Cheetah robot to run backward within a target speed.
+Cheetah-run_\_back, differently, hopes to control the Cheetah robot to run backward within a target speed.
 Cheetah-flip hopes the robot to move forward to the target speed by executing a rapid twist and jump.
-Similarly, Cheetah-flip__\__back hopes to control the robot to move backward by flipping.
+Similarly, Cheetah-flip_\_back hopes to control the robot to move backward by flipping.
     $\begin{split}\mathcal{R}_{\text{Cheetah}\_\text{run}}=&\text{tolerance}(\text{speed},(10,\infty)).\\ \mathcal{R}_{\text{Cheetah}\_\text{run$\_$back}}=&\text{tolerance}(-\text{speed},(10,\infty)).\\ \mathcal{R}_{\text{Cheetah}\_\text{flip}}=&\text{tolerance}(\text{angmomentum},(5,\infty)).\\ \mathcal{R}_{\text{Cheetah}\_\text{flip$\_$back}}=&\text{tolerance}(-\text{angmomentum},(5,\infty)).\end{split}$ (33)
-    Obviously, in this situation, for all states, the optimal actions of Cheetah-run and Cheetah-run__\__back are opposite, and TDR here is also huge.
-Algorithm Hypothesis Cartpole-balance Walker-stand&walk Cheetah-run&run__\__back
-&balance__\__sparse &prostrate&flip &flip&flip__\__back
-CURL $\Pi_{1}$ 994.5 $\pm$ 3.6 254.1 $\pm$ 9.2 229.7 $\pm$ 10.9
-CURL (w/ r) $\Pi_{3}$ 987.3 $\pm$ 12.9 265.0 $\pm$ 4.7 236.1 $\pm$ 5.8
-SAC+AE $\Pi_{1}$ 992.5 $\pm$ 2.6 256.9 $\pm$ 5.9 225.8 $\pm$ 10.1
-SAC+AE (w/ r) $\Pi_{3}$ 988.5 $\pm$ 7.0 257.2 $\pm$ 8.5 239.9 $\pm$ 12.4
-Table 4: Performance (mean $\pm$ std) in DMC of CURL, CURL (w/ r), SAC+AE, and SAC+AE (w/ r).
-Tasks Acrobot-Cartpole-Pendulum Walker-Cheetah-Hopper Cheetah-run__\__mass Walker-walk__\__mass
+    Obviously, in this situation, for all states, the optimal actions of Cheetah-run and Cheetah-run_\_back are opposite, and TDR here is also huge.
+Tasks Acrobot-Cartpole-Pendulum Walker-Cheetah-Hopper Cheetah-run_\_mass Walker-walk_\_mass
 Train&Test Train&Test Train Test Train Test
@@
 TAD 667.7 $\pm$ 6.4 554.7 $\pm$ 23.6 754.3 $\pm$ 22.9 738.2 $\pm$ 41.1 957.8 $\pm$ 35.1 963.1 $\pm$ 32.6
-[…]
+Table 4: Generalization performance (mean $\pm$ std) over different task distributions in image-based DMC of the best policy.
+Numbers greater than 95 $\%$ of the best performance for each environment are bold.
 C.2 Details of all task distributions
-Now we introduce the three task distributions in our experiments, which are designed based on existing tasks in DeepMind control suite for testing the generalization of trained agents.
-  • Cheetah__\__speed($\alpha,\beta$).
-This task distribution is designed in this paper with parameter $0\leq\beta\leq\alpha$, based on the task Cheetah__\__run in DeepMind control suite, and hopes the Cheetah robot can run with the target speed interval $(\alpha-\beta,\alpha+\beta)$.
+Now we introduce the three task distributions in our experiments, which are designed based on existing tasks in the DeepMind control suite for testing the generalization of trained agents.
+  • Cheetah_\_speed($\alpha,\beta$).
+This task distribution is designed in this paper with parameter $0\leq\beta\leq\alpha$, based on the task Cheetah_\_run in the DeepMind control suite, and hopes the Cheetah robot can run with the target speed interval $(\alpha-\beta,\alpha+\beta)$.
     $\begin{split}&\mathcal{R}_{\text{Cheetah}\_\text{speed}}(\alpha,\beta)=\text{tolerance}(\text{speed},(\alpha-\beta,\alpha+\beta)).\end{split}$ (34)
     We train the agents in tasks with parameters ($0.5,0.2$), ($1.5,0.2$), ($2.0,0.2$), ($3.0,0.2$) and test them in tasks with parameters ($1.0,0.2$), ($2,5,0.2$).
-  • Pendulum__\__angle($\alpha,\beta$).
-This task distribution is designed in this paper with parameter $-1\leq\alpha\leq\beta\leq 1$, based on the task Pendulumh__\__swingup in DeepMind control suite, and hopes the Pendulum robot can swing up within the target angle interval $(\arccos\alpha,\arccos\beta)$.
-    $\begin{split}\mathcal{R}_{\text{Pendulum}\_\text{angle}}(\alpha,\beta)=\text{tolerance}(\text{angle},(&\arccos\alpha,\\ &\arccos\beta)).\end{split}$ (34)
-    Training tasks are with parameters ($-0.95,-0.9$), ($-0.85,-0.8$), ($-0.8,-0.75$), ($-0.7,-0.65$) and test tasks are with parameters ($-0.9,-0.85$), ($-0,75,-0.7$).
-  • Walker__\__speed($\alpha,\beta$).
-This task distribution is designed in this paper with parameter $0\leq\beta\leq\alpha$, based on the task Walker__\__run in DeepMind control suite, and hopes the Walker robot can run within the target speed interval $(\alpha-\beta,\alpha+\beta)$.
+  • Pendulum_\_angle($\alpha,\beta$).
+This task distribution is designed in this paper with parameter $1\leq\alpha\leq\beta\leq 1$, based on the task Pendulumh_\_swingup in DeepMind control suite, and hopes the Pendulum robot can swing up within the target angle interval $(\arccos\alpha,\arccos\beta)$.
+    $\begin{split}\mathcal{R}_{\text{Pendulum}\_\text{angle}}(\alpha,\beta)=\text{tolerance}(\text{angle},(&\arccos\alpha,\arccos\beta)).\end{split}$ (35)
+    Training tasks are with parameters (0.95,0.9-0.95,-0.9), (0.85,0.8-0.85,-0.8), (0.8,0.75-0.8,-0.75), (0.7,0.65-0.7,-0.65) and test tasks are with parameters (0.9,0.85-0.9,-0.85), (0,75,0.7-0,75,-0.7).
+  • Walker_\_speed($\alpha,\beta$).
+This task distribution is designed in this paper with parameter $0\leq\beta\leq\alpha$, based on the task Walker_\_run in DeepMind control suite, and hopes the Walker robot can run within the target speed interval $(\alpha-\beta,\alpha+\beta)$.
     $\begin{split}&\mathcal{R}_{\text{Walker}\_\text{speed}}(\alpha,\beta)=\text{tolerance}(\text{speed},(\alpha-\beta,\alpha+\beta)).\end{split}$ (36)
@@
 All algorithms are trained on one NVIDIA GeForce RTX 2080 Ti.
-For each seed and each task setting, it will take around 3 days.
-As for the hyper-parameters, we follow previous works (Ha and Schmidhuber 2018; Hafner et al.
-2019b, a) and select 2 as the action repeat for all experiments following (Hafner et al.
-2019a).
+Each seed and each task setting will take around 3 days.
+As for the hyper-parameters, we follow previous works [3, 4, 5] and select 2 as the action repeat for all experiments following [5].
 C.3 Visualization results for task generalization
-Moreover, for each task sampled from the task distribution Cheetah__\__speed (here parameters $(3.0,0.2)$, $(2.0,0.2)$, $(1.5,0.2)$, $(0.5,0.2)$ are for training tasks and parameters $(2.5,0.2),(0.5,0.2)$ are for test tasks), we plot the speed of the agent as a function of the timestep in Fig.
-6.
-As depicted, for each task, the agent trained by TAD will quickly improve its speed until reaching the target speed and then keep its speed since the speed determines whether it has met the task requirements via utilizing historical information.
+Moreover, for each task sampled from the task distribution Cheetah_\_speed (here parameters $(3.0,0.2)$, $(2.0,0.2)$, $(1.5,0.2)$, $(0.5,0.2)$ are for training tasks and parameters $(2.5,0.2),(0.5,0.2)$ are for test tasks), we plot the speed of the agent as a function of the timestep in Fig.
+4.
+As depicted, for each task, the agent trained by TAD will quickly improve its speed until it reaches the target speed and then keep its speed since the speed determines whether it has met the task requirements via utilizing historical information.
 Consequently, TAD is aware of different tasks and can successfully generalize to unseen test tasks.
-We also provide videos of these trajectories in supplementary materials.
-Figure 6: Visualization of the trained TAD agent in the task distribution of Cheetah__\__speed. We plot the speed of the agent as a function of timesteps in all tasks.
+We also provide videos of these trajectories in the supplementary materials.
+Figure 4: Visualization of the trained TAD agent in the task distribution of Cheetah_\_speed. We plot the speed of the agent as a function of timesteps in all tasks.
+C.4 Implementation details
+TAD is implemented based on a widely adopted Dreamer repository https://github.com/yusukeurakami/dreamer-pytorch and we follow its default setups, including hyperparameters and network architecture.
+For the task model $p_{\theta}(\mathcal{M}|h_{t},s_{t})$, we choose to implement it with a simple 3-layer MLP, which maps current state $s_{t}$ and hidden state $s_{t}$ to a probability simplex $\triangle^{M-1}$ (assume we have $M$ tasks).
+C.5 TDR analyses in DMControl experiments
+We would like to illustrate the usage of TDR with a few examples in our experiments, including Walker-stand&\&walk and Walker-walk&\&prostrate.
+To simplify the analyses, we set the reward of each state in Walker-stand $\mathcal{M}_{1}$ to 1 if the height of the robot is higher than 1.2 and 0 otherwise.
+In Walker-prostrate $\mathcal{M}_{2}$, we set the reward of each state to 1 if the height is lower than 0.2 and 0 otherwise.
+In Walker-walk, $\mathcal{M}_{3}$, we set the reward of each state to 1 if the height is higher than 1.2 and the speed is greater than 1.0; otherwise, the reward is 0.
+We first estimate the optimal $Q^{*}$ function of these three tasks.
+For any state action pair $(s,a)$ in $\mathcal{M}_{1}$, if the height of the current state $s$ is higher than 1.2, the optimal policy is to maintain the height and we have $Q^{*}_{\mathcal{M}_{1}}=1+\gamma+\gamma^{2}+...=\frac{1}{1-\gamma}$.
+Otherwise, if the height of $s$ is lower than 1.2, the optimal policy is to make the agent stand in one step and maintain a height of at least 1.2.
+We can deduce that $Q^{*}_{\mathcal{M}_{1}}=\gamma+\gamma^{2}+...=\frac{\gamma}{1-\gamma}$.
+The analyses of $Q^{*}_{\mathcal{M}_{2}},Q^{*}_{\mathcal{M}_{3}}$ are similar and we conclude as follows:
+[…]
+Consider a task combination $\mathcal{T}_{1}$ that is distributed equally over Walker-stand and Walker-walk, i.e., $\mathbb{P}(\mathcal{M}_{1})=\mathbb{P}(\mathcal{M}_{3})=0.5$.
+If height of $s\geq 1.2$ and speed of $s\geq 1.0$, $Q^{*}_{\mathcal{M}_{1}}(s,a)=Q^{*}_{\mathcal{M}_{3}}(s,a)=\frac{1}{1-\gamma}$.
+Thus, we have:
+$\begin{split}D_{\text{TDR}}(\mathcal{T}_{1},s)=&\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}[\max_{a}Q_{\mathcal{M}}^{*}(s,a)]-\max_{a}\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}[Q_{\mathcal{M}}^{*}(s,a)]=\frac{1}{1-\gamma}-\frac{1}{1-\gamma}=0.\end{split}$ (38)
+Based on Theorem 3, for some policy $\pi\in\Pi_{1},\Pi_{2}$:
+$J_{\mathcal{T}}^{*}-\mathbb{E}_{\mathcal{M}\sim\mathcal{T}_{1}}\left[J_{\mathcal{M}}(\pi)\right]\geq\frac{1}{1-\gamma}\mathbb{E}_{s\sim d_{\mathcal{M},\pi}}[D_{\text{TDR}}(\mathcal{T}_{1},s)]=0.$ (39)
+Consequently, in the task distribution $\mathcal{T}_{1}$, there is no gap between the optimal performance achievable by the policies in $\Pi_{1},\Pi_{2}$ and the actual optimal value.
+A TDR of 0 indicates that policies in $\Pi_{1}$ and $\Pi_{2}$ are sufficient to achieve optimal performance in $\mathcal{T}_{1}$..
+In contrast, consider a task combination $\mathcal{T}_{2}$ that is distributed equally over Walker-walk and Walker-prostrate, i.e., $\mathbb{P}(\mathcal{M}_{2})=\mathbb{P}(\mathcal{M}_{3})=0.5$.
+For any $s$, $\frac{1}{2}(Q^{*}_{\mathcal{M}_{2}}(s,a)+Q^{*}_{\mathcal{M}_{3}}(s,a))\leq\frac{1}{2}(\frac{1}{1-\gamma}+\frac{\gamma}{1-\gamma})$.
+Thus, we have:
+$\begin{split}D_{\text{TDR}}(\mathcal{T}_{2},s)=&\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}[\max_{a}Q_{\mathcal{M}}^{*}(s,a)]-\max_{a}\mathbb{E}_{\mathcal{M}\sim\mathcal{T}_{2}}[Q_{\mathcal{M}}^{*}(s,a)]\geq\frac{1}{1-\gamma}-\frac{1}{2}\frac{1+\gamma}{1-\gamma}=0.5.\end{split}$ (40)
+Consequently, based on Theorem 3, for any policy $\pi\in\Pi_{1},\Pi_{2}$:
+$J_{\mathcal{T}}^{*}-\mathbb{E}_{\mathcal{M}\sim\mathcal{T}_{2}}\left[J_{\mathcal{M}}(\pi)\right]\geq\frac{1}{1-\gamma}\mathbb{E}_{s\sim d_{\mathcal{M},\pi}}[D_{\text{TDR}}(\mathcal{T}_{2},s)]=\frac{1}{2}\frac{1}{1-\gamma}.$ (41)
+In other words, our conclusion provides a non-trivial gap between the optimal performance achievable by the policies in $\Pi_{1},\Pi_{2}$ and the actual optimal value.
+A non-zero TDR shows that policies in $\Pi_{1},\Pi_{2}$ can only obtain half the optimal performance in $\mathcal{T}_{2}$.
+Our experimental results in Table 5 also confirm our analysis regarding the effect of TDR.
+While policies in $\Pi_{1},\Pi_{2}$ perform appealingly in Walker-stand&\&walk, their performance degenerates significantly in Walker-walk&\&prostrate.
+Algorithm Walker-stand&\&walk Walker-walk&\&prostrate
+CURL 563.4 $\pm$ 16.0 441.0 $\pm$ 7.7
+SAC+AE 683.8 $\pm$ 164.3 445.8 $\pm$ 2.6
+PlaNet 990.6 $\pm$ 3.8 445.4 $\pm$ 5.1
+Dreamer 992.3 $\pm$ 1.7 449.7 $\pm$ 12.5
+TAD 994.6 $\pm$ 0.1 855.3 $\pm$ 56.2
+Table 5: Performance (mean $\pm$ std) over different task combinations in DMC of the best policy. Numbers greater than 95 % of the best performance for each environment are bold.
+C.6 MSE loss of imagination images in DMControl experiments
+In Table 6, we have provided quantitative results on the models’ prediction error of imagined observations.
+As shown below, the prediction error of Dreamer gradually increases over the time horizon, while TAD exhibits no significant increasing trend.
+We attribute this phenomenon to TAD’s ability to access diverse trajectories sampled from different tasks during training.
+Algorithm Task step 10 step 15 step 20 step 25 step 30 step 35
+Dreamer Cheetah-run 19.2 18.8 28.7 29.3 30.1 32.5
+Cheetah-flip 18.6 24.3 34.8 37.5 33.5 31.1
+TAD (Ours) Cheetah-run 19.3 23.2 21.3 23.9 25.6 23.3
+Cheetah-flip 17.9 18.8 21.1 19.5 17.3 19.4
+Table 6: MSE Loss of imagination results of TAD and Dreamer
+C.7 Additional results with task parameters
+The core idea of TAD is to regularize the world model to distinguish tasks with $p_{\theta}(\mathcal{M}|h_{t},s_{t})$, with $\mathcal{M}$ being the task indicator.
+While the current implementation uses a set of discrete training tasks, we would like to clarify that our method is not limited to discrete spaces of tasks.
+For continuous spaces of tasks, when the task parameters are available, we can optimize TAD by directly predicting the continuous task parameters.
+When the task parameters are unavailable, we can switch to the discrete implementation, where we randomly sample $M$ tasks to construct a set of discrete training tasks and optimize TAD by predicting the task indices.
+Obviously, the task indices are less informative than task parameters, because they can only identify different tasks, while providing no information about the relationships or similarities among them.
+We have provided additional experiments of TAD with access to continuous task parameters in Table 7, labeled as TAD-Continuous.
+Results in the table indicate that our method using the discrete implementation (TAD-CE/SC) achieves performance comparable to its continuous counterpart.
+Algorithms Hypothesis Half-Cheetah-Fwd-Back(1e7) Half-Cheetah-Vel(1e7) Humanoid-Direc-2D(1e6)
+Train&Test Train Test Train Test
+PlaNet $\Pi_{2}$ 30.5 $\pm$ 42.9 -198.1 $\pm$ 1.9 -202.1 $\pm$ 1.8 215.9 $\pm$ 72.3 220.6 $\pm$ 75.3
+Dreamer $\Pi_{2}$ 127.4 $\pm$ 181.8 -151.4 $\pm$ 0.4 -169.4 $\pm$ 1.2 260.5 $\pm$ 48.9 263.5 $\pm$ 52.3
+RL2(zero-shot) $\Pi_{3}$ 1070.7 $\pm$ 109.7 -70.3 $\pm$ 6.7 191.9 $\pm$ 50.8
+RL2(few-shot) $\Pi_{3}$ 1006.9 $\pm$ 26.4 -146.9 $\pm$ 0.4 268.8 $\pm$ 30.2
+MAML(few-shot) 429.3 $\pm$ 81.4 -121.0 $\pm$ 37.1 205.3 $\pm$ 34.7
+VariBAD(zero-shot) $\Pi_{3}$ 1177.5 $\pm$ 94.9 -58.4 $\pm$ 20.6 260.3 $\pm$ 61.6
+TAD-CE (Ours) $\Pi_{3}$ 1455.8 $\pm$ 78.3 -49.3 $\pm$ 1.9 -47.1 $\pm$ 0.3 339.5 $\pm$ 78.7 335.5 $\pm$ 70.5
+TAD-SC (Ours) $\Pi_{3}$ 1541.5 $\pm$ 114.8 -50.5 $\pm$ 1.6 -49.6 $\pm$ 1.6 260.2 $\pm$ 185.0 249.0 $\pm$ 168.9
+TAD-Continuous (Ours) $\Pi_{3}$ 1539.9 $\pm$ 102.0 -47.8 $\pm$ 1.1 -47.5 $\pm$ 1.3 302.9 $\pm$ 36.3 304.6 $\pm$ 39.0
+[…]
 Appendix D Experimental details for MuJoCo
-We here introduce state-based control tasks, including Half-CheetahFwd-Back, Half-Cheetah-Vel, and Humanoid-Direc-2D, in detail, following the setting of previous meta RL works (Finn, Abbeel, and Levine 2017; Rakelly et al.
-2019).
+We here introduce state-based control tasks, including Half-CheetahFwd-Back, Half-Cheetah-Vel, and Humanoid-Direc-2D, in detail, following the setting of previous meta RL works [27, 29].
   • Half-Cheetah-Fwd-Back. This task distribution includes two tasks: moving forward and moving backward.
@@
 All algorithms are trained on one NVIDIA GeForce RTX 2080 Ti.
-For each seed and each task setting, it will take around 1 day.
-We select 1 as the action repeat for all experiments following.
+Each seed and each task setting will take around 1 day.
+We select 1 as the action repeat for all following experiments.
 $\beta$ 0.2 0.15 0.1
@@
 TAD 951.9 $\pm$ 3.3 876.9 $\pm$ 51.1 927.6 $\pm$ 2.6 800.1 $\pm$ 121.4 608.5 $\pm$ 321.4 491.6 $\pm$ 389.4
-Table 6: Average cumulative reward (mean $\pm$ one std) over different target region (smaller $\beta$ represents smaller target region and more sparse return) of the best policy trained by Dreamer and TAD in Cheetah__\__Speed.
+Table 8: Average cumulative reward (mean $\pm$ one std) over different target region (smaller $\beta$ represents smaller target region and more sparse return) of the best policy trained by Dreamer and TAD in Cheetah_\_Speed.
 For each $\beta$, we train agents in the train tasks and evaluate them in both train and test environments.
@@
 Dreamer 250.2 $\pm$ 9.6 3.0 $\pm$ 2.2 237.4 $\pm$ 14.9 28.9 $\pm$ 14.6 168.7 $\pm$ 96.2 157.2 $\pm$ 188.4
-TAD 951.9 $\pm$ 3.3 876.9 $\pm$ 51.1 841.9 $\pm$ 154.9 546.1 $\pm$ 291.6 $7$$\pm$ 170.1 716.8 $\pm$ 136.5
-Table 7: Average cumulative reward (mean $\pm$ one std) over different sparse rates of the best policy trained by Dreamer and TAD in Cheetah__\__Speed.
+TAD 951.9 $\pm$ 3.3 876.9 $\pm$ 51.1 841.9 $\pm$ 154.9 546.1 $\pm$ 291.6 777.7 $\pm$ 170.1 716.8 $\pm$ 136.5
+Table 9: Average cumulative reward (mean $\pm$ one std) over different sparse rates of the best policy trained by Dreamer and TAD in Cheetah_\_Speed.
 For each sparse rate, we train agents in the train tasks and evaluate them in both train and test environments.
@@
 Appendix E Ablation study
-E.1 Ablation Study on Reward Signals
-In this part, we do ablation studies on reward signals for model-free methods including CURL and SAC+AE, which both belong to $\Pi_{1}$.
-We design CURL (w/ r) and SAC+AE (w/ r) by directly adding reward signals into the observation based on CURL and SAC+AE, respectively.
-As shown in Table.
-4, the performance of CURL (w/ r) and SAC+AE (w/ r) is similar to CURL and SAC+AE.
-The major reason is that CURL (w/ r) and SAC+AE (w/ r)only utilize the observation and reward of the current timestep without historical information.
-Thus it is still difficult for them to distinguish different tasks.
-E.2 Dynamic Generalization
+E.1 Dynamic generalization
 As TAD utilizes all historical information to infer the environment, it can be directly applied to more general settings with different observations, dynamics, and/or actions.
 To evaluate the performance of TAD in these settings, we have conducted the following four experiments based on DMControl:
-  • Acrobot-Cartpole-Pendulum: includes 7 tasks of artpole-balance, cartpole-balance__\__sparse, cartpole-swingup, cartpole-swingup__\__sparse, acrobot-swingup, acrobot-swingup__\__sparse, and pendulum-swingup.
+  • Acrobot-Cartpole-Pendulum: includes 7 tasks of artpole-balance, cartpole-balance_\_sparse, cartpole-swingup, cartpole-swingup_\_sparse, acrobot-swingup, acrobot-swingup_\_sparse, and pendulum-swingup.
 All these tasks aim to control a rod-shaped robot, while they own different embodiments, dynamics, and observations.
@@
 The tasks own different embodiments, dynamics, actions, and observations.
-  • Cheetah-run-mass($m$): This task distribution is based on the task Cheetah-run and the mass of the robot is $m$ times that of the standard task.
+  • Cheetah-run-mass($m$): This task distribution is based on the task Cheetah-run, and the mass of the robot is $m$ times that of the standard task.
 Thus different tasks own different dynamics.
 We train the agents in tasks with $m=0.6,1.0,1.2,1.6$ and test them in tasks with parameters $m=0.8,1.4$.
-  • Walker-walk-mass($m$): This task distribution is based on the task Walker-walk and the mass of the robot is $m$ times that of the standard task.
+  • Walker-walk-mass($m$): This task distribution is based on the task Walker-walk, and the mass of the robot is $m$ times that of the standard task.
 Thus different tasks own different dynamics.
 We train the agents in tasks with $m=0.6,1.0,1.2,1.6$ and test them in tasks with parameters $m=0.8,1.4$.
-Then we test Dreamer and TAD in these four settings and report the results.
+Then, we test Dreamer and TAD in these four settings and report the results.
 TAD can achieve much greater performance and better convergence compared to Dreamer, as it can better infer the current task.
 This experiment indicates TAD’s potential in further handling dynamic generalization and even cross-embodiment tasks.
-E.3 Sparse Reward
+E.2 Sparse reward
 In this part, we will evaluate TAD in more challenging settings with sparse rewards.
-First, we evaluate Dreamer and TAD in Cheetah__\__speed with different $\beta$, which identifies the region of target speeds.
+First, we evaluate Dreamer and TAD in Cheetah_\_speed with different $\beta$, which identifies the region of target speeds.
 With smaller $\beta$, the reward signals are more sparse since the target intervals are smaller.
@@
 For each $\beta$, we take the training parameters ($0.5,\beta$), ($1.5,\beta$), ($2.0,\beta$), ($3.0,\beta$) and test them in tasks with parameters ($1.0,\beta$), ($2.5,\beta$).
-As shown in Table 6, with the decreasing of $\beta$, the performance of Dreamer and TAD decreases since reward signals are sparse so exploration here is much more difficult.
+As shown in Table 8, with the decreasing of $\beta$, the performance of Dreamer and TAD decreases since reward signals are sparse, so exploration here is much more difficult.
 However, our TAD still significantly outperforms Dreamer and shows strong generalization abilities, which shows that TAD can effectively utilize historical information, even sparse rewards.
-Moreover, we design Cheetah__\__speed__\__sparse based on Cheetah__\__speed.
-In Cheetah__\__speed__\__sparse($n$), we make the reward function sparse, i.e., the output reward is the same as Cheetah__\__speed every $n$ timesteps (in step $n-1,2n-1,...$) and 0 otherwise, of which the sparse rate (SR) is $(n-1)/n$.
+Moreover, we design Cheetah_\_speed_\_sparse based on Cheetah_\_speed.
+In Cheetah_\_speed_\_sparse($n$), we make the reward function sparse, i.e., the output reward is the same as Cheetah_\_speed every $n$ timesteps (in step n1,2n1,n-1,2n-1,...) and 0 otherwise, of which the sparse rate (SR) is (n1)/n(n-1)/n.
 We supplement experiments to evaluate the performance of Dreamer and TAD with $n=5$ (SR=0.8) and $n=10$ (SR=0.9).
 As shown in Table 9, with the increasing of SR, although the performance of TAD decreases since inferring task context from sparse reward is extremely difficult, TAD still shows strong performance in train tasks and generalizes well to unseen test tasks.
-Appendix F Ethics Issues and Broader Impact
-Designing agents that can generalize to unseen tasks is a major concern in reinforcement learning.
-This work focuses on task generalization in reinforcement learning and proposes a novel algorithm Reward Informed Dreamer.
-One of the potential negative impacts is that algorithms using deep neural networks, which lack interoperability and theoretical guarantee.
-If we hope to apply them in real-world applications, they may face security and robustness issues, and a possible way is to develop more explainable methods.
-There are no serious ethical issues as this is basic research.
-We hope our work can inspire more research on designing agents with stronger generalization abilities.